A Complete Bibliography of Lecture Notes in Mathematics (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.05

Title word cross-reference

(f, g) [NS12l, NS12x]. (g, f) [AK11b]. (H, G) [Bou10n]. (j = i − −2, i + 1case) [KKP16h]. (j = i − 1, icase) [KKP16i]. (k, n − k) [Isa11i]. (L, H) [FO14b]. (M, M) [DG14g]. (R^{D}, μ) [YYH13c]. (X , ν) [YYH13d]. (n − 1, 1) [Isa11c]. (n − 2, 2) [Isa11d]. (S^{n}, S^{n−1}) [KKP16p]. 0 [DG14c]. 1 [AK11h, Bou14b, DG14c, Kha10a, Zua13b]. 1 − D [Kha10p]. 1/4 [AH11o]. 1/H [Pra11a, Pra11b]. 1/J_{f} [HK14g]. 11 [DG14r]. 2 [DG14c, Kha10c, Kha10w, Kli14b, Sam14h]. 2 × 2 [Par10g]. 3 [DG14r, FKP13c, HKMR12c, NS12s, SGS15]. [a, b] [AK11a, AK11b]. A1 [Mor12a]. α [Ale12a, Has13]. B_{p,q}^{s,τ} (R^{n}) [YSY10k]. BES(3) [YY13k]. C∗ [BSG11]. C^{1} [ACT11a, ACT11b]. e [Rou14c, Rou14f]. E_{1} [DG14o, DG14p]. E_{2} [DG14i]. ℓ_{2} [Dod10a]. Ext(F) [JPT16m]. F [KKP16g, KKP16n, KKP16i, KKP16h, AK11j, NS12w]. f = f′(x, y) [NS12e]. f^{∗} [SAC^{+}14m]. F_{p,q}^{s,τ} (R^{n}) [YSY10k]. G [Bou10n, GQY14]. G = O(n + 1, 1) [KKP16q]. Γ [Bra14i]. γένoς [PP16a]. h [Bar11h]. H^{1} [Sti13d, YYH13h].

H^{1}(X , ν) [YYH13i]. H^{1}(μ) [YYH13j, YYH13l]. H_{1}(∂F) [NS12a, NS12r, NS12v]. H_{1}(∂F \ V g) [NS12v]. H_{n}^{(}E_{FIN} (); KZ^{−∞}) [FO14h]. I [FKP13i, Lau11]. J [Ves11f, Ves11s, Ves11t]. J(X; L, d) [Rei13a, Rei13i]. J_{f} [HK14g]. K [BCM^{+}11b, BSG11, Cor11, FO14a, FO14k, Mor12k, Sch11a, HMM^{+}13h]. KK [Mey11]. k ≤ (n − 2) [Isa11i]. L^{1} [BDR16a]. L^{2} [Ros11]. L^{∞} [Cap14]. L_{2} [Maj13u]. L_{p} [Def11b]. L_{q} [GM14]. M [CM11, DMS14e, DMS14g, DMS14h, GM14]. M(T) [Que10b]. A^{1} [Mor12j, Mor12b, Mor12c, Mor12d, Mor12l]. C^{n} [GZ12]. CR_{C} [Bou10m]. R^{4} [Frö12e]. R^{d} [BDR16g]. R^{n} [Par10g]. R^{n+2} [Frö12f]. G(F_{q}[t, t^{−1}]) [Wit14d]. G(F_{q}[t]) [Wit14c]. H̃_{−•} [Rei13n]. D [Sab13q]. F [AH11a]. F′ [Rei13b]. G_{−} [Rei13h, Rei13m, Rei13c]. H_{F} [JPT16c]. H_{K} [JPT16c]. I [Sab13a]. VC_{∞} [FO14d]. W [AH11b]. D_{l}^{μ} [KKP16k]. Γ_{C} [NS12m, NS12l]. RBMO(X , ν) [YYH13i]. M ≥ 2 [DG14c]. n [Dea14, YY13i, YY13q]. N(λ) [Par10k, Par10l]. O(n + 1, 1) [KKP16n]. p [Bou10o, Bou10p, Lew12, Sam14h, Sti13q]. p^{(0)} [Nås11c, Nås11l]. p^{(1)} [Nås11d, Nås11m]. P^{♯} [NS12b]. P_{j}^{♯} [NS12b]. ∂F [NS12b]. π_{1}^{A^{1}}(P^{n}) [Mor12b]. π_{1}^{A^{1}}(SL_{n}) [Mor12b]. Ψ [Nås11a]. ψ_{2} [GPV12]. q [AM12g, AM12h, AM12b, AM12c, AM12i, AM12j, AM12e, AM12l, AM12k, Nås11b]. Q_{(α,β)}^{w(x,D)} [Par10i]. S [AGJ^{+}13p, LE11e, LE11k]. sl_{2} [Rei13b, Rei13c]. SL_{2}(R) [Bou12b]. t = 1 [YY13o]. T_{π} [Rei13o]. ×Z [BM14]. U [Maj13c, Maj13i, Maj13l, Maj13r]. W [YY13i].

-adic [Sti13q]. -Algebraic [Mor12a]. -algebras [BSG11]. -Balls [FKP13c]. -Based [BDR16a]. -Biset [Bou10o]. -Bisets [Bou10n]. -Brownian [GQY14]. -Bundles [FKP13i]. -Centroid [GM14]. -Colorings [Kli14b]. -Convergence [Bra14i]. -Cosine [Ale12a]. -cosiness [Lau11]. -Coverings [Mor12b]. -Curves [CM11]. -D [Bou14b, Kha10c, Kha10a, Kha10w, Zua13b]. -Dense [Maj13u]. -Difference [AM12b, AM12i, AM12g]. -Estimates [GM14]. -Extremal [Cap14]. -Factorizations [AK11a]. -Factors [AK11b, AK11h, AK11j]. -Filtrations [Sab13a]. -Fractional [AM12h, AM12e, AM12l]. -Functional [AH11a, AH11b]. -Functions [Rou14c, Rou14f]. -Groups [Bou10p]. -Hermitian [Ves11s, Ves11t]. -Homology [Mor12d]. -Homotopy [Mor12c, Mor12d]. -Integral [AM12i]. -Invariant [Mor12j, Mor12l]. -Lefschetz [HMM^{+}13h]. -Leibniz [AM12c]. -Manifolds [HKMR12c, NS12s]. -Matrices [DMS14e, DMS14g, DMS14h]. -method [KKP16g, KKP16n]. -Mittag-Leffler [AM12j]. -Modules [Sab13q]. -Norm [GPV12]. -Numbers [LE11e, LE11k]. -Orthogonal [Ves11f]. -Pinching [AH11o]. -Rank [Sam14h]. -Regular [Bar11h]. -Sasakian [Dea14]. -Sets [Bou10n]. -Spaces [Def11b]. -Spherical [Isa11i, Isa11c, Isa11d]. -Statistics [Maj13c, Maj13i, Maj13l, Maj13r]. -Structures [Rei13b, Rei13c]. -Sturm [AM12k]. -system [KKP16i, KKP16h]. -Theories [Mor12k]. -Theory [BCM^{+}11b, BSG11, Cor11, FO14k, Mey11, Sch11a, FO14a]. -Type [DG14g].

-Vanishing [AGJ^{+}13p]. -Variation [Pra11a, Pra11b].

1-7 [DS10a]. 10-15 [CTSDV10a].

2007 [CTSDV10a, DS10a]. 2008 [ABB^{+}10b, BCMV11b, BMOR13b, Fas11, Kol11l, Le 11j]. 2009 [AT11i, ABH^{+}13b, BNQ^{+}12b, Eth11h, Kum15c, LLMS12b]. 2010 [ABCC^{+}12, Chu10, CDG^{+}13b, CCG^{+}11b, FFV11, Fla11f, Gia11d, KMM12, Kum14j, RSSS12b, Tia16c]. 2011 [ABIL13b, Ben13b, BDG^{+}14a, CKM^{+}13b, PBBD13b]. 2012 [CGGM14b, EH16b]. 2013 [BCG^{+}13b, Bur14c, Led14a]. 2015 [vdG16h].

Abel [PP16c, PP16b, PP16p, PP16r]. abelian [Sam14p, Sti13d, Sti13i, HY12i, Sam14a]. Abelianity [Sab13b]. Absolutely [Bou12b]. Abstract [DGH16a, DGH16p]. Accessibly [Her13g]. According [PP16a]. Acoustic [Amm12b, ABJW12]. Acting [Wit14e]. Actions [Bar11g, FO14f, Sea14]. Adapted [Lin10w, Lin10x, Lin10y]. Adapting [Lor10a]. Adaptive [Ber12, CGC15, NV12, Sie12b]. Adaptivity [BNQ^{+}12b]. Addition [PP16h]. Additive [Fla11g, RS10b]. Adic [Bos14a, Sti13q]. Adjoint [Par10g]. Adjoints [Neu10p]. Admissible [Sch14]. ADMs [LR12i]. Advanced [Bos14k, Die10l, SGS15]. Advances [AB12]. Affine [CT13, Isa11i, Mor12e, Sab13o, Sch10j]. Affinoid [Bos14b, Bos14c]. After [Aud11a, Bes16, DG14o, Ota15c, Sch11a]. Aggregation [OX13]. Albin [BDMY11b]. Alexander [Kum14a]. Algcurves [DP11]. Algebra [BF12b, BF12n, BF12o, Bro10h, HY12f, Koe14b, MT14, Que10b, Sch10o]. Algebraic [AT11a, AH11c, Bar11a, BCM^{+}11b, BF12k, Cor11, DP11, Dem13, Duc15a, FO14a, FO14k, FK11b, Mor12a, Mor12c, NS12a, Sch11a, Viv14b]. Algebras [BZ14, BF12c, BF12l, Bos14b, Bos14m, Bou10g, Con14, GW11c, Jak11, MV12b, Rei13m, BSG11]. Algorithm [NS12c, NS12p, NS12u, SAC^{+}14m]. Algorithmic [Ban14, Ceg13a, Ceg13b, Rob14c]. Algorithms [AGJ^{+}13v, DP11, DGQ12, Ken15c, Kli14b, NPS13, Wai15]. Almost [Ang14c, Ang14g, HR12, Sta12a, Sta12b, Sta12f, Sta12g, YSY10a]. Almost- [Ang14g]. Almost-Complex [Ang14c]. Along [Bra14k, HR12, Sab13r]. Alperin [Sam14j]. Alpha [LR12h]. Alternate [Neu10a]. Alternative [SS13b]. Amalgamated [Dod10b]. Amenable [Ben13c]. AMLE [Cap14]. Ampére [Mon14b, Bło13, Bou12a, Gue12a, PS13b]. Anabelian [Sti13f]. Analogues [Sti13e]. Analogy [Eis12]. Analysing [PAMC13]. Analysis [ABIL13b, BBB12, Bur14c, Die10a, EH16b, Frö12b, JR15, KMM12, Koe14b, LR12b, Led14a, Lit13b, Lor10h, MT14, MS16, MSU11h, Men13b, OZLK13f, OX13, Pos12j, Que10o, RS10b, Růž13a, Růž13b, SAC^{+}14i, SD16, SZ15, Tud11, Wai13c, Yse10g, dVBSV16, JPT16g]. Analytic [Duc15b, Lam15, MM13f, Nis14c, Pos12c, Pos12d, RTW15a, Rob14b, Sab13h, Tem15]. Analytical [Lin10a, Lin10b, Lin10c, Lin10d]. Analyticity [MSU11i].

Anderson [Bou14b, Bou14d]. Andrews [Bar11b]. Andronov [Lie15j, Lie15t]. Angeles [FFV11]. Angles [KO13b]. Angular [Yse10j]. Animal [PAMC13, Rey13a]. Anisotropic [HP13]. Anomalies [AGJK12]. Anomalous [AGJ^{+}13a]. Apnea [Bru13]. Apparent [DGH16b]. Appendix [BFMV14a, BCGG10a, KKP16a, SSM14a]. Application [ABMT13, DÉ13, Die10a, KKP16b, KS13, Lie15a, Lie15b, Lie15c, Lie15d, RL15]. Application-Oriented [Die10a]. Applications [ABIL13b, AT11g, AM12c, AH12a, BBD13c, Ban15c, Ban15d, Bar13, Bar11a, BBK13c, BNQ^{+}12b, BCMV11b, BG11, BGdC13, BY15, BL15b, BL15a, Bou10a, Bur14c, CM10a, CM10b, CKKR12b, CMN13b, Dem13, DGW14b, DRiSS10b, FK11a, FS13a, FKP13a, GLL11, JPT16g, Le 12, Lew12, Maj14e, NO13, OZLK13g, Ota15b, Pag13, RS10b, Riv12, Rud15, Sab13c, SG13, Sea14, SSM14i, Sie11, Sta12c, Tia16c, YYH13h]. Applying [Lam15]. Approach [BDR16a, BDR16i, BK11, Del12b, DH12, Die10c, Her13c, Kha10u, Kru14d, Neu10a, SS13b, YY13b]. Approaches [TP13]. Approaching [AGB15a]. Approximability [Yse10k]. Approximate [LR12b, LR12c]. Approximating [Sch15b]. Approximation [AGJ^{+}13d, AH12b, AN11, Bak13b, BSW13a, BSW13g, BCM15g, EH16b, Has13, KK14a, Kru14i, Nås11b, Nås11d, Nås11c, Nås11a, Nås11e, Ves11i, Ves11j, Voj10b]. Approximations [ABIL13b, AH12i, BV11, Haj12, MMN13g, MMN13k, MMN13n, MMN13o, Nås11o, Nås11n, Nås11r, Nou13b, YYH13a, YYH13k]. Arbitrary [Kok11b]. Arcs [FKP13d]. Arcsine [YY13m]. Argument [AH11i, Maj13t]. Arising [Que10e, VV13]. Aristotle [PP16a]. Arithmetic [FO14b, HS14a, Koe14b, MT14, SAC^{+}14e, Sti13u, Sti13y, Wit14e, CTSDV10a]. Arrays [MM13e]. Articles [BPR12a, Chu10]. Aspects [Ang14b, Ban14, Kha10n, KMM12, Led14a, Lie13]. Assembly [FO14d, Mey11]. Asset [Her13c]. Associated [BZ14, Bos14b, Bos14a, DGQ12, Her13k, Kum14k, NS12f, Pos12f, Que10f, RY14]. Associative [BZ14]. Associativity [BF12d]. Asymptotic [AN15a, BL15c, BL15d, BL15e, Cal13, DHI13a, DHI13b, FS13d, FS13e, FS13c, Hei13, Lie13, MS12, MMN13n, MMN13o, MMN13p, Pad11a, Sch10c, Spo13, Tud11]. Asymptotically [Hu11a, vdG16a]. Asymptotics [BL15f, Fra14, MMN13l]. Atomic [YYH13l, YSY10a]. Attenuating [ABJW12]. Attenuation [KS12]. Attractors [Mar11b]. Aubin [Lor10c]. Automata [Que10c, Que10d, Que10k, Que10l]. Automorphism [Lau13]. Automorphisms [Bed10, BCGG10h]. Autonomous [CKM^{+}13b, Maw13, NO13]. autres [MM13a]. Average [BGT14]. Aw [HJ13]. Axial [Kha10g, Kha10h].

B [HA13, KKP16p]. B.G. [Mor12e]. Back [AT11b, AK11c, AGB15b, AGJ^{+}13b, AN15b, AH11d, Ang14a, AM12a, AH12c, Aud11b, BDMY11a, Bak13a, Ban10a, Ban15b, BDR16b, Bar10a, Bar11c, Bec14c, Ben13a, BK12a, Toë11a, BFMV14b, BL15g, BF12a, Bos14d, BSW13b, Bou10b, Boy15a, Bra14a, Bro10a, BCGG10b, Bur14a, BCM15a, CM10c, CM10d, Ceg13c, CMN13a, DGH16c, DG14a, DHI13c, Def11a, DMS14a, DHHR11a, Die10b, Dod10c, DGW14a, EZ12a, Eth11a, EH16a, FFM11a, FO14c, FM12a, FV11a, Fla11a, Frö12a, FS13b, FKP13b, GOZ12a, GGS10a, Gia11a, GW11a, Gue13a, GZ14a, HMM^{+}13a, HK14a, Her13b, HPRY11a, HKMR12a, HY12a, Hu11b, Isa11e, JPT16a, Kaw13a, Ken15a, Kha10b, KK14b, KKP16c, Koe14a, Kok11a, Kol11a, Kor13a, Kru14a, Kum14b, Kum15a, KO13a, KKR12a, LE11c, LR12d, Le 11a, Léo12a, Lie15e, Lin10e, Lin10f, Lit13a, Lor10b, Maj13a, Maj14a, MV12a, Mas15a, MMA11a, MSU11a]. Back [MMN13a, Men13a, MM13b, Mon14a, MFR13, Mor12f, Nås11f, Nis14a, OZLK13a, Ota15a, Otw12a, Pad11b, Par10a, PS13a, PEP13a, PT10a, Pos12a, Pöt10a, Que10a, RS10a, Rei13d, Rob14a, Rou14a, Sab13d, SAC^{+}14a, Sal12a, Sam14b, Sch10a, Sch10d, Ser13a, Ser11a, SSM14b, Sie12a, Soi13a, Spe12a, Sta12d, Sti13a, SS13a, Tan11a, TBBF13a, Tia16a, Tuc13a, Viv14a, Voj10a, Wei14a, Wei14d, Wit14a, YYH13b, YY13a, Yse10a, YSY10b, Zua13a, vdG16b]. Backbone [KPR14]. Background [AH11e, BL15k, BF12b, GW11b, Nås11p, OZLK13b, Wei14e]. Backgrounds [AN15d]. Backward [QY12]. Baire [Dod10a]. Baker [BF12o]. Balance [Lie15d]. Balanced [BFMV14a]. Ball [KM12a]. Balls [FKP13c]. Banach [BHJKX12, Bai14, Dod10d, Dod10i, Dod10h, Que10b]. Band [Spe12b]. Baroreflex [Bro13, TBBF13b]. Barotropic [Pad11c]. Barrier [Pag13]. Based [BDR16a, BMOR13b, MMN13n]. Basic [AK11d, Bar11d, BFMV14c, DG14b, Die10m, Dod10e, Gia11d, Kaw13b, Kha10c, KKP16d, LE11d, Ser11b, Sti13b, Wit14b]. Basics [HMM^{+}13b, Yse10b]. Beginner [Van13a]. Behave [Kum14a]. Behavior [BFMV14c, Ish13, MS12, Par10k, Par10l]. Behaviour [AN15a, DG14k, FE13, Lin10a, Lin10b, Lin10c, Lin10d]. Beijing [Chu10]. Bellman [Lit13b]. Below [Sea14]. Bergman [BK12b, Bło14]. Berkovich [Fav15a, Jon15, Ota15b, Tem15]. Bernd [Viv14b]. Bernoulli [Bou14d, PP16l]. Besov [YSY10e, YSY10h]. Bessel [Bar10c, Bar10d, Bar10e]. Between [BF12l, Eis12, Kum15b]. Beyond [Lor10h, Rey13a, Aud11c]. Bianchi [Lie15a]. Bias [vdG16c]. Bicyclic [Sam14c]. Bifurcation [Ber13, BCD13, CKM^{+}13b, Lie15c, Lie15f, Lie15g, Lie15l, Lie15j, Lie15k, Lie15t, Lie15v, Lie15w]. Bijective [Que10c]. Bilinear [CM10e, CM10f, Kaw13j]. Bioinvasion [MJB13]. Biological [Lin13, OX13]. Biology [Bob15, DS13, Rud15]. Biomathematical [BBD13c]. Biomedical [Amm12b, SS13b]. Birational [PP16t, Sti13e]. Biset [Bou10d, Bou10c, Bou10o]. Bisets [Bou10n]. Black [ADFT11, Mar11b, SSM14d, Her13d]. Blocks [Sam14e, Sam14d]. Blow [Fla11d, Hu11a, Hu11c, Hu11d, Hu11e, Sab13n]. Blow-Up [Fla11d, Hu11a, Hu11c, Hu11e, Sab13n, Hu11d]. Blowing [MMA11b]. BM [YY13k]. Bochner [BC14]. Bodies [GPV12, GM14, GL12, Gru14, Kol14, MMN13j, NT12]. Body [BCM15f, BCM15e, Gal13, RSSS12b]. Bogdanov [Lie15c, Lie15g]. Böhm [AH11g]. Book [MMA11j, SSM14c]. Boolean [GW11c, Jeu15]. Borel [Cha14, Dod10h, Éme13]. Bose [Sei12]. Bottles [HKMR12e]. Bound [Bło14, Sam14f]. Boundaries [Pad11e]. Boundary [Bob15, BCD13, Bou12a, Bur14i, GGS10h, Jac13a, Lew12, MMN13f, MMN13h, MMN13i, MMN13m, MMN13l, MM13f, MM13h, NS12q, Neu10b, Pad11c, Pad11f, Pos12c]. Bounded [Dod10i, GGS10h, Sea14]. Boundedness [YYH13c, YYH13d]. Bounding [EL14, Kol11b]. Bounds [CS14, Hob11, Kaw13c, Kaw13d, Kol11e, Kol11f, Sch10n, Yse10c, vdG16n]. Bourgain [Bło14, Dod10f, Naz12]. Braid [Bro10e]. Branching [BH14, DR13, HR12]. Brascamp [Led14b]. Brauer [Sti13c]. Breaking [KKP16f, KKP16e, KKP16q, KKP16i, KKP16h]. Breathing [TP13]. Breeding [BH14]. Brendle [Bes16]. Bridge [Ort13, RY14]. Bridging [Kum15b]. Brief [Wai13a]. Brouwer [vdG16d]. Brownian [BDMY11b, BÉL14, Bur14c, Bur14b, Bur14k, DH12, GQY14, Haj12, HR12, Mar12, Ort13, Pic11b, Pra11a, Pra11b, RY14, Wu12, YY13b, YY13l, YY13o]. Bruhat [RTW15a]. Brún [DFO13]. BSDEs [MW12]. Bubbles [Pro13b]. Buildings [RTW15a, Wit14e]. Bundle [Frö12b]. Bundles [BFMV14a, CLR13, FKP13i, Mor12c, Pöt10d]. Burnside [Bou10e].

C [HA13]. C.I.M.E [ABB^{+}10b, BNQ^{+}12b, BCMV11b, CTSDV10a, DS10a, Fas11]. CAD [SD16]. CAGD [MS16]. Calabi [Bło12a]. Calculation [DGW14b]. Calculations [HY12g, Wei14f]. Calculi [AM12e, AM12l]. Calculus [AM12h, BT13, CLR13, Her13g, Her13m, INS12, Kru14d, Kru14g, Nou13b, Par10e, Par10j, Par10o, Ros11]. Calderón [MM13j]. Calibration [BR15]. Campanato [YSY10h]. Campbell [BF12o, Rou14c]. Canard [WMR13a]. Canards [SSM14d]. Canonical [Mor12h]. Capacities [DHHR11b]. Capturing [DGH16q]. Caputo [Die10a, Die10c, Die10j, Die10m, Die10l]. Car [Hel13]. Car-Following [Hel13]. Cardiovascular [BBK13c, ONO13]. careful [MM13a]. Caricature [DG14l]. Carleman [Le 12]. Carlo [Koe14b, DP14, SS14]. Cartan [Sam14g]. Case [BZ14, DG14d, FO14a, Hu11h, JPT16b, Kha10d, Le 11b, MM13j, Neu10e, Neu10f, Neu10v, Sab13p, Sab13o, Sch10j, Sch10k, Sti13y, Ves11r, WL13]. Cases [Die10l, SSM14l]. Catalytic [DR13]. Cataproducts [Sch10e]. Categories [Toë11b]. Cauchy [Lor10h, Nis14c, PP16d]. Caustic [Otw12g]. CBHD [BF12d, BF12j, BF12l, BF12m]. CDO [CJL11]. Cellular [RL15]. Central [BS13a, Gru14, WL13, Maj14f]. Central-Place [WL13]. Centred [vdG16k]. Centroid [GM14]. Cerf [HKMR12h]. Certain [Bou14d]. Cervonenkis [Maj13u]. Cetraro [ABB^{+}10b, ABIL13b, ABCC^{+}12, ABH^{+}13b, BNQ^{+}12b, BDG^{+}14a, BCMV11b, BMOR13b, CKM^{+}13b, CGGM14b, CTSDV10a, CDG^{+}13b, DS10a, EH16b, LLMS12b, PBBD13b, RSSS12b, Koe14b, Tia16c]. Chafee [DHI13e, DHI13j]. Chain [SS14]. Chaining [vdG16e]. Chains [BB14a, CCPP12, Kum14k]. Challenge [AT11a]. Challenges [BBKK13, BDG^{+}14a, FFM11b, ONO13]. Channel [Lie15c]. Chaos [Leh11, Mar12]. Characterising [BL13]. Characteristic [Bro10c, NS12b, NS12t, Sch10c, Sch10j, Sch10k, Sch10l]. Characteristics [Nis14f]. Characterizations [YSY10j]. Characters [BB14a, Sam14e, Bec14d]. Charge [Mar11b]. Chemical [SSM14g]. Chevalley [FG11]. Chiral [CFL11]. Choice [GMK13, PP16e]. Chronic [HA13]. Cibrario [Otw12d]. Cinquini [Otw12d]. Class [AN15f, Gia11f, Jak11, Maj13j]. Classes [BOGP12, BL15b, BL15a, BCGG10e, Dod10i, Isa11i, Maj13u, Sti13f, SS13b]. Classical [BBB12, Bur14j, CS14, Def11b, DHHR11c, Kha10e, MSU11b, Ser11b]. Classification [DGH16a, DGH16m, DGH16n, DGH16s, FO14b, HTKS13, PP16m, PP16o, Sam14t]. Classifications [DGH16d]. Clebsch [PP16e]. Clifford [BZ14, PP16f]. Climate [ZK13]. Climate-Driven [ZK13]. CLL [GW11d]. Cloaking [AGJ^{+}13a, AGJ^{+}13p]. Closedness [CS11]. Closure [Sch10j, Sch10k, Sch10l]. CLT [Ros11]. Cluster [Ben13p]. Clustering [BY15]. CM [HY12d]. Co [Pag13]. Co-monotony [Pag13]. Coagulation [Lau15]. Coalescing [DPPS14]. Coarse [Gia11b, Ben13b]. CoCoA [BD13b]. Codimension [Lie15h]. Codimension-One [Lie15h]. Coefficient [Mey11]. Coefficients [Has13, Nis14c, Sti13d]. Coherent [Bos14e]. Cohomological [Ang14b, Duc15a, Sti13h]. Cohomology [ÀFR13, Ang14c, Ang14d, Ang14e, Bec14e, Duc15b, HMM^{+}13c, Mon13]. Cokernels [FO14d]. Cold [Otw12b, Sei12]. Collapsing [NS12c]. Collected [Chu10]. Collective [FE13]. Collocated [AN15e]. Colored [FKP13h]. Colorings [Kli14b]. Combinatorial [Ban14, BB13, BFMV14d, Viv14b]. Combinatorics [DFO13]. Comments [HK14c, Maj14c, Nås11g]. Communities [Bri13]. Commutative [Bro10h, Def11b, Def11c, Par10m, Sch10o]. Compact [Bob11b, BNR11, BCGG10g, Kok11b, Que10d]. Compactification [Ota15c]. Compactifications [BFMV14e]. Compactness [AH11f, Lau15]. Comparing [JPT16c]. Comparison [BY15, Neu10c]. Compatible [EH16c]. Competitive [MJB13]. Complementability [BÉL14]. Complements [Bro10h]. Complete [AH11o, HMM^{+}13d]. Completion [Maj13b]. Complex [Ang14c, Ang14b, Ang14d, Ang14g, Bło13, Bou12a, Bro10e, CGGM14b, Con13, DS10b, Est15, Gue12a, Kum15c, Mor12j, Mor13b, Nor13, Rab11, Sch15b, Wil14]. Complex-Valued [Nor13]. Complexes [Bar11h, Sab13n]. Complexity [AT11i, Yse10c]. Component [AK11e, CM10p, CM10q]. Components [BFMV14h]. Composite [FS13d, FS13e, FS13c, FS13f]. Compressible [Pad11a]. Computation [BORM14, CM10e, CM10f, CM10g, CM10h]. Computational [BK11]. Computations [BGdC13]. Computing [DP11, DFO13, Sch11b, Soi13b]. Concave [CG14, EL14, KM12a, Kla14, Lat12]. Concentration [vdG16e, vdG16u]. Concepts [DG14l, Dod10e, KS13, Lie15s]. Concluding [Nås11g]. Concrete [AN15a]. Condensate [GW11d]. Condensation [Sei12]. Condition [DHHR11l, Léo12b, MSU11g, MW12, vdG16o].

Conditioned [Kor13b, Ser14]. Conditioning [BL15i, BL15h, YY13j]. Conditions [Bob15, Lor10g, MMN13l, Neu10b, Neu10i, Neu10j, Nis14e, Pro13a]. Conductance [Kum14i]. Cone [AH11g]. Cones [NS11]. conference [Chu10]. Confidence [vdG16f]. Configurations [DGH16d, Rei13e]. Conformal [KKP16d, KKP16e, MSU11e]. Congested [HTKS13]. Congruence [GW11e, GW11i]. Congruence-Permutable [GW11e]. Congruence-Preserving [GW11e]. Conjecture [Bar11b, Bar11g, Boi16, DÉ13, Ken14, KM14b, Kum14a, Sti13p, Sti13u, Sti13t, Sti13s, Sti13y, HKMR12d]. Conjectures [AGB15a, AGB15c, AGB15f, Sam14q, Sch10c]. Conjugacy [BCGG10e, MMA11c]. Connections [Bil14, Sab13i, Sab13h, Sab13p]. Consequences [Maj14i]. Conservation [Ban15d, Bre13]. Consistent [Hel13]. Consistently [Her13g]. Constrained [CWZ11, CS11]. Constructing [HPRY11b]. Construction [AH11g, BF12c, BSW13c, BSW13g, Dod10f, FG11, KKP16f, Maj14k, PP16l]. Constructions [Rei13l]. Containing [HKMR12e]. Context [CCPP12]. Continuity [HK14b, Lor10a, Lor10g, Ros11]. Continuous [AAFM12a, BDMY11b, Bou12b, DPPS14, JPT16f, Neu10d, Neu10e, Neu10f, Sti13d]. Continuum [AGJ^{+}13d]. Contour [Kor13b]. Contours [DGH16b]. Contraction [vdG16u]. Control [Bou14c, Bro12, Duf13, EZ12b, Gla12, HT13b, HM13, Kaw13b, Kaw13i, Le 12, MJB13, TBBF13b, TP13, Zua13b, ABCC^{+}12]. Controllability [Kaw13c, Kha10e, Kha10f, Kha10g, Kha10h, Kha10n, Kha10o, Kha10p, Kha10s, Kha10u, Kha10t]. Controls [Kha10f, Kha10a]. Controversy [HJ13, Aud11g]. Convection [Lin10g, Lin10h, Lin10i, Lin10j, Lin10k, Lin10l, Lin10y]. Convection-Diffusion [Lin10g, Lin10h, Lin10k, Lin10l]. Convergence [BV11, BF12d, Bra14b, Bra14i, Ceg13d, CWZ11, Gue13b, Kru14d, PS11, Pos12e, Yse10c]. Convex [AAFM12b, GPV12, GM14, GL12, Kol14, Kol11c, NT12, vdG16p, Gru14]. Convexity [CEK12, Neu10g]. Convolution [JPT16d, Rou14b]. Correspondence [Sab13c, Sab13h, Sab13q, Sab13p]. Cortical [Tuc13b]. Cosine [Ale12a]. cosiness [Lau11]. Cosmological [Lie15a]. Costs [GMK13, Sch14]. Cosymmetries [Lie15i]. Coulomb [Frö12b, Frö12e, Frö12f, Sol12]. Countability [Fav15a]. Counting [BCGG10c, Par10k, Par10l]. Coupled [CPSS13]. Coupling [Bur14l, Ken15b]. Coupling-from-the-Past [Ken15b]. Couplings [Bur14m]. Covariance [KKP16d]. Cover [Van13a]. Coverings [Mor12b, NS12d]. Covers [Sti13y]. Cox [VGS13]. CR [Isa11h]. CR-Hypersurfaces [Isa11h]. Cramér [Tud11]. Criterion [BFMV14f, Bra14h, Lau13]. Critical [Ben13c, Gia11c, Gia11d]. Crossings [Sab13r]. Crystallographic [FO14a, FO14j]. Cubature [CMN13b]. Culler [Ota15c]. Current [BDG^{+}14a]. Currents [Ber13]. Curtis [Bar11b]. Curvature [AH11c, AH11h, AH11m, FV11b, GG14, Gua14, Sea14, Wei14b, Zil14]. Curve [NS12e, NS12f]. Curves [BFMV14j, BFMV14q, Bru10, CM11, DGH16r, DGH16u, DP11, FK11b, Men13b, PP16g, PP16l, PP16o, PP16w, Sab13q, Wei14f, Wei14i, Wei14k, Wei14m, Wei14n, Wei14o]. Cuspidal [Sti13e]. Cutting [PP16u]. Cycle [DG14i, DG14j, Sti13f]. Cycles [HY12d, Sab13k, Sab13l]. Cyclic [NS12d]. Cylinders [NS12e]. Cylindrical [Rie11].

D [Bou14b, Kha10c, Kha10a, Kha10w, SGS15, Zua13b]. Dade [Bou10f]. Daganzo [HJ13]. Damped [Ves11a]. Damping [Kha10g, Ves11q]. Data [PAMC13, SZ15]. Decay [BGT14, Yse10f, Yse10l]. Decisions [vH14]. Decomposition [Boi16, BÉL14, FKP13c, KPR14, Ves11s, YY13o, Yse10j]. Decompositions [BD13b, DPPS14, Le 11c, YSY10a]. Deconvolution [LR12b, LR12c, LR12g, LR12h]. Decoupling [Lie15b]. Defect [Sam14a, Sam14h, Sam14o, Sam14p, Sam14u]. Defining [Isa11g]. Definite [JPT16f, JPT16g, Ves11b]. Definition [Her13e, NS12l]. Definitions [Sam14i, Wit14b]. Deformations [AGJ^{+}13q]. Degenerate [FM12b, FM12c, FM12d, FM12e, FM12g, Lie15j, Lie15k, Maj13c]. Degeneration [MMA11i]. degenerations [CT10]. dégénérescences [CT10]. Delay [AN15h, AN15g, HT13b]. Delocalized [Mou12]. Dense [Maj13u]. Density [DHHR11d]. Dependency [Ben13q]. Dependent [AGJ^{+}13e, BS13a, Bur14k, Kha10a, KPR14, Neu10a]. Depending [DPPS14]. Depression [Tuc13b]. Depth [Her13n]. Derivation [Hel13]. Derivative [KM12b, Mar12, vdG16k]. Derivatives [Yse10f]. Descartes [PP16g]. Descent [Neu10e, Neu10f, Sti13c]. Description [CM10k, CM10r, CM10s]. Descriptions [YY13q]. Descriptive [Dod10d]. Design [Sie12b]. Designs [Soi13b]. Destruction [Bri13]. d’Été [AT11i, Ben13b, Bur14c, Eth11h, Fla11f, Gia11d, Kol11l, Kum14j, Le 11j, vdG16h]. Detection [AGJ^{+}13c]. Deterministic [Kaw13i, Thi13]. Development [ONO13]. Developments [Rey13a]. Deviation [DHI13i]. Deviations [BBB12, Wai13a, AFV12]. DG [Toë11b]. DG-Categories [Toë11b]. Diagonal [YSY10a]. Diagonalisation [Ves11q]. Diagram [Maj13c, Maj13l, Maj14i]. Diagrams [FKP13d, GW11g]. Diffeomorphisms [HKMR12b, HKMR12c]. Difference [Ach13, AM12b, AM12i, BL15a, BL15d, BL15e, BL15h, BL15l, Lin10k, Lin10l, Pöt10e, Pöt10g, AM12g]. Differences [Neu10n]. Different [Bra14c, JPT16c, Pos12d]. Differentiable [AH11o, Bes16, DHHR11q]. Differential [ABCC^{+}12, BB14a, Bec14e, BDG^{+}14a, BL15b, BL15c, BL15d, BL15i, BL15m, CKM^{+}13b, Die10a, Die10d, Die10j, Die10k, Die10m, Die10l, Frö12b, FS13f, GQY14, Her13k, Kha10f, KKP16d, KKP16f, KKP16e, KKP16k, KKP16g, Lej12, Lor10c, Maw13, Mie11, Neu10o, Neu10s, NO13, OZLK13h, QY12, Rob14b, SS14, Sta12b, Sta12f, SS13b, Tuc13c, Wei14c, YSY10i, Bec14d]. Differentiation [AH12d]. Diffraction [AGJ^{+}13-29]. Diffusion [ACT11a, ACT11b, CP14, DHI13d, FM12b, FM12c, FM12d, FM12e, Her13h, HP13, Kha10d, Lin10g, Lin10h, Lin10i, Lin10j, Lin10k, Lin10l, Lin10y, Lin10z, Lin10-27, MC15, PEP13b, QY12, VV13, VA12]. Diffusions [CG14, Eth11f, Her13k, Jac13b]. Digital [CM10g, CM10h]. Dimension [BSKŠ15, Sab13s, Sal12b, Sti13h]. Dimensional [Eth11f, FKM14, FO14a, FO14j, FO14m, GG14, Jac13b, Jak11, KKP16b, OZLK13d, PS11, Sin16]. Dimensions [BBKK13, Jon15, MR14]. Diophantine [Mag14, SD10, Voj10b]. Direct [AGJK12, AGJ^{+}13d, Sch12]. directions [vdG16c]. Dirichlet [DHHR11e, GZ12, MMN13f, MMN13h, MMN13k, Otw12c]. Discontinuous [MR14]. Discrepancy [Ban14, Bil14, CS14, DP14, DGW14b, DGW14c, HS14a, Kli14b]. Discrete [Aba10, AN11, Haj12, KPR13, Pöt10c, Yuk13]. Discrete-Time [KPR13]. Discretely [Mas15b]. Discreteness [HK14j]. Discretisations [Lin10i, Lin10j]. Discretizations [EH16c]. Discussion [FKP13e, GW11f, Maj13g, Maj13k]. Disks [FKP13i, FKP13m]. Disorder [Gia11i, Gia11k, Gia11d]. Disordered [Gia11h, Kum14j]. Dispersal [LMP13a, OX13]. Dissipative [AN15a]. Distance [BM14, Lor10g, MUS11a]. Distortion [HK14e, HK14i]. Distribution [Bou14b, BGT14, DH12, GPV12, Lor10e, Mag14, Nås11b, Nås11d, Nås11c, Nås11m, Nås11l, Nås11o, Nås11n]. Distribution-Like [Lor10e]. Distributions [Mec12, Sab13c, Sch15b]. Diversity [Bri13]. Divides [Éme13]. Divisor [Sab13p, Sab13r]. Divisors [HY12d, MV12f]. Domain [MMN13f, MMN13h, MMN13i]. Domains [Bur14k, FO14f, FO14g, GGS10h, GZ12, MMN13g, MMN13m, MMN13l, MMN13o, MMN13p, MM13h]. Dominance [GGS10b]. Double [Lie15l, MM13c]. Double-Hopf [Lie15l]. doubling [YYH13h]. Down [MMA11b]. Drift [Eth11e]. Driven [Bai14, Bra14l, GQY14, Has13, ZK13]. Droplets [DG14g]. Dual [DG14b, Rab11, YYH13i, vdG16g]. Duality [HMM^{+}13l, Rei13a, Sti13i]. Dugald [Koe14b]. Duquesne [Kor13b]. Dvoretzky [Mec12]. Dyadic [Fla11b]. Dynamic [BORM14, BL15f, Fla11f]. Dynamical [Aba10, CCPP12, GKL12, KPR13, KP13c, KP13d, Pöt10c, Pöt10h, Que10f, Que10e, Que10g, Que10o, Que10m, dFS13, ABB^{+}10b]. Dynamics [Bed10, Car11, DHI13d, DHI13e, DHHR11r, DS10b, DGQ12, Est15, Jon15, Lie15c, ONO13, Par10b, Sch10b]. Dynkin [BF12o]. Dyson [BZ14].

École [Ben13b, Bur14c, Fla11f, Gia11d, Kum14j, Le 11j, AT11i, Eth11h, Kol11l, vdG16h]. Ecological [Bri13]. Ecology [LMP13a]. Economics [Her13c]. Eddy [LR12f]. Edge [Van13a, Van13b]. Editors [Koe14b, Viv14b, Wei14b]. Effect [HM13]. Effective [Kum14g]. Effects [Bri13, ZK13]. Eigenfunction [Yse10d]. Eigenfunctions [Bur14e]. Eigenspaces [Bro10b]. Eigenvalue [Bou14b, Bou14d, GGS10c, Ves11n, Ves11p]. Eigenvalues [Bur14e, DPPS14, LE11a, Par10k, Par10l, Par10i, Que10g]. Einstein [Gue13b, Sei12]. Eisenstein [HY12b]. Elastic [AN15e, Jac13a]. Electron [Riv12]. Electronic [Yse10e, Yse10k]. Element [Kru14f, Lin10m, Lin10n, NV12, Sie12b]. Elementary [Her13c, Her13e, Her13h, Her13i, Her13j, Her13k, Her13l, MMA11l, Sok13]. Elements [Bro10b, HMM^{+}13f]. Elimination [Rob14b, Rob14c]. Elliptic [BFMV14g, BFMV14r, CP14, FM12d, FM12e, FM12g, Frö12c, GGS10h, HKMR12c, HKMR12d, HKMR12e, Hu11i, LLMS12b, MM13h, Otw12c, Par10g, PP16n]. Elliptic-Hyperbolic [Otw12c]. Embedding [Dod10j, Hob11]. Embeddings [BS12, HPRY11b, HKMR12b, LE11a]. Emergence [DG14c, DG14g, DG14i, DG14m, DG14l, VV13]. Empirical [HTKS13, Kol11b, Kol11d, Kol11l, Kol11j, vdG16g, vdG16k]. Endomorphism [Bou10g]. Endomorphisms [DS10b]. Energetic [Mie11]. Energy [Ben13q, DHHR11e, Le 11d, Par10b]. Energy-Levels [Par10b]. Engineering [Her13d]. Entire [Sch10b]. Entropy [Kaw13i, Kaw13h, Léo12b]. Environment [CJL11]. Equation [Bło13, Bro12, DHI13e, DHI13j, DHHR11e, EZ12b, Fla11h, Kha10p, Kha10u, Kha10t, KM12b, Mon14b, Neu10v, Otw12d, Yse10e]. Equations [ABIL13b, ABCC^{+}12, AB12, AN15f, AN15g, AM12g, AM12h, AM12b, AM12i, Ban15c, BDR16a, BDR16c, BDR16d, BDR16g, BDR16i, Bar13, BDG^{+}14a, BG11, Bob15, BL15b, BL15a, BL15f, BL15d, Bou12a, CKM^{+}13b, Cré11, DHI13d, Deb13, Die10a, Die10d, Die10j, Die10m, Die10l, FM12b, FM12d, FS13f, GGS10b, GGS10h, GQY14, Gue12a, Hel13, Her13k, Hu11d, Hu11e, IS13, Ish13, Kha10e, Kha10f, Kha10a, Kru14h, Kru14i, Lam15, Lau15, Lej12, Lit13b, Lor10a, Lor10c, Lor10e, Mag14, Maw13, Neu10o, Neu10u, Neu10s, NO13, Otw12c, Pöt10e, Pöt10g, QY12, Rei13e, SAC^{+}14b, SS14, Ser13b, Sta12b, Sta12f, SS13b, SD10, Tuc13c, VV13, ZK13, BPR16]. equilibration [DG14p]. Equilibria [Lie15h]. Equilibrium [Kha10v]. Equivalence [DGH16q, Isa11i]. Equivalent [YSY10j]. Erasure [Le 11h]. ergodic [GKL12]. Ergodicity [DG14k, Deb13, Pro13a, vH14]. Error [Kru14f]. Escape [Kaw13d]. Essential [FKP13i, FKP13m, Sam14j]. Estimate [Bou14a]. Estimates [EL14, GGS10e, Gia11i, Gia11k, GM14, Kaw13g, KM12a, Kol14, Kru14f, Kum14f, Kum14g, Kum14e, LE11e, LE11i, Le 12, LLMS12b, Maj13q, Pos12i]. Estimation [ABMT13, BR15, CGC15, Maj13i, Mas15b, Mas15c, NPS13, TBBF13b, vdG16h]. Estimators [vdG16a]. Étale [Duc15b]. Étude [MM13a]. Euclidean [Frö12b]. Euler [PP16h]. Evaluation [Sti13g]. Events [Wai13c]. Every [Éme13]. Evidence [Sti13u]. Evolution [AN15f, AN15g, AH11h, Bra14i, DG14o, Hu11e, Kru14h, Kru14i, Lam15]. Evolutionary [Ban15c, Bob15]. Ewens [BNR11]. Exactly [Gia11f]. Exactness [Pro14]. Example [Fra14, Maj13o]. Examples [AGB15e, BPR12c, Bou10h, FS13g, JPT16e, Kaw13e, Kol11e, Lor10a, MMN13o, NS12f, NS12g, vdG16r]. Excess [Kol11e, Kol11f]. Excitability [WMR13a]. Excursion [AT11a, Her13c, Her13d, Her13e, YY13b, YY13d, YY13l]. Excursions [Wu12]. Existence [AH11n, AH11p, BDR16a, Die10d, FM12c, FM12d, FM12e, YY13c, Yse10f].

Exit [DHI13a]. Expanders [Ben13l]. Expanding [MSU11b, MUS11a, MSU11c, MSU11d, MSU11e]. Expansions [AGJ^{+}13t, AGJ^{+}13-27, FS13d, FS13e, FS13c, FS13f, Yse10d]. Expected [Kol11b]. Experimental [Bro13, Bru13]. Experiments [Kru14e]. Exponent [DHHR11t, LE11g]. Exponential [MW12, Tan11b, Ves11h, Yse10l]. Exponentials [MNU12]. Exponentiation [Wil14]. Exponents [DHHR11o, Kaw13c, Wei14j]. Exposition [Die10a]. Extended [AGJ^{+}13q]. Extending [Lor10c]. Extension [DG14j]. Extensions [GW11e, JPT16f, JPT16k, JPT16n, KK14d, KK14e, MV12b, SAC^{+}14j, iS10, JPT16g]. Exterior [Wei14c]. Extinction [Nås11e, Nås11h]. Extra [BFMV14h]. Extraction [Rum13]. Extrapolation [SSR15]. Extremal [Cap14]. Extremes [BFMV14c].

Factor [AK11f, Ben13q]. Factorization [KKP16o]. Factorizations [AK11a, AK11f]. Factors [AK11e, AK11b, AK11f, AK11h, AK11i, AK11j]. Facts [Sam14i]. Fagnano [PP16i]. Failure [Eis12]. Families [Ber13, BCGG10i, NS12f]. Family [Fai12]. Fano [Cao13, Gue13b]. Faster [HM13]. Fatou [Aud11a, Aud11c, Aud11j]. FC [DGH16v]. Features [DGH16a, DGH16l, DGH16t, DGH16u, DGH16v]. Feedback [AN15e, AN15g]. Feedbacks [AN15f]. Feller [BSW13c, BSW13d, BSW13i, BSW13k, RS10b]. Few [Sam14e, Ser14]. Feynman [Her13e, YY13d]. Fiber [NS12q, Pöt10d]. Fibers [FKP13g]. Fibrations [Di 14]. Field [Ach13, Ben14, Cha14, DG14g, DG14m, DG14l, GLL11, Le 11f, Mor12a]. Fields [Boy15d, Boy15e, BS13a, BS13b, Coh12, CKKR12b, Ken15c, Le 11j, Lej12, Maj14k, Spo13, SSR15, Sta13, Sti13h, Sti13t, Sti13w, vdD14]. Fifth [DFO13]. Filtered [DMS14b, Sab13l, Sab13m, Sab13s, Sab13r]. Filtering [ABMT13]. Filters [CM10l, CM10m]. Filtration [Rei13n]. Filtrations [DÉ13, Sab13a]. Final [AH11i, HK14c, Her13f]. Finance [BCG^{+}13b, CCG^{+}11b, Cré11, DRiSS10b]. Financial [Her13c, Her13d, Pro13b]. Fine [DHI13e, DHHR11f]. Finite [Ach13, Bar11a, Bar11d, Bar11j, Bar11m, BS12, Bou10d, Boy15b, Boy15d, Boy15g, Bro10c, Bro10f, HK14e, HK14i, Hu11e, Jak11, KKP16b, Kru14f, Le 11m, Léo12b, Lie13, Lin10k, Lin10l, Lin10m, Lin10n, Neu10n, NV12, OZLK13d, Par10h, Sam14d, Sie12b, Sti13w]. Finite-Dimensional [Jak11, KKP16b, OZLK13d]. Finite-Rank [Par10h]. Finitely [Bou12b]. Finiteness [Duc15a, Wit14e, Wit14c, Wit14d]. Fire [SG13]. First [ADFT11, Bar10c, Bar13, BZ12, BM14, BV11, GW11i, Nås11n, Pil14, Sok13, YY13b]. First-Order [Bar13, GW11i]. Fisher [DG14m, DG14n]. FitzHugh [Thi13]. Fixation [DG14g, DG14i, DG14m, DG14l, DG14o]. Fixed [Bar11e, Ceg13g, Hu11j, vdG16d]. Flat [FKM14]. Flatness [Sch10f]. Fleming [DG14q]. Flour [AT11i, Ben13b, Bur14c, Eth11h, Fla11f, Gia11d, Kol11l, Kum14j, Le 11j, vdG16h, AFV12, BHJKX12, BLSC12, DMZ12, GK12, GKL12, INS12]. Flow [AH11o, BEG13b, BG13, Cao13, EH16c, FFM11b, Gue13b, HJ13, Lie15c, Neu10h, SW13, Tia16b, Tia16c]. Flows [ABH^{+}13b, AH11a, ADFT11, Bai14, GK13, Haj12, HSL13, Sin16]. Fluctuations [BZ12, Nor13]. Fluid [BCM15e, CDG^{+}13b, DHHR11r, DGQ12, Fla11f, Gla12, Lie15c, Pad11g]. Fluids [Con13, Pad11a, Pad11c, Pad11e, Pad11f, Růž13a, Růž13b]. Fock [AN11]. Foliations [Bru10, Neu10i, Neu10j, PS13b]. Following [Hel13]. Foragers [WL13]. Foraging [BRVdL13, Mor13b]. Form [MMA11k]. Formal [BF12e, Bos14a, Bos14j, Rob14c, Rob14d, Sab13i]. Formalism [MUS11a]. Formation [MC15]. Forming [GML11]. Forms [Boy15d, Boy15e, Boy15f, KKP16e, Sam14s, vdG16n]. Formula [AT11d, Fai12, FO14k, Maj13c, Maj13l, Maj14i, Rou14c, YY13d]. Formulae [MMN13p, NS12t, Pic11b]. Formulation [BCAK13, DG14e, DG14d, Maj13d, Nås11k]. Formulations [Mie11]. Forward [Sab13m]. Foundations [DRiSS10b, Kum15c, Mol13]. Founded [Sie12b]. Four [FS13g]. Fourier [BR15, BGT14, Yse10g]. Fractal [MUS11a, MSU11e]. Fractional [AM12h, AM12b, AM12c, AM12i, AM12e, AM12l, Coh12, CKKR12b, Die10a, Die10d, Die10j, Die10m, Die10l, Gia11c, Mar11a, MS12, Pic11b, Pra11a, Pra11b, Sal12b, iS10]. Fragments [Sti13i]. Frames [Frö12b, Frö12e, Frö12f]. Framework [DHHR11g, Lor10h, Aud11f]. Fredholm [OZLK13g]. Free [AN11, Ben14, BF12c, BF12l, DH12, Le 11f, Pad11e]. French [CT10, DMLR11d, DMLR12, DMLR13, MM13a]. Frequency [AGJ^{+}13e]. Friction [BCM15d]. Friendly [Cor11]. Front [ABB^{+}10a, ABIL13a, AT11c, AK11g, ABBG^{+}12, AGB15d, ABH^{+}13a, Amm12a, AGJ^{+}13f, AGJ^{+}13g, AGJ^{+}13h, AGJ^{+}13i, AGJ^{+}13j, AGJ^{+}13k, AGJ^{+}13l, AGJ^{+}13m, AGJ^{+}13n, AN15c, AH11j, Ang14f, AM12d, AH12e, Aud11d, BBD13a, BBD13b, Ban10b, BMK15, BB14b, BDR16e, Bar10b, Bar11f, BBK13a, BBK13b, BCM^{+}11a, Bec14a, BCGC^{+}15, Ben13d, BCG^{+}13a, BNQ^{+}12a, BDG^{+}14b, BCMV11a, BD13a, BFMV14i, Bło12b, Bob11a, BL15j, BBST16, BDFT10, BF12f, BF12g, BF12h, Bos14f, Bos14g, Bos14h, BSW13e, Bou10i, Bou10j, Bou10k, Bou10l, Boy15c, Bra14d, Bro10d, BS16, BCGG10d, Bur14d, BMOR13a, BCM15b, CKM^{+}13a, CGGM14a, CM10i, CM10j, Ceg13e, CST14a, CST14b, CKKR12a, CTSDV10b, CDD^{+}14, CDG^{+}13a, CCG^{+}11a, DGH16e, DGH16f, DGH16g, DGH16h, DGH16i, DGH16j, DG14f, DGGK^{+}14, DHI13f, Def11d, Del12a, DMS14c, DHHR11h, DHHR11i, DHHR11j, DHHR11k, Die10e, Die10f, Die10g, Dod10g, DMLR11a]. Front [DMLR11b, DMLR11c, DMLR14, DFN15a, DFN15b, DG12a, DRiSS10a, Eth11b, FKM^{+}11, FO14e, FM12f, Fav15b, Fla11c, FK11c, Frö12d, FS13h, FKP13f, GGS10d, Gia11e, GW11h, HMM^{+}13e, HT13a, HK14d, Her13a, HS11a, HKMR12f, HY12c, Hu11f, Isa11a, JPT16h, Kaw13f, Kha10i, Kha10j, Kha10k, Kha10l, Kha10m, KM14a, Kli14a, KP13a, KP13b, KK14c, KKP16j, Kol11k, Kru14b, Kum14c, KWZ15, LE11b, LR12a, Le 11e, LLMS12a, LMP13b, LMP13c, Lie15m, Lie15n, Lie15o, Lie15p, Lie15q, Lin10o, Lin10p, Lin10q, Lin10r, Lin10s, Lin10t, Lor10d, Maj13e, Maj14b, MV12c, MMA11d, MMA11e, MMA11f, MUS11b, MMN13b, MMN13c, MMN13d, MMN13e, MM13d, Mor12g, Mor13a, Nås11i, NS12h, NS12i, NS12j, NS12k, Neu10k, Nis14b, Nou13a, OZLK13c, Otw12e, Pad11d, Par10c, PBBD13a, Pic11a, PP16j, PP16k, Pos12b, Pöt10b, Que10h, Rei13f]. Front [RTW15b, Rey13b, RSSS12a, Rob14e, Rou14d, Sab13e, Sab13f, Sab13g, SAC^{+}14c, Sam14k, Sam14l, Sam14m, Sam14n, Sch15a, Sch10g, SSM14e, Sta12e, Sti13j, Sti13k, Sti13l, Sti13m, Sti13n, Van13c, Ves11c, Wai13b, WMR13b, Wei14g, Wit14f, YYH13e, YYH13f, YYH13g, YY13e, YY13f, YY13g, YY13h, Yse10h, YSY10c, vdG16i, vdDKM^{+}14]. Fully [Gua14, IS13, CGGM14b]. Function [Bil14, CM11, DHHR11g, FKM14, MMN13f, MM13e, Par10k, Par10l, Par10n]. Functional [AH11a, AH11b, Car11, CP14, Gal13, KMM12, Lam15, Led14a, MM13f, Neu10r, NO13, Pag13, Par10e, Pos12c, Pos12d, SGS15, SZ15]. Functionals [AGJ^{+}13d, Bra14k, GPV12, Ort13]. Functions [AT11i, AM12j, AAFM12a, AAFM12b, Bar10c, Bar10d, Bar10e, Ben13f, Bos14c, CKKR12b, DMS14e, DHHR11d, DHHR11f, DHHR11q, Die10i, DG12b, Jeu15, JPT16f, JPT16g, JPT16i, Kum14a, KKR12b, LE11a, LE11k, Lew12, Lor10g, Maj13u, MMN13o, MMN13p, OZLK13d, OZLK13e, OZLK13f, PP16b, PP16n, Rob14b, Rou14c, Rou14f, Sch10b, Yse10k]. Functor [Bou10e, Bou10m]. Functors [Bou10d, Bou10c, Bou10o, Bou10q, GW11g]. Fundamental [Die10m, FO14f, FO14g, Her13c, Sti13o, Sti13u]. Further [Isa11f, Kum14d]. Fusion [Sam14f, Sam14j].

GAFA [Led14a]. Gain [Kha10g]. Galerkin [Kru14f]. Galton [Kor13b]. Games [Ach13, GLL11, Man12]. Gap [Kum15b]. Gas [BCM15c]. Gases [Riv12, Sei12, Sol12]. Gaussian [AT11d, AT11e, Ben14, BBB12, Eis12, Le 11f, Led14b, Leh11, Nou13b]. Gegenbauer [KKP16a]. General [BPR12a, CT13, DG14d, DG14g, DG14j, FS13d, Isa11g, JPT16b, Kaw13g, Kum14f, Le 11b, MV12d, Par10h, Que10k, Que10l, Ves11d, Ves11s, vdG16j]. Generalisations [LE11l]. Generalised [LE11a, LE11k]. Generalization [HMM^{+}13f]. Generalized [AGJ^{+}13e, AGJ^{+}13o, Bar10c, Bar10d, BOGP12, DHHR11l, MMA11g, MNU12, Růž13a, Růž13b]. Generalizing [KM12b]. Generates [Cha14]. Generators [BSW13d]. Generic [DGH16d]. Generically [DGH16t]. Genetic [Eth11e]. Genetics [Bul15, Eth11h]. Genus [Kok11b, PP16e, PP16t, PP16v]. Geodesic [Le 11g]. Geodesics [BK12b, Gue12a]. Geometric [BB14a, Bar10d, BFMV14j, BFMV14k, BY15, Bra14e, CEK12, DGH16t, DGH16u, DGH16v, DG12b, Frö12b, Gal13, Gut14, Kha10n, Mor12h, Pöt10c, Sti13b, Tia16c, CM10k, KMM12, Led14a]. Geometric-Functional [Gal13]. Geometrisation [Boi16]. Geometry [AT11h, AH11o, Ang14b, Bak13b, BPHZ15, Ben13b, Ben13o, Bło13, Bos14j, CGGM14b, CTSDV10a, Dem13, Duc15a, Frö12b, Frö12g, HS14b, JR15, Ken15c, Kid13, KKP16d, Kwo11, MUS11a, Mol13, Otw12i, Rab11, Rei13j, RTW15a, Spo13, Sti13f, Viv14b, Wei14b, Yuk13]. Germ [NS12w]. Getting [GW11i]. Gevrey [FS13e]. Gibbs [MUS11a, MSU11g]. Ginzburg [Neu10u, Neu10r]. Girsanov [Her13h, Léo12b]. GIT [BFMV14h, BFMV14k, BFMV14n]. Given [BY11, ABB^{+}10b, CTSDV10a, DS10a]. Glass [FFM11b, Fas11, FK11a, GML11, Sie11]. Global [Bom10, BCD13, BSW13f, Bra14f, Bra14l, Kha10o, Kha10p, Lej12, Pos12e]. Good [Sab13i, Sab13h, Sab13p]. Governed [Kha10f, Kha10a]. GPT [AGJ^{+}13p]. GPT- [AGJ^{+}13p]. GPTs [AGJ^{+}13z]. Gradient [AH11a, Neu10g]. Gradients [Neu10c, Neu10l, Neu10o, Neu10q, Neu10s, Neu10t]. Grain [CM10l, CM10m]. Graining [Gia11b]. Graph [BS12, Ben13h, DMS14d, NS12m, NS12l, Pos12j]. Graph-like [Pos12j]. Graphical [Wai15]. Graphs [AK11f, BZ12, Ben13e, Ben13f, Ben13g, Ben13i, Ben13j, Ben13m, Ben13k, FFR15, Kum14e, Kum14k, NS12d, NS12s, Pos12f, Pos12i]. Grassmannian [Mor12e]. Grassmannians [Ale12a]. Great [Aud11c, Aud11f, Aud11e]. Green [JPT16i, MMN13f, MMN13g, MMN13j, MMN13h, MMN13i, MMN13l, MMN13o, MMN13p]. Group [Bar11g, BSG11, Bou10f, BCGG10h, DFO13, NPS13, Riv12, Sea14]. Groupoid [Sti13o]. Groups [Ben13c, Bou10d, Bou10p, BNR11, Bro10c, Bro10e, Bro10f, CC11, DGH16q, FO14a, FO14g, FO14h, FO14j, FO14m, JPT16b, Lie13, Mor12e, Rou14e, Sam14a, Sam14d, Sam14c, Sam14h, Sam14o, Sam14p, Sam14r, Sam14t, Sam14u, Sch11b, Soi13b, Sti13u, Wit14e]. Growth [Ben13e, HR12, KS13]. Guide [AG13, BO13, GMK13, Van13a]. Guiding [OZLK13d, OZLK13e, OZLK13f]. Guts [FKP13g, FKP13h].

H [Koe14b]. Habitat [Bri13]. Habitats [HP13]. Hadamard [DMS14e]. Hamilton [ABIL13b, Bar13, BG11, Ish13, Lit13b]. Hamiltonians [Bou14d]. Hardy [LE11f, LE11g, KM14b, LE11e, YYH13j, YYH13i, YYH13h, YYH13l]. Hardy-Type [KM14b]. Harmonic [AH11k, Ben13f, JPT16g, Lew12, Par10d, Par10m]. Harmonics [AH12h, AH12i]. Harnack [Bur14i, Lew12]. Hausdorff [BF12o, Duq10, Rou14c, YSY10e]. Heat [Bur14f, Fra14, FK11a, Kum14f, Kum14g, Kum14e, Par10e, Sie11]. Heat-Semigroup [Par10e]. Hedging [CJL11, Tan11b]. held [ABB^{+}10b, CTSDV10a, DS10a]. Hepatitis [HA13]. Hermitian [Ves11d, Ves11s, Ves11t, Viv14c]. Hernández [Wei14b]. Hernández-Lamoneda [Wei14b]. Herrera [Wei14b]. Heterogeneous [DGQ12]. Hide [PP16w]. Hierarchy [DG14j]. High [FKM14]. High-Dimensional [FKM14]. Higher [DG14p, GGS10g, GGS10h, MM13h, Sch11a]. Higher-Order [MM13h]. Hilbert [Ceg13g, Kru14h, Kru14g, Neu10e, Neu10f, OZLK13e, Pos12c, Pos12d, Sab13c, Sab13h, Sab13q, Sab13p, Sab13o]. Hirzebruch [HY12d]. Historical [BF12i, FS13i]. History [Aud11g, Maj14c]. Hitting [LS14, Sok13]. Hodge [NS12r]. Holds [Kum14a]. Hole [Mar11b, MMN13i]. Holes [ADFT11, MMN13l, PP16f]. Holomorphic [Aba10, Ber13, ABB^{+}10b]. Holonomic [Sab13c, Sab13q]. Hom [Bou10r]. Homeomorphisms [HK14e]. Homogeneous [Gia11f, Jac13b, NS12g, NS12n, Rou14b, YSY10d]. Homological [Sch10c]. Homology [Ban10e, FO14h, Mor12d]. Homotopy [Ban10c, Bar11m, Bar11n, Mor12c, Mor12d]. Hopf [Ken14, Lie15l, Lie15j, Lie15t, Lie15w]. Horizontal [ACT11a, ACT11b]. Hörmander [Naz12, Par10o]. Hot [Bur14h]. Hulls [vdG16p]. Hybrid [FE13]. Hyperbolic [AB12, Ban15d, BS12, Ben13g, Ben14, Bre13, Lie15d, Nis14c, Otw12c]. Hyperbolicity [Nis14e]. Hypercontractive [CEL12]. Hyperelliptic [CM11]. Hypergeometric [Bar10e, BNR11]. Hypersurface [NS12w]. Hypersurfaces [Isa11h, Isa11i, Isa11j, Isa11k, Isa11c, Isa11d, Isa11b, Isa11l]. Hypoelliptic [Fra14].

Ideal [FKP13j]. Ideals [BGdC13, Mon13, Van13a, Van13b]. Idempotent [Lit13b]. Idempotent/Tropical [Lit13b]. Identification [AGJ^{+}13-28, FM12g, HT13b]. Identifying [FMS14]. Identities [KKP16k, KKP16o, YY13p]. Identity [AH11c, YYH13a, YYH13k]. II [Amm12b, BFMV14p, CKKR12b, DG14j, Jac13a, JPT16n, KK14e, Neu10j, Sie11]. IIC [Kum14d]. III [BSW13g]. Illuminated [DGH16t]. Illustrated [Bre13]. Image [CM10g, CM10h, CM10t, CM10u, JR15, Rum13]. Images [CM10k, DGH16l, HK14f, Nås11a]. Imaging [AGJK12, Amm12b, ABJW12, AGJ^{+}13d, AGJ^{+}13c, AGJ^{+}13q, AGJ^{+}13s, BMOR13b]. Immersed [BCM15f, BCM15e]. Implementations [AGJ^{+}13w]. Implications [JPT16m]. Improved [Bou14a]. Impulsive [Sta12b, Sta12f]. Including [vdG16e]. Inclusion [FMS14, MMN13h]. Inclusions [AGJ^{+}13d, Lor10i, MMN13f, MMN13j, MMN13l, OZLK13g, OZLK13h, Ves11p]. Incompressible [EH16c]. Incorporating [Mor13b]. Increments [Sta13]. indécis [MM13a]. Indecomposable [Jak11]. Indefinite [Ves11e, Ves11g]. Independence [Pil14]. Independent [Hob11, Maj13q, Mie11, Nis14g]. Indeterminate [BF12e]. Index [BCD13, GL12, Par10b]. Indexed [DÉ13]. Individual [FE13, LMP13a, WL13]. Induced [Par10h]. Industrial [FFM11b]. Inequalities [Bar10e, Car11, CEK12, Gua14, KM14b, Kol11l, Led14b, Lew12, MR14, Neu10g, vdG16k, vdG16q]. Inequality [BFMV14c, Bło14, CEL12, Lat14, Lor10g, NT12]. Infante [DHI13e, DHI13j]. Infection [HA13]. Inference [AT11g]. Infinite [Ben13p, MR14]. Infinitesimal [Her13g]. Infinitesimals [Her13m]. Infinity [Bou12b]. Influences [CEL12]. Information [Kum15b, Kum15c, SAC^{+}14i, vH14]. Inhomogeneous [BH14, CP14, YSY10e]. Injectivity [Sti13p]. Inner [KM12a, Neu10h, Neu10l]. Inputs [dFS13]. Instead [Lor10g]. Integrability [HK14g]. Integral [DHHR11e, Die10k, JR15, Kid13, LE11j, Maj14h, SSM14h, SSM14k, YY13i, AM12i]. Integrals [Ale12a, BOGP12, BV11, CS11, Her13e, Her13j, Maj13i, Maj13h, Maj14d, Maj14e, PP16h, PP16w, iS10].

Integrands [CS11]. Integrate [SG13]. Integration [AH12d, BL15c, BL15d, DP14, PP16d]. Integrative [TP13]. Integrodifference [ZK13]. Interactions [DGH16q]. Interior [PT10b]. Intermediate [Lau13]. Internal [Bou10r]. Interpolations [CEK12]. Interpretability [SAC^{+}14d]. Intersection [Ban10d, Ban10e, Mar12]. Intersections [HMM^{+}13d, HM13, HY12d]. Intertwining [Ale12a, KKP16l]. Interval [SAC^{+}14e, SAC^{+}14i, SAC^{+}14j]. Intervals [SAC^{+}14f, SAC^{+}14g, SAC^{+}14k, vdG16f]. Introducing [Lor10e]. Introduction [AT11f, AH11l, AH12i, Aud11h, BDR16f, Bar10f, Bar13, BFMV14l, Bob11b, BL15k, BPR12a, Bos14i, BEG13a, BEG13b, Bra14g, Bro10e, BS13b, CM10n, CM10o, Ceg13f, DGH16k, DG14h, Deb13, DHI13g, Def11e, DMS14f, DHHR11m, DHHR11n, Die10h, DS13, Est15, Eth11c, FO14i, Fla11d, FKP13k, Gia11g, Gia11h, Gue12b, HK14h, HY12e, Hu11g, IS13, Ish13, JPT16j, Kaw13i, Kaw13h, Ken15b, Kha10q, Kha10r, Kid13, KKP16m, Kol11g, Kru14c, Kum14h, LR12e, Lie15r, Lin10u, Lin10v, Lor10f, Maj13f, Man12, MSU11f, MM13g, Mor12i, Nås11j, NS12o, Neu10l, Nis14d, Otw12f, Par10f, Par10m, Pos12g, Rei13g, Riv12, Rob14f, Sab13j, Sch10h, Sea14, SSM14f, SSM14i, SW13, SZ15, Tem15, Var11, Ves11a, Wai13a, Wei14c, Wei14h, Yse10i, YSY10f, dVBSV16, vdG16l]. Introductory [Ben13h, FFR15, FS13g, Wai15]. Invariance [Her13d, Her13h, Kaw13i, Kaw13h, PP16t, Ser14]. Invariant [BFMV14j, HMM^{+}13g, Mor12j, Mor12l, NT12, NS11, Pöt10d, VA12]. Invariants [Bro10f, Isa11h, MMA11c, Sam14d]. Inverse [AGJ^{+}13-29, BCAK13, DMS14e, DMS14g, DMS14h, SBK^{+}13]. Inversion [VA12]. Investigation [Maj13g]. Involution [Rei13h]. Involving [Bar10e, DGH16r, DGH16u, DGH16v, Nås11o]. Irreducible [Ser11b]. Irregular [Sab13k]. Irregularities [BGT14]. Irrelevant [Gia11i]. Isogeometric [EH16c, MS16, SD16, dVBSV16, EH16b]. isolated [NS12q]. Isometric [Sea14]. Isometry [Bou14a]. Isomorphic [Del11]. Isomorphisms [AAFM12b]. Isoperimetric [Ben13e, Gua14]. Isothermal [FFM11b, Pad11e]. Isotropic [AH11m, GPV12, GM14]. Israel [KMM12, Led14a]. Issues [BDG^{+}14a, Duf13, DRiSS10b]. Itô [Her13k, Maj13h, Maj14d, Maj14e, Maj14h, Maj14i, YY13l, YY13q]. Italy [ABB^{+}10b, ABIL13b, ABCC^{+}12, ABH^{+}13b, BNQ^{+}12b, BDG^{+}14a, BCMV11b, BMOR13b, CKM^{+}13b, CGGM14b, CTSDV10a, CDG^{+}13b, DS10a, EH16b, Fas11, Koe14b, Kum15c, LLMS12b, PBBD13b, RSSS12b, Tia16c, Viv14b]. Iteration [DG14i, DG14j]. Iterative [Ceg13d, Ceg13g]. IV [Mas15b].

Jacobi [ABIL13b, Bar13, BG11, Boy15d, Boy15e, Boy15f, Ish13, Lit13b]. Jacobian [BFMV14e, Rei13i, Rei13j]. Jakob [PP16l]. Joint [GZ14b, Lor10h]. Jones [FKP13h]. Jordan [BL15m, BL15l, PP16m]. Julia [Aud11c, Aud11a]. July [ABB^{+}10b, DS10a]. Jump [Sok13]. June [Chu10].

Kac [YY13d]. Kähler [Ang14b, Bło13, BEG13b, BG13, Cao13, Gue12a, Gue13b, Kol12, SW13, Tia16b]. Kalman [ABMT13]. Kannan [AGB15a]. Kashiwara [Rou14g, Rou14e]. Keldysh [Otw12c]. Kernel [Bło14, Bur14f, Fra14, GGS10e, Kum14f, Kum14g, Kum14e]. Kernels [BNR11, MMN13g, MMN13l, Par10e]. Key [YSY10g]. Kind [Bar10c]. Kinematic [AT11d]. Kinetic [Ban15a, HA13]. Klein [CM11, HKMR12e]. KLS [KM14b]. Knowledge [BBKK13]. Kobayashi [Wei14i]. Kolmogorov [BC14]. Koszul [Con14]. Kronecker [KK14d].

Lagrangian [Con13]. Lamoneda [Wei14b]. Landau [Neu10u, Neu10r]. Langevin [Jac13a]. Langlands [Rei13a]. Laplace [DHHR11e, Sab13o]. Laplacian [Kok11b, LE11h, MMN13p, Sal12b]. Laplacians [Neu10p, Pos12f]. Larders [GW11i, GW11j]. Large [AFV12, Bak13b, BZ12, BBB12, BCGG10h, Ish13, LR12f, Wai13a]. Lasso [vdG16c, vdG16f, vdG16m, vdG16s]. Last [YY13j]. Lattices [Bar11l, Ben14, GW11e, GW11i]. Laurent [DÉ13]. Law [Kha10d, MY11, RY14, YY13p]. Laws [Ban15d, Bre13, Eis12, Lie15d, YY13j, YY13m]. Layer [AGJ^{+}13r, Lin10w, Lin10x, MM13c, MM13f, MM13h, MM13i, Lin10y]. Layer-Adapted [Lin10w, Lin10x, Lin10y]. Lebesgue [DHHR11o, DHHR11t]. Lectures [BCG^{+}13b, Bos14j, CTSDV10a, CCG^{+}11b, Dea14, DFO13, DS10a, HK14i, Nou13b, Toë11b, Wai15, Wil14, vdD14, ABB^{+}10b]. Leffler [AM12j, Die10i]. Lefschetz [Bar11e, HMM^{+}13c, HMM^{+}13f, HMM^{+}13g, HMM^{+}13i, HMM^{+}13h, HMM^{+}13j, HMM^{+}13l]. Legendre [PP16n]. Leibniz [AM12c, KM12b]. Lemma [GW11d]. Lemniscate [PP16i]. Lemniscatic [PP16h]. Length [CCPP12]. Lens [HKMR12g]. Lent [Bou11]. Leray [LR12g]. Leray-Deconvolution [LR12g]. Less [Lor10g, MM13a]. leurs [CT10]. Level [CM10e, CM10f, DG14p, Lau13, WL13, BMOR13b]. Levels [Par10b]. Levi [Mon14b]. Levico [Viv14b]. Lévy [BR15, BSW13g, Coh12, CKKR12b, CGC15, DHI13d, DRiSS10b, Her13l, KKR12b, Mas15b, Mas15c, PAMC13, Pro13a, Pro14, RS10b, Rey13a, Tan11b, YY13k, YY13m]. Lévy-Transformation [Pro14]. Lévy-Type [BSW13g]. Lie [BF12c, BF12l, CC11, Jak11, NS11, Rei13m, Rou14e, Ser11b]. Lieb [Led14b]. Life [KP13c, KP13d, ZK13]. Lifting [GW11d]. Light [Otw12g]. Like [Lor10e, DS10b, Pos12j]. Likelihood [HS14b]. Limit [BS13a, Gru14, Maj14g, Maj14e, Maj14f, Sta13, Yuk13]. Limits [Ben13i, Kum14j]. Line [BFMV14a, BFMV14g, Sab13o]. Linear [BL15b, BL15a, Bro10f, CH14, Di 14, GGS10b, GGS10f, GPV12, GZ14b, HY12f, Kaw13j, Kol11h, LE11j, Mor12e, Neu10e, Pöt10e, Tuc13c, Ves11a, vdG16a]. Linearization [Pöt10f]. Lines [Bur14g, CM10e, CM10f]. Links [FKP13l]. Liouville [AM12l, AM12k, Ben13j, Bro12, Die10d, Die10k]. Lipschitz [BDR16c, FKM14, Has13, MM13h]. Literature [Maj13k]. Littlewood [YYH13k]. Lizorkin [YSY10e, YSY10h]. LNM [Chu10]. Load [Kha10g, Kha10h]. Local [Aba10, ÀFR13, AH11b, Ben13i, Bou14d, Bra14b, Bra14h, Bra14i, BL13, DGH16l, HMM^{+}13b, HY12g, JR15, Kha10s, Lor10a, Mar12, Mar11a, Mon13, Ros11, Sab13l, Sab13m, Sab13s, Sab13r, Sch10k, Ser13b, Sti13q, Sti13t, YYH13l, YY13c, YY13l, Hel13]. Localization [DHI13h, Par10g]. Locus [BFMV14t]. Log [CG14, EL14, KM12a]. Log-Concave [CG14, EL14, KM12a]. Logarithmically [Kla14, Lat12]. Logarithmically-Concave [Kla14]. Logistic [Nås11h]. Long [AH11n, CH14, DG14k]. Long-Term [CH14]. Long-Time [AH11n, DG14k]. Loomis [NT12]. Loop [Le 11h, Le 11i]. Loops [Cha14, Le 11g, Le 11j, Le 11k, Wu12]. Loss [Bri13, vdG16j]. Lovász [AGB15a]. Low [GG14, Jon15, Kli14b, Kol11i]. Low-Discrepancy [Kli14b]. Lower [Bło14, DG14c, FO14k, GGS10i, Kaw13d, Sal12b, vdG16n]. Luis [Wei14b]. Lyapunov [Kaw13c, Sta12g, Wei14j].

Macaulay2 [ÀFR13, Van13b]. Machlup [CP14]. Macpherson [Koe14b]. Macroscale [Mor13b]. Macroscopic [Hel13]. Main [AGB15e, BF12j, DGH16s, DG14d, HY12h, Maj13d, Maj13m, Maj13p, Maj13s, NS12c, NS12p, NS12u, Rei13i]. Malliavin [BT13, INS12, Kru14d, Kru14g, Nou13b]. Manifold [Gue13b]. Manifolds [AH11f, Ang14c, Ang14d, Ang14g, Boi16, Bou12a, Cao13, Dea14, GG14, HKMR12b, HKMR12c, HKMR12d, HKMR12e, Lie15h, NS12s, Pos12h, SSM14h, SSM14k, Viv14c, Wei14b, Zil14]. Manin [Sti13c]. Manis [KK14e]. Manufacturing [Fas11]. Many [RSSS12b]. Map [CM10r, CM10s]. Maple [DP11]. Mappings [AH11k, Cap14, DGH16p, DS10b, HK14i, KO13b, MUS11a]. Maps [AAFM12a, Ber13, FO14d, MMA11h, MMA11i, MSU11d, Mey11, Rei13k, CM10k]. Margin [vdG16o]. Marginals [BY11, BDMY11b]. Marked [BSKŠ15, Bou11]. Markets [Ben13q]. Markov [BLSC12, CG14, CCPP12, Kum14k, Le 11b, Le 11j, Le 11m, SS14]. Markovian [BPR12c, CJL11]. Marks [SAC^{+}14g, SAC^{+}14h]. Martingale [MNU12, QY12]. Martingales [BY11, BDMY11b, BL13, HPRY11b, MR14]. Maslov [BCD13, Par10b]. Mass [Ves11r]. Match [Cor11]. Matching [Rum13]. Matchings [AK11h]. Material [AH11e, LE11d]. Materials [RL15, SGS15]. Mathematical [Amm12b, AGJ^{+}13s, Aud11c, Aud11e, BBKK13, BBK13c, BCG^{+}13b, Bur14c, BCM15d, CDG^{+}13b, CCG^{+}11b, DGQ12, Eth11h, FFM11b, GML11, HVM13, Her13e, KS13, Kum15c, LMP13a, OX13, Otw12h, Pro13b, SG13, Ves11a, Fas11]. Mathematically [Sie12b]. Mathematics [FFV11]. Matlab [FK11b]. Matrices [AGJ^{+}13t, BMS11, DMS14b, DMS14d, DMS14e, DMS14g, DMS14h, DMS14i, DPPS14, GZ14b, Nor13, Que10i, Spe12b, Ves11g, Ves11t, vdG16q]. Matrix [AGJ^{+}13u, KKP16g, KKP16n, KKP16o, Kol11i, Que10j, Ves11d, Ves11h, Ves11s, vdG16a]. Matrix-Valued [KKP16g, KKP16n, KKP16o]. Matter [ABB^{+}10a, ABIL13a, AT11b, AT11c, AK11c, AK11g, ABBG^{+}12, AGB15b, AGB15d, ABH^{+}13a, Amm12a, AGJ^{+}13b, AGJ^{+}13f, AGJ^{+}13g, AGJ^{+}13h, AGJ^{+}13i, AGJ^{+}13j, AGJ^{+}13k, AGJ^{+}13l, AGJ^{+}13m, AGJ^{+}13n, AN15b, AN15c, AH11d, AH11j, Ang14a, Ang14f, AM12a, AM12d, AH12c, AH12e, Aud11b, Aud11d, BBD13a, BBD13b, BDMY11a, Bak13a, Ban10a, Ban10b, BMK15, Ban15b, BB14b, BDR16b, BDR16e, Bar10a, Bar10b, Bar11c, Bar11f, BBK13a, BBK13b, BCM^{+}11a, Bec14a, Bec14c, BCGC^{+}15, Ben13a, Ben13d, BCG^{+}13a, BK12a, BNQ^{+}12a, Toë11a, BDG^{+}14b, BCMV11a, BD13a, BFMV14b, BFMV14i, Bło12b, Bob11a, BL15g, BL15j, BBST16, BDFT10, BF12a, BF12f, BF12g, BF12h, Bos14d, Bos14f, Bos14g, Bos14h, BSW13b, BSW13e, Bou10b, Bou10i, Bou10j, Bou10k, Bou10l, Boy15a, Boy15c, Bra14a, Bra14d, Bro10a, Bro10d, BS16, BCGG10b, BCGG10d, Bur14a, Bur14d, BMOR13a, BCM15a, BCM15b, CKM^{+}13a, CGGM14a]. Matter [CM10c, CM10d, CM10i, CM10j, Ceg13c, Ceg13e, CST14a, CST14b, CKKR12a, CTSDV10b, CDD^{+}14, CDG^{+}13a, CCG^{+}11a, CMN13a, DGH16c, DGH16e, DGH16f, DGH16g, DGH16h, DGH16i, DGH16j, DG14a, DG14f, DGGK^{+}14, DHI13c, DHI13f, Def11a, Def11d, Del12a, DMS14a, DMS14c, DHHR11a, DHHR11h, DHHR11i, DHHR11j, DHHR11k, Die10b, Die10e, Die10f, Die10g, Dod10c, Dod10g, DGW14a, DMLR11a, DMLR11b, DMLR11c, DMLR14, DFN15a, DFN15b, DG12a, DRiSS10a, EZ12a, Eth11a, Eth11b, EH16a, FFM11a, FKM^{+}11, FO14c, FO14e, FM12a, FM12f, Fav15b, FV11a, Fla11a, Fla11c, FK11c, Frö12a, Frö12d, FS13b, FS13h, FKP13b, FKP13f, GOZ12a, GGS10a, GGS10d, Gia11a, Gia11e, GW11a, GW11h, Gue13a, GZ14a, HMM^{+}13a, HMM^{+}13e, HT13a, HK14a, HK14d, Her13b, Her13a, HS11a, HPRY11a, HKMR12a, HKMR12f, HY12a, HY12c, Hu11b, Hu11f, Isa11e, Isa11a, JPT16a]. Matter [JPT16h, Kaw13a, Kaw13f, Ken15a, Kha10b, Kha10i, Kha10j, Kha10k, Kha10l, Kha10m, KM14a, Kli14a, KP13a, KP13b, KK14b, KK14c, KKP16c, KKP16j, Koe14a, Kok11a, Kol11a, Kol11k, Kor13a, Kru14a, Kru14b, Kum14b, Kum14c, Kum15a, KWZ15, KO13a, KKR12a, LE11c, LE11b, LR12d, LR12a, Le 11a, Le 11e, Léo12a, LLMS12a, LMP13b, LMP13c, Lie15e, Lie15m, Lie15n, Lie15o, Lie15p, Lie15q, Lin10e, Lin10f, Lin10o, Lin10p, Lin10q, Lin10r, Lin10s, Lin10t, Lit13a, Lor10b, Lor10d, Maj13a, Maj13e, Maj14a, Maj14b, MV12a, MV12c, Mas15a, MMA11a, MMA11d, MMA11e, MMA11f, MSU11a, MUS11b, MMN13a, MMN13b, MMN13c, MMN13d, MMN13e, Men13a, MM13b, MM13d, Mon14a, MFR13, Mor12f, Mor12g, Mor13a, Nås11f, Nås11i, NS12h, NS12i, NS12j, NS12k, Neu10k, Nis14a, Nis14b, Nou13a, OZLK13a, OZLK13c, Ota15a, Otw12a, Otw12e]. Matter [Pad11b, Pad11d, Par10a, Par10c, PS13a, PBBD13a, PEP13a, Pic11a, PT10a, PP16j, PP16k, Pos12a, Pos12b, Pöt10a, Pöt10b, Que10a, Que10h, RS10a, Rei13d, Rei13f, RTW15b, Rey13b, RSSS12a, Rob14a, Rob14e, Rou14a, Rou14d, Sab13d, Sab13e, Sab13f, Sab13g, SAC^{+}14a, SAC^{+}14c, Sal12a, Sam14b, Sam14k, Sam14l, Sam14m, Sam14n, Sch10a, Sch15a, Sch10d, Sch10g, Ser13a, Ser11a, SSM14b, SSM14e, Sie12a, Soi13a, Spe12a, Sta12d, Sta12e, Sti13a, Sti13j, Sti13k, Sti13l, Sti13m, Sti13n, SS13a, Tan11a, TBBF13a, Tia16a, Tuc13a, Van13c, Ves11c, Viv14a, Voj10a, Wai13b, WMR13b, Wei14a, Wei14d, Wei14g, Wit14a, Wit14f, YYH13b, YYH13e, YYH13f, YYH13g, YY13a, YY13e, YY13f, YY13g, YY13h, Yse10a, Yse10h, YSY10b, YSY10c, Zua13a, vdG16b, vdG16i, vdDKM^{+}14]. Matters [BSW13g, CKKR12b, DRiSS10b, Mas15b]. Mattila [JN14]. Maximal [BDR16d, DHHR11p, DG12b, FO14g, GG14, MR14, MV12e, MS11, Que10k, YYH13k]. Maximality [Wei14l]. Maximum [AH11r]. Mean [Ach13, DG14g, DG14m, DG14l, GLL11, MSU11c]. Mean-Field [DG14g, DG14m, DG14l]. Measurability [MSU11g]. Measure [Mec12]. Measure-Theoretic [Mec12]. Measurements [AGJ^{+}13c, AGJ^{+}13q, AGJ^{+}13z]. Measures [BNR11, Bou12b, CEL12, DGW14b, Duq10, Gua14, JN14, KM12a, Kla14, Kol14, Le 11i, Léo14, Maj14j, MUS11a, Que10i, YYH13h, YY13l]. Mechanics [CDG^{+}13b, Gla12, Pad11g, Spe12b, Yse10b]. Mechanisms [Enc13]. Media [ABJW12, BDR16a, BDR16g, BPR16, BDR16i, Kum14j, DMZ12]. Meet [YSY10h]. Melnikov [JN14]. memoirs [Aud11i]. memorial [Chu10]. Mercer [JPT16m]. Merging [BBKK13, CM10p, CM10q]. Meromorphic [Sab13i, Sab13h, Sab13p]. Meshes [Lin10w, Lin10x, Lin10y]. Meso [MMN13g, MMN13k]. Meso-Scale [MMN13g, MMN13k]. Mesoscale [Mor13b]. Message [Wai15]. Message-Passing [Wai15]. Metacyclic [Sam14o, Sam14r]. Metastability [DHI13h]. Method [Bil14, Bou11, Fle10, Gia11c, HKMR12h, Neu10d, OZLK13d, OZLK13e, OZLK13f, Rou14e, Sam14g, Sch15b, Sta12g, KKP16g, KKP16n, Rou14g]. Methods [Ach13, AGJK12, AGJ^{+}13s, AH12a, BCAK13, Bar11i, BR15, Ber12, Bul15, BMOR13b, Cal13, Ceg13d, Ceg13g, Ceg13h, CH14, CMN13b, DGH16m, DGH16n, DG14n, Hei13, Isa11g, Kru14f, Lie15s, Lin10m, Lin10n, NV12, PT10b, Rob14d, Rum13, SG13, SBK^{+}13, Sch12, SSM14g, SSM14i, Spo13, Wai13c, dVBSV16]. Metric [BS12, Ben13h, Lor10c, Lor10i, Mie11, Ves11e, vdG16p]. Metrics [Gue12a, Kol12, Men13b]. Micro [Mor13b, SGS15]. Micro- [Mor13b]. Micro-structure [SGS15]. Microscopic [Hel13]. Milman [Bło14, Naz12]. Milnor [Mor12k, NS12q]. Minimal [Bar11j, MMA11n, Neu10m, Sam14p]. Minimization [Bra14f, Bra14h, Bra14i, Bra14l, Kol11l, Kol11j]. Minimizers [Bra14b]. Minimizing [Bra14e, Bra14j, Bra14k]. minutieuse [MM13a]. Mirror [Bur14m]. Mittag [AM12j, Die10i]. Mittag-Leffler [Die10i]. Mixed [FMS14, MMN13i, MMN13m, MMN13l, NS12r, Sch10c, Yse10f]. Möbius [Bou12c]. Modal [SAC^{+}14k, SAC^{+}14i, SAC^{+}14j, Ves11i, Ves11j, Ves11q]. Model [Bou14b, DG14e, DG14g, DG14m, DG14l, DG14n, Duf13, HT13b, Her13d, Hob11, Kha10c, Kha10o, Kha10s, Kha10w, Koe14b, Kol11j, Kum14i, MT14, Mou12, Nås11b, Nås11d, Nås11c, Nås11a, Nås11e, Nås11h, Nås11k, Nås11m, Nås11l, Nås11n, Nås11q, Nås11r, Otw12b, TBBF13b, Ves11k, Wil14, vdD14]. Model-Independent [Hob11]. Modeling [Amm12b, BBD13c, BBK13c, BNQ^{+}12b, GK13, HVM13, HA13, SS13b, TP13, Tuc13b, ZK13]. Modelling [Ben13q, FE13, GML11, RL15, Rey13a, ABH^{+}13b]. Models [BBD13c, Bad13, Ban15a, Bar11j, BR15, BPHZ15, BY15, BCM15d, DG14q, DGQ12, DS13, Eth11h, Fas11, Fla11b, Fla11e, Fla11f, GGS10g, Gia11d, Gia11f, Gia11h, HJ13, Hel13, HP13, JPT16k, Ken15c, KS12, LR12b, LR12c, Lie15a, MC15, Mor13b, ONO13, SG13, SGS15, SD16, Tan11b, Tuc13c, Wai15]. Moderate [Sab13n]. Modified [Lat14]. Modules [Ale12b, Boy15b, Boy15d, Boy15g, Mon13, Sab13q]. Moduli [HY12i, Mar11b]. Modulus [Ros11]. moins [MM13a]. Molecular [YSY10a]. Molten [FFM11b]. Moment [Gia11c, Kla14]. Moments [Lat12, Leh11, MW12]. Monge [Bło13, Bou12a, Gue12a, Mon14b, PS13b]. Monodromies [NS12y]. Monodromy [MMA11m, NS12a]. Monomial [BGdC13, Mon13]. Monotone [BDR16d]. Monotonicity [Bur14f]. monotony [Pag13]. Monte [DP14, SS14]. Montecatini [Fas11]. Montel [Aud11c]. Montesinos [FKP13l]. Morgan [Ota15c]. Morphisms [Duc15a]. Morrey [YSY10h]. Motion [Bra14l, BÉL14, Bur14c, Bur14b, Bur14k, BCM15f, BCM15e, DH12, GQY14, Haj12, HR12, Mar12, Mar11a, Pic11b, Pra11a, Pra11b, Rum13, YY13l, YY13o]. Motivation [Maj13g]. Motivations [PP16c]. Move [ZK13]. Movement [FE13, HP13, LMP13a, PAMC13, Rey13a, WL13]. Movements [Bra14e, Bra14j, Bra14k]. MSR [AGJ^{+}13c, AGJ^{+}13q, AGJ^{+}13t, AGJ^{+}13z]. Muckenhoupt [DHHR11l]. Multi [Ben13q, Cha14, Die10j, MM13c, MM13f, MM13h, MM13i, PS11]. Multi-Dimensional [PS11]. Multi-Factor [Ben13q]. Multi-Layer [MM13c, MM13f, MM13h, MM13i]. Multi-Occupation [Cha14]. Multi-Term [Die10j]. Multicolor [HS14a]. Multifractal [MSU11h]. Multilevel [DG14j]. Multilocal [DGH16d]. Multiple [Maj13i, Maj13h, Maj14d, Maj14e, MMN13j]. Multiplicative [AAFM12a, Kha10f, Kha10u, Kha10t]. Multiplicity [Par10g, Que10l]. Multiplier [Tre14b]. Multiply [MMN13m]. Multiply-Perforated [MMN13m]. Multipolar [AGJ^{+}13t]. Multiscale [DG14e, BNQ^{+}12b]. Multisite [Enc13]. Multistatic [AGJ^{+}13u, AGJ^{+}13s]. Multistep [CH14]. Multitype [DG14e]. Must [BS12]. Mutants [DG14m, DG14l]. Mutation [DG14q, Eth11e]. Mutational [Lor10a, Lor10e, Lor10h, Lor10i]. Myrberg [HS11b].

Nagumo [Thi13]. Natural [Ban15a, Ban15c, DGH16l, LE11h]. Navier [Deb13, Gal13, Ser13b]. NCHO [Par10g]. Near [AGJ^{+}13p, Otw12g]. Near-Cloaking [AGJ^{+}13p]. Nearby [Sab13k, Sab13l]. Nearly [CT10]. Necessary [Nis14e]. Negative [KKR12b, Wei14b]. negligible [Éme13]. Neighbourhoods [Pos12h]. Network [BBRZ15, HSL13]. Networked [Kum15c]. Networking [Kum15b]. Networks [ABH^{+}13b, Est15, GK13, Lie15b, Lin13, Zua13b]. Neumann [Bur14e, Bur14f, GW11j, MMN13l]. Neurobiology [Tuc13c]. Neuronal [BBD13c]. Neurons [Tuc13c]. Neutral [DG14o]. Neutron [MK15]. Nevanlinna [Voj10b]. Newton [Kha10d, Neu10d, PP16o]. Newtonian [Růž13a, Růž13b]. Nice [Maj13j]. Nicely [Kum14a]. Nilmanifolds [Ang14e]. Nilpotent [Sti13r]. Nodal [Bur14g]. Noetherianity [Dra14]. Noise [DHI13d, DHI13i, Fla11g, Led14b, SBK^{+}13]. Non [Ang14b, Ben13c, CKM^{+}13b, DG14k, Éme13, FFM11b, FKP13d, Gru14, Has13, Hel13, KO13b, LE11j, Maj14f, MS16, Maw13, NS12q, Neu10h, Neu10l, NO13, Par10m, Riv12, Sam14p, Sti13d, Sti13i, Wei14b, YYH13h]. Non-abelian [Sam14p, Sti13d, Sti13i]. Non-Amenable [Ben13c]. Non-Autonomous [CKM^{+}13b, Maw13, NO13]. Non-central [Maj14f]. Non-Commutative [Par10m]. Non-convex [Gru14]. Non-doubling [YYH13h]. Non-ergodicity [DG14k]. Non-Inner [Neu10h, Neu10l]. Non-isolated [NS12q]. Non-Isothermal [FFM11b]. Non-Kähler [Ang14b]. Non-Linear [LE11j]. Non-Lipschitz [Has13]. Non-local [Hel13]. Non-negative [Wei14b]. Non-negligible [Éme13]. Non-Periodic [KO13b]. Non-prime [FKP13d]. Non-relativistic [Riv12]. Non-standard [MS16]. Nonabelian [Rei13i, Rei13j]. Nonamenable [Ben13j]. Nonautonomous [CPSS13, FM12d, KPR13, KP13c, KP13d, Pöt10c, Pöt10g, Pöt10h]. Noncollapsing [AH11b]. Noncommutative [BF12o, Def11b, Def11f, MV12f]. Nondegenerate [Nis14f]. Nonlinear [AGJ^{+}13v, Ban15d, BCD13, CGGM14b, DHI13d, FM12b, GGS10h, Gua14, IS13, LLMS12b, Neu10f, OZLK13f, OZLK13g, PT10b, Tuc13c, BCMV11b, DS10a]. Nonlinearities [BDR16c, BDR16d]. Nonlocal [VV13]. Nonnegative [Kha10p, Kha10v]. Nonsmooth [Dem10]. Norm [EL14, GPV12, Kol11i, MY11, vdG16j]. Norm-Penalty [vdG16j]. Normal [Frö12b, Frö12e, Frö12f, Nås11o, Sab13r]. Normalisation [BPR12b]. Normalized [BT13]. Norms [Kol11b, vdG16g]. Notation [BL15k]. Note [CLR13, GG14]. Notes [Bro12, FFR15, Tia16b]. Notion [Bos14n]. Notions [Ben13h]. NS [LR12h]. NS-Alpha- [LR12h]. NS-Omega-Deconvolution [LR12h]. Nuclear [Kol11i]. Null [HK14f]. Number [Bar11e, Boy15d, Boy15e, BCGG10e, Isa11i, Koe14c, MMA11l, PP16f]. Numbers [LE11e, LE11k, SSM14c]. Numerical [ABIL13b, AGJ^{+}13w, AGJ^{+}13x, AH12f, BDG^{+}14a, DRiSS10b, EH16b, HS11b, Kru14e, LR12b, MMN13n, PEP13b, RS10b, SAC^{+}14i, SS14, Wai13c, dVBSV16]. Numerics [BNQ^{+}12b, EZ12b].

Objects [GW11g]. Oblique [Ves11l]. Observed [Mas15b]. Obstacle [Sal12b]. Obstructions [Sti13c, Sti13q]. Obtained [BPR12b]. occasion [Chu10]. Occupation [Cha14, Wu12]. Ocone [BL13]. ODE [MC15]. Omega [LR12h]. One [BF12e, Eth11f, HKMR12e, Hu11h, Jac13b, Lie15h, MY11, Sab13s]. One-Dimensional [Jac13b]. One-Sided [HKMR12e]. Only [Kha10h, Kha10a]. Onsager [CP14]. Open [BSW13h, JPT16l, MMA11j, Sam14q, Sok13]. Open-Book [MMA11j]. Opening [HS11b]. Openness [HK14j]. Operation [BF12d, HM13, MMA11b]. Operations [KK14d, Sti13b]. Operator [DHHR11p, JPT16k, KM12b, MM13c, MM13i]. Operators [AH11c, Bou14c, Ceg13a, Ceg13b, DHHR11c, Die10a, Die10k, Fai12, JPT16m, KKP16d, KKP16f, KKP16e, KKP16l, KKP16k, KKP16g, KKP16q, KKP16i, KKP16h, LE11e, LE11f, LE11g, LR12c, Par10h, Pos12d, Que10n, Rud15, YYH13c, YYH13d, YYH13k, YSY10a, YSY10i]. Optical [Amm12b, Sch12]. Optics [CGGM14b, Gut14]. Optimal [AG13, BRVdL13, FV11b, Kru14f, Rey13a]. Optimality [SAC^{+}14d]. Optimisation [ABH^{+}13b]. Optimization [AGJ^{+}13v, BBRZ15, BB13, Bom10, Dem10, DS10a, GK13, Men13b, PT10b]. Option [Hob11, RS10b]. Options [Pag13]. Opto [Amm12b]. Opto-Acoustic [Amm12b]. Oracle [Kol11l]. Orbach [Kum14a]. Orbits [Mar11b]. Order [AN15f, AN15g, ADFT11, AAFM12b, Bar13, BV11, DG14c, FS12, GGS10g, GGS10h, GGS10i, GW11i, HJ13, MM13h, OZLK13h, Pil14, SSM14i]. Orders [MV12e]. Ordinals [DÉ13]. Ordinary [Lor10c, Maw13, Neu10o]. Organized [HSL13]. Oriented [Die10a, HP13, NS12s]. Original [Maj14h]. Ornstein [BOGP12]. Orthogonal [Bil14, Neu10p, Par10h, Ves11f]. Orthostatic [ONO13]. Oscillation [FKM14]. Oscillations [Ves11a]. Oscillator [Par10d]. Oscillators [CPSS13, Par10m]. Oscillatory [Lie15d, MS12]. Other [AM12e, DHHR11q, Fla11e, MMN13o]. Outline [Yse10i]. Outlook [Lie15u]. Ovals [BCGG10c]. Overdampedness [Ves11j]. Overrings [KK14f]. Overview [BF12i, Bur14h, DGH16o, JPT16l, Maj13k].

Package [DP11]. Packing [Duq10]. Pair [NS12x, NS12l]. Pairs [FO14b, Rab11, Ves11d]. Palais [HKMR12h]. Paley [YYH13k]. Panorama [DGW14c]. Paouris [Lat14]. Parabolic [Bur14i, FM12d, FM12e, FM12g, Hu11d, Hu11k, IS13, Kha10e, Kha10p, Kha10u, LLMS12b]. Paradigm [EH16b]. Parameter [ABMT13, BCAK13, FM12g, HT13b, TBBF13b]. Parameterized [Bra14l]. Parameters [DPPS14, Lie15f, Lor10a]. Parametric [Mas15c]. Paris [BCG^{+}13b, CCG^{+}11b]. Parity [AK11i]. Part [Par10k, Par10l, Pos12c, Pos12d, BFMV14o, BFMV14p]. Partial [ABCC^{+}12, Her13k, Kha10f, Maj13j, Tuc13c, vH14]. Partially [Jac13a]. Particle [Bou11]. Particles [BCM15c, Con13]. Partitioned [CH14]. Passage [BZ12, BM14, YY13j]. Passing [Wai15]. Past [Gal13, Ken15b]. Path [ACT11a, ACT11b, BSW13g, BSW13j, CT13, Gia11j, Her13e, Léo14, YY13o]. Paths [Bai14, GQY14, HR12, Le 11j, PAMC13, PP16d]. Patient [ONO13]. Pattern [MC15]. Patterns [Bad13, BORM14, Mor13b, Rey13a, VV13, WL13]. Paul [YY13m]. PBW [BF12l]. PDE [BCMV11b, BMOR13b, Rob14d]. PDEs [CGGM14b, EH16b, Fla11f, BORM14, DHHR11r, Gua14, Man12, Neu10a, Rob14c]. Peacocks [BPR12a, BPR12b, BPR12c, Pag13]. Penalization [Kol11c, Kol11i]. Penalized [Kol11j]. Penalty [vdG16j]. Horizontal [NS12y]. Shadow [DGH16r, DGH16u]. Thin [Boi16]. Tropical [Lit13b]. Percolation [BZ12, Ben13c, Ben13l, Ben13m, Ben13n, Ben13p, BM14, GK12]. Perforated [MMN13g, MMN13m]. Perforations [MMN13p]. Period [BMS11, Rei13k]. Periodic [KO13b, Sta12a, Sta12b, MMA11h, MMA11i]. Periodicity [Sta12f, Sta12g]. Permutable [GW11e]. Permutationally [NT12]. Permutations [PP16q]. Perspective [Bom10, LMP13a]. Perspectives [OX13]. Perturbation [Fla11f]. Perturbations [Ben13n, BL15m, BL15l, GGS10i, Pos12h, SSM14i]. Perturbed [FS13f, MMN13o, SSM14j]. Perverse [Sab13t]. Pest [PEP13b]. Pharmacodynamics [KS13]. Pharmacokinetics [KS13]. Phase [DG14o, DG14p, Mou12, Ves11m]. Phases [DG14c]. Phenomena [Gia11d]. Phenomenology [LR12b, LR12i]. Phenomenon [Gia11k, Kha10n]. Photoacoustic [ABJW12, GOZ12b]. Photoacoustics [KS12]. Physics [BBRZ15, FFV11, Her13e, Rud15]. Physiological [BBKK13, HVM13]. Physiology [BBK13c]. Pierre [Aud11j]. Pinching [AH11o]. Pinning [Gia11f, Gia11h, Mou12]. Pisier [Dod10f]. Pitman [YY13k]. Place [Sti13q, WL13]. Planar [Ben13e, Ben13o, BÉL14, DV13, Éme13, Lie15c]. Plane [Ben13g, DP11, NS12e, NS12f]. Planes [FO14f]. Plasma [Otw12b]. Plumbed [NS12s]. Plumber [Pos12i]. Plumbing [NS12s, NS12v]. Pluripotential [Dem13, PS13b, PBBD13b]. PM [KK14f]. PM-Spectra [KK14f]. Poincaré [Lie15j, Lie15t, Sch11b]. Point [Bad13, BB13, BM14, BY15, Bou11, BCM15c, Ceg13g, FO14m, Gia11c, Hei13, Hu11j, Kum14a, NS12a, PT10b, Sch15b, vdG16d]. Point-to-Point [BM14]. Points [Bar11e, BFMV14o, BFMV14p, Sti13u, CT10]. Poisson [Fai12, Le 11k, SBK^{+}13]. Poisson-Type [Fai12]. Poitou [Sti13i]. Polarization [AGJ^{+}13e, AGJ^{+}13o]. Polarized [BFMV14j]. Politecnico [Kum15c]. Polyharmonic [GGS10h]. Polyhedra [FKP13j]. Polyhedral [BMS11, Kok11b]. Polynomial [Bro10f, DS10b, FKP13h, NS12b, NS12t]. Polynomial-like [DS10b]. Polynomials [KKP16a, Maj14l]. Polystable [BFMV14o, BFMV14p]. Polytopes [Hug13]. Polytropic [Pad11f]. Population [Eth11h, PEP13b, WL13]. Porous [BDR16a, BDR16g, BPR16, BDR16i]. Portfolio [GMK13]. Portfolios [CWZ11]. Posedness [Kha10w, Ves11v, Nis14c]. Poset [GW11c, HMM^{+}13k]. Positional [PAMC13]. Positions [FMS14]. Positive [AH11m, JPT16f, JPT16g, Par10g, Sch10l, Zil14]. Positivity [BFMV14a, GGS10b, GGS10h, GGS10i, Le 11l]. Possible [vdG16c]. Potential [AGJ^{+}13r, BFMV14m, BH14, MM13c, MM13i]. Potentials [DMS14g, KW14, MM13f, MM13h]. Power [BF12e]. Practical [Kli14b]. Precision [vdG16a]. Predicting [ONO13]. Prediction [HTKS13, Kol11e]. Preface [MT14]. Preimages [HK14f]. Preliminaries [AGJ^{+}13y, Ang14g, AM12f, AH12g, Bar11k, BFMV14n, Bro10g, BCGG10f, Otw12h, YYH13m]. Preliminary [Bar10f]. Preparations [Nås11m, Nås11l]. Prerequisites [BF12k, Bro10h, YY13n]. Preserving [AH11m, DGH16a, FS12, GGS10h, GW11e]. presque [CT10]. Pressure [MSU11g, MSU11i]. Pressures [Hel13]. Prices [Hob11]. Pricing [Cré11, Her13c, RS10b, Tan11b]. Prime [CM11, MV12f, FKP13d]. Primer [BSW13i, NV12]. Primes [MV12d, MV12e]. Princeton [BCG^{+}13b, CCG^{+}11b]. Principal [KKP16q]. Principle [AH11r, Bur14i, Her13h, Ser14]. Prize [Aud11c, Aud11f, Aud11e]. Probabilistic [Bur14j, Del12b, Lie13, DFO13]. Probabilités [AT11i, Ben13b, Bur14c, DMLR11d, DMLR12, DMLR13, Eth11h, Fla11f, Gia11d, Kol11l, Kum14j, Le 11j, Tre14a, vdG16h]. Probabilities [DMLR11d, DMLR12, DMLR13]. Probability [BHJKX12, Del11, Gia11d, Her13i, Mec12, vdG16q]. Problem [BCAK13, Bou14c, Bou14a, Bur14h, FM12g, Gal13, GZ12, Hob11, Maj14g, MMN13f, MMN13h, MMN13i, Nis14c, Otw12c, Ves11n]. Problems [AGJ^{+}13-29, AM12k, BBRZ15, BCD13, BSW13h, Ceg13g, FM12c, FM12e, GGS10c, GGS10f, GGS10h, GGS10j, GZ14b, Gut14, KKP16p, Kol11e, Kol11l, Kol11j, LLMS12b, Lin10g, Lin10h, Lin10i, Lin10j, Lin10k, Lin10l, Lin10z, Lin10-27, Lor10h, Maj13g, Maj14c, Maw13, MMN13m, MMN13k, MM13f, MM13h, Neu10h, Neu10t, NPS13, OZLK13f, OX13, Rei13k, SAC^{+}14l, Sal12b, SBK^{+}13, SS13b, Lin10y]. Procedure [Gia11b]. Process [Has13, HS11b, Jac13a, KB15, Le 11k, MY11, Nås11p, Sok13, vdG16g]. Processes [AT11e, BZ14, BOGP12, BBB12, BY15, BSW13c, BSW13g, BSW13i, BSW13k, Bou11, CGC15, CT13, DV13, DR13, FK11a, GML11, Hei13, Her13l, Kol11b, Kol11d, Kor13b, KKR12b, KW14, Le 11b, Le 11m, LS14, MS12, Mas15b, Mas15c, Mie11, Mor13b, Pag13, QY12, RS10b, Rie11, Sch15b, Sie11, VGS13, VA12, MM13a]. processus [MM13a]. Product [Bou10r, BÉL14, Éme13, FKP13i, FKP13m, Neu10h, Neu10l]. Production [FK11a, Sie11]. Products [JPT16d, Maj13c, Que10j, Sam14r, Ves11g]. Profile [Ben13e]. Profiles [Lie15d]. Profinite [Sti13d]. Programming [Bom10, Fle10, Kol11h]. Progress [CKKR12b, DRiSS10b, OX13]. Projection [Ceg13b, Ceg13h, DGH16p]. Projections [DGH16b, DGH16r, Mec12, Neu10p, Par10h, Ves11l, Ves11f]. Projective [DS10b, Duc15a, FMS14, Otw12i, Rei13j]. Proof [AK11f, AH11o, BF12j, Duc15a, Kor13b, Maj13b, Maj13l, Maj13m, Maj13o, Maj13n, Maj13s, Maj14i, Naz12, NS12t, NS12u, PP16p, Ros11, Sok13, Tre14b]. Proofs [BF12k, BF12m, Bur14j, DG14n, DG14r, SSM14a]. Propagation [VV13]. Proper [BOGP12, Duc15a]. Properties [Bar10d, Bar11d, BFMV14a, BFMV14k, BSW13f, BSW13g, BSW13j, BG13, CT13, DHHR11f, DG12b, Fav15a, Gal13, Gia11j, HMM^{+}13g, HMM^{+}13i, HMM^{+}13h, HMM^{+}13j, Kaw13b, Kha10g, Kha10h, Léo14, Maj13h, MM13f, NS12m, Par10i, Rei13i, Rei13n, Ves11t, Wit14b, Wit14e, Wit14c, Wit14d]. Property [BY11, HMM^{+}13c, HMM^{+}13l, MNU12, Mor12e, Pil14, VA12]. Protoproducts [Sch10i]. Prüfer [KK14e]. Prym [Wei14k]. Pseudo [BFMV14m, MMA11h, MMA11i, RY14, YSY10i]. Pseudo-Brownian [RY14]. Pseudo-Differential [YSY10i]. Pseudo-periodic [MMA11h, MMA11i]. Pseudo-Stability [BFMV14m]. Pseudoconvex [Isa11k]. psh [DG12b]. Puiseux [PP16q]. Purely [Duc15a, MR14]. Push [Sab13m]. Push-Forward [Sab13m].

Quadratic [Bom10, Boy15b, Boy15d, Boy15g, Fle10, MW12, Sam14s, Ves11n, vdG16n]. Quadrature [AH12f]. Qualitative [Kha10u]. Quantitative [GOZ12b]. Quantized [MV12b, PS11]. Quantum [CFL11, Riv12, Sei12, Sol12, Yse10b, RSSS12b]. Quasi [DP14, Nås11b, Nås11h, Nås11n]. Quasi-Monte [DP14]. Quasi-Stationarity [Nås11h]. Quasi-stationary [Nås11b, Nås11n]. Quasipositive [GG14]. Quasisimple [Sam14t]. que [MM13a]. Questions [FFM11b, FKP13e, Gal13, Gla12, JPT16l]. Quillen [Bar11g, Sch11a]. Quotient [BFMV14h, BFMV14k, MMA11g, MMA11n]. Quotients [Bar11h].

R [Ken15b]. R. [Bes16]. Rademacher [Kol11b, Kol11d]. Radial [Yse10j]. Radial-Angular [Yse10j]. Radiative [FK11a, Sie11]. Radically [Her13c, Her13e, Her13h, Her13i, Her13j, Her13k, Her13l]. Radius [GZ14b]. Rafael [Wei14b]. Ramis [FS13j]. Random [AT11i, Bak13b, BB13, BSKŠ15, Ben13o, BY15, BH14, BS13a, BS13b, Cal13, DMZ12, Eth11e, FFR15, Hei13, Hug13, Jeu15, Ken15c, Kum14d, Kum14e, Kum14i, Maj13i, Maj13n, Maj13q, Maj14j, Maj14k, Man12, MY11, MSU11b, MUS11a, MSU11d, MSU11e, Mol13, Mou12, Nor13, RL15, Ser14, Spe12b, Spo13, SSR15, Sta13, VGS13, dFS13, Fla11f, Kum14j]. Randomised [NPS13]. Randomness [Ben13b]. Range [ZK13]. Ranges [GW11g]. Rank [GG14, Kol11i, Par10h, Sam14h]. Ranks [NS12v]. Rare [DG14m, DG14l, Wai13c]. Rascle [HJ13]. Rate [Hu11c, Mie11]. Rate-Independent [Mie11]. Rates [Kaw13d, PEP13b, Yse10c]. Rational [Bed10, Ber13, Bou10p, Sti13u, CT10]. rationnelles [CT10]. rationnels [CT10]. Raynaud [Bos14l]. Rays [BG11]. Re [DG14p]. Re-equilibration [DG14p]. Reachability [Kha10v]. Reaction [DHI13d, Kha10d, Lin10i, Lin10j, Lin10y, Lin10z, Lin10-27, MC15, VV13]. Reaction-Convection-Diffusion [Lin10i, Lin10j, Lin10y]. Reaction-Diffusion [DHI13d, Kha10d, Lin10z, Lin10-27, VV13]. Reaction-Diffusion-ODE [MC15]. Real [Ale12a, Boy15e, CGGM14b, CM11, MSU11i, Sab13n, Wil14]. Realizations [DGH16p]. Recognizing [FKP13m]. Reconstruction [AGJK12, AGJ^{+}13z, BMOR13b, Sch12]. Recovery [Kol11c, Kol11h, Kol11i, Kol11l]. Rectifiability [JN14]. Reduced [Bar11l]. Reduction [Bar11i, BSKŠ15, Maj13p, Sam14t, SSM14g, SSM14i, Sti13v, Ves11t]. Reductionist [TP13]. Refined [LE11i]. Reflected [Bur14k, Jac13a]. Reflecting [YY13k]. Reflection [Bro10c, Bro10e, Le 11l]. Reflectionless [JN14]. Reflections [BL13]. Refraction [Gut14]. Regimes [HM13]. Regular [AK11j, Bar11h, Bro10b, DHHR11d, GW11j, Sab13o]. Regularity [AH11n, CT13, Del12b, Lin12, Ser13b, YY13c, Yse10k, LLMS12b]. Regularization [Fla11g, KM12a, LR12g]. Regularizations [LR12h]. Regularizing [BG13]. Reinterpretation [PP16r]. Related [Ban15d, Gal13, Haj12, Lew12, Pos12i, SAC^{+}14l, YYH13k]. Relating [AGB15f, YY13i]. Relationship [BF12l]. Relative [Bec14e, FO14d]. relativistic [Riv12]. Relaxation [LR12j]. Relevant [Gia11k]. Reliability [Lin13]. Remark [GL12, Pra11a, Pra11b]. Remarks [CC11, FS13i, Gru14, Her13f, KM14b, Led14b]. Renewal [DR13]. Renormalization [Riv12]. Repellers [MSU11e]. Representation [BL15e, KKP16b, Pic11b, Rei13j, Rei13l, Ser11b, YY13k]. Representations [Bou10p, Boy15d, Boy15g, CC11, DG14b, Jak11, JPT16k, KKP16q, Le 11d, NS11, Ota15c, QY12, SSM14h, YY13i]. Representing [DGH16p]. Requiem [HJ13]. Resistance [Kum14g]. Resolvents [Ves11o]. Resonance [AGJ^{+}13a, Maw13]. Resonances [Ves11o]. Respect [GW11c, YY13j]. Respiration [Bru13]. Respiratory [BBK13c, Duf13, HT13b]. Response [AGJ^{+}13u, Lin13, ONO13]. Restricted [Bou14a]. Restrictive [Lor10g]. Result [Maj13m, Maj13p, Maj13s]. Results [AGJ^{+}13x, Bar10f, BFMV14d, DGH16s, DG14d, Deb13, Die10d, Die10m, Die10l, HY12h, Isa11f, Kum14d, Maj13d, Maj13k, Maj13r, Pos12e, Tan11b]. Resurrection [HJ13]. Reversal [AGJ^{+}13-29]. Reversible [Lie15c]. Review [Hu11i, Hu11j, Hu11k, Kru14g, SG13, Tan11b]. Revisited [Wei14i]. Revolution [GGS10k]. Rham [Sab13n]. Ricci [AH11o, BEG13b, BG13, Cao13, Gue13b, Sin16, SW13, Tia16b, Tia16c]. Riemann [AM12l, BK11, Bob11b, BCGG10c, BCGG10e, BCGG10g, BCGG10h, BCGG10i, DP11, Die10d, Die10k, FK11b, Kwo11, MMA11i, PP16t, PP16u, PP16s, Sab13c, Sab13h, Sab13q, Sab13p, Sab13o, Sab13t]. Riemann-Liouville [Die10d]. Riemannian [AH11f, AH11o, Kol12, Zil14]. Riesz [Que10j]. Rigid [Bos14e, Bos14j, Bos14l, Bos14n, Isa11j, MMN13j, Pad11c, Pad11f]. Rigidity [FKM14]. Rings [Bos14a, GW11j, HMM^{+}13b, HMM^{+}13c]. Risk [Kol11e, Kol11f, Kol11l, Kol11j, vdG16k]. RKHSs [JPT16c]. Rocco [Viv14b]. Roch [PP16s]. Rod [Kha10a]. Role [Mor13b, Rou14f]. Root [vdG16s]. Roots [PP16q]. Rost [Mor12j]. Rotating [Gal13]. Rotation [VA12]. Rotational [JR15]. Roth [Bil14]. Rough [Bai14, GQY14, Lej12]. Rowing [Kha10o]. RPF [MSU11j]. RPF-Theorem [MSU11j]. Rule [AM12c, KM12b]. Rules [WL13].

S. [Bes16, DÉ13]. Saint [AT11i, Ben13b, Bur14c, Eth11h, Fla11f, Gia11d, Kol11l, Kum14j, Le 11j, vdG16h, AFV12, BHJKX12, BLSC12, DMZ12, GK12, GKL12, INS12]. Saint-Flour [AT11i, Ben13b, Bur14c, Eth11h, Fla11f, Gia11d, Kol11l, Kum14j, Le 11j, vdG16h, AFV12, BHJKX12, BLSC12, DMZ12, GK12, GKL12, INS12]. Sample [BSW13g, BSW13j]. Sandra [Viv14b]. Sasakian [Dea14]. Satisfying [Fai12, Kha10d]. Scalar [BL15b, BL15a, GK13, KKP16k, MM13j, Ves11g]. Scalar-Valued [KKP16k, MM13j]. Scale [Bra14m, CKKR12b, KKR12b, Lau13, MMN13g, MMN13k]. Scaled [GW11c, HR12]. Scales [BL15f, Bra14c, LR12j, MM13j, Pos12c]. Scaling [BY11, BDMY11b, Bur14l, Kum14j]. scars [Aud11g]. Schemes [Bos14a, Duc15b, Lin10k, Lin10l]. Schmid [Mor12j]. Schoen [Bes16]. Scholes [Her13d]. School [ABB^{+}10b, BNQ^{+}12b, BCMV11b, CTSDV10a, DS10a, Fas11]. Schottky [CM11, HS11b, Sch11b]. Schrödinger [Yse10e, Bou12c, Bou14c, Kha10t]. Schur [HMM^{+}13l]. Sciences [Aud11c, Aud11e, Ban15a, Ban15c, KP13c, KP13d]. scientific [Aud11g]. SDE [Haj12, PS11]. SDEs [Has13]. Searching [Rey13a]. Second [AN15f, AN15g, BV11, HJ13, KM12b, OZLK13h]. Second-Order [HJ13, OZLK13h]. Section [Sti13p, Sti13u, Sti13t, Sti13s, Sti13y]. Sectional [Wei14b, Zil14]. Sections [Kol14, Sti13b, Sti13e, Sti13r, Sti13v, Sti13w, Sti13y, Sti13x]. Selected [Dod10d, Ser13b]. Selection [BCAK13, Bra14h, DG14q, Eth11g, Kol11j, Chu10]. Self [BT13, HSL13, HPRY11b, Hu11a, Maj14k, Par10g, VA12]. Self-Adjoint [Par10g]. Self-Organized [HSL13]. Self-Similar [HPRY11b, Hu11a, Maj14k, VA12]. Selfdecomposability [iS10]. Semi [BBB12, CG14]. Semi-Classical [BBB12]. Semicircle [Eis12]. Semiclassical [Par10j]. Semigroup [BDR16h, Par10e]. Semigroups [BSW13i, Rud15, BLSC12]. Semilinear [GGS10j, Hu11d, Kha10e, Kha10p, Kha10u, Kha10v, Kru14i]. Semimartingale [CWZ11, CS11, MW12, YY13c]. Séminaire [DMLR11d, DMLR12, DMLR13, Tre14a]. Seminar [KMM12, Led14a, DMLR11d, DMLR12, DMLR13]. Semistable [BFMV14o, BFMV14p, BFMV14t]. Separable [Dod10h]. September [CTSDV10a]. Sequence [BDMY11b, Bra14k]. Sequences [Que10f]. Sequential [Fle10]. Series [BF12d, BF12e, HY12b, KKP16q, Sch11b]. Set [BPR12c, BMOR13b, Dod10d, Éme13, FMS14, Sok13]. Set-Up [BPR12c]. Sets [AT11a, BB13, BSKŠ15, Bou10n, GZ14b, HK14f, Mol13]. Several [DS10b, MMN13f, Neu10q, YSY10j]. SFC [DGH16u]. Shade [DGH16r, DGH16u]. Shade/Shadow [DGH16r, DGH16u]. Shalen [Ota15c]. Shape [AGJ^{+}13q, Men13b]. Shapes [CM10g, CM10h, CM10t, CM10u]. Sheaf [Mor12j, Rei13m]. Sheaves [Bos14e, Mor12d, Mor12l, Sab13t]. Shell [EL14]. Shift [Gia11c, Ves11u]. Shifts [ZK13]. Shop [Pos12i]. Short [AH11p, BF12m, Ish13, Kru14g, Tre14b]. Short-Time [AH11p]. Sibuya [FS13j]. Sided [HKMR12e]. Sigma [Cha14]. Sigma-Field [Cha14]. Signal [Enc13]. Similar [HPRY11b, Hu11a, Maj14k, VA12]. Simonovits [AGB15a]. Simple [Bar11m, Bou10q, Kor13b, Lie13, Neu10r, Ves11p, YY13o]. Simulation [KB15, LR12f, Mar11a, Wai13c]. Simulations [MMN13n]. Simultaneous [Ves11q]. Single [Die10m, Die10l, MMN13h, MM13i]. Single-Term [Die10m, Die10l]. Singular [BFMV14q, Bou14b, Boy15f, MMA11j, SSM14i, SSM14j, Ves11r].

Singularities [DGH16a, DGH16m, DGH16n, DGH16p, Fla11e, NS12e, NS12f, NS12g, NS12n, Neu10r, Sin16]. Singularity [DGH16l, DGH16q, NS12q]. Singularly [FS13f, MMN13o, SSM14j]. SIS [Nås11b, Nås11d, Nås11c, Nås11a, Nås11e, Nås11h, Nås11m, Nås11l, Nås11n, Nås11q, Nås11r]. Site [Bou14b]. Size [PEP13b]. Skew [BÉL14]. Skew-Product [BÉL14]. Skorokhod [HPRY11b, Hob11]. Slepian [Led14b]. Slow [SSM14h, SSM14k]. Slower [HM13]. Slower-is-Faster-Effect [HM13]. SLP [HMM^{+}13d]. Smale [HKMR12d]. Small [AGJK12, AGJ^{+}13q, AGJ^{+}13-27, Bra14m, DHI13d, DHI13i, Fra14, KM12a, MMN13i, MMN13l, MMN13p, Sam14u]. Small-Ball [KM12a]. Small-Scale [Bra14m]. Smooth [AT11i, AAFM12a, DGH16b, DGH16r, Sab13p]. Smoothing [Gia11k]. Smoothness [MM13j]. Sobolev [DHHR11f, DHHR11n, DHHR11o, Neu10h, Neu10l, Neu10o, Neu10s, Neu10t]. Software [Sie12b]. Solution [DHI13i, DGQ12, Haj12]. Solutions [Bar13, BL15e, Has13, Hu11a, Hu11l, Ish13, KKP16p, Lej12, Lin10a, Lin10b, Lin10c, Lin10d, Lor10e, Mag14, MMN13k, Sta12a, Sta12b]. Solved [Gia11f]. Solving [AM12i, Isa11g]. Some [AT11h, AB12, AN15d, BOGP12, BZ12, BV11, Bil14, BPR12c, BF12m, BCGG10i, Bul15, CC11, Eth11h, FFR15, Fav15a, Gla12, LE11h, Le 12, Léo14, Maj13g, Maj13q, Maj13r, Maj14i, Maw13, Nås11a, Nås11o, Par10i, Par10j, Pil14, Pro13a, Rei13n, SAC^{+}14l, SS13b, Wai13c, Wai15, YY13p, vdG16r]. Source [AGJ^{+}13-29]. Sources [CCPP12]. Space [ACT11a, ACT11b, AN11, Bai14, BPHZ15, BS12, Dod10h, Gue12a, Hu11h, Kol12, Neu10e, Neu10f, Sti13y, Sti13x, Tud11, Ves11m, YYH13j, YYH13i, YYH13h, YYH13l]. Space-Time [BPHZ15]. Space-Valued [Bai14]. Spaces [Ban10d, Ban10e, Bar11a, Bar11d, Bar11m, Bos14b, Bos14e, Bos14l, Bos14n, Ceg13g, CT13, CS11, Def11b, Del11, DHHR11g, DHHR11n, DHHR11o, DHHR11q, DHHR11t, DS10b, Dod10b, Dod10d, Dod10i, Dod10h, Duc15b, FO14f, Fav15a, Frö12b, HKMR12g, HY12i, Jon15, Kru14h, Kru14g, LE11g, Le 11m, Lor10a, Lor10c, Lor10h, Lor10i, Mar11b, MM13e, Neu10h, Neu10l, OZLK13d, OZLK13e, Ota15b, Ota15c, Pos12c, Pos12d, Pos12i, Pos12j, Rou14b, Rou14g, Sab13n, Sea14, Tem15, YSY10d, YSY10e, YSY10k, BHJKX12]. Spacings [Bou14b, Bou14d]. Spanning [AK11k, Le 11h]. Sparse [Kol11c, Kol11h, Kol11l, vdG16n]. Sparsity [vdG16d, vdG16h, vdG16t]. Spatial [Bad13, Ban10e, DG14q, Eth11i, Ken15c, KB15, LMP13a, Lie15c, Spo13]. Spatially [KPR14]. SPDE [Bak13b]. Special [Die10l, Chu10]. Speciation [VV13]. Specific [ONO13, SSM14l]. Spectra [KK14f, Ves11b]. Spectral [AH12a, DH12, GZ14b, JPT16k, JPT16m, Kok11b, Maj14j, MK15, Par10k, Par10l, Par10m, Par10n, Pos12j, Que10k, Que10o, Que10l, Que10m, Que10n, Ves11s, Ves11t, Ves11u, MM13a]. spectrale [MM13a]. Spectrally [KKR12b]. Spectrum [Yse10l]. Sphere [AH11o, AH12d, AH12i, Bes16]. Spheres [KKP16e]. Spherical [AH12h, AH12i, Isa11i, Isa11k, Isa11c, Isa11d, Isa11b]. Spiking [Tuc13c]. Spine [DR13]. Split [FO14j]. Splitting [FO14a, FO14k].

Spots [Bur14h]. Spreading [Tuc13b]. Square [vdG16s]. Square-Root [vdG16s]. Squeeze [BS12]. Stability [BDG^{+}14a, BORM14, BFMV14f, BFMV14m, BFMV14r, Bra14m, Bra14n, CH14, FS12, GZ14b, Has13, Led14b, Pad11a, Pad11h, CKM^{+}13b]. Stabilization [AB12, AN15e, AN15f, AN15g, Zua13b]. Stabilizer [BFMV14s, Wei14l]. Stable [BFMV14o, BFMV14p, DGH16d, DGH16r, DV13, Duq10, Has13, KW14, LS14, Mar11a]. Standard [Dod10h, MS16, MMA11k]. Standardness [Lau11, Lau13]. Stanley [BD13b, Her13n]. Star [KK14d, Pos12i]. Starting [NS12a]. State [CT13, Gal13, Hu11l]. Statements [DGH16s]. States [HTKS13, Kha10v]. Stationarity [Nås11h]. Stationary [BBB12, Nås11d, Nås11c, Nås11m, Nås11l, SSR15, Nås11b, Nås11n]. Statistical [AGJ^{+}13u, AGJ^{+}13s, BBRZ15, Bul15, Spe12b]. Statistics [Bad13, Hei13, Ken15c, Maj13c, Maj13i, Maj13l, Spo13, Maj13r]. Steady [Gal13, Hu11l, Pad11a]. Steady-State [Gal13, Hu11l]. Steepest [Neu10e, Neu10f]. Stein [Sch15b, Tre14b]. Stereology [JR15, Kid13]. Stimulus [Lin13]. Stimulus-Response [Lin13]. Stochastic [BBD13c, BDR16a, BDR16g, BDR16i, BRVdL13, BPHZ15, Ben13q, BV11, CLR13, CS11, Deb13, DHI13j, DS13, Her13j, Her13m, Ken15c, Kru14h, Kru14i, MNU12, Mol13, Nås11h, Nås11p, QY12, Ros11, Rud15, SG13, SGS15, Spo13, SS13b, Thi13, Tuc13b, Tuc13c, Yuk13, BPR16]. Stokes [BCM15e, Deb13, Gal13, Sab13j, Sab13l, Sab13m, Sab13s, Sab13r, Sab13t, Ser13b]. Stokes-Filtered [Sab13l, Sab13m, Sab13s, Sab13r]. Stokes-Perverse [Sab13t]. Stopping [Sok13]. Strategies [Sch14]. Strategy [DG14l, Maj13s]. Stratification [BFMV14t, Rei13o]. Stratifications [DGH16t]. Stress [ONO13]. Stretch [BS12]. Strictly [Fra14]. Strictness [Sab13b]. String [Ban10e, Ban10f, Kha10g, Kha10h, Kha10v]. Strong [Bar11n, BPR12b, HMM^{+}13c, HMM^{+}13l, Kru14f, Kru14i, Nis14e, Sta13]. Strongly [Dod10i, Isa11k, Mor12j, Mor12l]. Structure [AGJ^{+}13u, Ben13k, Est15, Eth11i, MSU11e, NS12r, vdG16p, SGS15]. Structured [vdG16t]. Structures [AGJ^{+}13p, GW11i, HS14a, Rei13b, Rei13c, Sab13j]. Studies [Bro13, Bru13]. Study [FS13d, Nås11m, Nås11l, Neu10v, PEP13b, MM13a]. Stuff [Bos14k]. Sturm [AM12k]. Sturmfels [Viv14b]. Sub [vdG16c]. Sub-directions [vdG16c]. Subgroups [BFMV14s, Sam14j]. Subordinated [Maj14k]. Substitution [Que10g, Que10o]. Substitutions [Que10e]. Sufficient [Pro13a]. Suggestion [Lor10c]. Sum [Dod10a]. Summary [FO14l, Lie15u]. Summation [Def11b, Fai12]. Summer [ABB^{+}10b, BNQ^{+}12b, BCMV11b, CTSDV10a, DS10a, Fas11]. Sums [BT13, Maj13n, Maj13q, Maj13j]. Sup [Kol11b]. Sup-Norms [Kol11b]. Super [CC11, Kwo11]. Superalgebra [Ser11b]. Superalgebras [Jak11]. Superconductivity [Neu10u]. Supercritical [KPR14]. Supercurves [Rab11]. Superdimension [Ser11b]. Supergravity [ADFT11]. Supergroups [FG11, NS11]. Superirregularity [Bec14b]. Superlinear [Kha10e]. Superprocesses [KPR14]. Supersolutions [Lin12]. Superspaces [CFL11]. Supersymmetry [MS11, FFV11]. Supplementary [Neu10b, Neu10i, Neu10j]. Supported [Bou12b, Mon13]. Supremum [Maj13n, Maj13j]. Sure [HR12]. Surface [Bed10, DGH16r, Frö12g, NS12q]. Surfaces [BK11, Bob11b, BMS11, BCGG10c, BCGG10e, BCGG10g, BCGG10h, BCGG10i, DGH16b, DGH16t, DP11, FK11b, Frö12b, FKP13h, GGS10k, HY12i, Kok11b, Kwo11, MMA11i, Neu10m, PP16u, Rei13j, Sab13t]. Survey [Her13n]. SUSY [Spe12b]. Swans [SSM14d]. Swimming [Kha10c, Kha10n, Kha10o, Kha10s, Kha10w]. Symbols [BSW13d]. Symmetric [BF12n, CH14, Has13, Le 11b, Le 11m, Viv14c, Rou14g]. Symmetries [BCGG10c, BCGG10e, BCGG10g]. Symmetrization [Maj13t, vdG16u]. Symmetry [BCGG10h, BCGG10i, Dra14, GG14, Ken14, KKP16f, KKP16e, KKP16q, KKP16i, KKP16h]. Synchronous [Bur14m]. System [BCM15f, Duf13, HT13b, LE11j, SSM14i, KKP16i, KKP16h]. Systems [Aba10, AN15a, AN15h, AN15e, Ban15d, BBK13c, BL15c, BL15e, BL15i, BL15h, BL15m, BL15l, Frö12c, Gia11f, HVM13, Isa11g, Kaw13b, Kaw13i, Kaw13j, KPR13, KP13c, KP13d, Kum15c, Lie15d, MSU11b, MM13h, Nis14c, Nis14f, Nis14g, Pöt10c, Pöt10h, Que10f, Que10e, Que10g, Que10o, Que10m, RSSS12b, Rob14d, Sab13l, Sab13m, Sab13s, Sab13r, SAC^{+}14b, Sam14f, SSM14g, SSM14j, Thi13, Ves11a, Wei14c, dFS13, ABB^{+}10b, GKL12]. Syzygies [Con14].

Tacnodes [BFMV14g]. Tails [BFMV14f, BFMV14g, BFMV14r]. Takens [Lie15c, Lie15g]. Talagrand [CEL12]. Tale [YY13l]. Tanaka [Haj12]. Target [AGJ^{+}13q, AGJ^{+}13-28]. Tate [Bos14m, Sti13i]. Techniques [AK11f, AGJ^{+}13r, DHHR11s, DR13, Lam15, Lau15]. Teichmüller [Cap14, Wei14f, Wei14k, Wei14m, Wei14n, Wei14o]. Tensor [Bou10r, MMN13j, MMN13h, MMN13i]. Tensors [AGJ^{+}13e, AGJ^{+}13o]. Term [CH14, Die10j, Die10m, Die10l, Kha10d, PP16e]. Terme [Fas11, Viv14b]. Terminology [AK11d]. Terms [Kha10e, Sam14f, Sti13b]. Tessellations [Cal13, RL15, VGS13]. Testing [vdG16h]. Their [Bos14b, Bos14a, Bro10e, Con14, DGH16d, GM14, JPT16g, Maj13h, Pos12h, Rud15, SG13, Sam14d, YYH13h, CT10, Kum14j, RL15]. Theirs [Rei13e]. Themes [Pil14]. Theorem [AH11f, AH11o, Bes16, BC14, BFMV14m, Bło12a, BF12j, BF12m, BF12o, Dod10j, FS13j, Gru14, Her13h, JN14, Kor13b, Maj13b, MMA11l, MSU11j, Mec12, Naz12, PP16s, Sam14j, Sch10m, Tre14b, Tud11, vdG16d, BF12l]. Theorems [Bra14n, BS13a, Bur14j, DG14r, FY14, Her13c, Hu11j, KK14a, Maj13o, Maj13n, Maj14e, Maj14f, Mey11, Sta13, YSY10g, Yuk13]. Theoretic [Mec12, Rei13l]. Theoretical [HTKS13]. Theories [Hu11d, Hu11i, Hu11k, Mor12k, Pil14]. Theory [AK11f, AH12b, Ban10c, Ban10e, Ban10f, BDR16a, Bar13, BRVdL13, BCM^{+}11b, BSG11, Bil14, BFMV14j, CKM^{+}13b, Cap14, CS14, CKKR12b, Cor11, DGH16l, Def11c, Def11f, Dem13, DFO13, DP14, Die10m, Dod10d, DGW14c, DR13, DRiSS10b, FO14k, FS13e, HMM^{+}13b, HMM^{+}13g, HMM^{+}13k, Her13i, Her13k, Her13l, JPT16m, KKP16b, Koe14b, Koe14c, Kok11b, Kum14f, Kum15b, KKR12b, Le 12, Léo12b, MT14, MV12d, MV12f, MMA11l, Mey11, MM13j, MK15, Mol13, Par10m, PBBD13b, PS13b, Pöt10c, Pro13b, Que10m, Que10n, Rei13j, Sab13i, Sab13h, Sch11a, Ser13b, Voj10b, VV13, Wai13a, WMR13a, Wil14, YY13b, YY13d, YY13l, vdG16g, vdD14, FO14a, GK12, GKL12]. Thermodynamical [MUS11a]. Theta [Sch11b]. Thick [Boi16]. Thick/ Thin [Boi16]. Thin [EL14]. Thin-Shell [EL14]. Thomason [Sch11a]. Three [Boi16, FO14a, FO14j, FO14m, HKMR12d, HKMR12e, NS12w, Sin16]. Three-Dimensional [FO14a, FO14j, FO14m, Sin16]. Three-Manifolds [Boi16, HKMR12d, HKMR12e]. Thresholds [Nås11q]. Tight [Sch10j, Sch10k, Sch10l]. Tightness [BZ12]. Time [AGJ^{+}13-29, AH11n, AH11p, BPHZ15, BL15f, Bra14c, Bur14k, CP14, DG14k, Fra14, Hu11e, Ish13, Kha10a, KPR13, LR12j, Mar12, Mar11a, Nås11e, Neu10a, Ros11, Sok13, VA12, YY13o]. Time-Dependent [Kha10a, Neu10a]. Time-Inhomogeneous [CP14]. Time-Reversal [AGJ^{+}13-29]. Times [DHI13a, DHI13b, LS14, Wu12, YY13c, YY13j, YY13l]. Tits [RTW15a]. Toffalori [Koe14b]. Tomographies [Amm12b]. Tomography [AGJ^{+}13-29, GOZ12b, Sch12]. Tool [DG14b]. Tools [Duc15a, MS16, Par10j, SAC^{+}14i]. Topics [BCM^{+}11b, BF12o, Dod10d, Frö12b, Lew12, Pad11g, Pad11h, Ser13b, SD10, CDG^{+}13b]. Topographic [CM10k, CM10r, CM10s]. Topological [AT11g, AT11i, Bar11a, Bar11d, BCM^{+}11b, CM10r, CM10s, Cor11, DGH16n, MMA11b, MMA11m, PP16m, Sti13x]. Topology [AT11a, AT11h, Bar11a, CWZ11, CS11, Mor12a, NS12w, NS12x]. Torelli [Rei13k]. Tori [Bou14c]. Toric [Di 14]. Torino [Kum15c]. Torsors [Sti13s]. Torus [FKM14]. Totally [Boy15e]. Tour [Viv14c]. Traces [Nor13]. Tracking [AGJ^{+}13-28]. Trading [Sch14]. Traffic [HJ13, Hel13, HM13, HTKS13]. Tranches [CJL11]. Transaction [GMK13, Sch14]. Transcendental [Duc15a, PP16b]. Transcritical [Lie15k, Lie15v]. Transduction [Enc13]. Transfer [DHHR11s, FK11a, Sie11]. Transfers [Mor12h]. Transform [Ale12a]. Transformation [Pro13a, Pro14]. Transformations [BL15i, BL15h, BSW13k]. Transforms [AM12i, BGT14, FS12]. Transient [MY11]. Transition [BDR16h, DHI13b]. Transitions [DGH16d, DGH16r, DGH16u, DGH16v]. Transitive [BS12, Ben13k]. Transport [AG13, BG11, FV11b, Fla11h, HP13, MK15]. Tree [BM14, CM10e, CM10f, CM10g, CM10h, CM10t, CM10u]. Trees [AK11k, Bak13b, CM10p, CM10q, CCPP12, DRiSS10b, Duq10, Kor13b, Le 11h]. Triangle [Lor10g]. Trick [AH11q]. Tricomi [Neu10v]. Triebel [YSY10e, YSY10h]. Trigonometric [LE11a, LE11k, LE11l]. Triples [Pos12c]. Triplet [RY14]. Truncates [LR12j]. Truncation [Ban10e]. Tube [Isa11i, Isa11k, Isa11c, Isa11d, Isa11b, Isa11l]. Tubular [Pos12h]. Tug [Man12]. Tug-of-War [Man12]. Tumbling [Lie15a]. Tumor [KS13]. Turbulence [LR12b]. Tutorial [Bre13, MS16]. TV [BO13, SBK^{+}13]. Twin [Wit14e]. Twins [SAC^{+}14m]. Twist [KO13b]. Twisted [Wei14f, Wei14n, Wei14o]. Two [BB13, BPR12a, DG14m, DG14n, Eth11d, HPRY11b, Kum14a, Neu10c, Nis14g, Pos12d, YY13q]. Two-Point [Kum14a]. Type [BDMY11b, BSW13g, DG14g, DMS14d, Die10a, Fai12, FY14, FS13j, Gua14, KM14b, Lie15a, NT12, Otw12c, Que10k, JPT16n]. Types [Bar11m, Bar11n, BCGG10h, BCGG10i, DG14c, DG14m, DG14n, Eth11d].

UCLA [FFV11]. Uhlenbeck [AH11q, BOGP12]. Ultrametric [DMS14d, DMS14h, DMS14i]. Ultraproducts [Sch10m, Sch10o]. Ultrasensitivity [Enc13]. Ultrasound [AGJK12, Amm12b]. Unbounded [AN15f, AN15g, Lej12]. Unconditional [Lat12, NT12]. Undecidability [Koe14c]. undecided [MM13a]. Uniform [JN14, MMN13l, MMN13p, Nås11r, Sch10n]. Uniformisation [Bru10]. Uniformizations [HS11b]. Uniformizing [CM11]. Uniformly [CP14]. Unimodality [LS14]. Uniqueness [Ben13p, Die10d, Fla11d, Fla11e, MMA11n]. Unit [AH12i]. Unitary [DH12, Fai12, NS11, Que10n]. Units [Sti13g]. Universal [BFMV14e, Mey11, Tre14b]. Universe [Lie15a]. Unramified [Mor12k, Mor12l]. Unscented [ABMT13]. Unstressed [TBBF13b]. Upper [CS14, Kaw13c]. USA [FFV11]. Use [Sch10o]. User [AG13, GMK13]. Using [ÀFR13, AGJ^{+}13t, BD13b, CM11, Die10a, Kum14g, Van13b, vdG16f, Ken15b].

v [Cor11]. Validation [BBK13c, Duf13]. Valuation [MV12f]. Valuations [Ale12b, KK14e, MV12b]. Value [BCD13, GGS10h, MMN13f, MMN13h, MMN13i, MMN13m, MM13f]. Valued [Bai14, KKP16k, KKP16g, KKP16n, KKP16o, MM13j, Nor13, vdD14]. Vanishing [AGJ^{+}13p]. Vapnik [Maj13u]. Variable [CCPP12, DHHR11o, DHHR11t, Hu11h, Kha10g, Kha10h, LE11g]. Variables [DS10b, Maj13q, NS12w, Nis14g]. Variance [AGB15a]. Variants [LE11h]. Variation [Le 11d, Pra11a, Pra11b]. Variational [BDR16i, Bra14i, Neu10t, Rum13]. Variétés [CT10]. Varieties [Wei14k, CT10]. Various [PEP13b]. Vector [CLR13, EL14, Lej12, Lor10a, Lor10h, Mor12c]. Vectors [Lat12]. Verdera [JN14]. Vergne [Rou14g, Rou14e]. Verrès [Kum15c]. Vershik [Lau13]. Versus [Mor12h]. Vertex [Ben13k, GL12]. Vertical [NS12y]. Vertical/ Horizontal [NS12y]. Very [BPR12b]. Via [EL14, BR15, DGH16l, DR13, HPRY11b, NS12b, NS12v]. Vibrating [Kha10g, Kha10h, Kha10v]. View [Bos14l, DGH16r, JR15, Sch15b]. Views [DGH16v]. Viot [DG14q]. Viral [HA13]. Viscosity [Bar13, Ish13]. Viscous [BCM15d, EH16c]. Vitushkin [FY14]. Vitushkin-Type [FY14]. Vlasov [BCM15f, BCM15g]. Volatility [Ben13q, Her13d]. Volume [AGJ^{+}13-27, Lin10m, Lin10n, TBBF13b]. Volumes [FMS14]. vs [HTKS13, JPT16n].

Walk [BH14, Kum14d, MY11, Ser14]. Walks [Kum14j, Rey13a]. Walsh [Haj12]. War [Man12]. Watson [Kor13b]. Wave [EZ12b, Kha10a, Yse10k]. Wavelet [Ber12]. Waves [Zua13b]. Weak [AH11r, Kru14d, Kru14i, Lau15]. Weakly [BS13a]. Weighted [Kum14e, Kum14k, LE11e]. Weil [Boy15d, Boy15g]. Well [Kha10w, Nis14c, Ves11v]. Well-Posedness [Kha10w, Ves11v, Nis14c]. Weyl [HMM^{+}13l, Par10o]. Which [Bou12b, Éme13]. Whitney [MM13e, NT12]. Wick [Maj14l]. Wiener [Maj13h, Maj14d, Maj14e, Maj14h, Mar12, Rie11, Tud11, YY13l]. Wilking [AH11g]. Williams [YY13q]. Willmore [GGS10k]. Winding [DG14r]. Winding-Up [DG14r]. Windings [DV13]. Windows [AAFM12b]. Wireless [Kum15b]. Without [Duc15a, FKP13d, Pil14, Lie15f]. Witt [Mor12k]. Work [Maj13d, Maj13p, Maj13s]. Worked [vdG16r]. Worked-Out [vdG16r]. Works [PP16r]. Workshop [DFO13]. World [PP16g]. Worst [vdG16c]. Wright [DG14m, DG14n].

XL [Fla11f, Gia11d, Kum14j]. XLI [Ben13b]. XLIII [Bur14c, DMLR11d]. XLIV [DMLR12]. XLV [vdG16h, DMLR13]. XLVI [Tre14a]. XXXIX [AT11i, Eth11h]. XXXVIII [Kol11l, Le 11j].

Yau [Bło12a].

Zagier [HY12d]. Zero [Bro10c, Lie15w, MY11, Sch10j, Sch10k]. Zero-Hopf [Lie15w]. Zero-One [MY11]. Zeros [Ser14]. Zeta [Par10n]. Zippin [Dod10j]. Zoo [BO13]. Zooplankton [Mor13b]. Zygmund [MM13j].

References

Artstein-Avidan:2012:MMC

[AAFM12a] Shiri Artstein-Avidan, Dmitry Faifman, and Vitali Milman. On multiplicative maps of continuous and smooth functions. Lecture Notes in Mathematics, 2050:35–59, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

            springer.com/chapter/10.1007/978-3-642-29849-3_3/.

Artstein-Avidan:2012:OIC

[AAFM12b] Shiri Artstein-Avidan, Dan Florentin, and Vitali Milman. Order isomorphisms on convex functions in windows. Lecture Notes in Mathematics, 2050:61–122, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

            com/chapter/10.1007/978-3-642-29849-3_4/.

Alabau-Boussouira:2012:SRA

[AB12] Fatiha Alabau-Boussouira. On some recent advances on stabilization for hyperbolic equations. Lecture Notes in Mathematics, 2048:1–100, 2012. CODEN LNMAA2. ISBN 3-642-27892-2 (print), 3-642-27893-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-27893-8_1/.

Abate:2010:DHL

[Aba10] Marco Abate. Discrete holomorphic local dynamical systems. Lecture Notes in Mathematics, 1998:1–55, 2010. CO-DEN LNMAA2. ISBN 3-642-13170-0 (print), 3-642-13171-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-13171-4_1.pdf.

Abate:2010:FM

[ABB^{+}10a] Marco Abate, Eric Bedford, Marco Brunella, Tien-Cuong Dinh, Dierk Schleicher, and Nessim Sibony. Front matter. Lecture Notes in Mathematics, 1998:i–xiii, 2010. CODEN LNMAA2. ISBN 3-642-13170-0 (print), 3-642-13171-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-13171-4/1.pdf.

Abate:2010:HDS

[ABB^{+}10b] Marco Abate, Eric Bedford, Marco Brunella, Tien-Cuong Dinh, Dierk Schleicher, and Nessim Sibony, editors. Holomorphic dynamical systems: lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, July 7–12, 2008, volume 1998 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-13170-0 (print), 3-642-13171-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1998; QA331 .H62 2010; QA3 .L28 no. 1998. URL http://link.

springer.com/book/10.1007/978-3-642-13171-4;   http:/
/www.springerlink.com/content/978-3-642-13171-4.

Alabau-Boussouira:2012:FM

[ABBG^{+}12] Fatiha Alabau-Boussouira, Roger Brockett, Olivier Glass, Jérôme Le Rousseau, and Enrique Zuazua. Front matter. Lecture Notes in Mathematics, 2048:i–xiii, 2012. CODEN LNMAA2. ISBN 3-642-27892-2 (print), 3-642-27893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-27893-8/1.pdf.

Alabau-Boussouira:2012:CPD

[ABCC^{+}12] Fatiha Alabau-Boussouira, Piermarco Cannarsa, Jean-Michel Coron, et al., editors. Control of Partial Differential Equations: Cetraro, Italy 2010, volume 2048 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-27892-2 (print), 3-642-27893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2048; QA3 .L28 no. 2048;

QA374 .C56 2010. URL http://link.springer.com/book/
10.1007/978-3-642-27893-8;   http://www.springerlink.
com/content/978-3-642-27893-8.

Ambrosio:2013:FM

[ABH^{+}13a] Luigi Ambrosio, Alberto Bressan, Dirk Helbing, Axel Klar, and Enrique Zuazua. Front matter. Lecture Notes in Mathematics, 2062:i–xiv, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-32160-3/1.pdf.

Ambrosio:2013:MOF

[ABH^{+}13b] Luigi Ambrosio, Alberto Bressan, Dirk Helbing, Axel Klar, and Enrique Zuazua, editors. Modelling and Optimisation of Flows on Networks: Cetraro, Italy 2009, volume 2062 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2062; QA3 .L28 no. 2062; QA402 .M63 2013. URL http://link.

springer.com/book/10.1007/978-3-642-32160-3;   http:/
/www.springerlink.com/content/978-3-642-32160-3.

Achdou:2013:FM

[ABIL13a] Yves Achdou, Guy Barles, Hitoshi Ishii, and Grigory L. Litvinov. Front matter. Lecture Notes in Mathematics, 2074:i– xv, 2013. CODEN LNMAA2. ISBN 3-642-36432-2 (print), 3-642-36433-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-36433-4/1.pdf.

Achdou:2013:HJE

[ABIL13b] Yves Achdou, Guy Barles, Hitoshi Ishii, and Grigory L. Litvinov, editors. Hamilton–Jacobi Equations: Approximations, Numerical Analysis and Applications: Cetraro, Italy 2011, volume 2074 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-36432-2 (print), 3-642-36433-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2074; QA3 .L28 no. 2074; QA316 .C56 2011. URL http:

//link.springer.com/book/10.1007/978-3-642-36433-4;
http://www.springerlink.com/content/978-3-642-36433-4.

Ammari:2012:PIA

[ABJW12] Habib Ammari, Elie Bretin, Vincent Jugnon, and Abdul Wahab. Photoacoustic imaging for attenuating acoustic media. Lecture Notes in Mathematics, 2035:57–84, 2012. CODEN LNMAA2. ISBN 3-642-22989-1 (print), 3-642-22990-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-22990-9_3/.

Attarian:2013:AUK

[ABMT13] Adam Attarian, Jerry J. Batzel, Brett Matzuka, and Hien Tran. Application of the unscented Kalman filtering to parameter estimation. Lecture Notes in Mathematics, 2064:75– 88, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32882-4_4/.

Achdou:2013:FDM

[Ach13] Yves Achdou. Finite difference methods for mean field games. Lecture Notes in Mathematics, 2074:1–47, 2013. CODEN LN-MAA2. ISBN 3-642-36432-2 (print), 3-642-36433-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-36433-4_
1/.

Arnaudon:2011:HDC

[ACT11a] Marc Arnaudon, Koléhè Abdoulaye Coulibaly, and Anton Thalmaier. Horizontal diffusion in C^{1} path space. Lecture Notes in Mathematics, 2006:73–94, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-15217-7_2/.

Arnaudon:2011:HDP

[ACT11b] Marc Arnaudon, Koléhè Abdoulaye Coulibaly, and Anton Thalmaier. Horizontal diffusion in C^{1} path space. Lecture Notes in Mathematics, 2006:73–94, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-15217-7_2/.

Andrianopoli:2011:BHF

[ADFT11] L. Andrianopoli, R. D’Auria, S. Ferrara, and M. Trigiante. Black holes and first order flows in supergravity. Lecture Notes in Mathematics, 2027:17–43, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21744-9_2/.

AlvarezMontaner:2013:LCU

[ÀFR13] Josep Àlvarez Montaner and Oscar Fernández-Ramos. Local cohomology using Macaulay2. Lecture Notes in Mathematics, 2083: 179–185, 2013. CODEN LNMAA2. ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-38742-5_6/.

Azencott:2012:LDS

[AFV12] Robert Azencott, Mark Iosifovitch Freidlin, and S. R. Srinivasa Varadhan. Large deviations at Saint-Flour, volume ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-33199-8 (paperback). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 371 pp. LCCN QA3 .L28 no. ?? English translation of Grandes déviations et applications, Lecture Notes in Mathematics 774 (1980) ISBN 3-540-09741-4.

Ambrosio:2013:UGO

[AG13] Luigi Ambrosio and Nicola Gigli. A user’s guide to optimal transport. Lecture Notes in Mathematics, 2062:1–155, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32160-3_1/.

Alonso-Gutierrez:2015:AKL

[AGB15a] David Alonso-Gutiérrez and Jesús Bastero. Approaching the Kannan–Lovász–Simonovits and Variance Conjectures, volume 2131 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LNMAA2. ISBN 3-319-13262-8 (print), 3-319-13263-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 150

(est.) pp. LCCN ???? URL http://www.springerlink.com/
content/978-3-319-13263-1.

Alonso-Gutierrez:2015:BM

[AGB15b] David Alonso-Gutiérrez and Jesús Bastero. Back matter. Lecture Notes in Mathematics, 2131:137–150, 2015. CODEN LNMAA2. ISBN 3-319-13262-8 (print), 3-319-13263-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-319-13263-1/1.pdf.

Alonso-Gutierrez:2015:C

[AGB15c] David Alonso-Gutiérrez and Jesús Bastero. The conjectures. Lecture Notes in Mathematics, 2131:1–64, 2015. CODEN LNMAA2. ISBN 3-319-13262-8 (print), 3-319-13263-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-13263-1_1.
pdf.

Alonso-Gutierrez:2015:FM

[AGB15d] David Alonso-Gutiérrez and Jesús Bastero. Front matter. Lecture Notes in Mathematics, 2131:i–x, 2015. CODEN LNMAA2. ISBN 3-319-13262-8 (print), 3-319-13263-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-13263-1/1.pdf.

Alonso-Gutierrez:2015:ME

[AGB15e] David Alonso-Gutiérrez and Jesús Bastero. Main examples. Lecture Notes in Mathematics, 2131:65–101, 2015. CO-DEN LNMAA2. ISBN 3-319-13262-8 (print), 3-319-13263-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-13263-1_2.pdf.

Alonso-Gutierrez:2015:RC

[AGB15f] David Alonso-Gutiérrez and Jesús Bastero. Relating the conjectures. Lecture Notes in Mathematics, 2131:103–135, 2015. CODEN LNMAA2. ISBN 3-319-13262-8 (print), 3-319-13263-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-13263-1_3.pdf.

Ammari:2013:ARC

[AGJ^{+}13a] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Anomalous resonance cloaking. Lecture Notes in Mathematics, 2098:287–299, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02585-8_16/.

Ammari:2013:BM

[AGJ^{+}13b] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Back matter. Lecture Notes in Mathematics, 2098:351–382, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-319-02585-8/1.pdf.

Ammari:2013:DIM

[AGJ^{+}13c] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Detection and imaging from MSR measurements. Lecture Notes in Mathematics, 2098:189–202, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02585-8_9/.

Ammari:2013:DIF

[AGJ^{+}13d] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Direct imaging functionals for inclusions in the continuum approximation. Lecture Notes in Mathematics, 2098:173–188, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02585-8_8/.

Ammari:2013:FDG

[AGJ^{+}13e] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Frequency dependent generalized polarization tensors. Lecture Notes in Mathematics, 2098:133–142, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02585-8_5/.

Ammari:2013:FMa

[AGJ^{+}13f] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Front matter. Lecture Notes in Mathematics, 2098:i–xvii, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-319-02585-8/1.pdf.

Ammari:2013:FMb

[AGJ^{+}13g] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Front matter. Lecture Notes in Mathematics, 2098:1, 2013. CODEN LN-MAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-02585-8/
1/1.pdf.

Ammari:2013:FMc

[AGJ^{+}13h] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Front matter. Lecture Notes in Mathematics, 2098:95, 2013. CODEN LN-MAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-02585-8/
2/1.pdf.

Ammari:2013:FMd

[AGJ^{+}13i] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Front matter. Lecture Notes in Mathematics, 2098:143, 2013. CODEN LN-MAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-02585-8/
3/1.pdf.

Ammari:2013:FMe

[AGJ^{+}13j] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Front matter. Lecture Notes in Mathematics, 2098:171, 2013. CODEN LN-MAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-02585-8/
4/1.pdf.

Ammari:2013:FMf

[AGJ^{+}13k] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Front matter. Lecture Notes in Mathematics, 2098:203, 2013. CODEN LN-MAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-02585-8/
5/1.pdf.

Ammari:2013:FMg

[AGJ^{+}13l] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Front matter. Lecture Notes in Mathematics, 2098:227, 2013. CODEN LN-MAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-02585-8/
6/1.pdf.

Ammari:2013:FMh

[AGJ^{+}13m] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Front matter. Lecture Notes in Mathematics, 2098:267, 2013. CODEN LN-MAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-02585-8/
7/1.pdf.

Ammari:2013:FMi

[AGJ^{+}13n] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Front matter. Lecture Notes in Mathematics, 2098:301, 2013. CODEN LN-MAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-02585-8/
8/1.pdf.

Ammari:2013:GPT

[AGJ^{+}13o] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Generalized polarization tensors. Lecture Notes in Mathematics, 2098:115– 131, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02585-8_4/.

Ammari:2013:GVS

[AGJ^{+}13p] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. GPT- and Svanishing structures for near-cloaking. Lecture Notes in Mathematics, 2098:269–286, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02585-8_15/.

Ammari:2013:ISS

[AGJ^{+}13q] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Imaging small shape deformations of an extended target from MSR measurements. Lecture Notes in Mathematics, 2098:239–252, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02585-8_
13/.

Ammari:2013:LPT

[AGJ^{+}13r] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Layer potential techniques. Lecture Notes in Mathematics, 2098:51–94, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02585-8_2/.

Ammari:2013:MSM

[AGJ^{+}13s] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Mathematical and Statistical Methods for Multistatic Imaging, volume 2098 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CO-DEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 361 pp. LCCN QA3 .L28 no. 2098. URL http://link.

springer.com/book/10.1007/978-3-319-02585-8;   http:/
/www.springerlink.com/content/978-3-319-02585-8.

Ammari:2013:MMU

[AGJ^{+}13t] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. MSR matrices using multipolar expansions. Lecture Notes in Mathematics, 2098: 163–169, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02585-8_7/.

Ammari:2013:MRM

[AGJ^{+}13u] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Multistatic response matrix: Statistical structure. Lecture Notes in Mathematics, 2098:145–161, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02585-8_6/.

Ammari:2013:NOA

[AGJ^{+}13v] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Nonlinear optimization algorithms. Lecture Notes in Mathematics, 2098:253– 266, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02585-8_14/.

Ammari:2013:NI

[AGJ^{+}13w] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Numerical implementations. Lecture Notes in Mathematics, 2098:303–330, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02585-8_17/.

Ammari:2013:NR

[AGJ^{+}13x] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Numerical results. Lecture Notes in Mathematics, 2098:331–349, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02585-8_
18/.

Ammari:2013:P

[AGJ^{+}13y] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Preliminaries. Lecture Notes in Mathematics, 2098:3–50, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02585-8_1/.

Ammari:2013:RGM

[AGJ^{+}13z] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Reconstruction of GPTs from MSR measurements. Lecture Notes in Mathematics, 2098:205–210, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02585-8_10/.

Ammari:2013:SVE

[AGJ^{+}13-27] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Small volume expansions. Lecture Notes in Mathematics, 2098:97–113, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02585-8_3/.

Ammari:2013:TIT

[AGJ^{+}13-28] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Target identification and tracking. Lecture Notes in Mathematics, 2098:211– 226, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02585-8_11/.

Ammari:2013:TRD

[AGJ^{+}13-29] Habib Ammari, Josselin Garnier, Wenjia Jing, Hyeonbae Kang, Mikyoung Lim, Knut Sølna, and Han Wang. Time-reversal and diffraction tomography for inverse source problems. Lecture Notes in Mathematics, 2098:229–238, 2013. CODEN LNMAA2. ISBN 3-319-02584-8 (print), 3-319-02585-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02585-8_12/.

Ammari:2012:DRM

[AGJK12] Habib Ammari, Josselin Garnier, Vincent Jugnon, and Hyeonbae Kang. Direct reconstruction methods in ultrasound imaging of small anomalies. Lecture Notes in Mathematics, 2035: 31–55, 2012. CODEN LNMAA2. ISBN 3-642-22989-1 (print), 3-642-22990-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22990-9_2/.

Andrews:2011:FGF

[AH11a] Ben Andrews and Christopher Hopper. The F-functional and gradient flows. Lecture Notes in Mathematics, 2011:161–171, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-16286-2_10/.

Andrews:2011:FLN

[AH11b] Ben Andrews and Christopher Hopper. The W-functional and local noncollapsing. Lecture Notes in Mathematics, 2011:173– 191, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-16286-2_11/.

Andrews:2011:AIC

[AH11c] Ben Andrews and Christopher Hopper. An algebraic identity for curvature operators. Lecture Notes in Mathematics, 2011:193– 221, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-16286-2_12/.

Andrews:2011:BMb

[AH11d] Ben Andrews and Christopher Hopper. Back matter. Lecture Notes in Mathematics, 2011:287–296, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-16286-2/1.pdf.

Andrews:2011:BMa

[AH11e] Ben Andrews and Christopher Hopper. Background material. Lecture Notes in Mathematics, 2011:11–47, 2011. CODEN LN-MAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-16286-2_
2/.

Andrews:2011:CTR

[AH11f] Ben Andrews and Christopher Hopper. The compactness theorem for Riemannian manifolds. Lecture Notes in Mathematics, 2011:145–159, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-16286-2_9/.

Andrews:2011:CCB

[AH11g] Ben Andrews and Christopher Hopper. The cone construction of Böhm and Wilking. Lecture Notes in Mathematics, 2011:223– 233, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-16286-2_13/.

Andrews:2011:EC

[AH11h] Ben Andrews and Christopher Hopper. Evolution of the curvature. Lecture Notes in Mathematics, 2011:63–82, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-16286-2_

Andrews:2011:FA

[AH11i] Ben Andrews and Christopher Hopper. The final argument. Lecture Notes in Mathematics, 2011:259–269, 2011. CODEN LN-MAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-16286-2_
15/.

Andrews:2011:FM

[AH11j] Ben Andrews and Christopher Hopper. Front matter. Lecture Notes in Mathematics, 2011:i–xvii, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-16286-2/1.pdf.

Andrews:2011:HM

[AH11k] Ben Andrews and Christopher Hopper. Harmonic mappings. Lecture Notes in Mathematics, 2011:49–62, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-16286-2_3/.

Andrews:2011:I

[AH11l] Ben Andrews and Christopher Hopper. Introduction. Lecture Notes in Mathematics, 2011:1–9, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-16286-2_1/.

Andrews:2011:PPI

[AH11m] Ben Andrews and Christopher Hopper. Preserving positive isotropic curvature. Lecture Notes in Mathematics, 2011:235– 258, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-16286-2_14/.

Andrews:2011:RLT

[AH11n] Ben Andrews and Christopher Hopper. Regularity and longtime existence. Lecture Notes in Mathematics, 2011:137–143, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-16286-2_8/.

Andrews:2011:RFR

[AH11o] Ben Andrews and Christopher Hopper. The Ricci Flow in Riemannian Geometry: A Complete Proof of the Differentiable 1/4-Pinching Sphere Theorem, volume 2011 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 296 pp. LCCN

QA3 .L28 no. 2011. URL http://link.springer.com/book/
10.1007/978-3-642-16286-2;   http://www.springerlink.
com/content/978-3-642-16286-2.

Andrews:2011:STE

[AH11p] Ben Andrews and Christopher Hopper. Short-time existence. Lecture Notes in Mathematics, 2011:83–95, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-16286-2_5/.

Andrews:2011:UT

[AH11q] Ben Andrews and Christopher Hopper. Uhlenbeck’s trick. Lecture Notes in Mathematics, 2011:97–113, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-16286-2_6/.

Andrews:2011:WMP

[AH11r] Ben Andrews and Christopher Hopper. The weak maximum principle. Lecture Notes in Mathematics, 2011:115–135, 2011. CODEN LNMAA2. ISBN 3-642-16285-1 (print), 3-642-16286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-16286-2_7/.

Atkinson:2012:ASM

[AH12a] Kendall Atkinson and Weimin Han. Applications: Spectral methods. Lecture Notes in Mathematics, 2044:211–236, 2012. CODEN LNMAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-25983-8_6/.

Atkinson:2012:AT

[AH12b] Kendall Atkinson and Weimin Han. Approximation theory. Lecture Notes in Mathematics, 2044:131–163, 2012. CODEN LN-MAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25983-8_
4/.

Atkinson:2012:BM

[AH12c] Kendall Atkinson and Weimin Han. Back matter. Lecture Notes in Mathematics, 2044:237–244, 2012. CODEN LNMAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-25983-8/1.pdf.

Atkinson:2012:DIS

[AH12d] Kendall Atkinson and Weimin Han. Differentiation and integration over the sphere. Lecture Notes in Mathematics, 2044: 87–130, 2012. CODEN LNMAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-25983-8_3/.

Atkinson:2012:FM

[AH12e] Kendall Atkinson and Weimin Han. Front matter. Lecture Notes in Mathematics, 2044:i–ix, 2012. CODEN LNMAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-25983-8/1.pdf.

Atkinson:2012:NQ

[AH12f] Kendall Atkinson and Weimin Han. Numerical quadrature. Lecture Notes in Mathematics, 2044:165–210, 2012. CODEN LN-MAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25983-8_
5/.

Atkinson:2012:P

[AH12g] Kendall Atkinson and Weimin Han. Preliminaries. Lecture Notes in Mathematics, 2044:1–9, 2012. CODEN LNMAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25983-8_1/.

Atkinson:2012:SH

[AH12h] Kendall Atkinson and Weimin Han. Spherical harmonics. Lecture Notes in Mathematics, 2044:11–86, 2012. CODEN LNMAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25983-8_2/.

Atkinson:2012:SHA

[AH12i] Kendall Atkinson and Weimin Han. Spherical Harmonics and Approximations on the Unit Sphere: an Introduction, volume 2044 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-25982-0 (print), 3-642-25983-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 244 pp. LCCN QA3 .L28 no. 2044; QA406 .A85 2012.

URL http://link.springer.com/book/10.1007/978-3-642-
25983-8;   http://www.springerlink.com/content/978-3-
642-25983-8;   http://www.springerlink.com/content/u58550t8417n/

.

Akiyama:2011:F

[AK11a] Jin Akiyama and Mikio Kano. [a, b]-factorizations. Lecture Notes in Mathematics, 2031:193–218, 2011. CODEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21919-1_5/.

Akiyama:2011:FF

[AK11b] Jin Akiyama and Mikio Kano. (g, f)-factors and [a, b]-factors. Lecture Notes in Mathematics, 2031:143–191, 2011. CODEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21919-1_
4/.

Akiyama:2011:BM

[AK11c] Jin Akiyama and Mikio Kano. Back matter. Lecture Notes in Mathematics, 2031:337–356, 2011. CODEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-21919-1/1.pdf.

Akiyama:2011:BT

[AK11d] Jin Akiyama and Mikio Kano. Basic terminology. Lecture Notes in Mathematics, 2031:1–14, 2011. CODEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21919-1_1/.

Akiyama:2011:CF

[AK11e] Jin Akiyama and Mikio Kano. Component factors. Lecture Notes in Mathematics, 2031:253–293, 2011. CODEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21919-1_7/.

Akiyama:2011:FFG

[AK11f] Jin Akiyama and Mikio Kano. Factors and Factorizations of Graphs: Proof Techniques in Factor Theory, volume 2031 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CO-DEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 353 pp. LCCN QA3 .L28 no. 2031. URL http://link.

springer.com/book/10.1007/978-3-642-21919-1;   http:/
/www.springerlink.com/content/978-3-642-21919-1.

Akiyama:2011:FM

[AK11g] Jin Akiyama and Mikio Kano. Front matter. Lecture Notes in Mathematics, 2031:i–xii, 2011. CODEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-21919-1/1.pdf.

Akiyama:2011:MF

[AK11h] Jin Akiyama and Mikio Kano. Matchings and 1-factors. Lecture Notes in Mathematics, 2031:15–67, 2011. CODEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-21919-1_2/.

Akiyama:2011:PF

[AK11i] Jin Akiyama and Mikio Kano. Parity factors. Lecture Notes in Mathematics, 2031:219–251, 2011. CODEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21919-1_6/.

Akiyama:2011:RFF

[AK11j] Jin Akiyama and Mikio Kano. Regular factors and f-factors. Lecture Notes in Mathematics, 2031:69–141, 2011. CODEN LN-MAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21919-1_
3/.

Akiyama:2011:ST

[AK11k] Jin Akiyama and Mikio Kano. Spanning trees. Lecture Notes in Mathematics, 2031:295–336, 2011. CODEN LNMAA2. ISBN 3-642-21918-7 (print), 3-642-21919-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21919-1_8/.

Alesker:2012:CTI

[Ale12a] Semyon Alesker. The α-cosine transform and intertwining integrals on real Grassmannians. Lecture Notes in Mathematics, 2050:1–21, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-29849-3_1/;   http://link.springer.
com/content/pdf/bfm:978-3-642-29849-3/1.pdf.

Alesker:2012:MV

[Ale12b] Semyon Alesker. On modules over valuations. Lecture Notes in Mathematics, 2050:23–34, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29849-3_2/.

Annaby:2012:BM

[AM12a] Mahmoud H. Annaby and Zeinab S. Mansour. Back matter. Lecture Notes in Mathematics, 2056:295–318, 2012. CODEN LN-MAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-642-30898-7/
1.pdf.

Annaby:2012:FDE

[AM12b] Mahmoud H. Annaby and Zeinab S. Mansour. Fractional qdifference equations. Lecture Notes in Mathematics, 2056:223– 270, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30898-7_8/.

Annaby:2012:FLR

[AM12c] Mahmoud H. Annaby and Zeinab S. Mansour. Fractional q-Leibniz rule and applications. Lecture Notes in Mathematics, 2056:175–199, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30898-7_6/.

Annaby:2012:FM

[AM12d] Mahmoud H. Annaby and Zeinab S. Mansour. Front matter. Lecture Notes in Mathematics, 2056:i–xix, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-30898-7/1.pdf.

Annaby:2012:OFC

[AM12e] Mahmoud H. Annaby and Zeinab S. Mansour. Other qfractional calculi. Lecture Notes in Mathematics, 2056:147–173, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30898-7_5/.

Annaby:2012:P

[AM12f] Mahmoud H. Annaby and Zeinab S. Mansour. Preliminaries. Lecture Notes in Mathematics, 2056:1–39, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30898-7_1/.

Annaby:2012:DE

[AM12g] Mahmoud H. Annaby and Zeinab S. Mansour. q-Difference equations. Lecture Notes in Mathematics, 2056:41–71, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30898-7_
2/.

Annaby:2012:FCE

[AM12h] Mahmoud H. Annaby and Zeinab S. Mansour. q-Fractional Calculus and Equations, volume 2056 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-30897-X (paperback), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 318 pp. LCCN QA3 .L28 no.

2056. URL http://link.springer.com/book/10.1007/978-
3-642-30898-7;   http://www.springerlink.com/content/
978-3-642-30898-7.

Annaby:2012:ITS

[AM12i] Mahmoud H. Annaby and Zeinab S. Mansour. q-Integral transforms for solving fractional q-difference equations. Lecture Notes in Mathematics, 2056:271–293, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30898-7_9/.

Annaby:2012:MLF

[AM12j] Mahmoud H. Annaby and Zeinab S. Mansour. q-Mittag-Leffler functions. Lecture Notes in Mathematics, 2056:201–222, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30898-7_7/.

Annaby:2012:SLP

[AM12k] Mahmoud H. Annaby and Zeinab S. Mansour. q-Sturm– Liouville problems. Lecture Notes in Mathematics, 2056:73– 105, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30898-7_3/.

Annaby:2012:RLF

[AM12l] Mahmoud H. Annaby and Zeinab S. Mansour. Riemann–Liouville q-fractional calculi. Lecture Notes in Mathematics, 2056:107– 146, 2012. CODEN LNMAA2. ISBN 3-642-30897-X (print), 3-642-30898-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30898-7_4/.

Ammari:2012:FM

[Amm12a] Habib Ammari. Front matter. Lecture Notes in Mathematics, 2035:i–ix, 2012. CODEN LNMAA2. ISBN 3-642-22989-1 (print), 3-642-22990-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-22990-9/1.pdf.

Ammari:2012:MMB

[Amm12b] Habib Ammari. Mathematical Modeling in Biomedical Imaging II: Optical, Ultrasound, and Opto-Acoustic Tomographies, volume 2035 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-22989-1 (print), 3-642-22990-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 160 pp. LCCN QA3 .L28 no. 2035. URL http://link.

springer.com/book/10.1007/978-3-642-22990-9;   http:/
/www.springerlink.com/content/978-3-642-22990-9.

Attal:2011:DAF

[AN11] Stéphane Attal and Ion Nechita. Discrete approximation of the free Fock space. Lecture Notes in Mathematics, 2006:379– 394, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_16/.

Ammari:2015:ABC

[AN15a] Kaïs Ammari and Serge Nicaise. Asymptotic behaviour of concrete dissipative systems. Lecture Notes in Mathematics, 2124:73– 146, 2015. CODEN LNMAA2. ISBN 3-319-10899-9 (print), 3-319-10900-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10900-8_4.pdf.

Ammari:2015:BM

[AN15b] Kaïs Ammari and Serge Nicaise. Back matter. Lecture Notes in Mathematics, 2124:169–180, 2015. CODEN LNMAA2. ISBN 3-319-10899-9 (print), 3-319-10900-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-319-10900-8/1.pdf.

Ammari:2015:FM

[AN15c] Kaïs Ammari and Serge Nicaise. Front matter. Lecture Notes in Mathematics, 2124:i–xi, 2015. CODEN LNMAA2. ISBN 3-319-10899-9 (print), 3-319-10900-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-10900-8/1.pdf.

Ammari:2015:SB

[AN15d] Kaïs Ammari and Serge Nicaise. Some backgrounds. Lecture Notes in Mathematics, 2124:1–35, 2015. CODEN LNMAA2. ISBN 3-319-10899-9 (print), 3-319-10900-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-10900-8_1.
pdf.

Ammari:2015:SES

[AN15e] Kaïs Ammari and Serge Nicaise. Stabilization of Elastic Systems by Collocated Feedback, volume 2124 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LNMAA2. ISBN 3-319-10899-9 (print), 3-319-10900-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 180 (est.) pp. LCCN ???? URL

http://www.springerlink.com/content/978-3-319-10900-8.

Ammari:2015:SSOa

[AN15f] Kaïs Ammari and Serge Nicaise. Stabilization of second order evolution equations by a class of unbounded feedbacks. Lecture Notes in Mathematics, 2124:37–60, 2015. CODEN LNMAA2. ISBN 3-319-10899-9 (print), 3-319-10900-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-10900-8_2.pdf.

Ammari:2015:SSOb

[AN15g] Kaïs Ammari and Serge Nicaise. Stabilization of second order evolution equations with unbounded feedback with delay. Lecture Notes in Mathematics, 2124:61–71, 2015. CO-DEN LNMAA2. ISBN 3-319-10899-9 (print), 3-319-10900-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10900-8_3.pdf.

Ammari:2015:SD

[AN15h] Kaïs Ammari and Serge Nicaise. Systems with delay. Lecture Notes in Mathematics, 2124:147–168, 2015. CODEN LNMAA2. ISBN 3-319-10899-9 (print), 3-319-10900-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-10900-8_5.
pdf.

Angella:2014:BM

[Ang14a] Daniele Angella. Back matter. Lecture Notes in Mathematics, 2095:233–264, 2014. CODEN LNMAA2. ISBN 3-319-02440-X (print), 3-319-02441-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-02441-7/1.pdf.

Angella:2014:CACb

[Ang14b] Daniele Angella. Cohomological Aspects in Complex Non-Kähler Geometry, volume 2095 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-02440-X (print), 3-319-02441-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxvi + 262 pp. LCCN QA3 .L28 no. 2095.

URL http://link.springer.com/book/10.1007/978-3-319-
02441-7;   http://www.springerlink.com/content/978-3-
319-02441-7.

Angella:2014:CACa

[Ang14c] Daniele Angella. Cohomology of almost-complex manifolds. Lecture Notes in Mathematics, 2095:151–232, 2014. CODEN LN-MAA2. ISBN 3-319-02440-X (print), 3-319-02441-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02441-7_
4/.

Angella:2014:CCM

[Ang14d] Daniele Angella. Cohomology of complex manifolds. Lecture Notes in Mathematics, 2095:65–94, 2014. CODEN LNMAA2. ISBN 3-319-02440-X (print), 3-319-02441-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02441-7_2/.

Angella:2014:CN

[Ang14e] Daniele Angella. Cohomology of nilmanifolds. Lecture Notes in Mathematics, 2095:95–150, 2014. CODEN LNMAA2. ISBN 3-319-02440-X (print), 3-319-02441-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02441-7_3/.

Angella:2014:FM

[Ang14f] Daniele Angella. Front matter. Lecture Notes in Mathematics, 2095:i–xxv, 2014. CODEN LNMAA2. ISBN 3-319-02440-X (print), 3-319-02441-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-02441-7/1.pdf.

Angella:2014:PAC

[Ang14g] Daniele Angella. Preliminaries on (almost-)complex manifolds. Lecture Notes in Mathematics, 2095:1–63, 2014. CODEN LN-MAA2. ISBN 3-319-02440-X (print), 3-319-02441-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02441-7_
1/.

Adler:2011:ATE

[AT11a] Robert J. Adler and Jonathan E. Taylor. Algebraic topology of excursion sets: a new challenge. Lecture Notes in Mathematics, 2019:107–114, 2011. CODEN LNMAA2. ISBN 3-642-19579-2 (paperback), 3-642-19580-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-19580-8_6/.

Adler:2011:BM

[AT11b] Robert J. Adler and Jonathan E. Taylor. Back matter. Lecture Notes in Mathematics, 2019:115–122, 2011. CODEN LNMAA2. ISBN 3-642-19579-2 (paperback), 3-642-19580-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-19580-8/1.pdf.

Adler:2011:FM

[AT11c] Robert J. Adler and Jonathan E. Taylor. Front matter. Lecture Notes in Mathematics, 2019:i–viii, 2011. CODEN LNMAA2. ISBN 3-642-19579-2 (paperback), 3-642-19580-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-19580-8/1.pdf.

Adler:2011:GKF

[AT11d] Robert J. Adler and Jonathan E. Taylor. The Gaussian kinematic formula. Lecture Notes in Mathematics, 2019:59–85, 2011. CODEN LNMAA2. ISBN 3-642-19579-2 (paperback), 3-642-19580-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-19580-8_4/.

Adler:2011:GP

[AT11e] Robert J. Adler and Jonathan E. Taylor. Gaussian processes. Lecture Notes in Mathematics, 2019:13–35, 2011. CODEN LNMAA2. ISBN 3-642-19579-2 (paperback), 3-642-19580-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-19580-8_2/.

Adler:2011:I

[AT11f] Robert J. Adler and Jonathan E. Taylor. Introduction. Lecture Notes in Mathematics, 2019:1–12, 2011. CODEN LNMAA2. ISBN 3-642-19579-2 (paperback), 3-642-19580-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-19580-8_1/.

Adler:2011:ATI

[AT11g] Robert J. Adler and Jonathan E. Taylor. On applications: Topological inference. Lecture Notes in Mathematics, 2019:87–106, 2011. CODEN LNMAA2. ISBN 3-642-19579-2 (paperback), 3-642-19580-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-19580-8_5/.

Adler:2011:SGS

[AT11h] Robert J. Adler and Jonathan E. Taylor. Some geometry and some topology. Lecture Notes in Mathematics, 2019:37–58, 2011. CODEN LNMAA2. ISBN 3-642-19579-2 (paperback), 3-642-19580-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-19580-8_3/.

Adler:2011:TCS

[AT11i] Robert J. Adler and Jonathan E. Taylor. Topological Complexity of Smooth Random Functions: École d’Été de Probabilités de Saint-Flour XXXIX — 2009, volume 2019 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-19579-2 (paperback), 3-642-19580-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 122 pp. LCCN QA3 .L28 no. 2019; QA274.4 .A435 2011. URL http://link.

springer.com/book/10.1007/978-3-642-19580-8;   http:/
/www.springerlink.com/content/978-3-642-19580-8.

Audin:2011:AFJ

[Aud11a] Michèle Audin. After Fatou and Julia. Lecture Notes in Mathematics, 2014:115–133, 2011. CODEN LNMAA2. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-17854-2_5/.

Audin:2011:BM

[Aud11b] Michèle Audin. Back matter. Lecture Notes in Mathematics, 2014: 237–332, 2011. CODEN LNMAA2. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-17854-2/1.pdf.

Audin:2011:FJM

[Aud11c] Michèle Audin, editor. Fatou, Julia, Montel: The Great Prize of Mathematical Sciences of 1918, and Beyond, volume 2014 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 332 pp. LCCN QA3 .L28 no. 2014; QA26 .A8313 2011. URL http:

//link.springer.com/book/10.1007/978-3-642-17854-2;
http://www.springerlink.com/content/978-3-642-17854-2.

Audin:2011:FM

[Aud11d] Michèle Audin. Front matter. Lecture Notes in Mathematics, 2014:i–viii, 2011. CODEN LNMAA2. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-17854-2/1.pdf.

Audin:2011:GPM

[Aud11e] Michèle Audin. The great prize of mathematical sciences. Lecture Notes in Mathematics, 2014:59–89, 2011. CODEN LNMAA2.

ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-17854-2_3/.

Audin:2011:GPF

[Aud11f] Michèle Audin. The great prize, the framework. Lecture Notes in Mathematics, 2014:13–57, 2011. CODEN LNMAA2. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-17854-2_2/.

Audin:2011:HSS

[Aud11g] Michèle Audin. History’s scars — a scientific controversy … in 1965. Lecture Notes in Mathematics, 2014:193–235, 2011. CODEN LNMAA2. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-17854-2_7/.

Audin:2011:I

[Aud11h] Michèle Audin. Introduction. Lecture Notes in Mathematics, 2014:1–12, 2011. CODEN LNMAA2. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-17854-2_1/.

Audin:2011:M

[Aud11i] Michèle Audin. The memoirs. Lecture Notes in Mathematics, 2014:91–114, 2011. CODEN LNMAA2. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-17854-2_4/.

Audin:2011:PF

[Aud11j] Michèle Audin. On Pierre Fatou. Lecture Notes in Mathematics, 2014:135–192, 2011. CODEN LNMAA2. ISBN 3-642-17853-7 (paperback), 3-642-17854-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-17854-2_6/.

Baddeley:2013:SPP

[Bad13] Adrian Baddeley. Spatial point patterns: Models and statistics. Lecture Notes in Mathematics, 2068:49–114, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33305-7_
3/.

Bailleul:2014:FDB

[Bai14] Ismaël Bailleul. Flows driven by Banach space-valued rough paths. Lecture Notes in Mathematics, 2123:195–205, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_7.pdf.

Bakhtin:2013:BM

[Bak13a] Yuri Bakhtin. Back matter. Lecture Notes in Mathematics, 2068: 421–448, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-33305-7/1.pdf.

Bakhtin:2013:GLR

[Bak13b] Yuri Bakhtin. Geometry of large random trees: SPDE approximation. Lecture Notes in Mathematics, 2068:399–420, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33305-7_12/.

Banagl:2010:BM

[Ban10a] Markus Banagl. Back matter. Lecture Notes in Mathematics, 1997:211–223, 2010. CODEN LNMAA2. ISBN 3-642-12588-3 (print), 3-642-12589-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-12589-8/1.pdf.

Banagl:2010:FM

[Ban10b] Markus Banagl. Front matter. Lecture Notes in Mathematics, 1997:i–xvi, 2010. CODEN LNMAA2. ISBN 3-642-12588-3 (print), 3-642-12589-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-12589-8/1.pdf.

Banagl:2010:HT

[Ban10c] Markus Banagl. Homotopy theory. Lecture Notes in Mathematics, 1997:1–106, 2010. CODEN LNMAA2. ISBN 3-642-12588-3 (print), 3-642-12589-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12589-8_1.pdf.

Banagl:2010:IS

[Ban10d] Markus Banagl. Intersection spaces. Lecture Notes in Mathematics, 1997:107–189, 2010. CODEN LNMAA2. ISBN 3-642-12588-3 (print), 3-642-12589-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12589-8_2.pdf.

Banagl:2010:ISS

[Ban10e] Markus Banagl. Intersection Spaces, Spatial Homology Truncation, and String Theory, volume 1997 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-12588-3 (print), 3-642-12589-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 217 pp. LCCN QA3 .L28 no.

1997. URL http://link.springer.com/book/10.1007/978-
3-642-12589-8;   http://www.springerlink.com/content/
978-3-642-12589-8.

Banagl:2010:ST

[Ban10f] Markus Banagl. String theory. Lecture Notes in Mathematics, 1997:191–209, 2010. CODEN LNMAA2. ISBN 3-642-12588-3 (print), 3-642-12589-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12589-8_3.pdf.

Bansal:2014:AAC

[Ban14] Nikhil Bansal. Algorithmic aspects of combinatorial discrepancy. Lecture Notes in Mathematics, 2107:425–457, 2014. CO-DEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-04696-9_6.pdf.

Banasiak:2015:KMN

[Ban15a] Jacek Banasiak. Kinetic models in natural sciences. Lecture Notes in Mathematics, 2126:133–198, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11322-7_4.pdf.

Banda:2015:BM

[Ban15b] Mapundi Kondwani Banda. Back matter. Lecture Notes in Mathematics, 2126:495–496, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-319-11322-7/1.pdf.

Banda:2015:EEA

[Ban15c] Mapundi Kondwani Banda. Evolutionary Equations with Applications in Natural Sciences, volume 2126 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 496 (est.) pp. LCCN ???? URL

http://www.springerlink.com/content/978-3-319-11322-7.

Banda:2015:NHS

[Ban15d] Mapundi Kondwani Banda. Nonlinear hyperbolic systems of conservation laws and related applications. Lecture Notes in Mathematics, 2126:439–493, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11322-7_9.pdf.

Baricz:2010:BM

[Bar10a] Árpád Baricz. Back matter. Lecture Notes in Mathematics, 1994: 187–212, 2010. CODEN LNMAA2. ISBN 3-642-12229-9 (print), 3-642-12230-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-12230-9/1.pdf.

Baricz:2010:FM

[Bar10b] Árpád Baricz. Front matter. Lecture Notes in Mathematics, 1994: i–xiv, 2010. CODEN LNMAA2. ISBN 3-642-12229-9 (print), 3-642-12230-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-12230-9/1.pdf.

Baricz:2010:GBF

[Bar10c] Árpád Baricz. Generalized Bessel Functions of the First Kind, volume 1994 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-12229-9 (print), 3-642-12230-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 206 pp. LCCN QA3 .L28 no. 1994. URL http://link.

springer.com/book/10.1007/978-3-642-12230-9;   http:/
/www.springerlink.com/content/978-3-642-12230-9.

Baricz:2010:GPG

[Bar10d] Árpád Baricz. Geometric properties of generalized Bessel functions. Lecture Notes in Mathematics, 1994:23–69, 2010. CO-DEN LNMAA2. ISBN 3-642-12229-9 (print), 3-642-12230-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12230-9_2.pdf.

Baricz:2010:IIB

[Bar10e] Árpád Baricz. Inequalities involving Bessel and hypergeometric functions. Lecture Notes in Mathematics, 1994:71–186, 2010. CODEN LNMAA2. ISBN 3-642-12229-9 (print), 3-642-12230-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12230-9_3.pdf.

Baricz:2010:IPR

[Bar10f] Árpád Baricz. Introduction and preliminary results. Lecture Notes in Mathematics, 1994:1–22, 2010. CODEN LNMAA2. ISBN 3-642-12229-9 (print), 3-642-12230-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-12230-9_1.
pdf.

Barmak:2011:ATF

[Bar11a] Jonathan A. Barmak. Algebraic Topology of Finite Topological Spaces and Applications, volume 2032 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 170 pp. LCCN

QA3 .L28 no. 2032. URL http://link.springer.com/book/
10.1007/978-3-642-22003-6;   http://www.springerlink.
com/content/978-3-642-22003-6.

Barmak:2011:ACC

[Bar11b] Jonathan A. Barmak. The Andrews–Curtis conjecture. Lecture Notes in Mathematics, 2032:137–150, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-22003-6_11/.

Barmak:2011:BM

[Bar11c] Jonathan A. Barmak. Back matter. Lecture Notes in Mathematics, 2032:151–170, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-22003-6/1.pdf.

Barmak:2011:BTP

[Bar11d] Jonathan A. Barmak. Basic topological properties of finite spaces. Lecture Notes in Mathematics, 2032:19–35, 2011. CODEN LN-MAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22003-6_
2/.

Barmak:2011:FPL

[Bar11e] Jonathan A. Barmak. Fixed points and the Lefschetz number. Lecture Notes in Mathematics, 2032:129–135, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22003-6_
10/.

Barmak:2011:FM

[Bar11f] Jonathan A. Barmak. Front matter. Lecture Notes in Mathematics, 2032:i–xvii, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-22003-6/1.pdf.

Barmak:2011:GAC

[Bar11g] Jonathan A. Barmak. Group actions and a conjecture of Quillen. Lecture Notes in Mathematics, 2032:105–120, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22003-6_
8/.

Barmak:2011:RCQ

[Bar11h] Jonathan A. Barmak. h-regular complexes and quotients. Lecture Notes in Mathematics, 2032:93–104, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-22003-6_7/.

Barmak:2011:MR

[Bar11i] Jonathan A. Barmak. Methods of reduction. Lecture Notes in Mathematics, 2032:85–91, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-22003-6_6/.

Barmak:2011:MFM

[Bar11j] Jonathan A. Barmak. Minimal finite models. Lecture Notes in Mathematics, 2032:37–47, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-22003-6_3/.

Barmak:2011:P

[Bar11k] Jonathan A. Barmak. Preliminaries. Lecture Notes in Mathematics, 2032:1–18, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-22003-6_1/.

Barmak:2011:RL

[Bar11l] Jonathan A. Barmak. Reduced lattices. Lecture Notes in Mathematics, 2032:121–127, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-22003-6_9/.

Barmak:2011:SHTa

[Bar11m] Jonathan A. Barmak. Simple homotopy types and finite spaces. Lecture Notes in Mathematics, 2032:49–72, 2011. CODEN LN-MAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22003-6_
4/.

Barmak:2011:SHTb

[Bar11n] Jonathan A. Barmak. Strong homotopy types. Lecture Notes in Mathematics, 2032:73–84, 2011. CODEN LNMAA2. ISBN 3-642-22002-9 (print), 3-642-22003-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-22003-6_5/.

Barles:2013:ITV

[Bar13] Guy Barles. An introduction to the theory of viscosity solutions for first-order Hamilton–Jacobi equations and applications. Lecture Notes in Mathematics, 2074:49–109, 2013. CODEN LN-MAA2. ISBN 3-642-36432-2 (print), 3-642-36433-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-36433-4_
2/.

Barthe:2013:COT

[BB13] Franck Barthe and Charles Bordenave. Combinatorial optimization over two random point sets. Lecture Notes in Mathematics, 2078:483–535, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00321-4_19/.

Bar:2014:DCG

[BB14a] Christian Bär and Christian Becker. Differential characters and geometric chains. Lecture Notes in Mathematics, 2112:1–90, 2014. CODEN LNMAA2. ISBN 3-319-07033-9 (print), 3-319-07034-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-07034-6_1.pdf.

Bar:2014:FM

[BB14b] Christian Bär and Christian Becker. Front matter. Lecture Notes in Mathematics, 2112:i–viii, 2014. CODEN LNMAA2. ISBN 3-319-07033-9 (print), 3-319-07034-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-07034-6/1.pdf.

Bercu:2012:LDG

[BBB12] Bernard Bercu, Jean-François Bony, and Vincent Bruneau. Large deviations for Gaussian stationary processes and semiclassical analysis. Lecture Notes in Mathematics, 2046:409– 428, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27461-9_19/.

Bachar:2013:FMa

[BBD13a] Mostafa Bachar, Jerry Batzel, and Susanne Ditlevsen. Front matter. Lecture Notes in Mathematics, 2058:i–xvi, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-32157-3/
1.pdf.

Bachar:2013:FMb

[BBD13b] Mostafa Bachar, Jerry Batzel, and Susanne Ditlevsen. Front matter. Lecture Notes in Mathematics, 2058:1, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-32157-3/
1/1.pdf.

Bachar:2013:SBM

[BBD13c] Mostafa Bachar, Jerry Batzel, and Susanne Ditlevsen. Stochastic Biomathematical Models: with Applications to Neuronal Modeling, volume 2058 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 206 pp. LCCN QA3 .L28 no. 2058; QP363 .S785

2013. URL http://link.springer.com/book/10.1007/978-
3-642-32157-3;   http://www.springerlink.com/content/
978-3-642-32157-3.

Batzel:2013:FMa

[BBK13a] Jerry J. Batzel, Mostafa Bachar, and Franz Kappel. Front matter. Lecture Notes in Mathematics, 2064:i–xx, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-32882-4/
1.pdf.

Batzel:2013:FMb

[BBK13b] Jerry J. Batzel, Mostafa Bachar, and Franz Kappel. Front matter. Lecture Notes in Mathematics, 2064:1, 2013. CODEN LN-MAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-32882-4/
1/1.pdf.

Batzel:2013:MMV

[BBK13c] Jerry J. Batzel, Mostafa Bachar, and Franz Kappel. Mathematical Modeling and Validation in Physiology: Applications to the Cardiovascular and Respiratory Systems, volume 2064 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 254 pp. LCCN QA3 .L28 no. 2064; QA3 .L28 no. 2064; QP33.6.M36 M374 2013.

URL http://link.springer.com/book/10.1007/978-3-642-
32882-4;   http://www.springerlink.com/content/978-3-
642-32882-4.

Batzel:2013:MMP

[BBKK13] Jerry J. Batzel, Mostafa Bachar, John M. Karemaker, and Franz Kappel. Merging mathematical and physiological knowledge: Dimensions and challenges. Lecture Notes in Mathematics, 2064:3–19, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-32882-4_1/;   http://link.springer.
com/content/pdf/bfm:978-3-642-32882-4/1.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-642-32882-4/
1/1.pdf.

Baldassi:2015:SPN

[BBRZ15] Carlo Baldassi, Alfredo Braunstein, Abolfazl Ramezanpour, and Riccardo Zecchina. Statistical physics and network optimization problems. Lecture Notes in Mathematics, 2141:27–49, 2015. CODEN LNMAA2. ISBN 3-319-16966-1 (print), 3-319-16967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-16967-5_2.pdf.

Boileau:2016:FM

[BBST16] Michel Boileau, Gerard Besson, Carlo Sinestrari, and Gang Tian. Front matter. Lecture Notes in Mathematics, 2166:i– xi, 2016. CODEN LNMAA2. ISBN 3-319-42350-9 (print), 3-319-42351-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-42351-7/1.pdf.

Beznea:2014:BKT

[BC14] Lucian Beznea and Iulian Cîmpean. On Bochner–Kolmogorov theorem. Lecture Notes in Mathematics, 2123:61–70, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_3.pdf.

Banks:2013:PSM

[BCAK13] H. T. Banks, Ariel Cintrón-Arias, and Franz Kappel. Parameter selection methods in inverse problem formulation. Lecture Notes in Mathematics, 2064:43–73, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32882-4_3/.

Boscaggin:2013:MIG

[BCD13] Alberto Boscaggin, Anna Capietto, and Walter Dambrosio. The Maslov index and global bifurcation for nonlinear boundary value problems. Lecture Notes in Mathematics, 2065:1– 34, 2013. CODEN LNMAA2. ISBN 3-642-32905-5 (print), 3-642-32906-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32906-7_1/.

Benth:2013:FM

[BCG^{+}13a] Fred Espen Benth, Dan Crisan, Paolo Guasoni, Konstantinos Manolarakis, Johannes Muhle-Karbe, Colm Nee, and Philip Protter. Front matter. Lecture Notes in Mathematics, 2081:i– ix, 2013. CODEN LNMAA2. ISBN 3-319-00412-3 (print), 3-319-00413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-00413-6/1.pdf.

Benth:2013:PPL

[BCG^{+}13b] Fred Espen Benth, Dan Crisan, Paolo Guasoni, Konstantinos Manolarakis, Johannes Muhle-Karbe, Colm Nee, and Philip Protter, editors. Paris–Princeton Lectures on Mathematical Finance 2013, volume 2081 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-00412-3 (paperback), 3-319-00413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2081; QA3 L28 no. 2081;

HF5691 .P37 2013. URL http://link.springer.com/book/
10.1007/978-3-319-00413-6;   http://www.springerlink.
com/content/978-3-319-00413-6.

Belomestny:2015:FM

[BCGC^{+}15] Denis Belomestny, Fabienne Comte, Valentine Genon-Catalot, Hiroki Masuda, and Markus Reiß. Front matter. Lecture Notes in Mathematics, 2128:i–xv, 2015. CODEN LNMAA2. ISBN 3-319-12372-6 (print), 3-319-12373-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-12373-8/1.pdf.

Bujalance:2010:A

[BCGG10a] Emilio Bujalance, Francisco Javier Cirre, José Manuel Gamboa, and Grzegorz Gromadzki. Appendix. Lecture Notes in Mathematics, 2007:145–149, 2010. CODEN LNMAA2. ISBN 3-642-14827-1 (print), 3-642-14828-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-14828-6_6.pdf.

Bujalance:2010:BM

[BCGG10b] Emilio Bujalance, Francisco Javier Cirre, José Manuel Gamboa, and Grzegorz Gromadzki. Back matter. Lecture Notes in Mathematics, 2007:151–158, 2010. CODEN LNMAA2. ISBN 3-642-14827-1 (print), 3-642-14828-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-14828-6/1.pdf.

Bujalance:2010:COS

[BCGG10c] Emilio Bujalance, Francisco Javier Cirre, José Manuel Gamboa, and Grzegorz Gromadzki. Counting ovals of symmetries of Riemann surfaces. Lecture Notes in Mathematics, 2007:33–63, 2010. CODEN LNMAA2. ISBN 3-642-14827-1 (print), 3-642-14828-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14828-6_3.pdf.

Bujalance:2010:FM

[BCGG10d] Emilio Bujalance, Francisco Javier Cirre, José Manuel Gamboa, and Grzegorz Gromadzki. Front matter. Lecture Notes in Mathematics, 2007:i–xx, 2010. CODEN LNMAA2. ISBN 3-642-14827-1 (print), 3-642-14828-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-14828-6/1.pdf.

Bujalance:2010:NCC

[BCGG10e] Emilio Bujalance, Francisco Javier Cirre, José Manuel Gamboa, and Grzegorz Gromadzki. On the number of conjugacy classes of symmetries of Riemann surfaces. Lecture Notes in Mathematics, 2007:21–32, 2010. CODEN LNMAA2. ISBN 3-642-14827-1 (print), 3-642-14828-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-14828-6_2.pdf.

Bujalance:2010:P

[BCGG10f] Emilio Bujalance, Francisco Javier Cirre, José Manuel Gamboa, and Grzegorz Gromadzki. Preliminaries. Lecture Notes in Mathematics, 2007:1–20, 2010. CODEN LNMAA2. ISBN 3-642-14827-1 (print), 3-642-14828-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-14828-6_1.pdf.

Bujalance:2010:SCR

[BCGG10g] Emilio Bujalance, Francisco Javier Cirre, José Manuel Gamboa, and Grzegorz Gromadzki. Symmetries of Compact Riemann Surfaces, volume 2007 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-14827-1 (print), 3-642-14828-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 158 pp. LCCN QA3 .L28 no. 2007. URL http:

//link.springer.com/book/10.1007/978-3-642-14828-6;
http://www.springerlink.com/content/978-3-642-14828-6.

Bujalance:2010:STR

[BCGG10h] Emilio Bujalance, Francisco Javier Cirre, José Manuel Gamboa, and Grzegorz Gromadzki. Symmetry types of Riemann surfaces with a large group of automorphisms. Lecture Notes in Mathematics, 2007:91–143, 2010. CODEN LNMAA2. ISBN 3-642-14827-1 (print), 3-642-14828-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-14828-6_5.pdf.

Bujalance:2010:STS

[BCGG10i] Emilio Bujalance, Francisco Javier Cirre, José Manuel Gamboa, and Grzegorz Gromadzki. Symmetry types of some families of Riemann surfaces. Lecture Notes in Mathematics, 2007: 65–90, 2010. CODEN LNMAA2. ISBN 3-642-14827-1 (print), 3-642-14828-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-14828-6_4.pdf.

Baum:2011:FM

[BCM^{+}11a] Paul Frank Baum, Guillermo Cortiñas, Ralf Meyer, Rubén Sánchez-García, Marco Schlichting, and Bertrand Toën. Front matter. Lecture Notes in Mathematics, 2008:i–xvi, 2011. CODEN LNMAA2. ISBN 3-642-15707-6 (print), 3-642-15708-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-15708-0/
1.pdf.

Baum:2011:TAT

[BCM^{+}11b] Paul Frank Baum, Guillermo Cortiñas, Ralf Meyer, Rubén Sánchez-García, Marco Schlichting, and Bertrand Toën. Topics in Algebraic and Topological K-Theory, volume 2008 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-15707-6 (print), 3-642-15708-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 302 pp. LCCN

QA3 .L28 no. 2008. URL http://link.springer.com/book/
10.1007/978-3-642-15708-0;   http://www.springerlink.
com/content/978-3-642-15708-0.

Butta:2015:BM

[BCM15a] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Back matter. Lecture Notes in Mathematics, 2135:117–136, 2015. CODEN LNMAA2. ISBN 3-319-14758-7 (print), 3-319-14759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-319-14759-8/
1.pdf.

Butta:2015:FM

[BCM15b] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Front matter. Lecture Notes in Mathematics, 2135:i–xiv, 2015. CODEN LNMAA2. ISBN 3-319-14758-7 (print), 3-319-14759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-14759-8/
1.pdf.

Butta:2015:GPP

[BCM15c] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Gas of point particles. Lecture Notes in Mathematics, 2135:1–41, 2015. CODEN LNMAA2. ISBN 3-319-14758-7 (print), 3-319-14759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-14759-8_1.pdf.

Butta:2015:MMV

[BCM15d] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Mathematical Models of Viscous Friction, volume 2135 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LNMAA2. ISBN 3-319-14758-7 (print), 3-319-14759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 136 (est.) pp. LCCN ???? URL

http://www.springerlink.com/content/978-3-319-14759-8.

Butta:2015:MBIb

[BCM15e] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Motion of a body immersed in a Stokes fluid. Lecture Notes in Mathematics, 2135:101–116, 2015. CODEN LNMAA2. ISBN 3-319-14758-7 (print), 3-319-14759-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-14759-8_4.pdf.

Butta:2015:MBIa

[BCM15f] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Motion of a body immersed in a Vlasov system. Lecture Notes in Mathematics, 2135:63–100, 2015. CODEN LNMAA2. ISBN 3-319-14758-7 (print), 3-319-14759-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-14759-8_3.pdf.

Butta:2015:VA

[BCM15g] Paolo Buttà, Guido Cavallaro, and Carlo Marchioro. Vlasov approximation. Lecture Notes in Mathematics, 2135:43–61, 2015. CODEN LNMAA2. ISBN 3-319-14758-7 (print), 3-319-14759-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-14759-8_2.pdf.

Bianchini:2011:FM

[BCMV11a] Stefano Bianchini, Eric A. Carlen, Alexander Mielke, and Cédric Villani. Front matter. Lecture Notes in Mathematics, 2028:i– xiii, 2011. CODEN LNMAA2. ISBN 3-642-21718-4 (print), 3-642-21861-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-21861-3/1.pdf.

Bianchini:2011:NPA

[BCMV11b] Stefano Bianchini, Eric A. Carlen, Alexander Mielke, and Cédric Villani, editors. Nonlinear PDE’s and Applications: C.I.M.E. Summer School, Cetraro, Italy 2008, volume 2028 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-21718-4 (print), 3-642-21861-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2028;

QA374 .C53 2011. URL http://link.springer.com/book/
10.1007/978-3-642-21861-3;   http://www.springerlink.
com/content/978-3-642-21861-3.

Bigatti:2013:FM

[BD13a] Anna Maria Bigatti and Emanuela De Negri. Front matter. Lecture Notes in Mathematics, 2083:61, 2013. CODEN LN-MAA2. ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-38742-5/
2/1.pdf.

Bigatti:2013:SDU

[BD13b] Anna Maria Bigatti and Emanuela De Negri. Stanley decompositions using CoCoA. Lecture Notes in Mathematics, 2083: 47–59, 2013. CODEN LNMAA2. ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-38742-5_2/.

Bomze:2010:FM

[BDFT10] Immanuel M. Bomze, Vladimir F. Demyanov, Roger Fletcher, and Tamás Terlaky. Front matter. Lecture Notes in Mathematics, 1989:i–xiii, 2010. CODEN LNMAA2. ISBN 3-642-11338-9 (print), 3-642-11339-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-11339-0/1.pdf.

Beyn:2014:CCS

[BDG^{+}14a] Wolf-Jürgen Beyn, Luca Dieci, Nicola Guglielmi, Ernst Hairer, Jesús María Sanz-Serna, and Marino Zennaro, editors. Current Challenges in Stability Issues for Numerical Differential Equations: Cetraro, Italy 2011, volume 2082 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-01299-1 (print), 3-319-01300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2082; QA371 .C87 2014; QA3 .L28 no. 2082. URL http://link.

springer.com/book/10.1007/978-3-319-01300-8;   http:/
/www.springerlink.com/content/978-3-319-01300-8.

Beyn:2014:FM

[BDG^{+}14b] Wolf-Jürgen Beyn, Luca Dieci, Nicola Guglielmi, Ernst Hairer, Jesús María Sanz-Serna, and Marino Zennaro. Front matter. Lecture Notes in Mathematics, 2082:i–ix, 2014. CODEN LNMAA2. ISBN 3-319-01299-1 (print), 3-319-01300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-319-01300-8/1.pdf.

Baker:2011:BM

[BDMY11a] David Baker, Catherine Donati-Martin, and Marc Yor. Back matter. Lecture Notes in Mathematics, 2006:505–510, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-642-15217-7/
1.pdf.

Baker:2011:SAT

[BDMY11b] David Baker, Catherine Donati-Martin, and Marc Yor. A sequence of Albin type continuous martingales with Brownian marginals and scaling. Lecture Notes in Mathematics, 2006:441– 449, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_20/.

Barbu:2016:BAE

[BDR16a] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. L^{1}- based approach to existence theory for stochastic porous media equations. Lecture Notes in Mathematics, 2163:107–131, 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-41069-2_5.pdf.

Barbu:2016:BM

[BDR16b] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. Back matter. Lecture Notes in Mathematics, 2163:197–204, 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-41069-2/1.pdf.

Barbu:2016:ELN

[BDR16c] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. Equations with Lipschitz nonlinearities. Lecture Notes in Mathematics, 2163:19–47, 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-41069-2_2.pdf.

Barbu:2016:EMM

[BDR16d] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. Equations with maximal monotone nonlinearities. Lecture Notes in Mathematics, 2163:49–93, 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-41069-2_3.pdf.

Barbu:2016:FM

[BDR16e] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. Front matter. Lecture Notes in Mathematics, 2163:i–ix, 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-41069-2/
1.pdf.

Barbu:2016:I

[BDR16f] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. Introduction. Lecture Notes in Mathematics, 2163:1–18, 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-41069-2_1.pdf.

Barbu:2016:SPMa

[BDR16g] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. The stochastic porous media equations in R^{d}. Lecture Notes in Mathematics, 2163:133–165, 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-41069-2_6.pdf.

Barbu:2016:TS

[BDR16h] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. Transition semigroup. Lecture Notes in Mathematics, 2163:167–195, 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-41069-2_7.pdf.

Barbu:2016:VAS

[BDR16i] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner. Variational approach to stochastic porous media equations. Lecture Notes in Mathematics, 2163:95–106, 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-41069-2_4.
pdf.

Beck:2014:FM

[Bec14a] József Beck. Front matter. Lecture Notes in Mathematics, 2107: 317, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-04696-9/2/1.pdf.

Beck:2014:S

[Bec14b] József Beck. Superirregularity. Lecture Notes in Mathematics, 2107:221–316, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-04696-9_4.pdf.

Becker:2014:BM

[Bec14c] Christian Becker. Back matter. Lecture Notes in Mathematics, 2112:181–189, 2014. CODEN LNMAA2. ISBN 3-319-07033-9 (print), 3-319-07034-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-07034-6/1.pdf.

Becker:2014:DC

[Bec14d] Christian Becker. Differential Characters, volume 2112 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LN-MAA2. ISBN 3-319-07033-9 (print), 3-319-07034-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 189 (est.) pp.

LCCN ???? URL http://www.springerlink.com/content/
978-3-319-07034-6.

Becker:2014:RDC

[Bec14e] Christian Becker. Relative differential cohomology. Lecture Notes in Mathematics, 2112:91–180, 2014. CODEN LNMAA2. ISBN 3-319-07033-9 (print), 3-319-07034-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-07034-6_2.pdf.

Bedford:2010:DRS

[Bed10] Eric Bedford. Dynamics of rational surface automorphisms. Lecture Notes in Mathematics, 1998:57–104, 2010. CO-DEN LNMAA2. ISBN 3-642-13170-0 (print), 3-642-13171-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-13171-4_2.pdf.

Boucksom:2013:I

[BEG13a] Sébastien Boucksom, Philippe Eyssidieux, and Vincent Guedj. Introduction. Lecture Notes in Mathematics, 2086:1–6, 2013. CODEN LNMAA2. ISBN 3-319-00818-8 (print), 3-319-00819-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
319-00819-6_1/;   http://link.springer.com/content/pdf/
bfm:978-3-319-00819-6/1.pdf.

Boucksom:2013:IKR

[BEG13b] Sebastien Boucksom, Philippe Eyssidieux, and Vincent Guedj. An Introduction to the Kähler–Ricci Flow, volume 2086 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-00818-8 (print), 3-319-00819-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 333 pp. LCCN QA3 .L28 no. 2086; QA3 .L28 no. 2086; QA649 .I585 2013; QA670 .I587

2013. URL http://link.springer.com/book/10.1007/978-
3-319-00819-6;   http://www.springerlink.com/content/
978-3-319-00819-6.

Brossard:2014:SPD

[BÉL14] Jean Brossard, Michel Émery, and Christophe Leuridan. Skewproduct decomposition of planar Brownian motion and complementability. Lecture Notes in Mathematics, 2123:377–394, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_15.pdf.

Benjamini:2013:BM

[Ben13a] Itai Benjamini. Back matter. Lecture Notes in Mathematics, 2100:125–132, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-02576-6/1.pdf.

Benjamini:2013:CGR

[Ben13b] Itai Benjamini, editor. Coarse Geometry and Randomness: École d’Été de Probabilités de Saint-Flour XLI — 2011, volume 2100 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28

no. 2100; QA440-699. URL http://link.springer.com/book/
10.1007/978-3-319-02576-6;   http://www.springerlink.
com/content/978-3-319-02576-6.

Benjamini:2013:CPN

[Ben13c] Itai Benjamini. Critical percolation on non-amenable groups. Lecture Notes in Mathematics, 2100:63–68, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02576-6_8/.

Benjamini:2013:FM

[Ben13d] Itai Benjamini. Front matter. Lecture Notes in Mathematics, 2100:i–vii, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-02576-6/1.pdf.

Benjamini:2013:GIP

[Ben13e] Itai Benjamini. Growth and isoperimetric profile of planar graphs. Lecture Notes in Mathematics, 2100:59–61, 2013. CODEN LN-MAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02576-6_
7/.

Benjamini:2013:HFG

[Ben13f] Itai Benjamini. Harmonic functions on graphs. Lecture Notes in Mathematics, 2100:107–120, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02576-6_12/.

Benjamini:2013:HPH

[Ben13g] Itai Benjamini. The hyperbolic plane and hyperbolic graphs. Lecture Notes in Mathematics, 2100:23–31, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02576-6_3/.

Benjamini:2013:IGM

[Ben13h] Itai Benjamini. Introductory graph and metric notions. Lecture Notes in Mathematics, 2100:1–18, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02576-6_1/.

Benjamini:2013:LLG

[Ben13i] Itai Benjamini. Local limits of graphs. Lecture Notes in Mathematics, 2100:41–51, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02576-6_5/.

Benjamini:2013:NLG

[Ben13j] Itai Benjamini. Nonamenable Liouville graphs. Lecture Notes in Mathematics, 2100:121–124, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02576-6_13/.

Benjamini:2013:SVT

[Ben13k] Itai Benjamini. On the structure of vertex transitive graphs. Lecture Notes in Mathematics, 2100:19–21, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02576-6_2/.

Benjamini:2013:PE

[Ben13l] Itai Benjamini. Percolation on expanders. Lecture Notes in Mathematics, 2100:97–105, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02576-6_11/.

Benjamini:2013:PG

[Ben13m] Itai Benjamini. Percolation on graphs. Lecture Notes in Mathematics, 2100:33–40, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02576-6_4/.

Benjamini:2013:PP

[Ben13n] Itai Benjamini. Percolation perturbations. Lecture Notes in Mathematics, 2100:85–95, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02576-6_10/.

Benjamini:2013:RPG

[Ben13o] Itai Benjamini. Random planar geometry. Lecture Notes in Mathematics, 2100:53–58, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02576-6_6/.

Benjamini:2013:UIP

[Ben13p] Itai Benjamini. Uniqueness of the infinite percolation cluster. Lecture Notes in Mathematics, 2100:69–84, 2013. CODEN LNMAA2. ISBN 3-319-02575-9 (print), 3-319-02576-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02576-6_9/.

Benth:2013:SVD

[Ben13q] Fred Espen Benth. Stochastic volatility and dependency in energy markets: Multi-factor modelling. Lecture Notes in Mathematics, 2081:109–167, 2013. CODEN LNMAA2. ISBN 3-319-00412-3 (print), 3-319-00413-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00413-6_2/.

Benjamini:2014:GFF

[Ben14] Itai Benjamini. Gaussian free field on hyperbolic lattices. Lecture Notes in Mathematics, 2116:39–45, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-09477-9_2.
pdf.

Bertoluzza:2012:AWM

[Ber12] Silvia Bertoluzza. Adaptive wavelet methods. Lecture Notes in Mathematics, 2040:1–56, 2012. CODEN LNMAA2. ISBN 3-642-24078-X (print), 3-642-24079-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24079-9_1/.

Berteloot:2013:BCH

[Ber13] François Berteloot. Bifurcation currents in holomorphic families of rational maps. Lecture Notes in Mathematics, 2075:1– 93, 2013. CODEN LNMAA2. ISBN 3-642-36420-9 (print), 3-642-36421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-36421-1_1/.

Besson:2016:DST

[Bes16] Gérard Besson. The differentiable sphere theorem (after S. Brendle and R. Schoen). Lecture Notes in Mathematics, 2166:1–19, 2016. CODEN LNMAA2. ISBN 3-319-42350-9 (print), 3-319-42351-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42351-7_1.pdf.

Bonfiglioli:2012:BM

[BF12a] Andrea Bonfiglioli and Roberta Fulci. Back matter. Lecture Notes in Mathematics, 2034:523–539, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-22597-0/1.pdf.

Bonfiglioli:2012:BA

[BF12b] Andrea Bonfiglioli and Roberta Fulci. Background algebra. Lecture Notes in Mathematics, 2034:49–114, 2012. CODEN LN-MAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22597-0_
2/.

Bonfiglioli:2012:CFL

[BF12c] Andrea Bonfiglioli and Roberta Fulci. Construction of free Lie algebras. Lecture Notes in Mathematics, 2034:459–477, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22597-0_8/.

Bonfiglioli:2012:CCS

[BF12d] Andrea Bonfiglioli and Roberta Fulci. Convergence of the CBHD series and associativity of the CBHD operation. Lecture Notes in Mathematics, 2034:265–369, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-22597-0_5/.

Bonfiglioli:2012:FPS

[BF12e] Andrea Bonfiglioli and Roberta Fulci. Formal power series in one indeterminate. Lecture Notes in Mathematics, 2034:479– 499, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22597-0_9/.

Bonfiglioli:2012:FMa

[BF12f] Andrea Bonfiglioli and Roberta Fulci. Front matter. Lecture Notes in Mathematics, 2034:i–xxii, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-22597-0/1.pdf.

Bonfiglioli:2012:FMb

[BF12g] Andrea Bonfiglioli and Roberta Fulci. Front matter. Lecture Notes in Mathematics, 2034:47, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-22597-0/1/1.pdf.

Bonfiglioli:2012:FMc

[BF12h] Andrea Bonfiglioli and Roberta Fulci. Front matter. Lecture Notes in Mathematics, 2034:391, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-22597-0/2/1.pdf.

Bonfiglioli:2012:HO

[BF12i] Andrea Bonfiglioli and Roberta Fulci. Historical overview. Lecture Notes in Mathematics, 2034:1–45, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-22597-0_1/.

Bonfiglioli:2012:MPC

[BF12j] Andrea Bonfiglioli and Roberta Fulci. The main proof of the CBHD theorem. Lecture Notes in Mathematics, 2034:115–172, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22597-0_3/.

Bonfiglioli:2012:PAP

[BF12k] Andrea Bonfiglioli and Roberta Fulci. Proofs of the algebraic prerequisites. Lecture Notes in Mathematics, 2034:393– 457, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22597-0_7/.

Bonfiglioli:2012:RBC

[BF12l] Andrea Bonfiglioli and Roberta Fulci. Relationship between the CBHD Theorem, the PBW Theorem and the free Lie algebras. Lecture Notes in Mathematics, 2034:371–389, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22597-0_
6/.

Bonfiglioli:2012:SSP

[BF12m] Andrea Bonfiglioli and Roberta Fulci. Some “short” proofs of the CBHD theorem. Lecture Notes in Mathematics, 2034:173– 264, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22597-0_4/.

Bonfiglioli:2012:SA

[BF12n] Andrea Bonfiglioli and Roberta Fulci. Symmetric algebra. Lecture Notes in Mathematics, 2034:501–521, 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-22597-0_10/.

Bonfiglioli:2012:TNA

[BF12o] Andrea Bonfiglioli and Roberta Fulci. Topics in Noncommutative Algebra: The Theorem of Campbell, Baker, Hausdorff and Dynkin, volume 2034 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-22596-9 (print), 3-642-22597-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxii + 539 pp. LCCN QA3 .L28 no. 2034. URL http:

//link.springer.com/book/10.1007/978-3-642-22597-0;
http://www.springerlink.com/content/978-3-642-22597-0.

Bini:2014:APP

[BFMV14a] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Appendix: Positivity properties of balanced line bundles. Lecture Notes in Mathematics, 2122:197–203, 2014. CO-DEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11337-1_17.pdf.

Bini:2014:BM

[BFMV14b] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Back matter. Lecture Notes in Mathematics, 2122:205– 214, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-11337-1/1.pdf.

Bini:2014:BEB

[BFMV14c] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Behavior at the extremes of the basic inequality. Lecture Notes in Mathematics, 2122:81–90, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11337-1_7.pdf.

Bini:2014:CR

[BFMV14d] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Combinatorial results. Lecture Notes in Mathematics, 2122:27– 44, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11337-1_3.pdf.

Bini:2014:CUJ

[BFMV14e] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Compactifications of the universal Jacobian. Lecture Notes in Mathematics, 2122:171–195, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11337-1_16.pdf.

Bini:2014:CST

[BFMV14f] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. A criterion of stability for tails. Lecture Notes in Mathematics, 2122:91–105, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-11337-1_8.pdf.

Bini:2014:ETT

[BFMV14g] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Elliptic tails and tacnodes with a line. Lecture Notes in Mathematics, 2122:107–116, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11337-1_9.pdf.

Bini:2014:ECG

[BFMV14h] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Extra components of the GIT quotient. Lecture Notes in Mathematics, 2122:167–170, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11337-1_15.pdf.

Bini:2014:FM

[BFMV14i] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Front matter. Lecture Notes in Mathematics, 2122:i– x, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-11337-1/1.pdf.

Bini:2014:GIT

[BFMV14j] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Geometric Invariant Theory for Polarized Curves, volume 2122 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 214 (est.) pp.

LCCN ???? URL http://www.springerlink.com/content/
978-3-319-11337-1.

Bini:2014:GPG

[BFMV14k] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Geometric properties of the GIT quotient. Lecture Notes in Mathematics, 2122:155–165, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11337-1_14.pdf.

Bini:2014:I

[BFMV14l] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Introduction. Lecture Notes in Mathematics, 2122:1–16, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11337-1_1.pdf.

Bini:2014:PPS

[BFMV14m] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Potential pseudo-stability theorem. Lecture Notes in Mathematics, 2122:61–72, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-11337-1_5.pdf.

Bini:2014:PG

[BFMV14n] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Preliminaries on GIT. Lecture Notes in Mathematics, 2122:45– 59, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11337-1_4.pdf.

Bini:2014:SPSa

[BFMV14o] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Semistable, polystable and stable points (Part I). Lecture Notes in Mathematics, 2122:131–139, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11337-1_11.pdf.

Bini:2014:SPSb

[BFMV14p] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Semistable, polystable and stable points (Part II). Lecture Notes in Mathematics, 2122:149–154, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11337-1_13.pdf.

Bini:2014:SC

[BFMV14q] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Singular curves. Lecture Notes in Mathematics, 2122:17– 26, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11337-1_2.pdf.

Bini:2014:SET

[BFMV14r] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Stability of elliptic tails. Lecture Notes in Mathematics, 2122:141– 147, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11337-1_12.pdf.

Bini:2014:SS

[BFMV14s] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. Stabilizer subgroups. Lecture Notes in Mathematics, 2122:73–80, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11337-1_6.pdf.

Bini:2014:SSL

[BFMV14t] Gilberto Bini, Fabio Felici, Margarida Melo, and Filippo Viviani. A stratification of the semistable locus. Lecture Notes in Mathematics, 2122:117–130, 2014. CODEN LNMAA2. ISBN 3-319-11336-4 (print), 3-319-11337-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11337-1_10.pdf.

Bianchini:2011:TRA

[BG11] Stefano Bianchini and Matteo Gloyer. Transport rays and applications to Hamilton–Jacobi equations. Lecture Notes in Mathematics, 2028:1–15, 2011. CODEN LNMAA2. ISBN 3-642-21718-4 (print), 3-642-21861-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21861-3_1/.

Boucksom:2013:RPK

[BG13] Sébastien Boucksom and Vincent Guedj. Regularizing properties of the Kähler–Ricci flow. Lecture Notes in Mathematics, 2086: 189–237, 2013. CODEN LNMAA2. ISBN 3-319-00818-8 (print), 3-319-00819-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00819-6_4/.

Bigatti:2013:MIC

[BGdC13] Anna M. Bigatti, Philippe Gimenez, and Eduardo Sáenz de Cabezón. Monomial Ideals, Computations and Applications, volume 2083 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 194 pp. LCCN QA3 .L28 no. 2083. URL http:

//link.springer.com/book/10.1007/978-3-642-38742-5;
http://www.springerlink.com/content/978-3-642-38742-5.

Brandolini:2014:IDA

[BGT14] Luca Brandolini, Giacomo Gigante, and Giancarlo Travaglini. Irregularities of distribution and average decay of Fourier transforms. Lecture Notes in Mathematics, 2107:159–220, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-04696-9_3.pdf.

Bocharov:2014:BRW

[BH14] Sergey Bocharov and Simon C. Harris. Branching random walk in an inhomogeneous breeding potential. Lecture Notes in Mathematics, 2123:1–32, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-11970-0_1.pdf.

Badrikian:2012:PBS

[BHJKX12] A. Badrikian, Jørgen Hoffmann-Jørgensen, Jim Kuelbs, and Fernique Xavier. Probability in Banach spaces at Saint-Flour, volume ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-25276-1 (paperback). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 614 pp. LCCN QA3 .L28 no. ??; QA273 .B347 2012.

Bilyk:2014:ROF

[Bil14] Dmitriy Bilyk. Roth’s orthogonal function method in discrepancy theory and some new connections. Lecture Notes in Mathematics, 2107:71–158, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-04696-9_2.pdf.

Bobenko:2011:CAR

[BK11] Alexander I. Bobenko and Christian Klein. Computational Approach to Riemann Surfaces, volume 2013 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 257 pp. LCCN QA3 .L28 no.

2013. URL http://link.springer.com/book/10.1007/978-
3-642-17413-1;   http://www.springerlink.com/content/
978-3-642-17413-1.

Berman:2012:BM

[BK12a] Robert Berman and Julien Keller. Back matter. Lecture Notes in Mathematics, 2038:303–310, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-23669-3/1.pdf.

Berman:2012:BG

[BK12b] Robert Berman and Julien Keller. Bergman geodesics. Lecture Notes in Mathematics, 2038:283–302, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-23669-3_8/.

Brossard:2013:COL

[BL13] Jean Brossard and Christophe Leuridan. Characterising Ocone local martingales with reflections. Lecture Notes in Mathematics, 2078:167–180, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00321-4_6/.

Bodine:2015:ACSb

[BL15a] Sigrun Bodine and Donald A. Lutz. Applications to classes of scalar linear difference equations. Lecture Notes in Mathematics, 2129:295–368, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-18248-3_9.pdf.

Bodine:2015:ACSa

[BL15b] Sigrun Bodine and Donald A. Lutz. Applications to classes of scalar linear differential equations. Lecture Notes in Mathematics, 2129:237–294, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-18248-3_8.pdf.

Bodine:2015:AIDa

[BL15c] Sigrun Bodine and Donald A. Lutz. Asymptotic integration for differential systems. Lecture Notes in Mathematics, 2129:11– 67, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-319-18248-3_2.pdf.

Bodine:2015:AIDb

[BL15d] Sigrun Bodine and Donald A. Lutz. Asymptotic Integration of Differential and Difference Equations, volume 2129 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LN-MAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 404 (est.) pp.

LCCN ???? URL http://www.springerlink.com/content/
978-3-319-18248-3.

Bodine:2015:ARS

[BL15e] Sigrun Bodine and Donald A. Lutz. Asymptotic representation for solutions of difference systems. Lecture Notes in Mathematics, 2129:69–117, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-18248-3_3.pdf.

Bodine:2015:ADE

[BL15f] Sigrun Bodine and Donald A. Lutz. Asymptotics for dynamic equations on time scales. Lecture Notes in Mathematics, 2129: 369–391, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-319-18248-3_10.pdf.

Bodine:2015:BM

[BL15g] Sigrun Bodine and Donald A. Lutz. Back matter. Lecture Notes in Mathematics, 2129:393–404, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-319-18248-3/1.pdf.

Bodine:2015:CTDb

[BL15h] Sigrun Bodine and Donald A. Lutz. Conditioning transformations for difference systems. Lecture Notes in Mathematics, 2129:179–208, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-18248-3_5.pdf.

Bodine:2015:CTDa

[BL15i] Sigrun Bodine and Donald A. Lutz. Conditioning transformations for differential systems. Lecture Notes in Mathematics, 2129:119–177, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-18248-3_4.pdf.

Bodine:2015:FM

[BL15j] Sigrun Bodine and Donald A. Lutz. Front matter. Lecture Notes in Mathematics, 2129:i–xi, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-18248-3/1.pdf.

Bodine:2015:INB

[BL15k] Sigrun Bodine and Donald A. Lutz. Introduction, notation, and background. Lecture Notes in Mathematics, 2129:1–10, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-18248-3_1.pdf.

Bodine:2015:PJDb

[BL15l] Sigrun Bodine and Donald A. Lutz. Perturbations of Jordan difference systems. Lecture Notes in Mathematics, 2129:233– 236, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-319-18248-3_7.pdf.

Bodine:2015:PJDa

[BL15m] Sigrun Bodine and Donald A. Lutz. Perturbations of Jordan differential systems. Lecture Notes in Mathematics, 2129:209– 232, 2015. CODEN LNMAA2. ISBN 3-319-18247-1 (print), 3-319-18248-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-319-18248-3_6.pdf.

Blocki:2012:CYT

[Bło12a] Zbigniew Błocki. The Calabi–Yau theorem. Lecture Notes in Mathematics, 2038:201–227, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-23669-3_5/.

Blocki:2012:FM

[Bło12b] Zbigniew Błocki. Front matter. Lecture Notes in Mathematics, 2038:229, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-23669-3/4/1.pdf.

Blocki:2013:CMA

[Bło13] Zbigniew Błocki. The complex Monge–Ampère equation in Kähler geometry. Lecture Notes in Mathematics, 2075:95–141, 2013. CODEN LNMAA2. ISBN 3-642-36420-9 (print), 3-642-36421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-36421-1_2/.

Blocki:2014:LBB

[Bło14] Zbigniew Błocki. A lower bound for the Bergman kernel and the Bourgain–Milman inequality. Lecture Notes in Mathematics, 2116:53–63, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-09477-9_4.pdf.

Bakry:2012:MSS

[BLSC12] Dominique Bakry, Michel Ledoux, and L. Saloff-Coste. Markov semigroups at Saint-Flour, volume ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CO-DEN LNMAA2. ISBN 3-642-25937-5 (paperback). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 357 pp. LCCN QA3

.L28 no. ?? URL http://www.springerlink.com/content/

978-3-642-25983-8. Reprint of lectures originally published in the Lecture Notes in Mathematics volumes 1581 (1994), 1648 (199a6) and 1665 (1997).

Benjamini:2014:PPD

[BM14] Itai Benjamini and Pascal Maillard. Point-to-point distance in first passage percolation on (tree) ×Z. Lecture Notes in Mathematics, 2116:47–51, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-09477-9_3.pdf.

Banasiak:2015:FM

[BMK15] Jacek Banasiak and Mustapha Mokhtar-Kharroubi. Front matter. Lecture Notes in Mathematics, 2126:i–xi, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-11322-7/
1.pdf.

Burger:2013:FM

[BMOR13a] Martin Burger, Andrea C. G. Mennucci, Stanley Osher, and Martin Rumpf. Front matter. Lecture Notes in Mathematics, 2090: i–vii, 2013. CODEN LNMAA2. ISBN 3-319-01711-X (print), 3-319-01712-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-01712-9/1.pdf.

Burger:2013:LSP

[BMOR13b] Martin Burger, Andrea C. G. Mennucci, Stanley Osher, and Martin Rumpf, editors. Level Set and PDE Based Reconstruction Methods in Imaging: Cetraro, Italy 2008, volume 2090 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-01711-X (print), 3-319-01712-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2090.

URL http://link.springer.com/book/10.1007/978-3-319-
01712-9;   http://www.springerlink.com/content/978-3-
319-01712-9.

Bobenko:2011:PMP

[BMS11] Alexander I. Bobenko, Christian Mercat, and Markus Schmies. Period matrices of polyhedral surfaces. Lecture Notes in Mathematics, 2013:213–226, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-17413-1_7/.

Bertoluzza:2012:FM

[BNQ^{+}12a] Silvia Bertoluzza, Ricardo H. Nochetto, Alfio Quarteroni, Kunibert G. Siebert, and Andreas Veeser. Front matter. Lecture Notes in Mathematics, 2040:i–xii, 2012. CODEN LNMAA2. ISBN 3-642-24078-X (print), 3-642-24079-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-24079-9/1.pdf.

Bertoluzza:2012:MAM

[BNQ^{+}12b] Silvia Bertoluzza, Ricardo H. Nochetto, Alfio Quarteroni, Kunibert G. Siebert, and Andreas Veeser, editors. Multiscale and Adaptivity: Modeling, Numerics and Applications: C.I.M.E. Summer School, Cetraro, Italy 2009, volume 2040 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-24078-X (print), 3-642-24079-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2040; QA401 .M815 2012; QA3 .L28 no. 2040; TA342 .C56 2009. URL http:

//link.springer.com/book/10.1007/978-3-642-24079-9;
http://www.springerlink.com/content/978-3-642-24079-9.

Bourgade:2011:EMC

[BNR11] Paul Bourgade, Ashkan Nikeghbali, and Alain Rouault. Ewens measures on compact groups and hypergeometric kernels. Lecture Notes in Mathematics, 2006:351–377, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-15217-7_15/.

Burger:2013:GTZ

[BO13] Martin Burger and Stanley Osher. A guide to the TV zoo. Lecture Notes in Mathematics, 2090:1–70, 2013. CODEN LNMAA2. ISBN 3-319-01711-X (print), 3-319-01712-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01712-9_1/.

Bobenko:2011:FM

[Bob11a] Alexander I. Bobenko. Front matter. Lecture Notes in Mathematics, 2013:65, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-17413-1/2/1.pdf.

Bobenko:2011:ICR

[Bob11b] Alexander I. Bobenko. Introduction to compact Riemann surfaces. Lecture Notes in Mathematics, 2013:3–64, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-17413-1_
1/;   http://link.springer.com/content/pdf/bfm:978-3-
642-17413-1/1.pdf;   http://link.springer.com/content/
pdf/bfm:978-3-642-17413-1/1/1.pdf.

Bobrowski:2015:BCE

[Bob15] Adam Bobrowski. Boundary conditions in evolutionary equations in biology. Lecture Notes in Mathematics, 2126:47–92, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11322-7_2.pdf.

Basse-OConnor:2012:SCP

[BOGP12] Andreas Basse-O’Connor, Svend-Erik Graversen, and Jan Pedersen. Some classes of proper integrals and generalized Ornstein– Uhlenbeck processes. Lecture Notes in Mathematics, 2046:61– 74, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27461-9_3/.

Boileau:2016:TTD

[Boi16] Michel Boileau. Thick/thin decomposition of three-manifolds and the geometrisation conjecture. Lecture Notes in Mathematics, 2166:21–70, 2016. CODEN LNMAA2. ISBN 3-319-42350-9 (print), 3-319-42351-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-42351-7_2.pdf.

Bomze:2010:GOQ

[Bom10] Immanuel M. Bomze. Global optimization: a quadratic programming perspective. Lecture Notes in Mathematics, 1989:1– 53, 2010. CODEN LNMAA2. ISBN 3-642-11338-9 (print), 3-642-11339-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11339-0_1.pdf.

Beyn:2014:SCD

[BORM14] Wolf-Jürgen Beyn, Denny Otten, and Jens Rottmann-Matthes. Stability and computation of dynamic patterns in PDEs. Lecture Notes in Mathematics, 2082:89–172, 2014. CODEN LNMAA2. ISBN 3-319-01299-1 (print), 3-319-01300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01300-8_3/.

Bosch:2014:ART

[Bos14a] Siegfried Bosch. Adic rings and their associated formal schemes. Lecture Notes in Mathematics, 2105:151–173, 2014. CO-DEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-04417-0_7.pdf.

Bosch:2014:AAT

[Bos14b]  Siegfried Bosch.  Affinoid algebras and their associated

spaces. Lecture Notes in Mathematics, 2105:31–63, 2014. CO-DEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-04417-0_3.pdf.

Bosch:2014:AF

[Bos14c] Siegfried Bosch. Affinoid functions. Lecture Notes in Mathematics, 2105:65–91, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-04417-0_4.pdf.

Bosch:2014:BM

[Bos14d] Siegfried Bosch. Back matter. Lecture Notes in Mathematics, 2105:229–256, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-04417-0/1.pdf.

Bosch:2014:CSR

[Bos14e] Siegfried Bosch. Coherent sheaves on rigid spaces. Lecture Notes in Mathematics, 2105:117–147, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-04417-0_6.pdf.

Bosch:2014:FMa

[Bos14f] Siegfried Bosch. Front matter. Lecture Notes in Mathematics, 2105:i–viii, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-04417-0/1.pdf.

Bosch:2014:FMb

[Bos14g] Siegfried Bosch. Front matter. Lecture Notes in Mathematics, 2105:7, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-04417-0/1/1.pdf.

Bosch:2014:FMc

[Bos14h] Siegfried Bosch. Front matter. Lecture Notes in Mathematics, 2105:149, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-04417-0/2/1.pdf.

Bosch:2014:I

[Bos14i] Siegfried Bosch. Introduction. Lecture Notes in Mathematics, 2105:1–5, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-04417-0_1.pdf.

Bosch:2014:LFR

[Bos14j] Siegfried Bosch. Lectures on Formal and Rigid Geometry, volume 2105 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 256

(est.) pp. LCCN ???? URL http://www.springerlink.com/
content/978-3-319-04417-0.

Bosch:2014:MAS

[Bos14k] Siegfried Bosch. More advanced stuff. Lecture Notes in Mathematics, 2105:215–227, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-04417-0_9.pdf.

Bosch:2014:RVR

[Bos14l] Siegfried Bosch. Raynaud’s view on rigid spaces. Lecture Notes in Mathematics, 2105:175–214, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-04417-0_8.pdf.

Bosch:2014:TA

[Bos14m] Siegfried Bosch. Tate algebras. Lecture Notes in Mathematics, 2105:9–29, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-04417-0_2.pdf.

Bosch:2014:TNR

[Bos14n] Siegfried Bosch. Towards the notion of rigid spaces. Lecture Notes in Mathematics, 2105:93–116, 2014. CODEN LNMAA2. ISBN 3-319-04416-8 (print), 3-319-04417-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-04417-0_5.pdf.

Bouc:2010:A

[Bou10a] Serge Bouc. Applications. Lecture Notes in Mathematics, 1990: 215–240, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-11297-3_11.pdf.

Bouc:2010:BM

[Bou10b] Serge Bouc. Back matter. Lecture Notes in Mathematics, 1990: 293–299, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-11297-3/1.pdf.

Bouc:2010:BFa

[Bou10c] Serge Bouc. Biset functors. Lecture Notes in Mathematics, 1990:41–51, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11297-3_3.pdf.

Bouc:2010:BFF

[Bou10d] Serge Bouc. Biset Functors for Finite Groups, volume 1990 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CO-DEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 299 pp. LCCN QA3 .L28 no. 1990. URL http://link.

springer.com/book/10.1007/978-3-642-11297-3;   http:/
/www.springerlink.com/content/978-3-642-11297-3.

Bouc:2010:BFb

[Bou10e] Serge Bouc. The Burnside functor. Lecture Notes in Mathematics, 1990:75–95, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11297-3_5.pdf.

Bouc:2010:DG

[Bou10f] Serge Bouc. The Dade group. Lecture Notes in Mathematics, 1990:241–292, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11297-3_12.pdf.

Bouc:2010:EA

[Bou10g] Serge Bouc. Endomorphism algebras. Lecture Notes in Mathematics, 1990:97–119, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11297-3_6.pdf.

Bouc:2010:E

[Bou10h] Serge Bouc. Examples. Lecture Notes in Mathematics, 1990:1– 11, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11297-3_1.pdf.

Bouc:2010:FMa

[Bou10i] Serge Bouc. Front matter. Lecture Notes in Mathematics, 1990: I–IX, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-11297-3/1.pdf.

Bouc:2010:FMb

[Bou10j] Serge Bouc. Front matter. Lecture Notes in Mathematics, 1990: 14, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-11297-3/1/1.pdf.

Bouc:2010:FMc

[Bou10k] Serge Bouc. Front matter. Lecture Notes in Mathematics, 1990: 74, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-11297-3/2/1.pdf.

Bouc:2010:FMd

[Bou10l] Serge Bouc. Front matter. Lecture Notes in Mathematics, 1990: 154, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-11297-3/3/1.pdf.

Bouc:2010:F

[Bou10m] Serge Bouc. The functor CR_{C}. Lecture Notes in Mathematics, 1990:121–134, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11297-3_7.pdf.

Bouc:2010:SB

[Bou10n] Serge Bouc. G-sets and (H, G)-bisets. Lecture Notes in Mathematics, 1990:15–40, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11297-3_2.pdf.

Bouc:2010:BFc

[Bou10o] Serge Bouc. p-Biset functors. Lecture Notes in Mathematics, 1990:183–213, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11297-3_10.pdf.

Bouc:2010:RRG

[Bou10p] Serge Bouc. Rational representations of p-groups. Lecture Notes in Mathematics, 1990:155–181, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11297-3_9.pdf.

Bouc:2010:SF

[Bou10q] Serge Bouc. Simple functors. Lecture Notes in Mathematics, 1990:53–72, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11297-3_4.pdf.

Bouc:2010:TPI

[Bou10r] Serge Bouc. Tensor product and internal hom. Lecture Notes in Mathematics, 1990:135–152, 2010. CODEN LNMAA2. ISBN 3-642-11296-X (print), 3-642-11297-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11297-3_8.pdf.

Bouleau:2011:LPM

[Bou11] Nicolas Bouleau. The lent particle method for marked point processes. Lecture Notes in Mathematics, 2006:341–349, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
642-15217-7_14/.

Boucksom:2012:MAE

[Bou12a] Sébastien Boucksom. Monge–Ampère equations on complex manifolds with boundary. Lecture Notes in Mathematics, 2038:257– 282, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23669-3_7/.

Bourgain:2012:FSM

[Bou12b] Jean Bourgain. Finitely supported measures on SL_{2}(R) which are absolutely continuous at infinity. Lecture Notes in Mathematics, 2050:133–141, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-29849-3_7/.

Bourgain:2012:MS

[Bou12c] Jean Bourgain. Möbius Schrödinger. Lecture Notes in Mathematics, 2050:143–150, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-29849-3_8/.

Bourgain:2014:IER

[Bou14a] Jean Bourgain. An improved estimate in the restricted isometry problem. Lecture Notes in Mathematics, 2116:65–70, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_5.pdf.

Bourgain:2014:ESD

[Bou14b] Jean Bourgain. On eigenvalue spacings for the 1-D Anderson model with singular site distribution. Lecture Notes in Mathematics, 2116:71–83, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-09477-9_6.pdf.

Bourgain:2014:CPS

[Bou14c] Jean Bourgain. On the control problem for Schrödinger operators on tori. Lecture Notes in Mathematics, 2116:97–105, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_8.pdf.

Bourgain:2014:LES

[Bou14d] Jean Bourgain. On the local eigenvalue spacings for certain Anderson–Bernoulli Hamiltonians. Lecture Notes in Mathematics, 2116:85–96, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-09477-9_7.pdf.

Boylan:2015:BM

[Boy15a] Hatice Boylan. Back matter. Lecture Notes in Mathematics, 2130: 123–132, 2015. CODEN LNMAA2. ISBN 3-319-12915-5 (print), 3-319-12916-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-12916-7/1.pdf.

Boylan:2015:FQM

[Boy15b] Hatice Boylan. Finite quadratic modules. Lecture Notes in Mathematics, 2130:1–17, 2015. CODEN LNMAA2. ISBN 3-319-12915-5 (print), 3-319-12916-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-12916-7_1.pdf.

Boylan:2015:FM

[Boy15c] Hatice Boylan. Front matter. Lecture Notes in Mathematics, 2130:i–xix, 2015. CODEN LNMAA2. ISBN 3-319-12915-5 (print), 3-319-12916-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-12916-7/1.pdf.

Boylan:2015:JFF

[Boy15d] Hatice Boylan. Jacobi Forms, Finite Quadratic Modules and Weil Representations over Number Fields, volume 2130 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LN-MAA2. ISBN 3-319-12915-5 (print), 3-319-12916-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 132 (est.) pp.

LCCN ???? URL http://www.springerlink.com/content/
978-3-319-12916-7.

Boylan:2015:JFT

[Boy15e] Hatice Boylan. Jacobi forms over totally real number fields. Lecture Notes in Mathematics, 2130:65–101, 2015. CO-DEN LNMAA2. ISBN 3-319-12915-5 (print), 3-319-12916-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-12916-7_3.pdf.

Boylan:2015:SJF

[Boy15f] Hatice Boylan. Singular Jacobi forms. Lecture Notes in Mathematics, 2130:103–122, 2015. CODEN LNMAA2. ISBN 3-319-12915-5 (print), 3-319-12916-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-12916-7_4.pdf.

Boylan:2015:WRF

[Boy15g] Hatice Boylan. Weil representations of finite quadratic modules. Lecture Notes in Mathematics, 2130:19–64, 2015. CO-DEN LNMAA2. ISBN 3-319-12915-5 (print), 3-319-12916-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-12916-7_2.pdf.

Benes:2015:STM

[BPHZ15] Viktor Beneš, Michaela Prokešová, Kateřina Staňková Helisová, and Markéta Zikmundová. Space-time models in stochastic geometry. Lecture Notes in Mathematics, 2120:205–232, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10064-7_7.pdf.

Bogso:2012:PGI

[BPR12a] Antoine-Marie Bogso, Christophe Profeta, and Bernard Roynette. On peacocks: a general introduction to two articles. Lecture Notes in Mathematics, 2046:279–280, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27461-9_14.

Bogso:2012:PON

[BPR12b] Antoine-Marie Bogso, Christophe Profeta, and Bernard Roynette. Peacocks obtained by normalisation: Strong and very strong peacocks. Lecture Notes in Mathematics, 2046:317–374, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
642-27461-9_16/.

Bogso:2012:SEP

[BPR12c] Antoine-Marie Bogso, Christophe Profeta, and Bernard Roynette. Some examples of peacocks in a Markovian set-up. Lecture Notes in Mathematics, 2046:281–315, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27461-9_15/.

Barbu:2016:SPMb

[BPR16] Viorel Barbu, Giuseppe Da Prato, and Michael Röckner, editors. Stochastic Porous Media Equations, volume 2163 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. CODEN LNMAA2. ISBN 3-319-41068-7 (print), 3-319-41069-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 204 (est.) pp. LCCN ???? URL

http://www.springerlink.com/content/978-3-319-41069-2.

Belomestny:2015:ECL

[BR15] Denis Belomestny and Markus Reiß. Estimation and calibration of Lévy models via Fourier methods. Lecture Notes in Mathematics, 2128:1–76, 2015. CODEN LNMAA2. ISBN 3-319-12372-6 (print), 3-319-12373-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-12373-8_1.pdf.

Braides:2014:BM

[Bra14a] Andrea Braides. Back matter. Lecture Notes in Mathematics, 2094:173–176, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-01982-6/1.pdf.

Braides:2014:CLM

[Bra14b] Andrea Braides. Convergence of local minimizers. Lecture Notes in Mathematics, 2094:67–78, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01982-6_5/.

Braides:2014:DTS

[Bra14c] Andrea Braides. Different time scales. Lecture Notes in Mathematics, 2094:145–158, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01982-6_10/.

Braides:2014:FM

[Bra14d] Andrea Braides. Front matter. Lecture Notes in Mathematics, 2094:i–xi, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-01982-6/1.pdf.

Braides:2014:GMM

[Bra14e] Andrea Braides. Geometric minimizing movements. Lecture Notes in Mathematics, 2094:129–143, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01982-6_9/.

Braides:2014:GM

[Bra14f] Andrea Braides. Global minimization. Lecture Notes in Mathematics, 2094:7–24, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-01982-6_2/.

Braides:2014:I

[Bra14g] Andrea Braides. Introduction. Lecture Notes in Mathematics, 2094:1–6, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01982-6_1/.

Braides:2014:LMS

[Bra14h] Andrea Braides. Local minimization as a selection criterion. Lecture Notes in Mathematics, 2094:53–66, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01982-6_4/.

Braides:2014:LMV

[Bra14i] Andrea Braides. Local Minimization, Variational Evolution and Γ-Convergence, volume 2094 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 174 pp. LCCN QA3 .L28 no. 2094.

URL http://link.springer.com/book/10.1007/978-3-319-
01982-6;   http://www.springerlink.com/content/978-3-
319-01982-6.

Braides:2014:MM

[Bra14j] Andrea Braides. Minimizing movements. Lecture Notes in Mathematics, 2094:91–101, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01982-6_7/.

Braides:2014:MMA

[Bra14k] Andrea Braides. Minimizing movements along a sequence of functionals. Lecture Notes in Mathematics, 2094:103–128, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01982-6_8/.

Braides:2014:PMD

[Bra14l] Andrea Braides. Parameterized motion driven by global minimization. Lecture Notes in Mathematics, 2094:25–52, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
319-01982-6_3/.

Braides:2014:SSS

[Bra14m] Andrea Braides. Small-scale stability. Lecture Notes in Mathematics, 2094:79–89, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-01982-6_6/.

Braides:2014:ST

[Bra14n] Andrea Braides. Stability theorems. Lecture Notes in Mathematics, 2094:159–171, 2014. CODEN LNMAA2. ISBN 3-319-01981-3 (print), 3-319-01982-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-01982-6_11/.

Bressan:2013:HCL

[Bre13] Alberto Bressan. Hyperbolic conservation laws: an illustrated tutorial. Lecture Notes in Mathematics, 2062:157–245, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32160-3_2/.

Britton:2013:DDE

[Bri13] Nick F. Britton. Destruction and diversity: Effects of habitat loss on ecological communities. Lecture Notes in Mathematics, 2071:307–330, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-35497-7_11/.

Broue:2010:BM

[Bro10a] Michel Broué. Back matter. Lecture Notes in Mathematics, 1988: 119–138, 2010. CODEN LNMAA2. ISBN 3-642-11174-2 (print), 3-642-11175-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-11175-4/1.pdf.

Broue:2010:ERE

[Bro10b] Michel Broué. Eigenspaces and regular elements. Lecture Notes in Mathematics, 1988:97–118, 2010. CODEN LNMAA2. ISBN 3-642-11174-2 (print), 3-642-11175-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11175-4_5.pdf.

Broue:2010:FRG

[Bro10c] Michel Broué. Finite reflection groups in characteristic zero. Lecture Notes in Mathematics, 1988:57–96, 2010. CODEN LNMAA2. ISBN 3-642-11174-2 (print), 3-642-11175-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-11175-4_4.
pdf.

Broue:2010:FM

[Bro10d] Michel Broué. Front matter. Lecture Notes in Mathematics, 1988: I–XI, 2010. CODEN LNMAA2. ISBN 3-642-11174-2 (print), 3-642-11175-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-11175-4/1.pdf.

Broue:2010:ICR

[Bro10e] Michel Broué. Introduction to Complex Reflection Groups and Their Braid Groups, volume 1988 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-11174-2 (print), 3-642-11175-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 138 pp. LCCN QA3 .L28 no.

1988. URL http://link.springer.com/book/10.1007/978-
3-642-11175-4;   http://www.springerlink.com/content/
978-3-642-11175-4.

Broue:2010:PIF

[Bro10f] Michel Broué. Polynomial invariants of finite linear groups. Lecture Notes in Mathematics, 1988:35–56, 2010. CODEN LNMAA2.

ISBN 3-642-11174-2 (print), 3-642-11175-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-11175-4_3.
pdf.

Broue:2010:P

[Bro10g] Michel Broué. Preliminaries. Lecture Notes in Mathematics, 1988: 1–9, 2010. CODEN LNMAA2. ISBN 3-642-11174-2 (print), 3-642-11175-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11175-4_1.pdf.

Broue:2010:PCC

[Bro10h] Michel Broué. Prerequisites and complements in commutative algebra. Lecture Notes in Mathematics, 1988:11–33, 2010. CODEN LNMAA2. ISBN 3-642-11174-2 (print), 3-642-11175-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11175-4_2.pdf.

Brockett:2012:NCL

[Bro12] Roger Brockett. Notes on the control of the Liouville equation. Lecture Notes in Mathematics, 2048:101–129, 2012. CODEN LNMAA2. ISBN 3-642-27892-2 (print), 3-642-27893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27893-8_
2/.

Brown:2013:ESB

[Bro13] Clive M. Brown. Experimental studies of the baroreflex. Lecture Notes in Mathematics, 2064:163–176, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32882-4_9/.

Brunella:2010:UFC

[Bru10] Marco Brunella. Uniformisation of foliations by curves. Lecture Notes in Mathematics, 1998:105–163, 2010. CODEN LNMAA2. ISBN 3-642-13170-0 (print), 3-642-13171-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-13171-4_3.
pdf.

Bruce:2013:ESR

[Bru13] Eugene N. Bruce. Experimental studies of respiration and apnea. Lecture Notes in Mathematics, 2064:121–132, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32882-4_
7/.

Bartumeus:2013:SOF

[BRVdL13] Frederic Bartumeus, Ernesto P. Raposo, Gandhi M. Viswanathan, and Marcos G. E. da Luz. Stochastic optimal foraging theory. Lecture Notes in Mathematics, 2071:3–32, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-35497-7_
1/;   http://link.springer.com/content/pdf/bfm:978-3-
642-35497-7/1.pdf;   http://link.springer.com/content/
pdf/bfm:978-3-642-35497-7/1/1.pdf.

Benjamini:2012:FTG

[BS12] Itai Benjamini and Oded Schramm. Finite transitive graph embeddings into a hyperbolic metric space must stretch or squeeze. Lecture Notes in Mathematics, 2050:123–126, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29849-3_
5/.

Bulinski:2013:CLT

[BS13a] Alexander Bulinski and Evgeny Spodarev. Central limit theorems for weakly dependent random fields. Lecture Notes in Mathematics, 2068:337–383, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-33305-7_10/.

Bulinski:2013:IRF

[BS13b] Alexander Bulinski and Evgeny Spodarev. Introduction to random fields. Lecture Notes in Mathematics, 2068:277–335, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33305-7_9/.

Buffa:2016:FM

[BS16] Annalisa Buffa and Giancarlo Sangalli. Front matter. Lecture Notes in Mathematics, 2161:i–ix, 2016. CODEN LNMAA2. ISBN 3-319-42308-8 (print), 3-319-42309-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-42309-8/1.pdf.

Baum:2011:TGA

[BSG11] Paul F. Baum and Rubén J. Sánchez-García. K-theory for group C∗-algebras. Lecture Notes in Mathematics, 2008:1–43, 2011. CODEN LNMAA2. ISBN 3-642-15707-6 (print), 3-642-15708-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15708-0_1/.

Benes:2015:RMS

[BSKŠ15] Viktor Beneš, Jakub Stanĕk, Blažena Kratochvílová, and Ondřej Šedivý. Random marked sets and dimension reduction. Lecture Notes in Mathematics, 2120:171–203, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-10064-7_6.
pdf.

Bottcher:2013:A

[BSW13a] Björn Böttcher, René Schilling, and Jian Wang. Approximation. Lecture Notes in Mathematics, 2099:167–175, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02684-8_
7/.

Bottcher:2013:BM

[BSW13b] Björn Böttcher, René Schilling, and Jian Wang. Back matter. Lecture Notes in Mathematics, 2099:181–202, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-319-02684-8/
1.pdf.

Bottcher:2013:CFP

[BSW13c] Björn Böttcher, René Schilling, and Jian Wang. Construction of Feller processes. Lecture Notes in Mathematics, 2099:69– 98, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02684-8_3/.

Bottcher:2013:FGS

[BSW13d] Björn Böttcher, René Schilling, and Jian Wang. Feller generators and symbols. Lecture Notes in Mathematics, 2099:31– 67, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02684-8_2/.

Bottcher:2013:FM

[BSW13e] Björn Böttcher, René Schilling, and Jian Wang. Front matter. Lecture Notes in Mathematics, 2099:i–xviii, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-02684-8/
1.pdf.

Bottcher:2013:GP

[BSW13f] Björn Böttcher, René Schilling, and Jian Wang. Global properties. Lecture Notes in Mathematics, 2099:141–165, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02684-8_
6/.

Bottcher:2013:LMI

[BSW13g] Björn Böttcher, René Schilling, and Jian Wang. Lévy Matters III: Lévy-Type Processes: Construction, Approximation and Sample Path Properties, volume 2099 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 199 pp. LCCN QA3 .L28 no. 2099.

URL http://link.springer.com/book/10.1007/978-3-319-
02684-8;   http://www.springerlink.com/content/978-3-
319-02684-8.

Bottcher:2013:OP

[BSW13h] Björn Böttcher, René Schilling, and Jian Wang. Open problems. Lecture Notes in Mathematics, 2099:177–179, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02684-8_
8/.

Bottcher:2013:PFS

[BSW13i] Björn Böttcher, René Schilling, and Jian Wang. A primer on Feller semigroups and Feller processes. Lecture Notes in Mathematics, 2099:1–30, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02684-8_1/.

Bottcher:2013:SPP

[BSW13j] Björn Böttcher, René Schilling, and Jian Wang. Sample path properties. Lecture Notes in Mathematics, 2099:111–140, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02684-8_5/.

Bottcher:2013:TFP

[BSW13k] Björn Böttcher, René Schilling, and Jian Wang. Transformations of Feller processes. Lecture Notes in Mathematics, 2099: 99–110, 2013. CODEN LNMAA2. ISBN 3-319-02683-6 (print), 3-319-02684-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02684-8_4/.

Bourguin:2013:MCS

[BT13] Solesne Bourguin and Ciprian A. Tudor. Malliavin calculus and self normalized sums. Lecture Notes in Mathematics, 2078:323– 351, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00321-4_13/.

Bulinski:2015:SSM

[Bul15] Alexander Bulinski. Some statistical methods in genetics. Lecture Notes in Mathematics, 2120:293–320, 2015. CO-DEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10064-7_10.pdf.

Burdzy:2014:BMc

[Bur14a] Krzysztof Burdzy. Back matter. Lecture Notes in Mathematics, 2106:133–140, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/content/pdf/
bbm:978-3-319-04394-4/1.pdf.

Burdzy:2014:BMb

[Bur14b] Krzysztof Burdzy. Brownian motion. Lecture Notes in Mathematics, 2106:1–10, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-04394-4_1/.

Burdzy:2014:BMA

[Bur14c] Krzysztof Burdzy, editor. Brownian Motion and its Applications to Mathematical Analysis: École d’Été de Probabilités de Saint-Flour XLIII — 2013, volume 2106 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2106; QA3 L4 V.2106;

QA3 .L28 no. 2106. URL http://link.springer.com/book/
10.1007/978-3-319-04394-4;   http://www.springerlink.
com/content/978-3-319-04394-4.

Burdzy:2014:FM

[Bur14d] Krzysztof Burdzy. Front matter. Lecture Notes in Mathematics, 2106:i–xii, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-04394-4/1.pdf.

Burdzy:2014:NEE

[Bur14e] Krzysztof Burdzy. Neumann eigenfunctions and eigenvalues. Lecture Notes in Mathematics, 2106:31–39, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04394-4_4/.

Burdzy:2014:NHK

[Bur14f] Krzysztof Burdzy. Neumann heat kernel monotonicity. Lecture Notes in Mathematics, 2106:97–105, 2014. CODEN LNMAA2.

ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04394-4_9/.

Burdzy:2014:NL

[Bur14g] Krzysztof Burdzy. Nodal lines. Lecture Notes in Mathematics, 2106:89–96, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-04394-4_8/.

Burdzy:2014:OHS

[Bur14h] Krzysztof Burdzy. Overview of the “hot spots” problem. Lecture Notes in Mathematics, 2106:21–29, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04394-4_3/.

Burdzy:2014:PBH

[Bur14i] Krzysztof Burdzy. Parabolic boundary Harnack principle. Lecture Notes in Mathematics, 2106:63–75, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04394-4_6/.

Burdzy:2014:PPC

[Bur14j] Krzysztof Burdzy. Probabilistic proofs of classical theorems. Lecture Notes in Mathematics, 2106:11–19, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04394-4_2/.

Burdzy:2014:RBM

[Bur14k] Krzysztof Burdzy. Reflected Brownian motion in time dependent domains. Lecture Notes in Mathematics, 2106:107– 131, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-04394-4_10/.

Burdzy:2014:SC

[Bur14l] Krzysztof Burdzy. Scaling coupling. Lecture Notes in Mathematics, 2106:77–87, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-04394-4_7/.

Burdzy:2014:SMC

[Bur14m] Krzysztof Burdzy. Synchronous and mirror couplings. Lecture Notes in Mathematics, 2106:41–62, 2014. CODEN LNMAA2. ISBN 3-319-04393-5 (hardcover), 3-319-04394-3. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04394-4_5/.

Bergery:2011:CFS

[BV11] Blandine Bérard Bergery and Pierre Vallois. Convergence at first and second order of some approximations of stochastic integrals. Lecture Notes in Mathematics, 2006:241–268, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_10/.

Baker:2011:MGM

[BY11] David Baker and Marc Yor. On martingales with given marginals and the scaling property. Lecture Notes in Mathematics, 2006: 437–439, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_19/.

Blaszczyszyn:2015:CCP

[BY15] Bartłomiej Błaszczyszyn and Dhandapani Yogeshwaran. Clustering comparison of point processes, with applications to random geometric models. Lecture Notes in Mathematics, 2120:31–71, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10064-7_2.pdf.

Benjamini:2012:TFF

[BZ12] Itai Benjamini and Ofer Zeitouni. Tightness of fluctuations of first passage percolation on some large graphs. Lecture Notes in Mathematics, 2050:127–132, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29849-3_6/.

Bakry:2014:DPA

[BZ14] Dominique Bakry and Marguerite Zani. Dyson processes associated with associative algebras: The Clifford case. Lecture Notes in Mathematics, 2116:1–37, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-09477-9_1.pdf.

Calka:2013:AMR

[Cal13] Pierre Calka. Asymptotic methods for random tessellations. Lecture Notes in Mathematics, 2068:183–204, 2013. CODEN LN-MAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33305-7_
6/.

Cao:2013:KRF

[Cao13] Huai-Dong Cao. The Kähler–Ricci flow on Fano manifolds. Lecture Notes in Mathematics, 2086:239–297, 2013. CODEN LN-MAA2. ISBN 3-319-00818-8 (print), 3-319-00819-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-00819-6_
5/.

Capogna:2014:EMA

[Cap14] Luca Capogna. L^{∞}-extremal mappings in AMLE and Teichmüller theory. Lecture Notes in Mathematics, 2087:1–46, 2014. CODEN LNMAA2. ISBN 3-319-00941-9 (print), 3-319-00942-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00942-1_1/.

Carlen:2011:FID

[Car11] Eric A. Carlen. Functional inequalities and dynamics. Lecture Notes in Mathematics, 2028:17–85, 2011. CODEN LNMAA2. ISBN 3-642-21718-4 (print), 3-642-21861-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-21861-3_2/.

Carmeli:2011:RSL

[CC11] Claudio Carmeli and Gianni Cassinelli. Representations of super Lie groups: Some remarks. Lecture Notes in Mathematics, 2027:

45–67, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21744-9_3/.

Cousin:2011:FM

[CCG^{+}11a] Areski Cousin, Stéphane Crépey, Olivier Guéant, David Hobson, Monique Jeanblanc, Jean-Michel Lasry, Jean-Paul Laurent, Pierre-Louis Lions, and Peter Tankov. Front matter. Lecture Notes in Mathematics, 2003:i–x, 2011. CODEN LNMAA2. ISBN 3-642-14659-7 (print), 3-642-14660-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-14660-2/1.pdf.

Cousin:2011:PPL

[CCG^{+}11b] Areski Cousin, Stéphane Crépey, Olivier Guéant, David Hobson, Monique Jeanblanc, Jean-Michel Lasry, Jean-Paul Laurent, Pierre-Louis Lions, and Peter Tankov, editors. Paris–Princeton Lectures on Mathematical Finance 2010, volume 2003 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-14659-7 (print), 3-642-14660-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2003.

URL http://link.springer.com/book/10.1007/978-3-642-
14660-2;   http://www.springerlink.com/content/978-3-
642-14660-2.

Cenac:2012:CTV

[CCPP12] Peggy Cénac, Brigitte Chauvin, Frédéric Paccaut, and Nicolas Pouyanne. Context trees, variable length Markov chains and dynamical sources. Lecture Notes in Mathematics, 2046:1–39, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-27461-9_1/;   http://link.springer.
com/content/pdf/bfm:978-3-642-27461-9/1.pdf.

Conca:2014:FM

[CDD^{+}14] Aldo Conca, Sandra Di Rocco, Jan Draisma, June Huh, Bernd Sturmfels, and Filippo Viviani. Front matter. Lecture Notes in Mathematics, 2108:i–vii, 2014. CODEN LNMAA2. ISBN 3-319-04869-4 (print), 3-319-04870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-04870-3/1.pdf.

Constantin:2013:FM

[CDG^{+}13a] Peter Constantin, Arnaud Debussche, Giovanni P. Galdi, Michael Růžička, and Gregory Seregin. Front matter. Lecture Notes in Mathematics, 2073:i–ix, 2013. CODEN LNMAA2. ISBN 3-642-36296-6 (print), 3-642-36297-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-36297-2/1.pdf.

Constantin:2013:TMF

[CDG^{+}13b] Peter Constantin, Arnaud Debussche, Giovanni P. Galdi, Michael Růžička, and Gregory Seregin, editors. Topics in Mathematical Fluid Mechanics: Cetraro, Italy 2010, volume 2073 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-36296-6 (print), 3-642-36297-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2073; QA3 .L28 no. 2073; QA901 .T67 2013. URL http://link.

springer.com/book/10.1007/978-3-642-36297-2;   http:/
/www.springerlink.com/content/978-3-642-36297-2.

Cegielski:2013:AO

[Ceg13a] Andrzej Cegielski. Algorithmic operators. Lecture Notes in Mathematics, 2057:39–103, 2013. CODEN LNMAA2. ISBN 3-642-30900-3 (print), 3-642-30901-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30901-4_2/.

Cegielski:2013:APO

[Ceg13b] Andrzej Cegielski. Algorithmic projection operators. Lecture Notes in Mathematics, 2057:129–202, 2013. CODEN LNMAA2. ISBN 3-642-30900-3 (print), 3-642-30901-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30901-4_4/.

Cegielski:2013:BM

[Ceg13c] Andrzej Cegielski. Back matter. Lecture Notes in Mathematics, 2057:275–298, 2013. CODEN LNMAA2. ISBN 3-642-30900-3 (print), 3-642-30901-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-30901-4/1.pdf.

Cegielski:2013:CIM

[Ceg13d] Andrzej Cegielski. Convergence of iterative methods. Lecture Notes in Mathematics, 2057:105–127, 2013. CODEN LNMAA2. ISBN 3-642-30900-3 (print), 3-642-30901-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30901-4_3/.

Cegielski:2013:FM

[Ceg13e] Andrzej Cegielski. Front matter. Lecture Notes in Mathematics, 2057:i–xvi, 2013. CODEN LNMAA2. ISBN 3-642-30900-3 (print), 3-642-30901-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-30901-4/1.pdf.

Cegielski:2013:I

[Ceg13f] Andrzej Cegielski. Introduction. Lecture Notes in Mathematics, 2057:1–38, 2013. CODEN LNMAA2. ISBN 3-642-30900-3 (print), 3-642-30901-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30901-4_1/.

Cegielski:2013:IMF

[Ceg13g] Andrzej Cegielski. Iterative Methods for Fixed Point Problems in Hilbert Spaces, volume 2057 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-30900-3 (print), 3-642-30901-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 298 pp. LCCN QA3 .L28 no. 2057; QA297.8 .C44 2012; QA3 .L28 no. 2057. URL http://link.

springer.com/book/10.1007/978-3-642-30901-4;   http:/
/www.springerlink.com/content/978-3-642-30901-4.

Cegielski:2013:PM

[Ceg13h] Andrzej Cegielski. Projection methods. Lecture Notes in Mathematics, 2057:203–274, 2013. CODEN LNMAA2. ISBN 3-642-30900-3 (print), 3-642-30901-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30901-4_5/.

Cordero-Erausquin:2012:ICG

[CEK12] Dario Cordero-Erausquin and Bo’az Klartag. Interpolations, convexity and geometric inequalities. Lecture Notes in Mathematics, 2050:151–168, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-29849-3_9/.

Cordero-Erausquin:2012:HMT

[CEL12] Dario Cordero-Erausquin and Michel Ledoux. Hypercontractive measures, Talagrand’s inequality, and influences. Lecture Notes in Mathematics, 2050:169–189, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29849-3_10/.

Cervantes:2011:CQS

[CFL11] D. Cervantes, R. Fioresi, and M. A. Lledó. On chiral quantum superspaces. Lecture Notes in Mathematics, 2027:69–99, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21744-9_4/.

Cattiaux:2014:SLC

[CG14] P. Cattiaux and A. Guillin. Semi log-concave Markov diffusions. Lecture Notes in Mathematics, 2123:231–292, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_9.pdf.

Comte:2015:AEL

[CGC15] Fabienne Comte and Valentine Genon-Catalot. Adaptive estimation for Lévy processes. Lecture Notes in Mathematics, 2128:77– 177, 2015. CODEN LNMAA2. ISBN 3-319-12372-6 (print), 3-319-12373-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-12373-8_2.pdf.

Capogna:2014:FM

[CGGM14a] Luca Capogna, Pengfei Guan, Cristian E. Gutiérrez, and Annamaria Montanari. Front matter. Lecture Notes in Mathematics, 2087:i–xi, 2014. CODEN LNMAA2. ISBN 3-319-00941-9 (print), 3-319-00942-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-00942-1/1.pdf.

Capogna:2014:FNP

[CGGM14b] Luca Capogna, Pengfei Guan, Cristian E. Gutiérrez, and Annamaria Montanari, editors. Fully Nonlinear PDEs in Real and Complex Geometry and Optics: Cetraro, Italy 2012, volume 2087 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CO-DEN LNMAA2. ISBN 3-319-00941-9 (print), 3-319-00942-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 2087. URL http://link.springer.com/book/
10.1007/978-3-319-00942-1;   http://www.springerlink.
com/content/978-3-319-00942-1.

Console:2014:LTS

[CH14] Paola Console and Ernst Hairer. Long-term stability of symmetric partitioned linear multistep methods. Lecture Notes in Mathematics, 2082:1–37, 2014. CODEN LNMAA2. ISBN 3-319-01299-1 (print), 3-319-01300-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-01300-8_1/.

Chang:2014:MOF

[Cha14] Yinshan Chang. Multi-occupation field generates the Borel– sigma-field of loops. Lecture Notes in Mathematics, 2123:401–410, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_17.pdf.

Chung:2010:CAL

[Chu10] Kai Lai Chung, editor. Collected articles from LNM: a special selection on the occasion of the memorial conference for Kai Lai Chung, Beijing 13–16 June, 2010, Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-12694-4 (print), 3-642-12695-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA273 .C495 2010.

Cousin:2011:HCT

[CJL11] Areski Cousin, Monique Jeanblanc, and Jean-Paul Laurent. Hedging CDO tranches in a Markovian environment. Lecture Notes in Mathematics, 2003:1–61, 2011. CODEN LNMAA2. ISBN 3-642-14659-7 (print), 3-642-14660-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-14660-2_1/.

Cohen:2012:FM

[CKKR12a] Serge Cohen, Alexey Kuznetsov, Andreas E. Kyprianou, and Victor Rivero. Front matter. Lecture Notes in Mathematics, 2061: i–xii, 2012. CODEN LNMAA2. ISBN 3-642-31406-6 (print), 3-642-31407-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-31407-0/1.pdf.

Cohen:2012:LMI

[CKKR12b] Serge Cohen, Alexey Kuznetsov, Andreas E. Kyprianou, and Victor Rivero. Lévy Matters II: Recent Progress in Theory and Applications: Fractional Lévy Fields, and Scale Functions, volume 2061 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-31406-6 (print), 3-642-31407-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 186 pp. LCCN QA3 .L28 no. 2061; QA3 .L28 no. 2061;

QA274.73 .L488 2012. URL http://link.springer.com/book/
10.1007/978-3-642-31407-0;   http://www.springerlink.
com/content/978-3-642-31407-0.

Capietto:2013:FM

[CKM^{+}13a] Anna Capietto, Peter Kloeden, Jean Mawhin, Sylvia Novo, and Rafael Ortega. Front matter. Lecture Notes in Mathematics, 2065:i–ix, 2013. CODEN LNMAA2. ISBN 3-642-32905-5 (print), 3-642-32906-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-32906-7/1.pdf.

Capietto:2013:SBT

[CKM^{+}13b] Anna Capietto, Peter Kloeden, Jean Mawhin, Sylvia Novo, and Rafael Ortega, editors. Stability and Bifurcation Theory for Non-Autonomous Differential Equations: Cetraro, Italy 2011, volume 2065 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-32905-5 (print), 3-642-32906-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2065; QA372 .S73 2013. URL http://link.

springer.com/book/10.1007/978-3-642-32906-7;   http:/
/www.springerlink.com/content/978-3-642-32906-7.

Catuogno:2013:NSC

[CLR13] Pedro J. Catuogno, Diego S. Ledesma, and Paulo R. Ruffino. A note on stochastic calculus in vector bundles. Lecture Notes in Mathematics, 2078:353–364, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00321-4_14/.

Caselles:2010:Aa

[CM10a] Vicent Caselles and Pascal Monasse. Applications. Lecture Notes in Mathematics, 1984:155–171, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04611-7_8.pdf.

Caselles:2010:Ab

[CM10b] Vicent Caselles and Pascal Monasse. Applications. Lecture Notes in Mathematics, 1984:155–171, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04611-7_8.pdf.

Caselles:2010:BMa

[CM10c] Vicent Caselles and Pascal Monasse. Back matter. Lecture Notes in Mathematics, 1984:1–18, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-04611-7/1.pdf.

Caselles:2010:BMb

[CM10d] Vicent Caselles and Pascal Monasse. Back matter. Lecture Notes in Mathematics, 1984:1–18, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-04611-7/1.pdf.

Caselles:2010:CTBa

[CM10e] Vicent Caselles and Pascal Monasse. Computation of the tree of bilinear level lines. Lecture Notes in Mathematics, 1984:141–153, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04611-7_7.pdf.

Caselles:2010:CTBb

[CM10f] Vicent Caselles and Pascal Monasse. Computation of the tree of bilinear level lines. Lecture Notes in Mathematics, 1984:141–153, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04611-7_7.pdf.

Caselles:2010:CTSa

[CM10g] Vicent Caselles and Pascal Monasse. Computation of the tree of shapes of a digital image. Lecture Notes in Mathematics, 1984:115–140, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-04611-7_6.pdf.

Caselles:2010:CTSb

[CM10h] Vicent Caselles and Pascal Monasse. Computation of the tree of shapes of a digital image. Lecture Notes in Mathematics, 1984:115–140, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-04611-7_6.pdf.

Caselles:2010:FMa

[CM10i] Vicent Caselles and Pascal Monasse. Front matter. Lecture Notes in Mathematics, 1984:1–14, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-04611-7/1.pdf.

Caselles:2010:FMb

[CM10j] Vicent Caselles and Pascal Monasse. Front matter. Lecture Notes in Mathematics, 1984:1–14, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-04611-7/1.pdf.

Caselles:2010:GDI

[CM10k] Vicent Caselles and Pascal Monasse, editors. Geometric Description of Images as Topographic Maps, volume 1984 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LN-MAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvii + 186 pp. LCCN QA3 .L28 no. 1984; TA1637 .C374 2010. URL http:

//link.springer.com/book/10.1007/978-3-642-04611-7;
http://www.springerlink.com/content/978-3-642-04611-7.

Caselles:2010:GFa

[CM10l] Vicent Caselles and Pascal Monasse. Grain filters. Lecture Notes in Mathematics, 1984:35–73, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04611-7_3.pdf.

Caselles:2010:GFb

[CM10m] Vicent Caselles and Pascal Monasse. Grain filters. Lecture Notes in Mathematics, 1984:35–73, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04611-7_3.pdf.

Caselles:2010:Ia

[CM10n] Vicent Caselles and Pascal Monasse. Introduction. Lecture Notes in Mathematics, 1984:1–7, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04611-7_1.pdf.

Caselles:2010:Ib

[CM10o] Vicent Caselles and Pascal Monasse. Introduction. Lecture Notes in Mathematics, 1984:1–7, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04611-7_1.pdf.

Caselles:2010:MCTa

[CM10p] Vicent Caselles and Pascal Monasse. Merging the component trees. Lecture Notes in Mathematics, 1984:103–113, 2010. CO-DEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04611-7_5.pdf.

Caselles:2010:MCTb

[CM10q] Vicent Caselles and Pascal Monasse. Merging the component trees. Lecture Notes in Mathematics, 1984:103–113, 2010. CO-DEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04611-7_5.pdf.

Caselles:2010:TDTa

[CM10r] Vicent Caselles and Pascal Monasse. A topological description of the topographic map. Lecture Notes in Mathematics, 1984:75–102, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-04611-7_4.pdf.

Caselles:2010:TDTb

[CM10s] Vicent Caselles and Pascal Monasse. A topological description of the topographic map. Lecture Notes in Mathematics, 1984:75–102, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-04611-7_4.pdf.

Caselles:2010:TSIa

[CM10t] Vicent Caselles and Pascal Monasse. The tree of shapes of an image. Lecture Notes in Mathematics, 1984:9–34, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04611-7_2.pdf.

Caselles:2010:TSIb

[CM10u] Vicent Caselles and Pascal Monasse. The tree of shapes of an image. Lecture Notes in Mathematics, 1984:9–34, 2010. CODEN LNMAA2. ISBN 3-642-04610-X (print), 3-642-04611-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04611-7_2.pdf.

Crowdy:2011:URH

[CM11] Darren Crowdy and Jonathan S. Marshall. Uniformizing real hyperelliptic M-curves using the Schottky–Klein prime function. Lecture Notes in Mathematics, 2013:183–193, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-17413-1_
5/.

Crisan:2013:BM

[CMN13a] D. Crisan, K. Manolarakis, and C. Nee. Back matter. Lecture Notes in Mathematics, 2081:317–318, 2013. CODEN LNMAA2. ISBN 3-319-00412-3 (print), 3-319-00413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-319-00413-6/1.pdf.

Crisan:2013:CMA

[CMN13b] D. Crisan, K. Manolarakis, and C. Nee. Cubature methods and applications. Lecture Notes in Mathematics, 2081:203–316, 2013. CODEN LNMAA2. ISBN 3-319-00412-3 (print), 3-319-00413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00413-6_4/.

Cohen:2012:FLF

[Coh12] Serge Cohen. Fractional Lévy fields. Lecture Notes in Mathematics, 2061:1–95, 2012. CODEN LNMAA2. ISBN 3-642-31406-6 (print), 3-642-31407-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-31407-0_1/.

Constantin:2013:CFL

[Con13] Peter Constantin. Complex fluids and Lagrangian particles. Lecture Notes in Mathematics, 2073:1–21, 2013. CODEN LNMAA2. ISBN 3-642-36296-6 (print), 3-642-36297-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-36297-2_1/.

Conca:2014:KAT

[Con14] Aldo Conca. Koszul algebras and their syzygies. Lecture Notes in Mathematics, 2108:1–31, 2014. CODEN LNMAA2. ISBN 3-319-04869-4 (print), 3-319-04870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-04870-3_1.pdf.

Cortinas:2011:AVT

[Cor11] Guillermo Cortiñas. Algebraic v. topological K-theory: a friendly match. Lecture Notes in Mathematics, 2008:103–165, 2011. CODEN LNMAA2. ISBN 3-642-15707-6 (print), 3-642-15708-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15708-0_3/.

Coulibaly-Pasquier:2014:OMF

[CP14] Koléhè A. Coulibaly-Pasquier. Onsager–Machlup functional for uniformly elliptic time-inhomogeneous diffusion. Lecture Notes in Mathematics, 2123:105–123, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11970-0_5.pdf.

Clemson:2013:CNO

[CPSS13] Philip T. Clemson, Spase Petkoski, Tomislav Stankovski, and Aneta Stefanovska. Coupled nonautonomous oscillators. Lecture Notes in Mathematics, 2102:163–197, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-03080-7_5/.

Crepey:2011:APE

[Cré11] Stéphane Crépey. About the pricing equations in finance. Lecture Notes in Mathematics, 2003:63–203, 2011. CODEN LNMAA2. ISBN 3-642-14659-7 (print), 3-642-14660-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-14660-2_2/.

Czichowsky:2011:CSTb

[CS11] Christoph Czichowsky and Martin Schweizer. Closedness in the semimartingale topology for spaces of stochastic integrals with constrained integrands. Lecture Notes in Mathematics, 2006:413– 436, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_18/.

Chen:2014:UBC

[CS14] William Chen and Maxim Skriganov. Upper bounds in classical discrepancy theory. Lecture Notes in Mathematics, 2107:3–69, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-04696-9_1.pdf.

Chen:2014:FMa

[CST14a] William Chen, Anand Srivastav, and Giancarlo Travaglini. Front matter. Lecture Notes in Mathematics, 2107:i–xvi, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-04696-9/1.pdf.

Chen:2014:FMb

[CST14b] William Chen, Anand Srivastav, and Giancarlo Travaglini. Front matter. Lecture Notes in Mathematics, 2107:1, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-04696-9/
1/1.pdf.

Colliot-Thelene:2010:VPR

[CT10] Jean-Louis Colliot-Thélène. Variétés presque rationnelles, leurs points rationnels et leurs dégénérescences. (French) [nearly rational varieties, their rational points, and their degenerations]. Lecture Notes in Mathematics, 2009:1–44, 2010. CO-DEN LNMAA2. ISBN 3-642-15944-3 (print), 3-642-15945-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-15945-9_1.pdf.

Cuchiero:2013:PPR

[CT13] Christa Cuchiero and Josef Teichmann. Path properties and regularity of affine processes on general state spaces. Lecture Notes in Mathematics, 2078:201–244, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00321-4_8/.

Colliot-Thelene:2010:AGL

[CTSDV10a] Jean-Louis Colliot-Thélène, Peter Swinnerton-Dyer, and Paul Vojta, editors. Arithmetic Geometry: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, September 10-15, 2007, volume 2009 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-15944-3 (print), 3-642-15945-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2009. URL http://link.

springer.com/book/10.1007/978-3-642-15945-9;   http:/
/www.springerlink.com/content/978-3-642-15945-9.

Colliot-Thelene:2010:FM

[CTSDV10b] Jean-Louis Colliot-Thélène, Peter Swinnerton-Dyer, and Paul Vojta. Front matter. Lecture Notes in Mathematics, 2009:i– xi, 2010. CODEN LNMAA2. ISBN 3-642-15944-3 (print), 3-642-15945-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-15945-9/1.pdf.

Czichowsky:2011:CSTa

[CWZ11] Christoph Czichowsky, Nicholas Westray, and Harry Zheng. Convergence in the semimartingale topology and constrained portfolios. Lecture Notes in Mathematics, 2006:395–412, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-15217-7_
17/.

Dellacherie:2013:FIO

[DÉ13] Claude Dellacherie and Michel Émery. Filtrations indexed by ordinals; application to a conjecture of S. Laurent. Lecture Notes in Mathematics, 2078:141–157, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00321-4_4/.

Dearricott:2014:LSM

[Dea14] Owen Dearricott. Lectures on n-Sasakian manifolds. Lecture Notes in Mathematics, 2110:57–109, 2014. CODEN LNMAA2. ISBN 3-319-06372-3 (print), 3-319-06373-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-06373-7_4.
pdf.

Debussche:2013:ERS

[Deb13] Arnaud Debussche. Ergodicity results for the stochastic Navier– Stokes equations: an introduction. Lecture Notes in Mathematics, 2073:23–108, 2013. CODEN LNMAA2. ISBN 3-642-36296-6 (print), 3-642-36297-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-36297-2_2/.

Defant:2011:BM

[Def11a] Andreas Defant. Back matter. Lecture Notes in Mathematics, 2021:159–173, 2011. CODEN LNMAA2. ISBN 3-642-20437-6 (print), 3-642-20438-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-20438-8/1.pdf.

Defant:2011:CSC

[Def11b] Andreas Defant. Classical Summation in Commutative and Noncommutative L_{p}-Spaces, volume 2021 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-20437-6 (print), 3-642-20438-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 171 pp. LCCN QA3 .L28 no.

2021. URL http://link.springer.com/book/10.1007/978-
3-642-20438-8;   http://www.springerlink.com/content/
978-3-642-20438-8.

Defant:2011:CT

[Def11c] Andreas Defant. Commutative theory. Lecture Notes in Mathematics, 2021:15–78, 2011. CODEN LNMAA2. ISBN 3-642-20437-6 (print), 3-642-20438-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-20438-8_2/.

Defant:2011:FM

[Def11d] Andreas Defant. Front matter. Lecture Notes in Mathematics, 2021:i–viii, 2011. CODEN LNMAA2. ISBN 3-642-20437-6 (print), 3-642-20438-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-20438-8/1.pdf.

Defant:2011:I

[Def11e] Andreas Defant. Introduction. Lecture Notes in Mathematics, 2021:1–13, 2011. CODEN LNMAA2. ISBN 3-642-20437-6 (print), 3-642-20438-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-20438-8_1/.

Defant:2011:NT

[Def11f] Andreas Defant. Noncommutative theory. Lecture Notes in Mathematics, 2021:79–158, 2011. CODEN LNMAA2. ISBN 3-642-20437-6 (print), 3-642-20438-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-20438-8_3/.

Dellacherie:2011:IPS

[Del11] Claude Dellacherie. On isomorphic probability spaces. Lecture Notes in Mathematics, 2006:187–189, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-15217-7_6/.

Delarue:2012:FM

[Del12a] François Delarue. Front matter. Lecture Notes in Mathematics, 2038:199, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-23669-3/3/1.pdf.

Delarue:2012:PAR

[Del12b] François Delarue. Probabilistic approach to regularity. Lecture Notes in Mathematics, 2038:55–198, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-23669-3_4/.

Demyanov:2010:NO

[Dem10] Vladimir F. Demyanov. Nonsmooth optimization. Lecture Notes in Mathematics, 1989:55–163, 2010. CODEN LNMAA2. ISBN 3-642-11338-9 (print), 3-642-11339-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11339-0_2.pdf.

Demailly:2013:APT

[Dem13] Jean-Pierre Demailly. Applications of pluripotential theory to algebraic geometry. Lecture Notes in Mathematics, 2075:143– 263, 2013. CODEN LNMAA2. ISBN 3-642-36420-9 (print), 3-642-36421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-36421-1_3/.

Ducros:2015:FMa

[DFN15a] Antoine Ducros, Charles Favre, and Johannes Nicaise. Front matter. Lecture Notes in Mathematics, 2119:i–xix, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-11029-5/
1.pdf.

Ducros:2015:FMb

[DFN15b] Antoine Ducros, Charles Favre, and Johannes Nicaise. Front matter. Lecture Notes in Mathematics, 2119:1, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-11029-5/
1/1.pdf.

Detinko:2013:PGT

[DFO13] Alla Detinko, Dane Flannery, and Eamonn O’Brien, editors. Probabilistic Group Theory, Combinatorics, and Computing: Lectures from the Fifth de Brún Workshop, volume 2070 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 1-4471-4813-4 (print), 1-4471-4814-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2070; QA3 .L28 no. 2070; QA174 .P76 2013. URL http://link.

springer.com/book/10.1007/978-1-4471-4814-2;   http:/
/www.springerlink.com/content/978-1-4471-4814-2.

deFreitas:2013:RDS

[dFS13] Michael Marcondes de Freitas and Eduardo D. Sontag. Random dynamical systems with inputs. Lecture Notes in Mathematics, 2102:41–87, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-03080-7_2/.

Dujardin:2012:FM

[DG12a] Romain Dujardin and Vincent Guedj. Front matter. Lecture Notes in Mathematics, 2038:53, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-23669-3/2/1.pdf.

Dujardin:2012:GPM

[DG12b] Romain Dujardin and Vincent Guedj. Geometric properties of maximal psh functions. Lecture Notes in Mathematics, 2038: 33–52, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23669-3_3/.

Dawson:2014:BM

[DG14a] Donald A. Dawson and Andreas Greven. Back matter. Lecture Notes in Mathematics, 2092:841–858, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-319-02153-9/1.pdf.

Dawson:2014:BTD

[DG14b] Donald A. Dawson and Andreas Greven. A basic tool: Dual representations. Lecture Notes in Mathematics, 2092:105–145, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02153-9_5/.

Dawson:2014:ELO

[DG14c] Donald A. Dawson and Andreas Greven. Emergence with M ≥ 2 lower order types (phases 0, 1, 2). Lecture Notes in Mathematics, 2092:377–714, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02153-9_9/.

Dawson:2014:FMR

[DG14d] Donald A. Dawson and Andreas Greven. Formulation of the main results in the general case. Lecture Notes in Mathematics, 2092: 55–104, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02153-9_4/.

Dawson:2014:FMM

[DG14e] Donald A. Dawson and Andreas Greven. Formulation of the multitype and multiscale model. Lecture Notes in Mathematics, 2092: 39–53, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02153-9_3/.

Dawson:2014:FM

[DG14f] Donald A. Dawson and Andreas Greven. Front matter. Lecture Notes in Mathematics, 2092:i–xvii, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-319-02153-9/1.pdf.

Dawson:2014:GTM

[DG14g] Donald A. Dawson and Andreas Greven. The general (M, M)- type mean-field model: Emergence, fixation and droplets. Lecture Notes in Mathematics, 2092:715–780, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02153-9_10/.

Dawson:2014:I

[DG14h] Donald A. Dawson and Andreas Greven. Introduction. Lecture Notes in Mathematics, 2092:1–10, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-02153-9_1/.

Dawson:2014:ICE

[DG14i] Donald A. Dawson and Andreas Greven. Iteration of the cycle I: Emergence and fixation on E_{2}. Lecture Notes in Mathematics, 2092:811–828, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02153-9_13/.

Dawson:2014:ICI

[DG14j] Donald A. Dawson and Andreas Greven. Iteration of the cycle II: Extension to the general multilevel hierarchy. Lecture Notes in Mathematics, 2092:829–837, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02153-9_14/.

Dawson:2014:LTB

[DG14k] Donald A. Dawson and Andreas Greven. Long-time behaviour: Ergodicity and non-ergodicity. Lecture Notes in Mathematics, 2092:147–159, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02153-9_6/.

Dawson:2014:MFEb

[DG14l] Donald A. Dawson and Andreas Greven. Mean-field emergence and fixation of rare mutants: Concepts, strategy and a caricature model. Lecture Notes in Mathematics, 2092:161–165, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02153-9_7/.

Dawson:2014:MFEa

[DG14m] Donald A. Dawson and Andreas Greven. Mean-field emergence and fixation of rare mutants in the Fisher–Wright model with two types. Lecture Notes in Mathematics, 2092:11–38, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02153-9_2/.

Dawson:2014:MPF

[DG14n] Donald A. Dawson and Andreas Greven. Methods and proofs for the Fisher–Wright model with two types. Lecture Notes in Mathematics, 2092:167–375, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02153-9_8/.

Dawson:2014:NEA

[DG14o] Donald A. Dawson and Andreas Greven. Neutral evolution on E_{1} after fixation (phase 3). Lecture Notes in Mathematics, 2092: 781–786, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02153-9_11/.

Dawson:2014:REH

[DG14p] Donald A. Dawson and Andreas Greven. Re-equilibration on higher level E_{1} (phase 4). Lecture Notes in Mathematics, 2092: 787–810, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02153-9_12/.

Dawson:2014:SFV

[DG14q] Donald A. Dawson and Andreas Greven. Spatial Fleming–Viot Models with Selection and Mutation, volume 2092 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 856 pp. LCCN

QA3 .L28 no. 2092. URL http://link.springer.com/book/
10.1007/978-3-319-02153-9;   http://www.springerlink.
com/content/978-3-319-02153-9.

Dawson:2014:WPT

[DG14r] Donald A. Dawson and Andreas Greven. Winding-up: Proofs of the theorems 3–11. Lecture Notes in Mathematics, 2092: 839, 2014. CODEN LNMAA2. ISBN 3-319-02152-4 (print), 3-319-02153-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02153-9_15.

Dearricott:2014:FM

[DGGK^{+}14] Owen Dearricott, Fernando Galaz-García, Lee Kennard, Catherine Searle, Gregor Weingart, and Wolfgang Ziller. Front matter. Lecture Notes in Mathematics, 2110:i–vii, 2014. CODEN LNMAA2. ISBN 3-319-06372-3 (print), 3-319-06373-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-06373-7/
1.pdf.

Damon:2016:ACS

[DGH16a] James Damon, Peter Giblin, and Gareth Haslinger. Abstract classification of singularities preserving features. Lecture Notes in Mathematics, 2165:35–39, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-41471-3_4.pdf.

Damon:2016:ACP

[DGH16b] James Damon, Peter Giblin, and Gareth Haslinger. Apparent contours for projections of smooth surfaces. Lecture Notes in Mathematics, 2165:23–33, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-41471-3_3.pdf.

Damon:2016:BM

[DGH16c] James Damon, Peter Giblin, and Gareth Haslinger. Back matter. Lecture Notes in Mathematics, 2165:253–257, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-319-41471-3/
1.pdf.

Damon:2016:CSM

[DGH16d] James Damon, Peter Giblin, and Gareth Haslinger. Classifications of stable multilocal configurations and their generic transitions. Lecture Notes in Mathematics, 2165:243–252, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-41471-3_14.pdf.

Damon:2016:FMa

[DGH16e] James Damon, Peter Giblin, and Gareth Haslinger. Front matter. Lecture Notes in Mathematics, 2165:i–x, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-41471-3/
1.pdf.

Damon:2016:FMb

[DGH16f] James Damon, Peter Giblin, and Gareth Haslinger. Front matter. Lecture Notes in Mathematics, 2165:1, 2016. CODEN LN-MAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-41471-3/
1/1.pdf.

Damon:2016:FMc

[DGH16g] James Damon, Peter Giblin, and Gareth Haslinger. Front matter. Lecture Notes in Mathematics, 2165:21, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-41471-3/
2/1.pdf.

Damon:2016:FMd

[DGH16h] James Damon, Peter Giblin, and Gareth Haslinger. Front matter. Lecture Notes in Mathematics, 2165:115, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-41471-3/
3/1.pdf.

Damon:2016:FMe

[DGH16i] James Damon, Peter Giblin, and Gareth Haslinger. Front matter. Lecture Notes in Mathematics, 2165:179, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-41471-3/
4/1.pdf.

Damon:2016:FMf

[DGH16j] James Damon, Peter Giblin, and Gareth Haslinger. Front matter. Lecture Notes in Mathematics, 2165:213, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-41471-3/
5/1.pdf.

Damon:2016:I

[DGH16k] James Damon, Peter Giblin, and Gareth Haslinger. Introduction. Lecture Notes in Mathematics, 2165:3–10, 2016. CO-DEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-41471-3_1.pdf.

Damon:2016:LFN

[DGH16l] James Damon, Peter Giblin, and Gareth Haslinger, editors. Local Features in Natural Images via Singularity Theory, volume 2165 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 257

(est.) pp. LCCN ???? URL http://www.springerlink.com/
content/978-3-319-41471-3.

Damon:2016:MCS

[DGH16m] James Damon, Peter Giblin, and Gareth Haslinger. Methods for classification of singularities. Lecture Notes in Mathematics, 2165:73–99, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-41471-3_6.pdf.

Damon:2016:MTC

[DGH16n] James Damon, Peter Giblin, and Gareth Haslinger. Methods for topological classification of singularities. Lecture Notes in Mathematics, 2165:101–114, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-41471-3_7.pdf.

Damon:2016:O

[DGH16o] James Damon, Peter Giblin, and Gareth Haslinger. Overview. Lecture Notes in Mathematics, 2165:11–20, 2016. CODEN LN-MAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-319-41471-
3_2.pdf.

Damon:2016:RAM

[DGH16p] James Damon, Peter Giblin, and Gareth Haslinger. Realizations of abstract mappings representing projection singularities. Lecture Notes in Mathematics, 2165:135–155, 2016.

CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-41471-3_9.pdf.

Damon:2016:SEG

[DGH16q] James Damon, Peter Giblin, and Gareth Haslinger. Singularity equivalence groups capturing interactions. Lecture Notes in Mathematics, 2165:41–71, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-41471-3_5.pdf.

Damon:2016:SVP

[DGH16r] James Damon, Peter Giblin, and Gareth Haslinger. Stable view projections and transitions involving shade/shadow curves on a smooth surface (SC). Lecture Notes in Mathematics, 2165:181– 191, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-41471-3_11.pdf.

Damon:2016:SMC

[DGH16s] James Damon, Peter Giblin, and Gareth Haslinger. Statements of the main classification results. Lecture Notes in Mathematics, 2165:157–177, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-41471-3_10.pdf.

Damon:2016:SGI

[DGH16t] James Damon, Peter Giblin, and Gareth Haslinger. Stratifications of generically illuminated surfaces with geometric features. Lecture Notes in Mathematics, 2165:117–134, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-41471-3_8.pdf.

Damon:2016:TIG

[DGH16u] James Damon, Peter Giblin, and Gareth Haslinger. Transitions involving geometric features and shade/shadow curves (SFC). Lecture Notes in Mathematics, 2165:215–241, 2016.

CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-41471-3_13.pdf.

Damon:2016:TIV

[DGH16v] James Damon, Peter Giblin, and Gareth Haslinger. Transitions involving views of geometric features (FC). Lecture Notes in Mathematics, 2165:193–212, 2016. CODEN LNMAA2. ISBN 3-319-41470-4 (print), 3-319-41471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-41471-3_12.pdf.

Discacciati:2012:HMM

[DGQ12] Marco Discacciati, Paola Gervasio, and Alfio Quarteroni. Heterogeneous mathematical models in fluid dynamics and associated solution algorithms. Lecture Notes in Mathematics, 2040: 57–123, 2012. CODEN LNMAA2. ISBN 3-642-24078-X (print), 3-642-24079-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-24079-9_2/.

Doerr:2014:BM

[DGW14a] Carola Doerr, Michael Gnewuch, and Magnus Wahlström. Back matter. Lecture Notes in Mathematics, 2107:679–698, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-04696-9/1.pdf.

Doerr:2014:CDM

[DGW14b] Carola Doerr, Michael Gnewuch, and Magnus Wahlström. Calculation of discrepancy measures and applications. Lecture Notes in Mathematics, 2107:621–678, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-04696-9_10.pdf.

Doerr:2014:PDT

[DGW14c] Carola Doerr, Michael Gnewuch, and Magnus Wahlström. A Panorama of Discrepancy Theory, volume 2107 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 698 (est.) pp. LCCN ???? URL

http://www.springerlink.com/content/978-3-319-04696-9.

Demni:2012:SDF

[DH12] Nizar Demni and Taoufik Hmidi. Spectral distribution of the free unitary Brownian motion: Another approach. Lecture Notes in Mathematics, 2046:191–206, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27461-9_9/.

Diening:2011:BM

[DHHR11a] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Back matter. Lecture Notes in Mathematics, 2017: 483–509, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-18363-8/1.pdf.

Diening:2011:C

[DHHR11b] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Capacities. Lecture Notes in Mathematics, 2017:315– 338, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-18363-8_10/.

Diening:2011:CO

[DHHR11c] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Classical operators. Lecture Notes in Mathematics, 2017: 199–212, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-18363-8_6/.

Diening:2011:DRF

[DHHR11d] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Density of regular functions. Lecture Notes in Mathematics, 2017:289–314, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18363-8_9/.

Diening:2011:DEI

[DHHR11e] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Dirichlet energy integral and Laplace equation. Lecture Notes in Mathematics, 2017:401–436, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-18363-8_13/.

Diening:2011:FPS

[DHHR11f] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Fine properties of Sobolev functions. Lecture Notes in Mathematics, 2017:339–366, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18363-8_11/.

Diening:2011:FFS

[DHHR11g] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. A framework for function spaces. Lecture Notes in Mathematics, 2017:21–68, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18363-8_2/.

Diening:2011:FMa

[DHHR11h] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Ruzicka. Front matter. Lecture Notes in Mathematics, 2017:i– ix, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-18363-8/1.pdf.

Diening:2011:FMb

[DHHR11i] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Front matter. Lecture Notes in Mathematics, 2017: 19, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-18363-8/1/1.pdf.

Diening:2011:FMc

[DHHR11j] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Front matter. Lecture Notes in Mathematics, 2017:

245, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-18363-8/2/1.pdf.

Diening:2011:FMd

[DHHR11k] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Front matter. Lecture Notes in Mathematics, 2017: 399, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-18363-8/3/1.pdf.

Diening:2011:GMC

[DHHR11l] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. The generalized Muckenhoupt condition. Lecture Notes in Mathematics, 2017:143–197, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18363-8_5/.

Diening:2011:I

[DHHR11m] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Introduction. Lecture Notes in Mathematics, 2017:1– 17, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-18363-8_1/.

Diening:2011:ISS

[DHHR11n] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Introduction to Sobolev spaces. Lecture Notes in Mathematics, 2017:247–288, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18363-8_8/.

Diening:2011:LSS

[DHHR11o] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Lebesgue and Sobolev Spaces with Variable Exponents, volume 2017 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 509 pp. LCCN QA3 .L28 no. 2017. URL http://link.

springer.com/book/10.1007/978-3-642-18363-8;   http:/
/www.springerlink.com/content/978-3-642-18363-8.

Diening:2011:MO

[DHHR11p] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. The maximal operator. Lecture Notes in Mathematics, 2017:99–141, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18363-8_4/.

Diening:2011:OSD

[DHHR11q] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Other spaces of differentiable functions. Lecture Notes in Mathematics, 2017:367–398, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18363-8_12/.

Diening:2011:PFD

[DHHR11r] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. PDEs and fluid dynamics. Lecture Notes in Mathematics, 2017:437–481, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18363-8_14/.

Diening:2011:TT

[DHHR11s] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Transfer techniques. Lecture Notes in Mathematics, 2017:213–244, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18363-8_7/.

Diening:2011:VEL

[DHHR11t] Lars Diening, Petteri Harjulehto, Peter Hästö, and Michael Růžička. Variable exponent Lebesgue spaces. Lecture Notes in Mathematics, 2017:69–97, 2011. CODEN LNMAA2. ISBN 3-642-18362-X (print), 3-642-18363-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18363-8_3/.

Debussche:2013:AET

[DHI13a] Arnaud Debussche, Michael Högele, and Peter Imkeller. Asymptotic exit times. Lecture Notes in Mathematics, 2085:87–120, 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00828-8_5/.

Debussche:2013:ATT

[DHI13b] Arnaud Debussche, Michael Högele, and Peter Imkeller. Asymptotic transition times. Lecture Notes in Mathematics, 2085:121– 130, 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00828-8_6/.

Debussche:2013:BM

[DHI13c] Arnaud Debussche, Michael Högele, and Peter Imkeller. Back matter. Lecture Notes in Mathematics, 2085:151–165, 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/bbm:978-3-
319-00828-8/1.pdf.

Debussche:2013:DNR

[DHI13d] Arnaud Debussche, Michael Högele, and Peter Imkeller. The Dynamics of Nonlinear Reaction-Diffusion Equations with Small Lévy Noise, volume 2085 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 163 pp. LCCN QA3 .L28 no. 2085.

URL http://link.springer.com/book/10.1007/978-3-319-
00828-8;   http://www.springerlink.com/content/978-3-
319-00828-8.

Debussche:2013:FDC

[DHI13e] Arnaud Debussche, Michael Högele, and Peter Imkeller. The fine dynamics of the Chafee–Infante equation. Lecture Notes in Mathematics, 2085:11–43, 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00828-8_2/.

Debussche:2013:FM

[DHI13f] Arnaud Debussche, Michael Högele, and Peter Imkeller. Front matter. Lecture Notes in Mathematics, 2085:i–xiii, 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-00828-8/
1.pdf.

Debussche:2013:I

[DHI13g] Arnaud Debussche, Michael Högele, and Peter Imkeller. Introduction. Lecture Notes in Mathematics, 2085:1–10, 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-00828-8_
1/.

Debussche:2013:LM

[DHI13h] Arnaud Debussche, Michael Högele, and Peter Imkeller. Localization and metastability. Lecture Notes in Mathematics, 2085: 131–149, 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00828-8_7/.

Debussche:2013:SDS

[DHI13i] Arnaud Debussche, Michael Högele, and Peter Imkeller. The small deviation of the small noise solution. Lecture Notes in Mathematics, 2085:69–85, 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00828-8_4/.

Debussche:2013:SCI

[DHI13j] Arnaud Debussche, Michael Högele, and Peter Imkeller. The stochastic Chafee–Infante equation. Lecture Notes in Mathematics, 2085:45–68, 2013. CODEN LNMAA2. ISBN 3-319-00827-7 (print), 3-319-00828-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00828-8_3/.

DiRocco:2014:LTF

[Di 14] Sandra Di Rocco. Linear toric fibrations. Lecture Notes in Mathematics, 2108:119–147, 2014. CODEN LNMAA2. ISBN 3-319-04869-4 (print), 3-319-04870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-04870-3_4.pdf.

Diethelm:2010:AFD

[Die10a] Kai Diethelm. The Analysis of Fractional Differential Equations: An Application-Oriented Exposition Using Differential Operators of Caputo Type, volume 2004 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 247 pp. LCCN QA3 .L28 no.

2004. URL http://link.springer.com/book/10.1007/978-
3-642-14574-2;   http://www.springerlink.com/content/
978-3-642-14574-2.

Diethelm:2010:BM

[Die10b] Kai Diethelm. Back matter. Lecture Notes in Mathematics, 2004: 187–253, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-14574-2/1.pdf.

Diethelm:2010:CA

[Die10c] Kai Diethelm. Caputo’s approach. Lecture Notes in Mathematics, 2004:49–65, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-14574-2_3.pdf.

Diethelm:2010:EUR

[Die10d] Kai Diethelm. Existence and uniqueness results for Riemann-Liouville fractional differential equations. Lecture Notes in Mathematics, 2004:77–83, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-14574-2_5.pdf.

Diethelm:2010:FMa

[Die10e] Kai Diethelm. Front matter. Lecture Notes in Mathematics, 2004: i–viii, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-14574-2/1.pdf.

Diethelm:2010:FMb

[Die10f] Kai Diethelm. Front matter. Lecture Notes in Mathematics, 2004:1, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-14574-2/1/1.pdf.

Diethelm:2010:FMc

[Die10g] Kai Diethelm. Front matter. Lecture Notes in Mathematics, 2004: 75, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-14574-2/2/1.pdf.

Diethelm:2010:I

[Die10h] Kai Diethelm. Introduction. Lecture Notes in Mathematics, 2004: 3–12, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-14574-2_1.pdf.

Diethelm:2010:MLF

[Die10i] Kai Diethelm. Mittag-Leffler functions. Lecture Notes in Mathematics, 2004:67–73, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-14574-2_4.pdf.

Diethelm:2010:MTC

[Die10j] Kai Diethelm. Multi-term Caputo fractional differential equations. Lecture Notes in Mathematics, 2004:167–186, 2010. CO-DEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14574-2_8.pdf.

Diethelm:2010:RLD

[Die10k] Kai Diethelm. Riemann–Liouville differential and integral operators. Lecture Notes in Mathematics, 2004:13–47, 2010. CO-DEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14574-2_2.pdf.

Diethelm:2010:STCb

[Die10l] Kai Diethelm. Single-term Caputo fractional differential equations: Advanced results for special cases. Lecture Notes in Mathematics, 2004:133–166, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-14574-2_7.pdf.

Diethelm:2010:STCa

[Die10m] Kai Diethelm. Single-term Caputo fractional differential equations: Basic theory and fundamental results. Lecture Notes in Mathematics, 2004:85–132, 2010. CODEN LNMAA2. ISBN 3-642-14573-6 (print), 3-642-14574-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-14574-2_6.pdf.

Donati-Martin:2011:FMa

[DMLR11a] Catherine Donati-Martin, Antoine Lejay, and Alain Rouault. Front matter. Lecture Notes in Mathematics, 2006:i–xi, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/bfm:978-3-
642-15217-7/1.pdf.

Donati-Martin:2011:FMb

[DMLR11b] Catherine Donati-Martin, Antoine Lejay, and Alain Rouault. Front matter. Lecture Notes in Mathematics, 2006:1, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/bfm:978-3-
642-15217-7/1/1.pdf.

Donati-Martin:2011:FMc

[DMLR11c] Catherine Donati-Martin, Antoine Lejay, and Alain Rouault. Front matter. Lecture Notes in Mathematics, 2006:71, 2011.

CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/bfm:978-3-
642-15217-7/2/1.pdf.

Donati-Martin:2011:SPX

[DMLR11d] Catherine Donati-Martin, Antoine Lejay, and Alain Rouault, editors. Séminaire de Probabilités XLIII. (French) [Seminar on Probabilities XLIII], volume 2006 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2006; QA273.A1 S46

2011. URL http://link.springer.com/book/10.1007/978-
3-642-15217-7;   http://www.springerlink.com/content/
978-3-642-15217-7.

Donati-Martin:2012:SPX

[DMLR12] Catherine Donati-Martin, Antoine Lejay, and Alain Rouault, editors. Séminaire de Probabilités XLIV. (French) [Seminar on Probabilities XLIV], volume 2046 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2046; QA273.A1 S44

2010. URL http://link.springer.com/book/10.1007/978-
3-642-27461-9;   http://www.springerlink.com/content/
978-3-642-27461-9.

Donati-Martin:2013:SPX

[DMLR13] Catherine Donati-Martin, Antoine Lejay, and Alain Rouault, editors. Séminaire de Probabilités XLV. (French) [Seminar on Probabilities XLV], volume 2078 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (paperback), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2078; QA3 .L28 no. 2078; QA273.A1 S44 2011. URL http://link.

springer.com/book/10.1007/978-3-319-00321-4;   http:/
/www.springerlink.com/content/978-3-319-00321-4.

Donati-Martin:2014:FM

[DMLR14] Catherine Donati-Martin, Antoine Lejay, and Alain Rouault. Front matter. Lecture Notes in Mathematics, 2123:i–viii, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-11970-0/1.pdf.

Dellacherie:2014:BM

[DMS14a] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Back matter. Lecture Notes in Mathematics, 2118:215–238, 2014. CODEN LNMAA2. ISBN 3-319-10297-4 (print), 3-319-10298-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-10298-6/1.pdf.

Dellacherie:2014:FMb

[DMS14b] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Filtered matrices. Lecture Notes in Mathematics, 2118:119–163, 2014. CODEN LNMAA2. ISBN 3-319-10297-4 (print), 3-319-10298-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10298-6_5.pdf.

Dellacherie:2014:FMa

[DMS14c] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Front matter. Lecture Notes in Mathematics, 2118:i– x, 2014. CODEN LNMAA2. ISBN 3-319-10297-4 (print), 3-319-10298-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-10298-6/1.pdf.

Dellacherie:2014:GUT

[DMS14d] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Graph of ultrametric type matrices. Lecture Notes in Mathematics, 2118:85–117, 2014. CODEN LNMAA2. ISBN 3-319-10297-4 (print), 3-319-10298-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-10298-6_4.pdf.

Dellacherie:2014:HFI

[DMS14e] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Hadamard functions of inverse M-matrices. Lecture Notes in Mathematics, 2118:165–213, 2014. CODEN LNMAA2. ISBN 3-319-10297-4 (print), 3-319-10298-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-10298-6_6.pdf.

Dellacherie:2014:I

[DMS14f] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Introduction. Lecture Notes in Mathematics, 2118:1–3, 2014. CODEN LNMAA2. ISBN 3-319-10297-4 (print), 3-319-10298-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10298-6_1.pdf.

Dellacherie:2014:IMP

[DMS14g] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Inverse M-matrices and potentials. Lecture Notes in Mathematics, 2118:5–55, 2014. CODEN LNMAA2. ISBN 3-319-10297-4 (print), 3-319-10298-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-10298-6_2.pdf.

Dellacherie:2014:IMU

[DMS14h] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Inverse M-Matrices and Ultrametric Matrices, volume 2118 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LN-MAA2. ISBN 3-319-10297-4 (print), 3-319-10298-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 238 (est.) pp.

LCCN ???? URL http://www.springerlink.com/content/
978-3-319-10298-6.

Dellacherie:2014:UM

[DMS14i] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Ultrametric matrices. Lecture Notes in Mathematics, 2118:57–84, 2014. CODEN LNMAA2. ISBN 3-319-10297-4 (print), 3-319-10298-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10298-6_3.pdf.

DenHollander:2012:RMS

[DMZ12] Frank Den Hollander, Stanislav Alekseevič Molčanov, and Ofer Zeitouni. Random media at Saint-Flour. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-32948-9 (paperback). vii + 1 + 564 pp. LCCN

QA273.43. URL http://www.springerlink.com/content/978-

3-642-32948-7. Reprint of lectures originally published in the Lecture Notes in Mathematics volumes 1581 (1994), 1837 (2004) and 1974 (2009).

Dodos:2010:BS

[Dod10a] Pandelis Dodos. The ℓ_{2} Baire sum. Lecture Notes in Mathematics, 1993:37–56, 2010. CODEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12153-1_3.pdf.

Dodos:2010:AS

[Dod10b] Pandelis Dodos. Amalgamated spaces. Lecture Notes in Mathematics, 1993:57–70, 2010. CODEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12153-1_4.pdf.

Dodos:2010:BM

[Dod10c] Pandelis Dodos. Back matter. Lecture Notes in Mathematics, 1993:127–167, 2010. CODEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-12153-1/1.pdf.

Dodos:2010:BSD

[Dod10d] Pandelis Dodos. Banach Spaces and Descriptive Set Theory: Selected Topics, volume 1993 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 161 pp. LCCN QA3 .L28 no. 1993.

URL http://link.springer.com/book/10.1007/978-3-642-
12153-1;   http://www.springerlink.com/content/978-3-
642-12153-1.

Dodos:2010:BC

[Dod10e] Pandelis Dodos. Basic concepts. Lecture Notes in Mathematics, 1993:1–8, 2010. CODEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12153-1_1.pdf.

Dodos:2010:BPC

[Dod10f] Pandelis Dodos. The Bourgain–Pisier construction. Lecture Notes in Mathematics, 1993:89–114, 2010. CODEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12153-1_6.pdf.

Dodos:2010:FM

[Dod10g] Pandelis Dodos. Front matter. Lecture Notes in Mathematics, 1993:i–xi, 2010. CODEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-12153-1/1.pdf.

Dodos:2010:SBS

[Dod10h] Pandelis Dodos. The standard Borel space of all separable Banach spaces. Lecture Notes in Mathematics, 1993:9–35, 2010. CODEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12153-1_2.pdf.

Dodos:2010:SBC

[Dod10i] Pandelis Dodos. Strongly bounded classes of Banach spaces. Lecture Notes in Mathematics, 1993:115–126, 2010. CO-DEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12153-1_7.pdf.

Dodos:2010:ZET

[Dod10j] Pandelis Dodos. Zippin’s embedding theorem. Lecture Notes in Mathematics, 1993:71–88, 2010. CODEN LNMAA2. ISBN 3-642-12152-7 (print), 3-642-12153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12153-1_5.pdf.

Deconinck:2011:CPA

[DP11] Bernard Deconinck and Matthew S. Patterson. Computing with plane algebraic curves and Riemann surfaces: The algorithms of the Maple package “Algcurves”. Lecture Notes in Mathematics, 2013:67–123, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-17413-1_2/.

Dick:2014:DTQ

[DP14] Josef Dick and Friedrich Pillichshammer. Discrepancy theory and quasi-Monte Carlo integration. Lecture Notes in Mathematics, 2107:539–619, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-04696-9_9.pdf.

Dieci:2014:CDC

[DPPS14] Luca Dieci, Alessandra Papini, Alessandro Pugliese, and Alessandro Spadoni. Continuous decompositions and coalescing eigenvalues for matrices depending on parameters. Lecture Notes in Mathematics, 2082:173–264, 2014. CODEN LNMAA2. ISBN 3-319-01299-1 (print), 3-319-01300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01300-8_4/.

Doring:2013:CBP

[DR13] Leif Döring and Matthew I. Roberts. Catalytic branching processes via spine techniques and renewal theory. Lecture Notes in Mathematics, 2078:305–322, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00321-4_12/.

Draisma:2014:NS

[Dra14] Jan Draisma. Noetherianity up to symmetry. Lecture Notes in Mathematics, 2108:33–61, 2014. CODEN LNMAA2. ISBN 3-319-04869-4 (print), 3-319-04870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-04870-3_2.pdf.

Duquesne:2010:FM

[DRiSS10a] Thomas Duquesne, Oleg Reichmann, Ken iti Sato, and Christoph Schwab. Front matter. Lecture Notes in Mathematics, 2001: i–xiv, 2010. CODEN LNMAA2. ISBN 3-642-14006-8 (print), 3-642-14007-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-14007-5/1.pdf.

Duquesne:2010:LMR

[DRiSS10b] Thomas Duquesne, Oleg Reichmann, Ken iti Sato, and Christoph Schwab. Lévy Matters I: Recent Progress in Theory and Applications: Foundations, Trees and Numerical Issues in Finance, volume 2001 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-14006-8 (print), 3-642-14007-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 198 pp. LCCN QA3 .L28 no. 2001; QA274.73 .R43 2010.

URL http://link.springer.com/book/10.1007/978-3-642-
14007-5;   http://www.springerlink.com/content/978-3-
642-14007-5.

DiPillo:2010:NOL

[DS10a] Gianni Di Pillo and Fabio Schoen, editors. Nonlinear Optimization: Lectures given at the C.I.M.E. Summer School held in Cetraro, Italy, July 1-7, 2007, volume 1989 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-11338-9 (print), 3-642-11339-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 1989.

URL http://link.springer.com/book/10.1007/978-3-642-
11339-0;   http://www.springerlink.com/content/978-3-
642-11339-0.

Dinh:2010:DSC

[DS10b] Tien-Cuong Dinh and Nessim Sibony. Dynamics in several complex variables: Endomorphisms of projective spaces and polynomial-like mappings. Lecture Notes in Mathematics, 1998: 165–294, 2010. CODEN LNMAA2. ISBN 3-642-13170-0 (print), 3-642-13171-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-13171-4_4.pdf.

Ditlevsen:2013:ISM

[DS13] Susanne Ditlevsen and Adeline Samson. Introduction to stochastic models in biology. Lecture Notes in Mathematics, 2058:3–35, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-32157-3_1/;   http://link.springer.
com/content/pdf/bfm:978-3-642-32157-3/1.pdf;   http:/

/link.springer.com/content/pdf/bfm:978-3-642-32157-3/
1/1.pdf.

Ducros:2015:CFP

[Duc15a] Antoine Ducros. Cohomological finiteness of proper morphisms in algebraic geometry: a purely transcendental proof, without projective tools. Lecture Notes in Mathematics, 2119:135–140, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11029-5_4.pdf.

Ducros:2015:ECS

[Duc15b] Antoine Ducros. Étale cohomology of schemes and analytic spaces. Lecture Notes in Mathematics, 2119:67–118, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11029-5_2.pdf.

Duffin:2013:MVC

[Duf13] James Duffin. Model validation and control issues in the respiratory system. Lecture Notes in Mathematics, 2064:133–162, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32882-4_8/.

Duquesne:2010:PHM

[Duq10] Thomas Duquesne. Packing and Hausdorff measures of stable trees. Lecture Notes in Mathematics, 2001:93–136, 2010. CODEN LNMAA2. ISBN 3-642-14006-8 (print), 3-642-14007-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14007-5_2.pdf.

Doney:2013:WPS

[DV13] R. A. Doney and S. Vakeroudis. Windings of planar stable processes. Lecture Notes in Mathematics, 2078:277–300, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00321-4_10/.

daVeiga:2016:INA

[dVBSV16] L. Beirão da Veiga, A. Buffa, G. Sangalli, and R. Vázquez. An introduction to the numerical analysis of isogeometric methods. Lecture Notes in Mathematics, 2161:87–154, 2016. CO-DEN LNMAA2. ISBN 3-319-42308-8 (print), 3-319-42309-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42309-8_3.pdf.

Evans:2016:BM

[EH16a] John A. Evans and Thomas J. R. Hughes. Back matter. Lecture Notes in Mathematics, 2161:195–196, 2016. CODEN LNMAA2. ISBN 3-319-42308-8 (print), 3-319-42309-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-319-42309-8/1.pdf.

Evans:2016:IAN

[EH16b] John A. Evans and Thomas J. R. Hughes, editors. IsoGeometric Analysis: a New Paradigm in the Numerical Approximation of PDEs: Cetraro, Italy 2012, volume 2161 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. CODEN LNMAA2. ISBN 3-319-42308-8 (print), 3-319-42309-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 196 (est.) pp. LCCN ???? URL

http://www.springerlink.com/content/978-3-319-42309-8.

Evans:2016:ICD

[EH16c] John A. Evans and Thomas J. R. Hughes. Isogeometric compatible discretizations for viscous incompressible flow. Lecture Notes in Mathematics, 2161:155–193, 2016. CODEN LNMAA2. ISBN 3-319-42308-8 (print), 3-319-42309-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-42309-8_4.pdf.

Eisenbaum:2012:AFA

[Eis12] Nathalie Eisenbaum. Another failure in the analogy between Gaussian and semicircle laws. Lecture Notes in Mathematics, 2046:207–213, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-27461-9_10/.

Eldan:2014:BNL

[EL14] Ronen Eldan and Joseph Lehec. Bounding the norm of a logconcave vector via thin-Shell estimates. Lecture Notes in Mathematics, 2116:107–122, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-09477-9_9.pdf.

Emery:2013:PBS

[Éme13] Michel Émery. A planar Borel set which divides every nonnegligible Borel product. Lecture Notes in Mathematics, 2078: 159–165, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00321-4_5/.

Enciso:2013:MMU

[Enc13] Germán A. Enciso. Multisite mechanisms for ultrasensitivity in signal Transduction. Lecture Notes in Mathematics, 2102:199– 224, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-03080-7_6/.

Estrada:2015:ICN

[Est15] Ernesto Estrada. Introduction to complex networks: Structure and dynamics. Lecture Notes in Mathematics, 2126:93–131, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11322-7_3.pdf.

Etheridge:2011:BM

[Eth11a] Alison Etheridge. Back matter. Lecture Notes in Mathematics, 2012:109–119, 2011. CODEN LNMAA2. ISBN 3-642-16631-8 (paperback), 3-642-16632-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-16632-7/1.pdf.

Etheridge:2011:FM

[Eth11b] Alison Etheridge. Front matter. Lecture Notes in Mathematics, 2012:i–viii, 2011. CODEN LNMAA2. ISBN 3-642-16631-8 (paperback), 3-642-16632-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-16632-7/1.pdf.

Etheridge:2011:I

[Eth11c] Alison Etheridge. Introduction. Lecture Notes in Mathematics, 2012:1–3, 2011. CODEN LNMAA2. ISBN 3-642-16631-8 (paperback), 3-642-16632-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-16632-7_1/.

Etheridge:2011:MTT

[Eth11d] Alison Etheridge. More than two types. Lecture Notes in Mathematics, 2012:53–64, 2011. CODEN LNMAA2. ISBN 3-642-16631-8 (paperback), 3-642-16632-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-16632-7_4/.

Etheridge:2011:MRG

[Eth11e] Alison Etheridge. Mutation and random genetic drift. Lecture Notes in Mathematics, 2012:5–32, 2011. CODEN LNMAA2. ISBN 3-642-16631-8 (paperback), 3-642-16632-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-16632-7_2/.

Etheridge:2011:ODD

[Eth11f] Alison Etheridge. One dimensional diffusions. Lecture Notes in Mathematics, 2012:33–51, 2011. CODEN LNMAA2. ISBN 3-642-16631-8 (paperback), 3-642-16632-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-16632-7_3/.

Etheridge:2011:S

[Eth11g] Alison Etheridge. Selection. Lecture Notes in Mathematics, 2012: 65–87, 2011. CODEN LNMAA2. ISBN 3-642-16631-8 (paperback), 3-642-16632-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-16632-7_5/.

Etheridge:2011:SMM

[Eth11h] Alison Etheridge. Some Mathematical Models from Population Genetics: École d’Été de Probabilités de Saint-Flour XXXIX — 2009, volume 2012 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-16631-8 (paperback), 3-642-16632-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 119 pp. LCCN QA3 .L28 no. 2012; QA3 .L28 no. 2012; QH455 .E84 2011. URL http://link.

springer.com/book/10.1007/978-3-642-16632-7;   http:/
/www.springerlink.com/content/978-3-642-16632-7.

Etheridge:2011:SS

[Eth11i] Alison Etheridge. Spatial structure. Lecture Notes in Mathematics, 2012:89–107, 2011. CODEN LNMAA2. ISBN 3-642-16631-8 (paperback), 3-642-16632-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-16632-7_6/.

Ervedoza:2012:BM

[EZ12a] Sylvain Ervedoza and Enrique Zuazua. Back matter. Lecture Notes in Mathematics, 2048:341–344, 2012. CODEN LNMAA2. ISBN 3-642-27892-2 (print), 3-642-27893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-27893-8/1.pdf.

Ervedoza:2012:WEC

[EZ12b] Sylvain Ervedoza and Enrique Zuazua. The wave equation: Control and numerics. Lecture Notes in Mathematics, 2048:245– 339, 2012. CODEN LNMAA2. ISBN 3-642-27892-2 (print), 3-642-27893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27893-8_5/.

Faifman:2012:FUO

[Fai12] Dmitry Faifman. A family of unitary operators satisfying a Poisson-type summation formula. Lecture Notes in Mathematics, 2050:191–204, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-29849-3_11/.

Fasano:2011:MMM

[Fas11] Antonio Fasano, editor. Mathematical Models in the Manufacturing of Glass: C.I.M.E. Summer School, Montecatini Terme, Italy 2008, volume 2010 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-15966-4 (print), 3-642-15967-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2010. URL http://link.

springer.com/book/10.1007/978-3-642-15967-1;   http:/
/www.springerlink.com/content/978-3-642-15967-1.

Favre:2015:CPS

[Fav15a] Charles Favre. Countability properties of some Berkovich spaces. Lecture Notes in Mathematics, 2119:119–132, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11029-5_3.pdf.

Favre:2015:FM

[Fav15b] Charles Favre. Front matter. Lecture Notes in Mathematics, 2119:133, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-11029-5/2/1.pdf.

Franz:2013:HMI

[FE13] Benjamin Franz and Radek Erban. Hybrid modelling of individual movement and collective behaviour. Lecture Notes in Mathematics, 2071:129–157, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35497-7_5/.

Farina:2011:BM

[FFM11a] Angiolo Farina, Antonio Fasano, and Andro Mikelić. Back matter. Lecture Notes in Mathematics, 2010:225–227, 2011. CODEN LNMAA2. ISBN 3-642-15966-4 (print), 3-642-15967-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-642-15967-1/
1.pdf.

Farina:2011:NIF

[FFM11b] Angiolo Farina, Antonio Fasano, and Andro Mikelić. Nonisothermal flow of molten glass: Mathematical challenges and industrial questions. Lecture Notes in Mathematics, 2010:173– 224, 2011. CODEN LNMAA2. ISBN 3-642-15966-4 (print), 3-642-15967-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15967-1_4/.

Fagnani:2015:SIN

[FFR15] Fabio Fagnani, Sophie M. Fosson, and Chiara Ravazzi. Some introductory notes on random graphs. Lecture Notes in Mathematics, 2141:1–26, 2015. CODEN LNMAA2. ISBN 3-319-16966-1 (print), 3-319-16967-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-16967-5_1.pdf.

Ferrara:2011:SMP

[FFV11] Sergio Ferrara, Rita Fioresi, and V. S. Varadarajan, editors. Supersymmetry in Mathematics and Physics: UCLA Los Angeles, USA 2010, volume 2027 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2027; QC174.17.S9 S982 2011; QA3 .L28 no. 2027. URL http://link.

springer.com/book/10.1007/978-3-642-21744-9;   http:/
/www.springerlink.com/content/978-3-642-21744-9.

Fioresi:2011:CCS

[FG11] R. Fioresi and F. Gavarini. On the construction of Chevalley supergroups. Lecture Notes in Mathematics, 2027:101–123, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21744-9_5/.

Frank:2011:RHT

[FK11a] Martin Frank and Axel Klar. Radiative heat transfer and applications for glass production processes. Lecture Notes in Mathematics, 2010:57–134, 2011. CODEN LNMAA2. ISBN 3-642-15966-4 (print), 3-642-15967-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-15967-1_2/.

Frauendiener:2011:ACR

[FK11b] Jörg Frauendiener and Christian Klein. Algebraic curves and Riemann surfaces in Matlab. Lecture Notes in Mathematics, 2013: 125–162, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-17413-1_3/.

Frauendiener:2011:FM

[FK11c] Jörg Frauendiener and Christian Klein. Front matter. Lecture Notes in Mathematics, 2013:163, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-17413-1/3/1.pdf.

Farina:2011:FM

[FKM^{+}11] Angiolo Farina, Axel Klar, Robert M. M. Mattheij, Andro Mikelić, and Norbert Siedow. Front matter. Lecture Notes in Mathematics, 2010:i–xi, 2011. CODEN LNMAA2. ISBN 3-642-15966-4 (print), 3-642-15967-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-15967-1/1.pdf.

Faifman:2014:ORL

[FKM14] Dmitry Faifman, Bo’az Klartag, and Vitali Milman. On the oscillation rigidity of a Lipschitz function on a high-dimensional flat torus. Lecture Notes in Mathematics, 2116:123–131, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_10.pdf.

Futer:2013:A

[FKP13a] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Applications. Lecture Notes in Mathematics, 2069:139–154, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
642-33302-6_9/.

Futer:2013:BM

[FKP13b] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Back matter. Lecture Notes in Mathematics, 2069:163–170, 2013.

CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/bbm:978-3-
642-33302-6/1.pdf.

Futer:2013:DB

[FKP13c] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Decomposition into 3-balls. Lecture Notes in Mathematics, 2069:17– 33, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33302-6_2/.

Futer:2013:DNP

[FKP13d] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Diagrams without non-prime arcs. Lecture Notes in Mathematics, 2069: 109–118, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33302-6_7/.

Futer:2013:DQ

[FKP13e] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Discussion and questions. Lecture Notes in Mathematics, 2069:155– 161, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33302-6_10/.

Futer:2013:FM

[FKP13f] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Front matter. Lecture Notes in Mathematics, 2069:i–x, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-33302-6/
1.pdf.

Futer:2013:GF

[FKP13g] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Guts and fibers. Lecture Notes in Mathematics, 2069:73–90, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33302-6_

Futer:2013:GSC

[FKP13h] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Guts of Surfaces and the Colored Jones Polynomial, volume 2069 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CO-DEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 170 pp. LCCN QA3 .L28 no. 2069. URL http://link.

springer.com/book/10.1007/978-3-642-33302-6;   http:/
/www.springerlink.com/content/978-3-642-33302-6.

Futer:2013:BEP

[FKP13i] David Futer, Efstratia Kalfagianni, and Jessica Purcell. I-bundles and essential product disks. Lecture Notes in Mathematics, 2069: 53–72, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33302-6_4/.

Futer:2013:IP

[FKP13j] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Ideal polyhedra. Lecture Notes in Mathematics, 2069:35–51, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
642-33302-6_3/.

Futer:2013:I

[FKP13k] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Introduction. Lecture Notes in Mathematics, 2069:1–15, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33302-6_
1/.

Futer:2013:ML

[FKP13l] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Montesinos links. Lecture Notes in Mathematics, 2069:119–138, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33302-6_8/.

Futer:2013:REP

[FKP13m] David Futer, Efstratia Kalfagianni, and Jessica Purcell. Recognizing essential product disks. Lecture Notes in Mathematics, 2069: 91–108, 2013. CODEN LNMAA2. ISBN 3-642-33301-X (print), 3-642-33302-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33302-6_6/.

Flandoli:2011:BM

[Fla11a] Franco Flandoli. Back matter. Lecture Notes in Mathematics, 2015:161–176, 2011. CODEN LNMAA2. ISBN 3-642-18230-5 (print), 3-642-18231-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-18231-0/1.pdf.

Flandoli:2011:DM

[Fla11b] Franco Flandoli. Dyadic models. Lecture Notes in Mathematics, 2015:71–99, 2011. CODEN LNMAA2. ISBN 3-642-18230-5 (print), 3-642-18231-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18231-0_3/.

Flandoli:2011:FM

[Fla11c] Franco Flandoli. Front matter. Lecture Notes in Mathematics, 2015:i–ix, 2011. CODEN LNMAA2. ISBN 3-642-18230-5 (print), 3-642-18231-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-18231-0/1.pdf.

Flandoli:2011:IUB

[Fla11d] Franco Flandoli. Introduction to uniqueness and blow-up. Lecture Notes in Mathematics, 2015:1–16, 2011. CODEN LNMAA2. ISBN 3-642-18230-5 (print), 3-642-18231-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-18231-0_1/.

Flandoli:2011:OMU

[Fla11e] Franco Flandoli. Other models: Uniqueness and singularities. Lecture Notes in Mathematics, 2015:133–159, 2011. CODEN LNMAA2. ISBN 3-642-18230-5 (print), 3-642-18231-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-18231-0_

Flandoli:2011:RPP

[Fla11f] Franco Flandoli, editor. Random Perturbation of PDEs and Fluid Dynamic Models: École d’Été de Probabilités de Saint-Flour XL — 2010, volume 2015 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-18230-5 (print), 3-642-18231-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2015. URL http:

//link.springer.com/book/10.1007/978-3-642-18231-0;
http://www.springerlink.com/content/978-3-642-18231-0.

Flandoli:2011:RAN

[Fla11g] Franco Flandoli. Regularization by additive noise. Lecture Notes in Mathematics, 2015:17–69, 2011. CODEN LNMAA2. ISBN 3-642-18230-5 (print), 3-642-18231-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18231-0_2/.

Flandoli:2011:TE

[Fla11h] Franco Flandoli. Transport equation. Lecture Notes in Mathematics, 2015:101–131, 2011. CODEN LNMAA2. ISBN 3-642-18230-5 (print), 3-642-18231-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18231-0_4/.

Fletcher:2010:SQP

[Fle10] Roger Fletcher. The sequential quadratic programming method. Lecture Notes in Mathematics, 1989:165–214, 2010. CO-DEN LNMAA2. ISBN 3-642-11338-9 (print), 3-642-11339-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11339-0_3.pdf.

Favini:2012:BM

[FM12a] Angelo Favini and Gabriela Marinoschi. Back matter. Lecture Notes in Mathematics, 2049:135–143, 2012. CODEN LNMAA2. ISBN 3-642-28284-9 (print), 3-642-28285-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-28285-0/1.pdf.

Favini:2012:DND

[FM12b] Angelo Favini and Gabriela Marinoschi. Degenerate Nonlinear Diffusion Equations, volume 2049 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-28284-9 (print), 3-642-28285-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxii + 143 pp. LCCN QA3 .L28 no.

2049. URL http://link.springer.com/book/10.1007/978-
3-642-28285-0;   http://www.springerlink.com/content/
978-3-642-28285-0.

Favini:2012:EDD

[FM12c] Angelo Favini and Gabriela Marinoschi. Existence for diffusion degenerate problems. Lecture Notes in Mathematics, 2049:57– 90, 2012. CODEN LNMAA2. ISBN 3-642-28284-9 (print), 3-642-28285-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-28285-0_2/.

Favini:2012:ENP

[FM12d] Angelo Favini and Gabriela Marinoschi. Existence for nonautonomous parabolic–elliptic degenerate diffusion equations. Lecture Notes in Mathematics, 2049:91–108, 2012. CODEN LN-MAA2. ISBN 3-642-28284-9 (print), 3-642-28285-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28285-0_
3/.

Favini:2012:EPE

[FM12e] Angelo Favini and Gabriela Marinoschi. Existence for parabolic– elliptic degenerate diffusion problems. Lecture Notes in Mathematics, 2049:1–56, 2012. CODEN LNMAA2. ISBN 3-642-28284-9 (print), 3-642-28285-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-28285-0_1/.

Favini:2012:FM

[FM12f] Angelo Favini and Gabriela Marinoschi. Front matter. Lecture Notes in Mathematics, 2049:i–xxi, 2012. CODEN LNMAA2. ISBN 3-642-28284-9 (print), 3-642-28285-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-28285-0/1.pdf.

Favini:2012:PIP

[FM12g] Angelo Favini and Gabriela Marinoschi. Parameter identification in a parabolic–elliptic degenerate problem. Lecture Notes in Mathematics, 2049:109–133, 2012. CODEN LNMAA2. ISBN 3-642-28284-9 (print), 3-642-28285-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28285-0_4/.

Florentin:2014:ISI

[FMS14] Dan Florentin, Vitali Milman, and Alexander Segal. Identifying set inclusion by projective positions and mixed volumes. Lecture Notes in Mathematics, 2116:133–145, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_11.pdf.

Farley:2014:ATC

[FO14a] Daniel Scott Farley and Ivonne Johanna Ortiz. Algebraic K-theory of Crystallographic Groups: The Three-Dimensional Splitting Case, volume 2113 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 150 (est.) pp. LCCN ???? URL http://www.

springerlink.com/content/978-3-319-08153-3.

Farley:2014:ACP

[FO14b] Daniel Scott Farley and Ivonne Johanna Ortiz. Arithmetic classification of pairs (L, H). Lecture Notes in Mathematics, 2113:23– 39, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-08153-3_3.pdf.

Farley:2014:BM

[FO14c] Daniel Scott Farley and Ivonne Johanna Ortiz. Back matter. Lecture Notes in Mathematics, 2113:143–150, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-319-08153-3/
1.pdf.

Farley:2014:CRA

[FO14d] Daniel Scott Farley and Ivonne Johanna Ortiz. Cokernels of the relative assembly maps for VC_{∞}. Lecture Notes in Mathematics, 2113:119–136, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-08153-3_9.pdf.

Farley:2014:FM

[FO14e] Daniel Scott Farley and Ivonne Johanna Ortiz. Front matter. Lecture Notes in Mathematics, 2113:i–x, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-319-08153-3/1.pdf.

Farley:2014:FDA

[FO14f] Daniel Scott Farley and Ivonne Johanna Ortiz. Fundamental domains for actions on spaces of planes. Lecture Notes in Mathematics, 2113:99–117, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-08153-3_8.pdf.

Farley:2014:FDM

[FO14g] Daniel Scott Farley and Ivonne Johanna Ortiz. Fundamental domains for the maximal groups. Lecture Notes in Mathematics, 2113:59–79, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-08153-3_6.pdf.

Farley:2014:HGH

[FO14h] Daniel Scott Farley and Ivonne Johanna Ortiz. The homology groups H_{n}^{(}E_{FIN} (); KZ^{−∞}). Lecture Notes in Mathematics, 2113:81–98, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-08153-3_7.pdf.

Farley:2014:I

[FO14i] Daniel Scott Farley and Ivonne Johanna Ortiz. Introduction. Lecture Notes in Mathematics, 2113:1–8, 2014. CO-DEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-08153-3_1.pdf.

Farley:2014:STD

[FO14j] Daniel Scott Farley and Ivonne Johanna Ortiz. The split threedimensional crystallographic groups. Lecture Notes in Mathematics, 2113:41–43, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-08153-3_4.pdf.

Farley:2014:SFL

[FO14k] Daniel Scott Farley and Ivonne Johanna Ortiz. A splitting formula for lower algebraic K-theory. Lecture Notes in Mathematics, 2113:45–57, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-08153-3_5.pdf.

Farley:2014:S

[FO14l] Daniel Scott Farley and Ivonne Johanna Ortiz. Summary. Lecture Notes in Mathematics, 2113:137–141, 2014. CO-DEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-08153-3_10.pdf.

Farley:2014:TDP

[FO14m] Daniel Scott Farley and Ivonne Johanna Ortiz. Three-dimensional point groups. Lecture Notes in Mathematics, 2113:9–21, 2014. CODEN LNMAA2. ISBN 3-319-08152-7 (print), 3-319-08153-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-08153-3_2.pdf.

Franchi:2014:STA

[Fra14] Jacques Franchi. Small time asymptotics for an example of strictly hypoelliptic heat kernel. Lecture Notes in Mathematics, 2123:71–103, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-11970-0_4.pdf.

Frohlich:2012:BM

[Frö12a] Steffen Fröhlich. Back matter. Lecture Notes in Mathematics, 2053:107–117, 2012. CODEN LNMAA2. ISBN 3-642-29845-1 (print), 3-642-29846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-29846-2/1.pdf.

Frohlich:2012:CFN

[Frö12b] Steffen Fröhlich. Coulomb Frames in the Normal Bundle of Surfaces in Euclidean Spaces: Topics from Differential Geometry and Geometric Analysis of Surfaces, volume 2053 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-29845-1 (print), 3-642-29846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 117 pp. LCCN QA3 .L28 no.

2053. URL http://link.springer.com/book/10.1007/978-
3-642-29846-2;   http://www.springerlink.com/content/
978-3-642-29846-2.

Frohlich:2012:ES

[Frö12c] Steffen Fröhlich. Elliptic systems. Lecture Notes in Mathematics, 2053:31–52, 2012. CODEN LNMAA2. ISBN 3-642-29845-1 (print), 3-642-29846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-29846-2_2/.

Frohlich:2012:FM

[Frö12d] Steffen Fröhlich. Front matter. Lecture Notes in Mathematics, 2053:i–xiv, 2012. CODEN LNMAA2. ISBN 3-642-29845-1 (print), 3-642-29846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-29846-2/1.pdf.

Frohlich:2012:NCFa

[Frö12e] Steffen Fröhlich. Normal Coulomb frames in R^{4}. Lecture Notes in Mathematics, 2053:53–73, 2012. CODEN LNMAA2. ISBN 3-642-29845-1 (print), 3-642-29846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29846-2_3/.

Frohlich:2012:NCFb

[Frö12f] Steffen Fröhlich. Normal Coulomb frames in R^{n+2}. Lecture Notes in Mathematics, 2053:75–105, 2012. CODEN LNMAA2. ISBN 3-642-29845-1 (print), 3-642-29846-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29846-2_4/.

Frohlich:2012:SG

[Frö12g] Steffen Fröhlich. Surface geometry. Lecture Notes in Mathematics, 2053:1–29, 2012. CODEN LNMAA2. ISBN 3-642-29845-1 (print), 3-642-29846-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-29846-2_1/.

Florentin:2012:SOP

[FS12] Dan Florentin and Alexander Segal. Stability of order preserving transforms. Lecture Notes in Mathematics, 2050:205– 225, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29849-3_12/.

Fruchard:2013:A

[FS13a] Augustin Fruchard and Reinhard Schäfke. Applications. Lecture Notes in Mathematics, 2066:119–150, 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34035-2_6/.

Fruchard:2013:BM

[FS13b] Augustin Fruchard and Reinhard Schäfke. Back matter. Lecture Notes in Mathematics, 2066:155–161, 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-34035-2/1.pdf.

Fruchard:2013:CAEc

[FS13c] Augustin Fruchard and Reinhard Schäfke. Composite Asymptotic Expansions, volume 2066 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 161 pp. LCCN QA3 .L28 no. 2066.

URL http://link.springer.com/book/10.1007/978-3-642-
34035-2;   http://www.springerlink.com/content/978-3-
642-34035-2.

Fruchard:2013:CAEa

[FS13d] Augustin Fruchard and Reinhard Schäfke. Composite asymptotic expansions: General study. Lecture Notes in Mathematics, 2066:

17–41, 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34035-2_2/.

Fruchard:2013:CAEb

[FS13e] Augustin Fruchard and Reinhard Schäfke. Composite asymptotic expansions: Gevrey theory. Lecture Notes in Mathematics, 2066: 43–61, 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34035-2_3/.

Fruchard:2013:CES

[FS13f] Augustin Fruchard and Reinhard Schäfke. Composite expansions and singularly perturbed differential equations. Lecture Notes in Mathematics, 2066:81–118, 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34035-2_5/.

Fruchard:2013:FIE

[FS13g] Augustin Fruchard and Reinhard Schäfke. Four introductory examples. Lecture Notes in Mathematics, 2066:1–15, 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34035-2_1/.

Fruchard:2013:FM

[FS13h] Augustin Fruchard and Reinhard Schäfke. Front matter. Lecture Notes in Mathematics, 2066:i–x, 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-34035-2/1.pdf.

Fruchard:2013:HR

[FS13i] Augustin Fruchard and Reinhard Schäfke. Historical remarks. Lecture Notes in Mathematics, 2066:151–153, 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-34035-2_

Fruchard:2013:TRS

[FS13j] Augustin Fruchard and Reinhard Schäfke. A theorem of Ramis–Sibuya type. Lecture Notes in Mathematics, 2066:63– 80, 2013. CODEN LNMAA2. ISBN 3-642-34034-2 (print), 3-642-34035-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34035-2_4/.

Figalli:2011:BM

[FV11a] Alessio Figalli and Cédric Villani. Back matter. Lecture Notes in Mathematics, 2028:219–224, 2011. CODEN LNMAA2. ISBN 3-642-21718-4 (print), 3-642-21861-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-21861-3/1.pdf.

Figalli:2011:OTC

[FV11b] Alessio Figalli and Cédric Villani. Optimal transport and curvature. Lecture Notes in Mathematics, 2028:171–217, 2011. CODEN LNMAA2. ISBN 3-642-21718-4 (print), 3-642-21861-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21861-3_
4/.

Friedland:2014:VTT

[FY14] Omer Friedland and Yosef Yomdin. Vitushkin-type theorems. Lecture Notes in Mathematics, 2116:147–157, 2014. CO-DEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_12.pdf.

Galdi:2013:SSN

[Gal13] Giovanni P. Galdi. Steady-state Navier–Stokes problem past a rotating body: Geometric-functional properties and related questions. Lecture Notes in Mathematics, 2073:109–197, 2013. CODEN LNMAA2. ISBN 3-642-36296-6 (print), 3-642-36297-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-36297-2_3/.

Galaz-Garcia:2014:NMS

[GG14] Fernando Galaz-Garcia. A note on maximal symmetry rank, quasipositive curvature, and low dimensional manifolds. Lecture Notes in Mathematics, 2110:45–55, 2014. CODEN LNMAA2. ISBN 3-319-06372-3 (print), 3-319-06373-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-06373-7_3.
pdf.

Gazzola:2010:BM

[GGS10a] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Back matter. Lecture Notes in Mathematics, 1991:393–429, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-12245-3/1.pdf.

Gazzola:2010:DPL

[GGS10b] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Dominance of positivity in linear equations. Lecture Notes in Mathematics, 1991:187–226, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12245-3_6.pdf.

Gazzola:2010:EP

[GGS10c] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Eigenvalue problems. Lecture Notes in Mathematics, 1991:61– 98, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12245-3_3.pdf.

Gazzola:2010:FM

[GGS10d] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Front matter. Lecture Notes in Mathematics, 1991:i–xviii, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-12245-3/1.pdf.

Gazzola:2010:KE

[GGS10e] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Kernel estimates. Lecture Notes in Mathematics, 1991:99–146, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12245-3_4.pdf.

Gazzola:2010:LP

[GGS10f] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Linear problems. Lecture Notes in Mathematics, 1991:27–60, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12245-3_2.pdf.

Gazzola:2010:MHO

[GGS10g] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Models of higher order. Lecture Notes in Mathematics, 1991:1– 25, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12245-3_1.pdf.

Gazzola:2010:PBV

[GGS10h] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Polyharmonic Boundary Value Problems: Positivity Preserving and Nonlinear Higher Order Elliptic Equations in Bounded Domains, volume 1991 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 423 pp. LCCN QA3 .L28 no. 1991. URL http:

//link.springer.com/book/10.1007/978-3-642-12245-3;
http://www.springerlink.com/content/978-3-642-12245-3.

Gazzola:2010:PLO

[GGS10i] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Positivity and lower order perturbations. Lecture Notes in Mathematics, 1991:147–185, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12245-3_5.pdf.

Gazzola:2010:SP

[GGS10j] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Semilinear problems. Lecture Notes in Mathematics, 1991:227– 370, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12245-3_7.pdf.

Gazzola:2010:WSR

[GGS10k] Filippo Gazzola, Hans-Christoph Grunau, and Guido Sweers. Willmore surfaces of revolution. Lecture Notes in Mathematics, 1991:371–392, 2010. CODEN LNMAA2. ISBN 3-642-12244-2 (print), 3-642-12245-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12245-3_8.pdf.

Giacomin:2011:BM

[Gia11a] Giambattista Giacomin. Back matter. Lecture Notes in Mathematics, 2025:113–130, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-21156-0/1.pdf.

Giacomin:2011:CGP

[Gia11b] Giambattista Giacomin. The coarse graining procedure. Lecture Notes in Mathematics, 2025:91–99, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21156-0_7/.

Giacomin:2011:CPS

[Gia11c] Giambattista Giacomin. Critical point shift: The fractional moment method. Lecture Notes in Mathematics, 2025:63– 90, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21156-0_6/.

Giacomin:2011:DCP

[Gia11d] Giambattista Giacomin, editor. Disorder and Critical Phenomena Through Basic Probability Models: École d’Été de Probabilités de Saint-Flour XL — 2010, volume 2025 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2025; QC173.4.O73 G53 2011. URL http://link.

springer.com/book/10.1007/978-3-642-21156-0;   http:/
/www.springerlink.com/content/978-3-642-21156-0.

Giacomin:2011:FM

[Gia11e] Giambattista Giacomin. Front matter. Lecture Notes in Mathematics, 2025:i–xi, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/content/pdf/
bfm:978-3-642-21156-0/1.pdf.

Giacomin:2011:HPS

[Gia11f] Giambattista Giacomin. Homogeneous pinning systems: a class of exactly solved models. Lecture Notes in Mathematics, 2025: 5–27, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21156-0_2/.

Giacomin:2011:I

[Gia11g] Giambattista Giacomin. Introduction. Lecture Notes in Mathematics, 2025:1–4, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21156-0_1/.

Giacomin:2011:IDP

[Gia11h] Giambattista Giacomin. Introduction to disordered pinning models. Lecture Notes in Mathematics, 2025:29–40, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-21156-0_3/.

Giacomin:2011:IDE

[Gia11i] Giambattista Giacomin. Irrelevant disorder estimates. Lecture Notes in Mathematics, 2025:41–50, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21156-0_4/.

Giacomin:2011:PP

[Gia11j] Giambattista Giacomin. Path properties. Lecture Notes in Mathematics, 2025:101–112, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21156-0_8/.

Giacomin:2011:RDE

[Gia11k] Giambattista Giacomin. Relevant disorder estimates: The smoothing phenomenon. Lecture Notes in Mathematics, 2025: 51–61, 2011. CODEN LNMAA2. ISBN 3-642-21155-0 (paperback), 3-642-21156-9. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21156-0_5/.

Grimmett:2012:PTS

[GK12] Geoffrey R. Grimmett and Harry Kesten. Percolation theory at Saint-Flour, volume 2053. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-32508-4 (paperback). ISSN 0075-8434 (print), 1617-9692 (electronic). xxvii + 303 pp. LCCN QA3 .L28

no. 2053. URL http://www.springerlink.com/content/978-

3-642-29846-2. Reprint of a lecture originally published in the Lecture Notes in Mathematics volume 1180 (1986) and corrected reprint of a lecture originally published in the Lecture Notes in Mathematics volume 1665 (1997)”.

Gottlich:2013:MOS

[GK13] Simone Göttlich and Axel Klar. Modeling and optimization of scalar flows on networks. Lecture Notes in Mathematics, 2062: 395–461, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32160-3_8/.

Guivarch:2012:DSE

[GKL12] Yves Guivarc’h, John F. C. Kingman, and François Ledrappier. Dynamical systems and ergodic theory at Saint-Flour. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-25965-0 (paperback). vii + 224 pp.

LCCN QA313 .D963 2012. URL http://www.springerlink.
com/content/978-3-642-23669-3. Reprint of lectures origi-

nally published in the Lecture Notes in Mathematics volumes 539 (1976), 774 (1980) and 1097 (1984).

Gluskin:2012:RVI

[GL12] Efim D. Gluskin and Alexander E. Litvak. A remark on vertex index of the convex bodies. Lecture Notes in Mathematics, 2050: 255–265, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29849-3_14/.

Glass:2012:SQC

[Gla12] Olivier Glass. Some questions of control in fluid mechanics. Lecture Notes in Mathematics, 2048:131–206, 2012. CODEN LN-MAA2. ISBN 3-642-27892-2 (print), 3-642-27893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27893-8_
3/.

Gueant:2011:MFG

[GLL11] Olivier Guéant, Jean-Michel Lasry, and Pierre-Louis Lions. Mean field games and applications. Lecture Notes in Mathematics, 2003: 205–266, 2011. CODEN LNMAA2. ISBN 3-642-14659-7 (print), 3-642-14660-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-14660-2_3/.

Giannopoulos:2014:EIC

[GM14] Apostolos Giannopoulos and Emanuel Milman. M-estimates for isotropic convex bodies and their L_{q}-centroid bodies. Lecture Notes in Mathematics, 2116:159–182, 2014. CO-DEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_13.pdf.

Guasoni:2013:PCT

[GMK13] Paolo Guasoni and Johannes Muhle-Karbe. Portfolio choice with transaction costs: a user’s guide. Lecture Notes in Mathematics, 2081:169–201, 2013. CODEN LNMAA2. ISBN 3-319-00412-3 (print), 3-319-00413-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00413-6_3/.

Groot:2011:MMG

[GML11] J. A. W. M. Groot, Robert M. M. Mattheij, and K. Y. Laevsky. Mathematical modelling of glass forming processes. Lecture Notes in Mathematics, 2010:1–56, 2011. CODEN LNMAA2. ISBN 3-642-15966-4 (print), 3-642-15967-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-15967-1_1/.

Gao:2012:BM

[GOZ12a] Hao Gao, Stanley Osher, and Hongkai Zhao. Back matter. Lecture Notes in Mathematics, 2035:159–160, 2012. CODEN LNMAA2. ISBN 3-642-22989-1 (print), 3-642-22990-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-22990-9/1.pdf.

Gao:2012:QPT

[GOZ12b] Hao Gao, Stanley Osher, and Hongkai Zhao. Quantitative photoacoustic tomography. Lecture Notes in Mathematics, 2035:131– 158, 2012. CODEN LNMAA2. ISBN 3-642-22989-1 (print), 3-642-22990-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22990-9_5/.

Giannopoulos:2012:DNL

[GPV12] Apostolos Giannopoulos, Grigoris Paouris, and Petros Valettas. On the distribution of the ψ_{2}-norm of linear functionals on isotropic convex bodies. Lecture Notes in Mathematics, 2050: 227–253, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29849-3_13/.

Geng:2014:BMR

[GQY14] Xi Geng, Zhongmin Qian, and Danyu Yang. G-Brownian motion as rough paths and differential equations driven by G-Brownian motion. Lecture Notes in Mathematics, 2123:125–193, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_6.pdf.

Grupel:2014:RCL

[Gru14] Uri Grupel. Remarks on the central limit theorem for nonconvex bodies. Lecture Notes in Mathematics, 2116:183–198, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_14.pdf.

Guan:2014:CMI

[Gua14] Pengfei Guan. Curvature measures, isoperimetric type inequalities and fully nonlinear PDEs. Lecture Notes in Mathematics, 2087:47–94, 2014. CODEN LNMAA2. ISBN 3-319-00941-9 (print), 3-319-00942-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00942-1_2/.

Guedj:2012:CMA

[Gue12a]  Vincent Guedj.  Complex Monge–Ampère Equations and

Geodesics in the Space of Kähler Metrics, volume 2038 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 310 pp. LCCN

QA3 .L28 no. 2038. URL http://link.springer.com/book/
10.1007/978-3-642-23669-3;   http://www.springerlink.
com/content/978-3-642-23669-3.

Guedj:2012:I

[Gue12b] Vincent Guedj. Introduction. Lecture Notes in Mathematics, 2038:1–10, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-23669-3_1/;   http://link.springer.
com/content/pdf/bfm:978-3-642-23669-3/1.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-642-23669-3/
1/1.pdf.

Guedj:2013:BM

[Gue13a] Vincent Guedj. Back matter. Lecture Notes in Mathematics, 2086:335–336, 2013. CODEN LNMAA2. ISBN 3-319-00818-8 (print), 3-319-00819-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-00819-6/1.pdf.

Guedj:2013:CKR

[Gue13b] Vincent Guedj. Convergence of the Kähler–Ricci flow on a Kähler–Einstein Fano manifold. Lecture Notes in Mathematics, 2086:299–333, 2013. CODEN LNMAA2. ISBN 3-319-00818-8 (print), 3-319-00819-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00819-6_6/.

Gutierrez:2014:RPG

[Gut14] Cristian E. Gutiérrez. Refraction problems in geometric optics. Lecture Notes in Mathematics, 2087:95–150, 2014. CODEN LN-MAA2. ISBN 3-319-00941-9 (print), 3-319-00942-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-00942-1_
3/.

Gillibert:2011:BM

[GW11a] Pierre Gillibert and Friedrich Wehrung. Back matter. Lecture Notes in Mathematics, 2029:143–158, 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-21774-6/1.pdf.

Gillibert:2011:B

[GW11b] Pierre Gillibert and Friedrich Wehrung. Background. Lecture Notes in Mathematics, 2029:1–34, 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-21774-6_1/.

Gillibert:2011:BAS

[GW11c] Pierre Gillibert and Friedrich Wehrung. Boolean algebras that are scaled with respect to a poset. Lecture Notes in Mathematics, 2029:35–50, 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21774-6_2/.

Gillibert:2011:CLL

[GW11d] Pierre Gillibert and Friedrich Wehrung. The condensate lifting lemma (CLL). Lecture Notes in Mathematics, 2029:51– 79, 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21774-6_3/.

Gillibert:2011:CPC

[GW11e] Pierre Gillibert and Friedrich Wehrung. Congruence-permutable, congruence-preserving extensions of lattices. Lecture Notes in Mathematics, 2029:117–129, 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21774-6_5/.

Gillibert:2011:D

[GW11f] Pierre Gillibert and Friedrich Wehrung. Discussion. Lecture Notes in Mathematics, 2029:139–141, 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21774-6_7/.

Gillibert:2011:ODR

[GW11g] Pierre Gillibert and Friedrich Wehrung. From Objects to Diagrams for Ranges of Functors, volume 2029 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 158 pp. LCCN QA3 .L28 no.

2029. URL http://link.springer.com/book/10.1007/978-
3-642-21774-6;   http://www.springerlink.com/content/
978-3-642-21774-6.

Gillibert:2011:FM

[GW11h] Pierre Gillibert and Friedrich Wehrung. Front matter. Lecture Notes in Mathematics, 2029:i–x, 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-21774-6/1.pdf.

Gillibert:2011:GLC

[GW11i] Pierre Gillibert and Friedrich Wehrung. Getting larders from congruence lattices of first-order structures. Lecture Notes in Mathematics, 2029:81–116, 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21774-6_4/.

Gillibert:2011:LNR

[GW11j] Pierre Gillibert and Friedrich Wehrung. Larders from von Neumann regular rings. Lecture Notes in Mathematics, 2029:131– 138, 2011. CODEN LNMAA2. ISBN 3-642-21773-7 (print), 3-642-21774-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21774-6_6/.

Guedj:2012:DPD

[GZ12] Vincent Guedj and Ahmed Zeriahi. Dirichlet problem in domains of C^{n}. Lecture Notes in Mathematics, 2038:13–32, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23669-3_2/.

Guglielmi:2014:BM

[GZ14a] Nicola Guglielmi and Marino Zennaro. Back matter. Lecture Notes in Mathematics, 2082:315–316, 2014. CODEN LNMAA2. ISBN 3-319-01299-1 (print), 3-319-01300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-319-01300-8/1.pdf.

Guglielmi:2014:SLP

[GZ14b] Nicola Guglielmi and Marino Zennaro. Stability of linear problems: Joint spectral radius of sets of matrices. Lecture Notes in Mathematics, 2082:265–313, 2014. CODEN LNMAA2. ISBN 3-319-01299-1 (print), 3-319-01300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01300-8_5/.

Herrmann:2013:VKM

[HA13] Eva Herrmann and Yusuke Asai. Viral kinetic modeling of chronic hepatitis C and B infection. Lecture Notes in Mathematics, 2102: 251–268, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-03080-7_8/.

Hajri:2012:DAS

[Haj12] Hatem Hajri. Discrete approximations to solution flows of Tanaka’s SDE related to Walsh Brownian motion. Lecture Notes in Mathematics, 2046:167–190, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27461-9_8/.

Hashimoto:2013:ASS

[Has13] Hiroya Hashimoto. Approximation and stability of solutions of SDEs driven by a symmetric α stable process with non-Lipschitz coefficients. Lecture Notes in Mathematics, 2078:181– 199, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00321-4_7/.

Heinrich:2013:AMS

[Hei13] Lothar Heinrich. Asymptotic methods in statistics of random point processes. Lecture Notes in Mathematics, 2068:115–150, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33305-7_4/.

Helbing:2013:DNL

[Hel13] Dirk Helbing. Derivation of non-local macroscopic traffic equations and consistent traffic pressures from microscopic carfollowing models. Lecture Notes in Mathematics, 2062:247– 269, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32160-3_3/.

Herzberg:2013:FM

[Her13a] Frederik Herzberg. Front matter. Lecture Notes in Mathematics, 2067:i–xviii, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-33149-7/1.pdf.

Herzberg:2013:BM

[Her13b] Frederik S. Herzberg. Back matter. Lecture Notes in Mathematics, 2067:95–112, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-33149-7/1.pdf.

Herzberg:2013:EFEa

[Her13c] Frederik S. Herzberg. Excursion to financial economics: a radically elementary approach to the fundamental theorems of asset pricing. Lecture Notes in Mathematics, 2067:45–53, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33149-7_5/.

Herzberg:2013:EFEb

[Her13d] Frederik S. Herzberg. Excursion to financial engineering: Volatility invariance in the Black–Scholes model. Lecture Notes in Mathematics, 2067:55–59, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33149-7_6/.

Herzberg:2013:EMP

[Her13e] Frederik S. Herzberg. Excursion to mathematical physics: a radically elementary definition of Feynman path integrals. Lecture Notes in Mathematics, 2067:71–75, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33149-7_8/.

Herzberg:2013:FR

[Her13f] Frederik S. Herzberg. Final remarks. Lecture Notes in Mathematics, 2067:93, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-33149-7_10.

Herzberg:2013:ICC

[Her13g] Frederik S. Herzberg. Infinitesimal calculus, consistently and accessibly. Lecture Notes in Mathematics, 2067:1–5, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33149-7_
1/.

Herzberg:2013:REG

[Her13h] Frederik S. Herzberg. The radically elementary Girsanov theorem and the diffusion invariance principle. Lecture Notes in Mathematics, 2067:35–44, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-33149-7_4/.

Herzberg:2013:REP

[Her13i] Frederik S. Herzberg. Radically elementary probability theory. Lecture Notes in Mathematics, 2067:7–17, 2013. CODEN LN-MAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33149-7_
2/.

Herzberg:2013:RES

[Her13j] Frederik S. Herzberg. Radically elementary stochastic integrals. Lecture Notes in Mathematics, 2067:19–34, 2013. CODEN LN-MAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33149-7_
3/.

Herzberg:2013:RETa

[Her13k] Frederik S. Herzberg. A radically elementary theory of Itô diffusions and associated partial differential equations. Lecture Notes in Mathematics, 2067:61–70, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33149-7_7/.

Herzberg:2013:RETb

[Her13l] Frederik S. Herzberg. A radically elementary theory of Lévy processes. Lecture Notes in Mathematics, 2067:77–92, 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33149-7_
9/.

Herzberg:2013:SCI

[Her13m] Frederik S. Herzberg. Stochastic Calculus with Infinitesimals, volume 2067 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-33148-3 (print), 3-642-33149-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xix + 112 pp. LCCN QA3 .L28 no. 2067. URL http://link.

springer.com/book/10.1007/978-3-642-33149-7;   http:/
/www.springerlink.com/content/978-3-642-33149-7.

Herzog:2013:SSD

[Her13n] Jürgen Herzog. A survey on Stanley depth. Lecture Notes in Mathematics, 2083:3–45, 2013. CODEN LNMAA2. ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-38742-5_1/;   http://link.springer.
com/content/pdf/bfm:978-3-642-38742-5/1.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-642-38742-5/
1/1.pdf.

Helbing:2013:CAD

[HJ13] Dirk Helbing and Anders Johansson. On the controversy around Daganzo’s requiem for and Aw–Rascle’s resurrection of secondorder traffic flow models. Lecture Notes in Mathematics, 2062: 271–302, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32160-3_4/.

Hencl:2014:BM

[HK14a] Stanislav Hencl and Pekka Koskela. Back matter. Lecture Notes in Mathematics, 2096:139–178, 2014. CODEN LNMAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-319-03173-6/1.pdf.

Hencl:2014:C

[HK14b] Stanislav Hencl and Pekka Koskela. Continuity. Lecture Notes in Mathematics, 2096:17–39, 2014. CODEN LNMAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-03173-6_2/.

Hencl:2014:FC

[HK14c] Stanislav Hencl and Pekka Koskela. Final comments. Lecture Notes in Mathematics, 2096:123–138, 2014. CODEN LNMAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-03173-6_7/.

Hencl:2014:FM

[HK14d] Stanislav Hencl and Pekka Koskela. Front matter. Lecture Notes in Mathematics, 2096:i–xi, 2014. CODEN LNMAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-03173-6/1.pdf.

Hencl:2014:HFD

[HK14e] Stanislav Hencl and Pekka Koskela. Homeomorphisms of finite distortion. Lecture Notes in Mathematics, 2096:81–105, 2014. CODEN LNMAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-03173-6_5/.

Hencl:2014:IPN

[HK14f] Stanislav Hencl and Pekka Koskela. Images and preimages of null sets. Lecture Notes in Mathematics, 2096:63–79, 2014. CODEN LNMAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-03173-6_4/.

Hencl:2014:Ib

[HK14g] Stanislav Hencl and Pekka Koskela. Integrability of J_{f} and 1/J_{f}. Lecture Notes in Mathematics, 2096:107–121, 2014. CODEN LNMAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-03173-6_
6/.

Hencl:2014:Ia

[HK14h] Stanislav Hencl and Pekka Koskela. Introduction. Lecture Notes in Mathematics, 2096:1–15, 2014. CODEN LNMAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-03173-6_1/.

Hencl:2014:LMF

[HK14i] Stanislav Hencl and Pekka Koskela. Lectures on Mappings of Finite Distortion, volume 2096 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 176 pp. LCCN QA3 .L28 no. 2096.

URL http://link.springer.com/book/10.1007/978-3-319-
03173-6;   http://www.springerlink.com/content/978-3-
319-03173-6.

Hencl:2014:OD

[HK14j] Stanislav Hencl and Pekka Koskela. Openness and discreteness. Lecture Notes in Mathematics, 2096:41–61, 2014. CODEN LN-MAA2. ISBN 3-319-03172-4 (print), 3-319-03173-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-03173-6_
3/.

Hong:2012:BM

[HKMR12a] Sungbok Hong, John Kalliongis, Darryl McCullough, and J. Hyam Rubinstein. Back matter. Lecture Notes in Mathematics, 2055:145–155, 2012. CODEN LNMAA2. ISBN 3-642-31563-1 (print), 3-642-31564-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-31564-0/1.pdf.

Hong:2012:DEMa

[HKMR12b] Sungbok Hong, John Kalliongis, Darryl McCullough, and J. Hyam Rubinstein. Diffeomorphisms and embeddings of manifolds. Lecture Notes in Mathematics, 2055:9–17, 2012. CODEN LNMAA2. ISBN 3-642-31563-1 (print), 3-642-31564-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31564-0_
2/.

Hong:2012:DEMb

[HKMR12c] Sungbok Hong, John Kalliongis, Darryl McCullough, and J. Hyam Rubinstein. Diffeomorphisms of Elliptic 3-Manifolds, volume 2055 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-31563-1 (print), 3-642-31564-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 155 pp. LCCN QA3 .L28 no. 2055. URL http:

//link.springer.com/book/10.1007/978-3-642-31564-0;
http://www.springerlink.com/content/978-3-642-31564-0.

Hong:2012:ETMa

[HKMR12d] Sungbok Hong, John Kalliongis, Darryl McCullough, and J. Hyam Rubinstein. Elliptic three-manifolds and the Smale Conjecture. Lecture Notes in Mathematics, 2055:1–7, 2012. CODEN LNMAA2. ISBN 3-642-31563-1 (print), 3-642-31564-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31564-0_
1/.

Hong:2012:ETMb

[HKMR12e] Sungbok Hong, John Kalliongis, Darryl McCullough, and J. Hyam Rubinstein. Elliptic three-manifolds containing onesided Klein bottles. Lecture Notes in Mathematics, 2055:53– 83, 2012. CODEN LNMAA2. ISBN 3-642-31563-1 (print), 3-642-31564-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31564-0_4/.

Hong:2012:FM

[HKMR12f] Sungbok Hong, John Kalliongis, Darryl McCullough, and J. Hyam Rubinstein. Front matter. Lecture Notes in Mathematics, 2055:i–x, 2012. CODEN LNMAA2. ISBN 3-642-31563-1 (print), 3-642-31564-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-31564-0/1.pdf.

Hong:2012:LS

[HKMR12g] Sungbok Hong, John Kalliongis, Darryl McCullough, and J. Hyam Rubinstein. Lens spaces. Lecture Notes in Mathematics, 2055:85–144, 2012. CODEN LNMAA2. ISBN 3-642-31563-1 (print), 3-642-31564-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-31564-0_5/.

Hong:2012:MCP

[HKMR12h] Sungbok Hong, John Kalliongis, Darryl McCullough, and J. Hyam Rubinstein. The method of Cerf and Palais. Lecture Notes in Mathematics, 2055:19–51, 2012. CODEN LN-MAA2. ISBN 3-642-31563-1 (print), 3-642-31564-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31564-0_

Helbing:2013:ORS

[HM13] Dirk Helbing and Amin Mazloumian. Operation regimes and slower-is-faster-effect in the control of traffic intersections. Lecture Notes in Mathematics, 2062:357–394, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32160-3_7/.

Harima:2013:BM

[HMM^{+}13a] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. Back matter. Lecture Notes in Mathematics, 2080:235–252, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-38206-2/1.pdf.

Harima:2013:BTL

[HMM^{+}13b] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. Basics on the theory of local rings. Lecture Notes in Mathematics, 2080:39– 95, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-38206-2_2/.

Harima:2013:CRS

[HMM^{+}13c] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. Cohomology rings and the strong Lefschetz property. Lecture Notes in Mathematics, 2080:189–199, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-38206-2_7/.

Harima:2013:CIS

[HMM^{+}13d] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. Complete intersections with the SLP. Lecture Notes in Mathematics, 2080:141– 156, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-38206-2_4/.

Harima:2013:FM

[HMM^{+}13e] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. Front matter. Lecture Notes in Mathematics, 2080:i–xix, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-38206-2/1.pdf.

Harima:2013:GLE

[HMM^{+}13f] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. A generalization of Lefschetz elements. Lecture Notes in Mathematics, 2080:157– 170, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-38206-2_5/.

Harima:2013:ITL

[HMM^{+}13g] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. Invariant theory and Lefschetz properties. Lecture Notes in Mathematics, 2080: 201–209, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-38206-2_8/.

Harima:2013:LPb

[HMM^{+}13h] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. k-Lefschetz properties. Lecture Notes in Mathematics, 2080:171–188, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-38206-2_
6/.

Harima:2013:LPa

[HMM^{+}13i] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. Lefschetz properties. Lecture Notes in Mathematics, 2080:97–140, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-38206-2_

Harima:2013:LPc

[HMM^{+}13j] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. The Lefschetz Properties, volume 2080 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 250 pp. LCCN QA3 .L28 no. 2080. URL http:

//link.springer.com/book/10.1007/978-3-642-38206-2;
http://www.springerlink.com/content/978-3-642-38206-2.

Harima:2013:PT

[HMM^{+}13k] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. Poset theory. Lecture Notes in Mathematics, 2080:1–38, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-38206-2_1/.

Harima:2013:SLP

[HMM^{+}13l] Tadahito Harima, Toshiaki Maeno, Hideaki Morita, Yasuhide Numata, Akihito Wachi, and Junzo Watanabe. The strong Lefschetz property and the Schur–Weyl duality. Lecture Notes in Mathematics, 2080:211–234, 2013. CODEN LNMAA2. ISBN 3-642-38205-3 (print), 3-642-38206-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-38206-2_9/.

Hobson:2011:SEP

[Hob11] David Hobson. The Skorokhod embedding problem and modelindependent bounds for option prices. Lecture Notes in Mathematics, 2003:267–318, 2011. CODEN LNMAA2. ISBN 3-642-14659-7 (print), 3-642-14660-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-14660-2_4/.

Hillen:2013:TAD

[HP13] Thomas Hillen and Kevin J. Painter. Transport and anisotropic diffusion models for movement in oriented habitats. Lecture Notes in Mathematics, 2071:177–222, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35497-7_7/.

Hirsch:2011:BM

[HPRY11a] Francis Hirsch, Christophe Profeta, Bernard Roynette, and Marc Yor. Back matter. Lecture Notes in Mathematics, 2006:505– 510, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-15217-7/1.pdf.

Hirsch:2011:CSS

[HPRY11b] Francis Hirsch, Christophe Profeta, Bernard Roynette, and Marc Yor. Constructing self-similar martingales via two Skorokhod embeddings. Lecture Notes in Mathematics, 2006:451– 503, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_21/.

Harris:2012:BBM

[HR12] Simon C. Harris and Matthew I. Roberts. Branching Brownian motion: Almost sure growth along scaled paths. Lecture Notes in Mathematics, 2046:375–399, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27461-9_17/.

Hidalgo:2011:FM

[HS11a] Rubén A. Hidalgo and Mika Seppälä. Front matter. Lecture Notes in Mathematics, 2013:211, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-17413-1/4/1.pdf.

Hidalgo:2011:NSU

[HS11b] Rubén A. Hidalgo and Mika Seppälä. Numerical Schottky uniformizations: Myrberg’s opening process. Lecture Notes in Mathematics, 2013:195–209, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-17413-1_6/.

Hebbinghaus:2014:MDA

[HS14a] Nils Hebbinghaus and Anand Srivastav. Multicolor discrepancy of arithmetic structures. Lecture Notes in Mathematics, 2107:319– 424, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-04696-9_5.pdf.

Huh:2014:LG

[HS14b] June Huh and Bernd Sturmfels. Likelihood geometry. Lecture Notes in Mathematics, 2108:63–117, 2014. CODEN LNMAA2. ISBN 3-319-04869-4 (print), 3-319-04870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-04870-3_3.
pdf.

Helbing:2013:SON

[HSL13] Dirk Helbing, Jan Siegmeier, and Stefan Lämmer. Self-organized network flows. Lecture Notes in Mathematics, 2062:335–355, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32160-3_6/.

Hartung:2013:FM

[HT13a] Ferenc Hartung and Janos Turi. Front matter. Lecture Notes in Mathematics, 2064:119, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-32882-4/2/1.pdf.

Hartung:2013:PIR

[HT13b] Ferenc Hartung and Janos Turi. Parameter identification in a respiratory control system model with delay. Lecture Notes in Mathematics, 2064:105–118, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32882-4_6/.

Helbing:2013:TVE

[HTKS13] Dirk Helbing, Martin Treiber, Arne Kesting, and Martin Schönhof. Theoretical vs. empirical classification and prediction of congested traffic states. Lecture Notes in Mathematics, 2062: 303–333, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32160-3_5/.

Hu:2011:ASS

[Hu11a] Bei Hu. Asymptotically self-similar blow-up solutions. Lecture Notes in Mathematics, 2018:85–95, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-18460-4_8/.

Hu:2011:BM

[Hu11b] Bei Hu. Back matter. Lecture Notes in Mathematics, 2018: 127, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-18460-4/1.pdf.

Hu:2011:BR

[Hu11c] Bei Hu. Blow-up rate. Lecture Notes in Mathematics, 2018:65– 83, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-18460-4_7/.

Hu:2011:BTS

[Hu11d] Bei Hu. Blow-up Theories for Semilinear Parabolic Equations, volume 2018 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 125 pp. LCCN QA3 .L28 no. 2018. URL http:

//link.springer.com/book/10.1007/978-3-642-18460-4;
http://www.springerlink.com/content/978-3-642-18460-4.

Hu:2011:FTB

[Hu11e] Bei Hu. Finite time blow-up for evolution equations. Lecture Notes in Mathematics, 2018:33–46, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-18460-4_5/.

Hu:2011:FM

[Hu11f] Bei Hu. Front matter. Lecture Notes in Mathematics, 2018: i–x, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-18460-4/1.pdf.

Hu:2011:I

[Hu11g] Bei Hu. Introduction. Lecture Notes in Mathematics, 2018:1– 5, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-18460-4_1/.

Hu:2011:OSV

[Hu11h] Bei Hu. One space variable case. Lecture Notes in Mathematics, 2018:97–118, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18460-4_9/.

Hu:2011:RET

[Hu11i] Bei Hu. A review of elliptic theories. Lecture Notes in Mathematics, 2018:7–18, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18460-4_2/.

Hu:2011:RFP

[Hu11j] Bei Hu. A review of fixed point theorems. Lecture Notes in Mathematics, 2018:29–31, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18460-4_4/.

Hu:2011:RPT

[Hu11k] Bei Hu. A review of parabolic theories. Lecture Notes in Mathematics, 2018:19–27, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18460-4_3/.

Hu:2011:SSS

[Hu11l] Bei Hu. Steady-state solutions. Lecture Notes in Mathematics, 2018:47–63, 2011. CODEN LNMAA2. ISBN 3-642-18459-6 (print), 3-642-18460-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18460-4_6/.

Hug:2013:RP

[Hug13] Daniel Hug. Random polytopes. Lecture Notes in Mathematics, 2068:205–238, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-33305-7_7/.

Heldt:2013:MMP

[HVM13] Thomas Heldt, George C. Verghese, and Roger G. Mark. Mathematical modeling of physiological systems. Lecture Notes in Mathematics, 2064:21–41, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32882-4_2/.

Howard:2012:BM

[HY12a] Benjamin Howard and Tonghai Yang. Back matter. Lecture Notes in Mathematics, 2041:135–140, 2012. CODEN LNMAA2. ISBN 3-642-23978-1 (print), 3-642-23979-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-23979-3/1.pdf.

Howard:2012:ES

[HY12b] Benjamin Howard and Tonghai Yang. Eisenstein series. Lecture Notes in Mathematics, 2041:43–63, 2012. CODEN LNMAA2. ISBN 3-642-23978-1 (print), 3-642-23979-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-23979-3_4/.

Howard:2012:FM

[HY12c] Benjamin Howard and Tonghai Yang. Front matter. Lecture Notes in Mathematics, 2041:i–viii, 2012. CODEN LNMAA2. ISBN 3-642-23978-1 (print), 3-642-23979-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-23979-3/1.pdf.

Howard:2012:IHZ

[HY12d] Benjamin Howard and Tonghai Yang. Intersections of Hirzebruch– Zagier Divisors and CM Cycles, volume 2041 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-23978-1 (print), 3-642-23979-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 140 pp. LCCN QA3 .L28 no.

2041. URL http://link.springer.com/book/10.1007/978-
3-642-23979-3;   http://www.springerlink.com/content/
978-3-642-23979-3.

Howard:2012:I

[HY12e] Benjamin Howard and Tonghai Yang. Introduction. Lecture Notes in Mathematics, 2041:1–9, 2012. CODEN LNMAA2. ISBN 3-642-23978-1 (print), 3-642-23979-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-23979-3_1/.

Howard:2012:LA

[HY12f] Benjamin Howard and Tonghai Yang. Linear algebra. Lecture Notes in Mathematics, 2041:11–24, 2012. CODEN LNMAA2. ISBN 3-642-23978-1 (print), 3-642-23979-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-23979-3_2/.

Howard:2012:LC

[HY12g] Benjamin Howard and Tonghai Yang. Local calculations. Lecture Notes in Mathematics, 2041:85–133, 2012. CODEN LNMAA2. ISBN 3-642-23978-1 (print), 3-642-23979-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-23979-3_6/.

Howard:2012:MR

[HY12h] Benjamin Howard and Tonghai Yang. The main results. Lecture Notes in Mathematics, 2041:65–84, 2012. CODEN LNMAA2. ISBN 3-642-23978-1 (print), 3-642-23979-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-23979-3_5/.

Howard:2012:MSA

[HY12i] Benjamin Howard and Tonghai Yang. Moduli spaces of Abelian surfaces. Lecture Notes in Mathematics, 2041:25–41, 2012. CODEN LNMAA2. ISBN 3-642-23978-1 (print), 3-642-23979-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23979-3_3/.

Ikeda:2012:MCS

[INS12] Nobuyuki Ikeda, David Nualart, and Daniel W. Stroock. Malliavin Calculus at Saint-Flour. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-25931-6. vii + 346 pp. LCCN QA274.2 .I443 2012. URL http://

www.springerlink.com/content/978-3-642-30898-7. Reprint

of lectures originally published in the Lecture Notes in Mathematics volumes 976 (1983), 1427 (1990) and 1690 (1998).

Sato:2010:FIE

[iS10] Ken iti Sato. Fractional integrals and extensions of selfdecomposability. Lecture Notes in Mathematics, 2001:1–91, 2010. CODEN LNMAA2. ISBN 3-642-14006-8 (print), 3-642-14007-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14007-5_1.pdf.

Imbert:2013:IFN

[IS13] Cyril Imbert and Luis Silvestre. An introduction to fully nonlinear parabolic equations. Lecture Notes in Mathematics, 2086: 7–88, 2013. CODEN LNMAA2. ISBN 3-319-00818-8 (print), 3-319-00819-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00819-6_2/.

Isaev:2011:FM

[Isa11a] Alexander Isaev. Front matter. Lecture Notes in Mathematics, 2020:i–xii, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-19783-3/1.pdf.

Isaev:2011:STHc

[Isa11b] Alexander Isaev. Spherical Tube Hypersurfaces, volume 2020 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CO-DEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 220 pp. LCCN QA3 .L28 no. 2020. URL http://link.

springer.com/book/10.1007/978-3-642-19783-3;   http:/
/www.springerlink.com/content/978-3-642-19783-3.

Isaev:2011:STHa

[Isa11c] Prof. Alexander Isaev. (n − 1, 1)-spherical tube hypersurfaces. Lecture Notes in Mathematics, 2020:97–121, 2011. CODEN LN-MAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-19783-3_
6/.

Isaev:2011:STHb

[Isa11d] Prof. Alexander Isaev. (n − 2, 2)-spherical tube hypersurfaces. Lecture Notes in Mathematics, 2020:123–184, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-19783-3_
7/.

Isaev:2011:BM

[Isa11e] Prof. Alexander Isaev. Back matter. Lecture Notes in Mathematics, 2020:213–220, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-19783-3/1.pdf.

Isaev:2011:FR

[Isa11f] Prof. Alexander Isaev. Further results. Lecture Notes in Mathematics, 2020:195–212, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-19783-3_9/.

Isaev:2011:GMS

[Isa11g] Prof. Alexander Isaev. General methods for solving defining systems. Lecture Notes in Mathematics, 2020:55–82, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-19783-3_

Isaev:2011:ICH

[Isa11h] Prof. Alexander Isaev. Invariants of CR-hypersurfaces. Lecture Notes in Mathematics, 2020:1–33, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-19783-3_1/.

Isaev:2011:NAE

[Isa11i] Prof. Alexander Isaev. Number of affine equivalence classes of (k, n − k)-spherical tube hypersurfaces for k ≤ (n − 2). Lecture Notes in Mathematics, 2020:185–194, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-19783-3_8/.

Isaev:2011:RH

[Isa11j] Prof. Alexander Isaev. Rigid hypersurfaces. Lecture Notes in Mathematics, 2020:35–40, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-19783-3_2/.

Isaev:2011:SPS

[Isa11k] Prof. Alexander Isaev. Strongly pseudoconvex spherical tube hypersurfaces. Lecture Notes in Mathematics, 2020:83–96, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-19783-3_5/.

Isaev:2011:TH

[Isa11l] Prof. Alexander Isaev. Tube hypersurfaces. Lecture Notes in Mathematics, 2020:41–53, 2011. CODEN LNMAA2. ISBN 3-642-19782-5 (print), 3-642-19783-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-19783-3_3/.

Ishii:2013:SIV

[Ish13] Hitoshi Ishii. A short introduction to viscosity solutions and the large time behavior of solutions of Hamilton–Jacobi equations. Lecture Notes in Mathematics, 2074:111–249, 2013. CODEN LNMAA2. ISBN 3-642-36432-2 (print), 3-642-36433-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-36433-4_
3/.

Jacob:2013:LPR

[Jac13a] Emmanuel Jacob. Langevin process reflected on a partially elastic boundary II. Lecture Notes in Mathematics, 2078:245– 275, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00321-4_9/.

Jacobsen:2013:ODH

[Jac13b] Martin Jacobsen. One-dimensional homogeneous diffusions. Lecture Notes in Mathematics, 2058:37–55, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32157-3_2/.

Jakobsen:2011:IFD

[Jak11] Hans Plesner Jakobsen. Indecomposable finite-dimensional representations of a class of Lie algebras and Lie superalgebras. Lecture Notes in Mathematics, 2027:125–138, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-21744-9_6/.

Jeulin:2015:BRF

[Jeu15] Dominique Jeulin. Boolean random functions. Lecture Notes in Mathematics, 2120:143–169, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-10064-7_5.pdf.

Jaye:2014:RMM

[JN14] Benjamin Jaye and Fedor Nazarov. Reflectionless measures and the Mattila–Melnikov–Verdera uniform rectifiability theorem. Lecture Notes in Mathematics, 2116:199–229, 2014. CO-DEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_15.pdf.

Jonsson:2015:DBS

[Jon15] Mattias Jonsson. Dynamics on Berkovich spaces in low dimensions. Lecture Notes in Mathematics, 2119:205–366, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11029-5_6.pdf.

Jorgensen:2016:BM

[JPT16a] Palle Jorgensen, Steen Pedersen, and Feng Tian. Back matter. Lecture Notes in Mathematics, 2160:219–233, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-319-39780-1/
1.pdf.

Jorgensen:2016:CMG

[JPT16b] Palle Jorgensen, Steen Pedersen, and Feng Tian. The case of more general groups. Lecture Notes in Mathematics, 2160:47–66, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-39780-1_3.pdf.

Jorgensen:2016:CDR

[JPT16c] Palle Jorgensen, Steen Pedersen, and Feng Tian. Comparing the different RKHSs H_{F} and H_{K}. Lecture Notes in Mathematics, 2160:171–191, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-39780-1_8.pdf.

Jorgensen:2016:CP

[JPT16d] Palle Jorgensen, Steen Pedersen, and Feng Tian. Convolution products. Lecture Notes in Mathematics, 2160:193–195, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-39780-1_9.pdf.

Jorgensen:2016:E

[JPT16e] Palle Jorgensen, Steen Pedersen, and Feng Tian. Examples. Lecture Notes in Mathematics, 2160:67–92, 2016. CO-DEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-39780-1_4.pdf.

Jorgensen:2016:ECP

[JPT16f] Palle Jorgensen, Steen Pedersen, and Feng Tian. Extensions of continuous positive definite functions. Lecture Notes in Mathematics, 2160:17–46, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-39780-1_2.pdf.

Jorgensen:2016:EPD

[JPT16g] Palle Jorgensen, Steen Pedersen, and Feng Tian, editors. Extensions of Positive Definite Functions: Applications and Their Harmonic Analysis, volume 2160 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 233 (est.) pp. LCCN ???? URL http://www.

springerlink.com/content/978-3-319-39780-1.

Jorgensen:2016:FM

[JPT16h] Palle Jorgensen, Steen Pedersen, and Feng Tian. Front matter. Lecture Notes in Mathematics, 2160:i–xxvi, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-39780-1/
1.pdf.

Jorgensen:2016:GF

[JPT16i] Palle Jorgensen, Steen Pedersen, and Feng Tian. Green’s functions. Lecture Notes in Mathematics, 2160:151–169, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-39780-1_7.pdf.

Jorgensen:2016:I

[JPT16j] Palle Jorgensen, Steen Pedersen, and Feng Tian. Introduction. Lecture Notes in Mathematics, 2160:1–16, 2016. CO-DEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-39780-1_1.pdf.

Jorgensen:2016:MSR

[JPT16k] Palle Jorgensen, Steen Pedersen, and Feng Tian. Models for, and spectral representations of, operator extensions. Lecture Notes in Mathematics, 2160:197–216, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-39780-1_10.pdf.

Jorgensen:2016:OOQ

[JPT16l] Palle Jorgensen, Steen Pedersen, and Feng Tian. Overview and open questions. Lecture Notes in Mathematics, 2160:217– 218, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-319-39780-1_11.pdf.

Jorgensen:2016:STM

[JPT16m] Palle Jorgensen, Steen Pedersen, and Feng Tian. Spectral theory for Mercer operators, and implications for Ext(F). Lecture Notes in Mathematics, 2160:115–150, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-39780-1_6.pdf.

Jorgensen:2016:TVT

[JPT16n] Palle Jorgensen, Steen Pedersen, and Feng Tian. Type I vs. Type II extensions. Lecture Notes in Mathematics, 2160:93–113, 2016. CODEN LNMAA2. ISBN 3-319-39779-6 (print), 3-319-39780-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-39780-1_5.pdf.

Jensen:2015:RIG

[JR15] Eva B. Vedel Jensen and Allan Rasmusson. Rotational integral geometry and local stereology — with a view to image analysis. Lecture Notes in Mathematics, 2120:233–255, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10064-7_8.pdf.

Kawan:2013:BM

[Kaw13a] Christoph Kawan. Back matter. Lecture Notes in Mathematics, 2089:221–272, 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-01288-9/1.pdf.

Kawan:2013:BPC

[Kaw13b] Christoph Kawan. Basic properties of control systems. Lecture Notes in Mathematics, 2089:1–42, 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01288-9_1/.

Kawan:2013:CLE

[Kaw13c] Christoph Kawan. Controllability, Lyapunov exponents, and upper bounds. Lecture Notes in Mathematics, 2089:121–150, 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01288-9_5/.

Kawan:2013:ERL

[Kaw13d] Christoph Kawan. Escape rates and lower bounds. Lecture Notes in Mathematics, 2089:151–175, 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01288-9_6/.

Kawan:2013:E

[Kaw13e] Christoph Kawan. Examples. Lecture Notes in Mathematics, 2089:177–220, 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-01288-9_7/.

Kawan:2013:FM

[Kaw13f] Christoph Kawan. Front matter. Lecture Notes in Mathematics, 2089:i–xxii, 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-01288-9/1.pdf.

Kawan:2013:GE

[Kaw13g] Christoph Kawan. General estimates. Lecture Notes in Mathematics, 2089:107–120, 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01288-9_4/.

Kawan:2013:IIE

[Kaw13h] Christoph Kawan. Introduction to invariance entropy. Lecture Notes in Mathematics, 2089:43–87, 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01288-9_2/.

Kawan:2013:IED

[Kaw13i] Christoph Kawan. Invariance Entropy for Deterministic Control Systems: An Introduction, volume 2089 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxii + 270 pp. LCCN

QA3 .L28 no. 2089. URL http://link.springer.com/book/
10.1007/978-3-319-01288-9;   http://www.springerlink.
com/content/978-3-319-01288-9.

Kawan:2013:LBS

[Kaw13j] Christoph Kawan. Linear and bilinear systems. Lecture Notes in Mathematics, 2089:89–105, 2013. CODEN LNMAA2. ISBN 3-319-01287-8 (print), 3-319-01288-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01288-9_3/.

Kroese:2015:SPS

[KB15] Dirk P. Kroese and Zdravko I. Botev. Spatial process simulation. Lecture Notes in Mathematics, 2120:369–404, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10064-7_12.pdf.

Kennard:2014:HCS

[Ken14] Lee Kennard. On the Hopf conjecture with symmetry. Lecture Notes in Mathematics, 2110:111–116, 2014. CODEN LNMAA2. ISBN 3-319-06372-3 (print), 3-319-06373-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-06373-7_5.
pdf.

Kendall:2015:BM

[Ken15a] Wilfrid S. Kendall. Back matter. Lecture Notes in Mathematics, 2120:441–466, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-10064-7/1.pdf.

Kendall:2015:ICP

[Ken15b] Wilfrid S. Kendall. Introduction to coupling-from-the-past using R. Lecture Notes in Mathematics, 2120:405–439, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10064-7_13.pdf.

Kendall:2015:SGS

[Ken15c] Wilfrid S. Kendall. Stochastic Geometry, Spatial Statistics and Random Fields: Models and Algorithms, volume 2120 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LN-MAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 466 (est.) pp.

LCCN ???? URL http://www.springerlink.com/content/
978-3-319-10064-7.

Khapalov:2010:DWR

[Kha10a] Alexander Y. Khapalov. The 1-D wave and rod equations governed by controls that are time-dependent only. Lecture Notes in Mathematics, 1995:147–156, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12413-6_9.pdf.

Khapalov:2010:BM

[Kha10b] Alexander Y. Khapalov. Back matter. Lecture Notes in Mathematics, 1995:275–290, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-12413-6/1.pdf.

Khapalov:2010:BDS

[Kha10c] Alexander Y. Khapalov. A “basic” 2-D swimming model. Lecture Notes in Mathematics, 1995:165–170, 2010. CO-DEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12413-6_11.pdf.

Khapalov:2010:CRD

[Kha10d] Alexander Y. Khapalov. The case of the reaction-diffusion term satisfying Newton’s law. Lecture Notes in Mathematics, 1995:49– 65, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12413-6_4.pdf.

Khapalov:2010:CCS

[Kha10e] Alexander Y. Khapalov. Classical controllability for the semilinear parabolic equations with superlinear terms. Lecture Notes in Mathematics, 1995:67–80, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12413-6_5.pdf.

Khapalov:2010:CPD

[Kha10f] Alexander Y. Khapalov. Controllability of Partial Differential Equations Governed by Multiplicative Controls, volume 1995 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CO-DEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 284 pp. LCCN QA3 .L28 no. 1995. URL http://link.

springer.com/book/10.1007/978-3-642-12413-6;   http:/
/www.springerlink.com/content/978-3-642-12413-6.

Khapalov:2010:CPVa

[Kha10g] Alexander Y. Khapalov. Controllability properties of a vibrating string with variable axial load and damping gain. Lecture Notes in Mathematics, 1995:83–104, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12413-6_6.pdf.

Khapalov:2010:CPVb

[Kha10h] Alexander Y. Khapalov. Controllability properties of a vibrating string with variable axial load only. Lecture Notes in Mathematics, 1995:105–119, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12413-6_7.pdf.

Khapalov:2010:FMa

[Kha10i] Alexander Y. Khapalov. Front matter. Lecture Notes in Mathematics, 1995:i–xv, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-12413-6/1.pdf.

Khapalov:2010:FMb

[Kha10j] Alexander Y. Khapalov. Front matter. Lecture Notes in Mathematics, 1995:14, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-12413-6/1/1.pdf.

Khapalov:2010:FMc

[Kha10k] Alexander Y. Khapalov. Front matter. Lecture Notes in Mathematics, 1995:82, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-12413-6/2/1.pdf.

Khapalov:2010:FMd

[Kha10l] Alexander Y. Khapalov. Front matter. Lecture Notes in Mathematics, 1995:158, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-12413-6/3/1.pdf.

Khapalov:2010:FMe

[Kha10m] Alexander Y. Khapalov. Front matter. Lecture Notes in Mathematics, 1995:264, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-12413-6/4/1.pdf.

Khapalov:2010:GAC

[Kha10n] Alexander Y. Khapalov. Geometric aspects of controllability for a swimming phenomenon. Lecture Notes in Mathematics, 1995:195–217, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12413-6_13.pdf.

Khapalov:2010:GCR

[Kha10o] Alexander Y. Khapalov. Global controllability for a “rowing” swimming model. Lecture Notes in Mathematics, 1995:237–262, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12413-6_15.pdf.

Khapalov:2010:GNC

[Kha10p] Alexander Y. Khapalov. Global nonnegative controllability of the 1 − D semilinear parabolic equation. Lecture Notes in Mathematics, 1995:15–31, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12413-6_2.pdf.

Khapalov:2010:Ia

[Kha10q] Alexander Y. Khapalov. Introduction. Lecture Notes in Mathematics, 1995:1–12, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12413-6_1.pdf.

Khapalov:2010:Ib

[Kha10r] Alexander Y. Khapalov. Introduction. Lecture Notes in Mathematics, 1995:159–164, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12413-6_10.pdf.

Khapalov:2010:LCS

[Kha10s] Alexander Y. Khapalov. Local controllability for a swimming model. Lecture Notes in Mathematics, 1995:219–236, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12413-6_14.pdf.

Khapalov:2010:MCSb

[Kha10t] Alexander Y. Khapalov. Multiplicative controllability for the Schrödinger equation. Lecture Notes in Mathematics, 1995:265– 274, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12413-6_16.pdf.

Khapalov:2010:MCSa

[Kha10u] Alexander Y. Khapalov. Multiplicative controllability of the semilinear parabolic equation: a qualitative approach. Lecture Notes in Mathematics, 1995:33–48, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12413-6_3.pdf.

Khapalov:2010:RNE

[Kha10v] Alexander Y. Khapalov. Reachability of nonnegative equilibrium states for the semilinear vibrating string. Lecture Notes in Mathematics, 1995:121–145, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12413-6_8.pdf.

Khapalov:2010:WPD

[Kha10w] Alexander Y. Khapalov. The well-posedness of a 2-D swimming model. Lecture Notes in Mathematics, 1995:171–193, 2010. CODEN LNMAA2. ISBN 3-642-12412-7 (print), 3-642-12413-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12413-6_12.pdf.

Kiderlen:2013:IIG

[Kid13] Markus Kiderlen. Introduction into integral geometry and stereology. Lecture Notes in Mathematics, 2068:21–48, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33305-7_
2/.

Knebusch:2014:AT

[KK14a] Manfred Knebusch and Tobias Kaiser. Approximation theorems. Lecture Notes in Mathematics, 2103:59–121, 2014. CODEN LN-MAA2. ISBN 3-319-03211-9 (print), 3-319-03212-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-03212-2_
2/.

Knebusch:2014:BM

[KK14b] Manfred Knebusch and Tobias Kaiser. Back matter. Lecture Notes in Mathematics, 2103:179–192, 2014. CODEN LNMAA2. ISBN 3-319-03211-9 (print), 3-319-03212-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-319-03212-2/1.pdf.

Knebusch:2014:FM

[KK14c] Manfred Knebusch and Tobias Kaiser. Front matter. Lecture Notes in Mathematics, 2103:i–xii, 2014. CODEN LNMAA2. ISBN 3-319-03211-9 (print), 3-319-03212-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-03212-2/1.pdf.

Knebusch:2014:KES

[KK14d] Manfred Knebusch and Tobias Kaiser. Kronecker extensions and star operations. Lecture Notes in Mathematics, 2103:123– 178, 2014. CODEN LNMAA2. ISBN 3-319-03211-9 (print), 3-319-03212-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-03212-2_3/.

Knebusch:2014:MVP

[KK14e] Manfred Knebusch and Tobias Kaiser. Manis Valuations and Prüfer Extensions II, volume 2103 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-03211-9 (print), 3-319-03212-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 190 pp. LCCN QA3 .L28 no.

2103. URL http://link.springer.com/book/10.1007/978-
3-319-03212-2;   http://www.springerlink.com/content/
978-3-319-03212-2.

Knebusch:2014:OPS

[KK14f] Manfred Knebusch and Tobias Kaiser. Overrings and PMspectra. Lecture Notes in Mathematics, 2103:1–57, 2014. CODEN LNMAA2. ISBN 3-319-03211-9 (print), 3-319-03212-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-03212-2_
1/.

Kobayashi:2016:AGP

[KKP16a] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Appendix: Gegenbauer polynomials. Lecture Notes in Mathematics, 2170:173–184, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-981-10-2657-7_14.pdf.

Kobayashi:2016:AFD

[KKP16b] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Application of finite-dimensional representation theory. Lecture Notes in Mathematics, 2170:51–65, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-981-10-2657-7_5.
pdf.

Kobayashi:2016:BM

[KKP16c] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Back matter. Lecture Notes in Mathematics, 2170:185–192, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/bbm:978-981-
10-2657-7/1.pdf.

Kobayashi:2016:BOD

[KKP16d] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Basic operators in differential geometry and conformal covariance. Lecture Notes in Mathematics, 2170:93–109, 2016. CO-DEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
981-10-2657-7_8.pdf.

Kobayashi:2016:CSB

[KKP16e] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner, editors. Conformal Symmetry Breaking Operators for Differential Forms on Spheres, volume 2170 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 192 (est.) pp. LCCN ???? URL http://

www.springerlink.com/content/978-981-10-2657-7.

Kobayashi:2016:CDS

[KKP16f] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Construction of differential symmetry breaking operators. Lecture Notes in Mathematics, 2170:121–129, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/content/pdf/10.1007/978-981-
10-2657-7_10.pdf.

Kobayashi:2016:MMV

[KKP16g] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. F - method for matrix-valued differential operators. Lecture Notes in Mathematics, 2170:31–39, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-981-10-2657-7_3.pdf.

Kobayashi:2016:SSBb

[KKP16h] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. F -system for symmetry breaking operators (j = i − −2, i + 1case). Lecture Notes in Mathematics, 2170:87–91, 2016. CO-DEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
981-10-2657-7_7.pdf.

Kobayashi:2016:SSBa

[KKP16i] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. F - system for symmetry breaking operators (j = i−1, icase). Lecture Notes in Mathematics, 2170:67–85, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-981-10-2657-7_6.
pdf.

Kobayashi:2016:FM

[KKP16j] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Front matter. Lecture Notes in Mathematics, 2170:i–ix, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/bfm:978-981-
10-2657-7/1.pdf.

Kobayashi:2016:ISV

[KKP16k] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Identities of scalar-valued differential operators d_{l}^{μ}. Lecture Notes in Mathematics, 2170:111–119, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-981-10-2657-7_9.pdf.

Kobayashi:2016:IO

[KKP16l] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Intertwining operators. Lecture Notes in Mathematics, 2170:141– 153, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-981-10-2657-7_12.pdf.

Kobayashi:2016:I

[KKP16m] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Introduction. Lecture Notes in Mathematics, 2170:1–12, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
981-10-2657-7_1.pdf.

Kobayashi:2016:MVFa

[KKP16n] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Matrix-valued F -method for O(n+1, 1). Lecture Notes in Mathematics, 2170:41–49, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-981-10-2657-7_4.pdf.

Kobayashi:2016:MVFb

[KKP16o] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Matrix-valued factorization identities. Lecture Notes in Mathematics, 2170:155–172, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-981-10-2657-7_13.pdf.

Kobayashi:2016:SPB

[KKP16p] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Solutions to problems A and B for (S^{n}, S^{n−1}). Lecture Notes in Mathematics, 2170:131–139, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-981-10-2657-7_11.pdf.

Kobayashi:2016:SBO

[KKP16q] Toshiyuki Kobayashi, Toshihisa Kubo, and Michael Pevzner. Symmetry breaking operators and principal series representations of G = O(n + 1, 1). Lecture Notes in Mathematics, 2170:13–30, 2016. CODEN LNMAA2. ISBN 981-10-2656-4 (print), 981-10-2657-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
981-10-2657-7_2.pdf.

Kuznetsov:2012:BM

[KKR12a] Alexey Kuznetsov, Andreas E. Kyprianou, and Victor Rivero. Back matter. Lecture Notes in Mathematics, 2061:187–188, 2012. CODEN LNMAA2. ISBN 3-642-31406-6 (print), 3-642-31407-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-31407-0/1.pdf.

Kuznetsov:2012:TSF

[KKR12b] Alexey Kuznetsov, Andreas E. Kyprianou, and Victor Rivero. The theory of scale functions for spectrally negative Lévy processes. Lecture Notes in Mathematics, 2061:97–186, 2012. CO-DEN LNMAA2. ISBN 3-642-31406-6 (print), 3-642-31407-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
642-31407-0_2/.

Klartag:2014:LCM

[Kla14] Bo’az Klartag. Logarithmically-concave moment measures I. Lecture Notes in Mathematics, 2116:231–260, 2014. CO-DEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_16.pdf.

Kliemann:2014:FM

[Kli14a] Lasse Kliemann. Front matter. Lecture Notes in Mathematics, 2107:485, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-04696-9/3/1.pdf.

Kliemann:2014:PAL

[Kli14b] Lasse Kliemann. Practical algorithms for low-discrepancy 2colorings. Lecture Notes in Mathematics, 2107:459–484, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-04696-9_7.pdf.

Klartag:2012:IRL

[KM12a] Bo’az Klartag and Emanuel Milman. Inner regularization of logconcave measures and small-ball estimates. Lecture Notes in Mathematics, 2050:267–278, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29849-3_15/.

Konig:2012:OEG

[KM12b] Hermann König and Vitali Milman. An operator equation generalizing the Leibniz rule for the second derivative. Lecture Notes in Mathematics, 2050:279–299, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29849-3_16/.

Klartag:2014:FM

[KM14a] Bo’az Klartag and Emanuel Milman. Front matter. Lecture Notes in Mathematics, 2116:i–ix, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-09477-9/1.pdf.

Kolesnikov:2014:RKC

[KM14b] Alexander V. Kolesnikov and Emanuel Milman. Remarks on the KLS conjecture and Hardy-type inequalities. Lecture Notes in Mathematics, 2116:273–292, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-09477-9_18.pdf.

Klartag:2012:GAF

[KMM12] Bo’az Klartag, Shahar Mendelson, and Vitali D. Milman, editors. Geometric Aspects of Functional Analysis: Israel Seminar 2006–2010, volume 2050 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (paperback), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2050. URL http:

//link.springer.com/book/10.1007/978-3-642-29849-3;
http://www.springerlink.com/content/978-3-642-29849-3.

Kunze:2013:BM

[KO13a] Markus Kunze and Rafael Ortega. Back matter. Lecture Notes in Mathematics, 2065:301–303, 2013. CODEN LNMAA2. ISBN 3-642-32905-5 (print), 3-642-32906-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-32906-7/1.pdf.

Kunze:2013:TMN

[KO13b] Markus Kunze and Rafael Ortega. Twist mappings with nonperiodic angles. Lecture Notes in Mathematics, 2065:265–300, 2013. CODEN LNMAA2. ISBN 3-642-32905-5 (print), 3-642-32906-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32906-7_5/.

Koenigsmann:2014:BM

[Koe14a] Jochen Koenigsmann. Back matter. Lecture Notes in Mathematics, 2111:197–198, 2014. CODEN LNMAA2. ISBN 3-642-54935-7 (print), 3-642-54936-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-54936-6/1.pdf.

Koenigsmann:2014:MTA

[Koe14b] Jochen Koenigsmann, editor. Model Theory in Algebra, Analysis and Arithmetic: Cetraro, Italy 2012, Editors: H. Dugald Macpherson, Carlo Toffalori, volume 2111 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-642-54935-7 (print), 3-642-54936-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-54936-6.

Koenigsmann:2014:UNT

[Koe14c] Jochen Koenigsmann. Undecidability in number theory. Lecture Notes in Mathematics, 2111:159–195, 2014. CODEN LNMAA2. ISBN 3-642-54935-7 (print), 3-642-54936-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-54936-6_5.
pdf.

Kokotov:2011:BM

[Kok11a] Alexey Kokotov. Back matter. Lecture Notes in Mathematics, 2013:255–257, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-17413-1/1.pdf.

Kokotov:2011:STL

[Kok11b] Alexey Kokotov. On the spectral theory of the Laplacian on compact polyhedral surfaces of arbitrary genus. Lecture Notes in Mathematics, 2013:227–253, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-17413-1_8/.

Koltchinskii:2011:BM

[Kol11a] Prof. Vladimir Koltchinskii. Back matter. Lecture Notes in Mathematics, 2033:235–254, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-22147-7/1.pdf.

Koltchinskii:2011:BES

[Kol11b] Prof. Vladimir Koltchinskii. Bounding expected sup-norms of empirical and Rademacher processes. Lecture Notes in Mathematics, 2033:33–57, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-22147-7_3/.

Koltchinskii:2011:CPS

[Kol11c] Prof. Vladimir Koltchinskii. Convex penalization in sparse recovery. Lecture Notes in Mathematics, 2033:151–189, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22147-7_
8/.

Koltchinskii:2011:ERP

[Kol11d] Prof. Vladimir Koltchinskii. Empirical and Rademacher processes. Lecture Notes in Mathematics, 2033:17–32, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22147-7_
2/.

Koltchinskii:2011:EER

[Kol11e] Prof. Vladimir Koltchinskii. Examples of excess risk bounds in prediction problems. Lecture Notes in Mathematics, 2033:81– 97, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22147-7_5/.

Koltchinskii:2011:ERB

[Kol11f] Prof. Vladimir Koltchinskii. Excess risk bounds. Lecture Notes in Mathematics, 2033:59–79, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-22147-7_4/.

Koltchinskii:2011:I

[Kol11g] Prof. Vladimir Koltchinskii. Introduction. Lecture Notes in Mathematics, 2033:1–16, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-22147-7_1/.

Koltchinskii:2011:LPS

[Kol11h] Prof. Vladimir Koltchinskii. Linear programming in sparse recovery. Lecture Notes in Mathematics, 2033:121–149, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22147-7_
7/.

Koltchinskii:2011:LRM

[Kol11i] Prof. Vladimir Koltchinskii. Low rank matrix recovery: Nuclear norm penalization. Lecture Notes in Mathematics, 2033: 191–234, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22147-7_9/.

Koltchinskii:2011:PER

[Kol11j] Prof. Vladimir Koltchinskii. Penalized empirical risk minimization and model selection problems. Lecture Notes in Mathematics, 2033:99–119, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-22147-7_6/.

Koltchinskii:2011:FM

[Kol11k] Vladimir Koltchinskii. Front matter. Lecture Notes in Mathematics, 2033:i–ix, 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-22147-7/1.pdf.

Koltchinskii:2011:OIE

[Kol11l] Vladimir Koltchinskii. Oracle Inequalities in Empirical Risk Minimization and Sparse Recovery Problems: École d’Été de Probabilités de Saint-Flour XXXVIII — 2008, volume 2033 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-22146-7 (print), 3-642-22147-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 254 pp. LCCN QA3 .L28 no. 2033;

QA278.2 .K65 2011.  URL http://link.springer.com/
book/10.1007/978-3-642-22147-7;   http://www.loc.gov/
catdir/enhancements/fy1206/2011934366-d.html;   http:
//www.springerlink.com/content/978-3-642-22147-7.

Kolev:2012:RSK

[Kol12] Boris Kolev. The Riemannian space of Kähler metrics. Lecture Notes in Mathematics, 2038:231–255, 2012. CODEN LNMAA2. ISBN 3-642-23668-5 (print), 3-642-23669-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-23669-3_6/.

Koldobsky:2014:EMS

[Kol14] Alexander Koldobsky. Estimates for measures of sections of convex bodies. Lecture Notes in Mathematics, 2116:261–271, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09477-9_17.pdf.

Kortchemski:2013:BM

[Kor13a] Igor Kortchemski. Back matter. Lecture Notes in Mathematics, 2078:559–560, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-00321-4/1.pdf.

Kortchemski:2013:SPD

[Kor13b] Igor Kortchemski. A simple proof of Duquesne’s theorem on contour processes of conditioned Galton–Watson trees. Lecture Notes in Mathematics, 2078:537–558, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00321-4_20/.

Kloeden:2013:FMa

[KP13a] Peter E. Kloeden and Christian Pötzsche. Front matter. Lecture Notes in Mathematics, 2102:i–xviii, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-319-03080-7/1.pdf.

Kloeden:2013:FMb

[KP13b] Peter E. Kloeden and Christian Pötzsche. Front matter. Lecture Notes in Mathematics, 2102:1, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-03080-7/1/1.pdf.

Kloeden:2013:NDSa

[KP13c] Peter E. Kloeden and Christian Pötzsche. Nonautonomous dynamical systems in the life sciences. Lecture Notes in Mathematics, 2102:3–39, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-03080-7_1/;   http://link.springer.
com/content/pdf/bfm:978-3-319-03080-7/1.pdf;   http:/
/link.springer.com/content/pdf/bfm:978-3-319-03080-7/
1/1.pdf.

Kloeden:2013:NDSb

[KP13d] Peter E. Kloeden and Christian Pötzsche. Nonautonomous Dynamical Systems in the Life Sciences, volume 2102 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 312 pp. LCCN QA3 .L28 no. 2102; QA3 .L28 no. 2102; QH324.2. URL http:

//link.springer.com/book/10.1007/978-3-319-03080-7;
http://www.springerlink.com/content/978-3-319-03080-7.

Kloeden:2013:DTN

[KPR13] P. E. Kloeden, C. Pötzsche, and M. Rasmussen. Discrete-time nonautonomous dynamical systems. Lecture Notes in Mathematics, 2065:35–102, 2013. CODEN LNMAA2. ISBN 3-642-32905-5 (print), 3-642-32906-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-32906-7_2/.

Kyprianou:2014:BDS

[KPR14] A. E. Kyprianou, J-L. Pérez, and Y.-X. Ren. The backbone decomposition for spatially dependent supercritical superprocesses. Lecture Notes in Mathematics, 2123:33–59, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_2.pdf.

Kruse:2014:BM

[Kru14a] Raphael Kruse. Back matter. Lecture Notes in Mathematics, 2093:155–180, 2014. CODEN LNMAA2. ISBN 3-319-02230-X (print), 3-319-02231-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-02231-4/1.pdf.

Kruse:2014:FM

[Kru14b] Raphael Kruse. Front matter. Lecture Notes in Mathematics, 2093:i–xiv, 2014. CODEN LNMAA2. ISBN 3-319-02230-X (print), 3-319-02231-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-02231-4/1.pdf.

Kruse:2014:I

[Kru14c] Raphael Kruse. Introduction. Lecture Notes in Mathematics, 2093:1–10, 2014. CODEN LNMAA2. ISBN 3-319-02230-X (print), 3-319-02231-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02231-4_1/.

Kruse:2014:MCA

[Kru14d] Raphael Kruse. A Malliavin calculus approach to weak convergence. Lecture Notes in Mathematics, 2093:109–127, 2014. CODEN LNMAA2. ISBN 3-319-02230-X (print), 3-319-02231-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02231-4_5/.

Kruse:2014:NE

[Kru14e] Raphael Kruse. Numerical experiments. Lecture Notes in Mathematics, 2093:129–153, 2014. CODEN LNMAA2. ISBN 3-319-02230-X (print), 3-319-02231-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02231-4_6/.

Kruse:2014:OSE

[Kru14f] Raphael Kruse. Optimal strong error estimates for Galerkin finite element methods. Lecture Notes in Mathematics, 2093:51– 84, 2014. CODEN LNMAA2. ISBN 3-319-02230-X (print), 3-319-02231-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02231-4_3/.

Kruse:2014:SRM

[Kru14g] Raphael Kruse. A short review of the Malliavin calculus in Hilbert spaces. Lecture Notes in Mathematics, 2093:85–108, 2014. CODEN LNMAA2. ISBN 3-319-02230-X (print), 3-319-02231-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02231-4_4/.

Kruse:2014:SEE

[Kru14h] Raphael Kruse. Stochastic evolution equations in Hilbert spaces. Lecture Notes in Mathematics, 2093:11–49, 2014. CODEN LN-MAA2. ISBN 3-319-02230-X (print), 3-319-02231-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02231-4_
2/.

Kruse:2014:SWA

[Kru14i] Raphael Kruse. Strong and Weak Approximation of Semilinear Stochastic Evolution Equations, volume 2093 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-02230-X (print), 3-319-02231-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 177 pp. LCCN QA3 .L28 no.

2093. URL http://link.springer.com/book/10.1007/978-
3-319-02231-4;   http://www.springerlink.com/content/
978-3-319-02231-4.

Kowar:2012:AMP

[KS12] Richard Kowar and Otmar Scherzer. Attenuation models in photoacoustics. Lecture Notes in Mathematics, 2035:85–130, 2012. CODEN LNMAA2. ISBN 3-642-22989-1 (print), 3-642-22990-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22990-9_4/.

Koch:2013:MCP

[KS13] Gilbert Koch and Johannes Schropp. Mathematical concepts in pharmacokinetics and pharmacodynamics with application to tumor growth. Lecture Notes in Mathematics, 2102:225– 250, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-03080-7_7/.

Kumagai:2014:AOC

[Kum14a] Takashi Kumagai. Alexander–Orbach conjecture holds when twopoint functions behave nicely. Lecture Notes in Mathematics, 2101:65–77, 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-03152-1_6/.

Kumagai:2014:BM

[Kum14b] Takashi Kumagai. Back matter. Lecture Notes in Mathematics, 2101:135–150, 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692

(electronic). URL http://link.springer.com/content/pdf/
bbm:978-3-319-03152-1/1.pdf.

Kumagai:2014:FM

[Kum14c] Takashi Kumagai. Front matter. Lecture Notes in Mathematics, 2101:i–x, 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-03152-1/1.pdf.

Kumagai:2014:FRR

[Kum14d] Takashi Kumagai. Further results for random walk on IIC. Lecture Notes in Mathematics, 2101:79–93, 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-03152-1_7/.

Kumagai:2014:HKEc

[Kum14e] Takashi Kumagai. Heat kernel estimates for random weighted graphs. Lecture Notes in Mathematics, 2101:59–64, 2014. CO-DEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-03152-1_
5/.

Kumagai:2014:HKEa

[Kum14f] Takashi Kumagai. Heat kernel estimates: General theory. Lecture Notes in Mathematics, 2101:21–41, 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-03152-1_3/.

Kumagai:2014:HKEb

[Kum14g] Takashi Kumagai. Heat kernel estimates using effective resistance. Lecture Notes in Mathematics, 2101:43–58, 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-03152-1_4/.

Kumagai:2014:I

[Kum14h] Takashi Kumagai. Introduction. Lecture Notes in Mathematics, 2101:1–2, 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-03152-1_1.

Kumagai:2014:RCM

[Kum14i] Takashi Kumagai. Random conductance model. Lecture Notes in Mathematics, 2101:95–134, 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-03152-1_8/.

Kumagai:2014:RWD

[Kum14j] Takashi Kumagai, editor. Random Walks on Disordered Media and their Scaling Limits: École d’Été de Probabilités de Saint-Flour XL — 2010, volume 2101 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2101. URL http://link.

springer.com/book/10.1007/978-3-319-03152-1;   http:/
/www.springerlink.com/content/978-3-319-03152-1.

Kumagai:2014:WGA

[Kum14k] Takashi Kumagai. Weighted graphs and the associated Markov chains. Lecture Notes in Mathematics, 2101:3–19, 2014. CODEN LNMAA2. ISBN 3-319-03151-1 (hardcover), 3-319-03152-X. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-03152-1_2/.

Kumar:2015:BM

[Kum15a] P. R. Kumar. Back matter. Lecture Notes in Mathematics, 2141: 137–138, 2015. CODEN LNMAA2. ISBN 3-319-16966-1 (print), 3-319-16967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-16967-5/1.pdf.

Kumar:2015:BGB

[Kum15b] P. R. Kumar. Bridging the gap between information theory and wireless networking. Lecture Notes in Mathematics, 2141:109– 135, 2015. CODEN LNMAA2. ISBN 3-319-16966-1 (print), 3-319-16967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-319-16967-5_4.pdf.

Kumar:2015:MFC

[Kum15c] P. R. Kumar, editor. Mathematical Foundations of Complex Networked Information Systems: Politecnico di Torino, Verrès, Italy 2009, volume 2141 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LNMAA2. ISBN 3-319-16966-1 (print), 3-319-16967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 138 (est.) pp. LCCN ???? URL http://www.

springerlink.com/content/978-3-319-16967-5.

Kyprianou:2014:PSP

[KW14] A. E. Kyprianou and A. R. Watson. Potentials of stable processes. Lecture Notes in Mathematics, 2123:333–343, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_12.pdf.

Kwok:2011:GSR

[Kwo11] Stephen Kwok. On the geometry of super Riemann surfaces. Lecture Notes in Mathematics, 2027:139–154, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21744-9_
7/.

Kumar:2015:FM

[KWZ15] P. R. Kumar, Martin J. Wainwright, and Riccardo Zecchina. Front matter. Lecture Notes in Mathematics, 2141:i–vii, 2015. CODEN LNMAA2. ISBN 3-319-16966-1 (print), 3-319-16967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-16967-5/1.pdf.

Lamb:2015:AFA

[Lam15] Wilson Lamb. Applying functional analytic techniques to evolution equations. Lecture Notes in Mathematics, 2126:1–46, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11322-7_1.pdf.

Latala:2012:MUL

[Lat12] Rafał Latała. Moments of unconditional logarithmically concave vectors. Lecture Notes in Mathematics, 2050:301–315, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29849-3_17/.

Latala:2014:MPI

[Lat14] Rafał Latała. Modified Paouris inequality. Lecture Notes in Mathematics, 2116:293–307, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-09477-9_19.pdf.

Laurent:2011:SC

[Lau11] Stéphane Laurent. On standardness and I-cosiness. Lecture Notes in Mathematics, 2006:127–186, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-15217-7_5/.

Laurent:2013:VIL

[Lau13] Stéphane Laurent. Vershik’s intermediate level standardness criterion and the scale of an automorphism. Lecture Notes in Mathematics, 2078:123–139, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00321-4_3/.

Laurencot:2015:WCT

[Lau15] Philippe Laurençot. Weak compactness techniques and coagulation equations. Lecture Notes in Mathematics, 2126:199–253, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11322-7_5.pdf.

LeJan:2011:BM

[Le 11a] Yves Le Jan. Back matter. Lecture Notes in Mathematics, 2026: 115–124, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-21216-1/1.pdf.

LeJan:2011:CGS

[Le 11b] Yves Le Jan. The case of general symmetric Markov processes. Lecture Notes in Mathematics, 2026:99–113, 2011. CODEN LN-MAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-21216-1_10/.

LeJan:2011:D

[Le 11c] Yves Le Jan. Decompositions. Lecture Notes in Mathematics, 2026:65–73, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21216-1_7/.

LeJan:2011:EVR

[Le 11d] Yves Le Jan. Energy variation and representations. Lecture Notes in Mathematics, 2026:57–64, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21216-1_6/.

LeJan:2011:FM

[Le 11e] Yves Le Jan. Front matter. Lecture Notes in Mathematics, 2026: i–viii, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-21216-1/1.pdf.

LeJan:2011:GFF

[Le 11f] Yves Le Jan. The Gaussian free field. Lecture Notes in Mathematics, 2026:47–56, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21216-1_5/.

LeJan:2011:GL

[Le 11g] Yves Le Jan. Geodesic loops. Lecture Notes in Mathematics, 2026: 29–34, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21216-1_3/.

LeJan:2011:LES

[Le 11h] Yves Le Jan. Loop erasure and spanning trees. Lecture Notes in Mathematics, 2026:75–89, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21216-1_8/.

LeJan:2011:LM

[Le 11i] Yves Le Jan. Loop measures. Lecture Notes in Mathematics, 2026:13–28, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21216-1_2/.

LeJan:2011:MPL

[Le 11j] Yves Le Jan, editor. Markov Paths, Loops and Fields: École d’Été de Probabilités de Saint-Flour XXXVIII — 2008, volume 2026 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no.

2026; QA274.7 .L46 2011. URL http://link.springer.com/
book/10.1007/978-3-642-21216-1;   http://site.ebrary.
com/id/10483931;   http://www.springerlink.com/content/
978-3-642-21215-4/;   http://www.springerlink.com/content/
978-3-642-21216-1.

LeJan:2011:PPL

[Le 11k] Yves Le Jan. Poisson process of loops. Lecture Notes in Mathematics, 2026:35–45, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21216-1_4/.

LeJan:2011:RP

[Le 11l] Yves Le Jan. Reflection positivity. Lecture Notes in Mathematics, 2026:91–97, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21216-1_9/.

LeJan:2011:SMP

[Le 11m] Yves Le Jan. Symmetric Markov processes on finite spaces. Lecture Notes in Mathematics, 2026:1–12, 2011. CODEN LNMAA2. ISBN 3-642-21215-8 (paperback), 3-642-21216-6. ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21216-1_1/.

LeRousseau:2012:CES

[Le 12] Jérôme Le Rousseau. Carleman estimates and some applications to control theory. Lecture Notes in Mathematics, 2048: 207–243, 2012. CODEN LNMAA2. ISBN 3-642-27892-2 (print), 3-642-27893-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27893-8_4/.

Lang:2011:EEG

[LE11a] Jan Lang and David Edmunds. Eigenvalues, Embeddings and Generalised Trigonometric Functions, volume 2016 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 220 pp. LCCN

QA3 .L28 no. 2016. URL http://link.springer.com/book/
10.1007/978-3-642-18429-1;   http://www.springerlink.
com/content/978-3-642-18429-1.

Lang:2011:FM

[LE11b] Jan Lang and David Edmunds. Front matter. Lecture Notes in Mathematics, 2016:i–xi, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-18429-1/1.pdf.

Lang:2011:BMb

[LE11c] Prof. Jan Lang and Prof. David Edmunds. Back matter. Lecture Notes in Mathematics, 2016:211–220, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-18429-1/1.pdf.

Lang:2011:BMa

[LE11d] Prof. Jan Lang and Prof. David Edmunds. Basic material. Lecture Notes in Mathematics, 2016:1–31, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-18429-1_1/.

Lang:2011:ENW

[LE11e] Prof. Jan Lang and Prof. David Edmunds. Estimates of snumbers of weighted Hardy operators. Lecture Notes in Mathematics, 2016:105–128, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-18429-1_6/.

Lang:2011:HO

[LE11f] Prof. Jan Lang and Prof. David Edmunds. Hardy operators. Lecture Notes in Mathematics, 2016:65–71, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-18429-1_4/.

Lang:2011:HOV

[LE11g] Prof. Jan Lang and Prof. David Edmunds. Hardy operators on variable exponent spaces. Lecture Notes in Mathematics, 2016:

183–209, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-18429-1_9/.

Lang:2011:LSN

[LE11h] Prof. Jan Lang and Prof. David Edmunds. The Laplacian and some natural variants. Lecture Notes in Mathematics, 2016:49– 63, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-18429-1_3/.

Lang:2011:MRE

[LE11i] Prof. Jan Lang and Prof. David Edmunds. More refined estimates. Lecture Notes in Mathematics, 2016:129–151, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-18429-1_
7/.

Lang:2011:NLI

[LE11j] Prof. Jan Lang and Prof. David Edmunds. A non-linear integral system. Lecture Notes in Mathematics, 2016:153–182, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-18429-1_8/.

Lang:2011:NGT

[LE11k] Prof. Jan Lang and Prof. David Edmunds. s-numbers and generalised trigonometric functions. Lecture Notes in Mathematics, 2016:73–104, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-18429-1_5/.

Lang:2011:TG

[LE11l] Prof. Jan Lang and Prof. David Edmunds. Trigonometric generalisations. Lecture Notes in Mathematics, 2016:33–48, 2011. CODEN LNMAA2. ISBN 3-642-18267-4 (print), 3-642-18429-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-18429-1_2/.

Ledoux:2014:GAF

[Led14a] Michel Ledoux, editor. Geometric Aspects of Functional Analysis: Israel Seminar (GAFA) 2011–2013, volume 2116 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 333 (est.) pp. LCCN ???? URL

http://www.springerlink.com/content/978-3-319-09477-9.

Ledoux:2014:RGN

[Led14b] Michel Ledoux. Remarks on Gaussian noise stability, Brascamp– Lieb and Slepian inequalities. Lecture Notes in Mathematics, 2116:309–333, 2014. CODEN LNMAA2. ISBN 3-319-09476-9 (print), 3-319-09477-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-09477-9_20.pdf.

Lehec:2011:MGC

[Leh11] Joseph Lehec. Moments of the Gaussian chaos. Lecture Notes in Mathematics, 2006:327–340, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-15217-7_13/.

Lejay:2012:GSR

[Lej12] Antoine Lejay. Global solutions to rough differential equations with unbounded vector fields. Lecture Notes in Mathematics, 2046:215–246, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-27461-9_11/.

Leonard:2012:BM

[Léo12a] Christian Léonard. Back matter. Lecture Notes in Mathematics, 2046:467–469, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-27461-9/1.pdf.

Leonard:2012:GTU

[Léo12b] Christian Léonard. Girsanov theory under a finite entropy condition. Lecture Notes in Mathematics, 2046:429–465, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27461-9_
20/.

Leonard:2014:SPP

[Léo14] Christian Léonard. Some properties of path measures. Lecture Notes in Mathematics, 2123:207–230, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-11970-0_8.
pdf.

Lewis:2012:ABH

[Lew12] J. Lewis. Applications of boundary Harnack inequalities for p harmonic functions and related topics. Lecture Notes in Mathematics, 2045:1–72, 2012. CODEN LNMAA2. ISBN 3-642-27144-8 (print), 3-642-27145-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-27145-8_1/.

Liebeck:2013:PAA

[Lie13] Martin W. Liebeck. Probabilistic and asymptotic aspects of finite simple groups. Lecture Notes in Mathematics, 2070:1–34, 2013. CODEN LNMAA2. ISBN 1-4471-4813-4 (print), 1-4471-4814-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-1-4471-4814-2_1/;   http://link.springer.
com/content/pdf/bfm:978-1-4471-4814-2/1.pdf.

Liebscher:2015:ACM

[Lie15a] Stefan Liebscher. Application: Cosmological models of Bianchi type, the tumbling universe. Lecture Notes in Mathematics, 2117:115–118, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-10777-6_13.pdf.

Liebscher:2015:ADN

[Lie15b] Stefan Liebscher. Application: Decoupling in networks. Lecture Notes in Mathematics, 2117:43–47, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-10777-6_6.
pdf.

Liebscher:2015:AFF

[Lie15c] Stefan Liebscher. Application: Fluid flow in a planar channel, spatial dynamics with reversible Bogdanov–Takens bifurcation. Lecture Notes in Mathematics, 2117:119–128, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10777-6_14.pdf.

Liebscher:2015:AOP

[Lie15d] Stefan Liebscher. Application: Oscillatory profiles in systems of hyperbolic balance laws. Lecture Notes in Mathematics, 2117:49– 54, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10777-6_7.pdf.

Liebscher:2015:BM

[Lie15e] Stefan Liebscher. Back matter. Lecture Notes in Mathematics, 2117:139–144, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-10777-6/1.pdf.

Liebscher:2015:BP

[Lie15f] Stefan Liebscher. Bifurcation without Parameters, volume 2117 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 144 (est.) pp.

LCCN ???? URL http://www.springerlink.com/content/
978-3-319-10777-6.

Liebscher:2015:BTB

[Lie15g] Stefan Liebscher. Bogdanov–Takens bifurcation. Lecture Notes in Mathematics, 2117:81–102, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-10777-6_10.pdf.

Liebscher:2015:COM

[Lie15h] Stefan Liebscher. Codimension-one manifolds of equilibria. Lecture Notes in Mathematics, 2117:131–133, 2015. CO-DEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10777-6_15.pdf.

Liebscher:2015:C

[Lie15i] Stefan Liebscher. Cosymmetries. Lecture Notes in Mathematics, 2117:21–23, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-10777-6_3.pdf.

Liebscher:2015:DPA

[Lie15j] Stefan Liebscher. Degenerate Poincaré–Andronov–Hopf bifurcation. Lecture Notes in Mathematics, 2117:67–79, 2015. CO-DEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10777-6_9.pdf.

Liebscher:2015:DTB

[Lie15k] Stefan Liebscher. Degenerate transcritical bifurcation. Lecture Notes in Mathematics, 2117:57–65, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-10777-6_8.
pdf.

Liebscher:2015:DHB

[Lie15l] Stefan Liebscher. Double-Hopf bifurcation. Lecture Notes in Mathematics, 2117:109–113, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-10777-6_12.pdf.

Liebscher:2015:FMa

[Lie15m] Stefan Liebscher. Front matter. Lecture Notes in Mathematics, 2117:i–xii, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-10777-6/1.pdf.

Liebscher:2015:FMb

[Lie15n] Stefan Liebscher. Front matter. Lecture Notes in Mathematics, 2117:1, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-10777-6/1/1.pdf.

Liebscher:2015:FMc

[Lie15o] Stefan Liebscher. Front matter. Lecture Notes in Mathematics, 2117:25, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-10777-6/2/1.pdf.

Liebscher:2015:FMd

[Lie15p] Stefan Liebscher. Front matter. Lecture Notes in Mathematics, 2117:55, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-10777-6/3/1.pdf.

Liebscher:2015:FMe

[Lie15q] Stefan Liebscher. Front matter. Lecture Notes in Mathematics, 2117:129, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-10777-6/4/1.pdf.

Liebscher:2015:I

[Lie15r] Stefan Liebscher. Introduction. Lecture Notes in Mathematics, 2117:3–12, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-10777-6_1.pdf.

Liebscher:2015:MC

[Lie15s] Stefan Liebscher. Methods and concepts. Lecture Notes in Mathematics, 2117:13–19, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-10777-6_2.pdf.

Liebscher:2015:PAH

[Lie15t] Stefan Liebscher. Poincaré–Andronov–Hopf bifurcation. Lecture Notes in Mathematics, 2117:35–41, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-10777-6_5.
pdf.

Liebscher:2015:SO

[Lie15u] Stefan Liebscher. Summary and outlook. Lecture Notes in Mathematics, 2117:135–137, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-10777-6_16.pdf.

Liebscher:2015:TB

[Lie15v] Stefan Liebscher. Transcritical bifurcation. Lecture Notes in Mathematics, 2117:27–34, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-10777-6_4.pdf.

Liebscher:2015:ZHB

[Lie15w] Stefan Liebscher. Zero-Hopf bifurcation. Lecture Notes in Mathematics, 2117:103–108, 2015. CODEN LNMAA2. ISBN 3-319-10776-3 (print), 3-319-10777-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-10777-6_11.pdf.

Linss:2010:ABSa

[Lin10a] Torsten Linß. The analytical behaviour of solutions. Lecture Notes in Mathematics, 1985:33–76, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-05134-0_3.
pdf.

Linss:2010:ABSb

[Lin10b] Torsten Linß. The analytical behaviour of solutions. Lecture Notes in Mathematics, 1985:33–76, 2010. CODEN LNMAA2.

ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-05134-0_3.
pdf.

Linss:2010:ABSc

[Lin10c] Torsten Linß. The analytical behaviour of solutions. Lecture Notes in Mathematics, 1985:235–246, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-05134-0_7.
pdf.

Linss:2010:ABSd

[Lin10d] Torsten Linß. The analytical behaviour of solutions. Lecture Notes in Mathematics, 1985:235–246, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-05134-0_7.
pdf.

Linss:2010:BMa

[Lin10e] Torsten Linß. Back matter. Lecture Notes in Mathematics, 1985: 1–18, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-05134-0/1.pdf.

Linss:2010:BMb

[Lin10f] Torsten Linß. Back matter. Lecture Notes in Mathematics, 1985: 1–18, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-05134-0/1.pdf.

Linss:2010:CDPa

[Lin10g] Torsten Linß. Convection-diffusion problems. Lecture Notes in Mathematics, 1985:257–307, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-05134-0_9.pdf.

Linss:2010:CDPb

[Lin10h] Torsten Linß. Convection-diffusion problems. Lecture Notes in Mathematics, 1985:257–307, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-05134-0_9.pdf.

Linss:2010:DRCa

[Lin10i] Torsten Linß. Discretisations of reaction-convection-diffusion problems. Lecture Notes in Mathematics, 1985:183–231, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05134-0_6.pdf.

Linss:2010:DRCb

[Lin10j] Torsten Linß. Discretisations of reaction-convection-diffusion problems. Lecture Notes in Mathematics, 1985:183–231, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05134-0_6.pdf.

Linss:2010:FDSa

[Lin10k] Torsten Linß. Finite difference schemes for convection-diffusion problems. Lecture Notes in Mathematics, 1985:77–149, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05134-0_4.pdf.

Linss:2010:FDSb

[Lin10l] Torsten Linß. Finite difference schemes for convection-diffusion problems. Lecture Notes in Mathematics, 1985:77–149, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05134-0_4.pdf.

Linss:2010:FEFa

[Lin10m] Torsten Linß. Finite element and finite volume methods. Lecture Notes in Mathematics, 1985:151–182, 2010. CO-DEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05134-0_5.pdf.

Linss:2010:FEFb

[Lin10n] Torsten Linß. Finite element and finite volume methods. Lecture Notes in Mathematics, 1985:151–182, 2010. CO-DEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05134-0_5.pdf.

Linss:2010:FMa

[Lin10o] Torsten Linß. Front matter. Lecture Notes in Mathematics, 1985: 1–10, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-05134-0/1.pdf.

Linss:2010:FMb

[Lin10p] Torsten Linß. Front matter. Lecture Notes in Mathematics, 1985: 1–10, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-05134-0/1.pdf.

Linss:2010:FMc

[Lin10q] Torsten Linß. Front matter. Lecture Notes in Mathematics, 1985: 31, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-05134-0/1/1.pdf.

Linss:2010:FMd

[Lin10r] Torsten Linß. Front matter. Lecture Notes in Mathematics, 1985: 31, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-05134-0/1/1.pdf.

Linss:2010:FMe

[Lin10s] Torsten Linß. Front matter. Lecture Notes in Mathematics, 1985: 233, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-05134-0/2/1.pdf.

Linss:2010:FMf

[Lin10t] Torsten Linß. Front matter. Lecture Notes in Mathematics, 1985: 233, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-05134-0/2/1.pdf.

Linss:2010:Ia

[Lin10u] Torsten Linß. Introduction. Lecture Notes in Mathematics, 1985: 1–4, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05134-0_1.pdf.

Linss:2010:Ib

[Lin10v] Torsten Linß. Introduction. Lecture Notes in Mathematics, 1985: 1–4, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-05134-0_1.pdf.

Linss:2010:LAMa

[Lin10w] Torsten Linß. Layer-adapted meshes. Lecture Notes in Mathematics, 1985:5–29, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-05134-0_2.pdf.

Linss:2010:LAMb

[Lin10x] Torsten Linß. Layer-adapted meshes. Lecture Notes in Mathematics, 1985:5–29, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-05134-0_2.pdf.

Linss:2010:LAMc

[Lin10y] Torsten Linß, editor. Layer-Adapted Meshes for Reaction-Convection-Diffusion Problems, volume 1985 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 320 pp. LCCN QA3 .L28 no.

1985. URL http://link.springer.com/book/10.1007/978-
3-642-05134-0;   http://www.springerlink.com/content/
978-3-642-05134-0.

Linss:2010:RDPa

[Lin10z] Torsten Linß. Reaction-diffusion problems. Lecture Notes in Mathematics, 1985:247–256, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-05134-0_8.pdf.

Linss:2010:RDPb

[Lin10-27] Torsten Linß. Reaction-diffusion problems. Lecture Notes in Mathematics, 1985:247–256, 2010. CODEN LNMAA2. ISBN 3-642-05133-2 (print), 3-642-05134-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-05134-0_8.pdf.

Lindqvist:2012:RS

[Lin12] Peter Lindqvist. Regularity of supersolutions. Lecture Notes in Mathematics, 2045:73–131, 2012. CODEN LNMAA2. ISBN 3-642-27144-8 (print), 3-642-27145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27145-8_2/.

Lin:2013:SRR

[Lin13] Kevin K. Lin. Stimulus-response reliability of biological networks. Lecture Notes in Mathematics, 2102:135–161, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-03080-7_
4/.

Litvinov:2013:BM

[Lit13a] Grigory L. Litvinov. Back matter. Lecture Notes in Mathematics, 2074:303–304, 2013. CODEN LNMAA2. ISBN 3-642-36432-2 (print), 3-642-36433-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-36433-4/1.pdf.

Litvinov:2013:ITA

[Lit13b] Grigory L. Litvinov. Idempotent/tropical analysis, the Hamilton– Jacobi and Bellman equations. Lecture Notes in Mathematics, 2074:251–301, 2013. CODEN LNMAA2. ISBN 3-642-36432-2 (print), 3-642-36433-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-36433-4_4/.

Lewis:2012:FM

[LLMS12a] John Lewis, Peter Lindqvist, Juan J. Manfredi, and Sandro Salsa. Front matter. Lecture Notes in Mathematics, 2045:i– xi, 2012. CODEN LNMAA2. ISBN 3-642-27144-8 (print), 3-642-27145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-27145-8/1.pdf.

Lewis:2012:REN

[LLMS12b] John Lewis, Peter Lindqvist, Juan J. Manfredi, and Sandro Salsa, editors. Regularity Estimates for Nonlinear Elliptic and Parabolic Problems: Cetraro, Italy 2009, volume 2045 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-27144-8 (print), 3-642-27145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2045; QA3 .L28 no. 2045; QA377 .R448 2012. URL http://link.

springer.com/book/10.1007/978-3-642-27145-8;   http:/
/www.springerlink.com/content/978-3-642-27145-8.

Lewis:2013:DIM

[LMP13a] Mark A. Lewis, Philip K. Maini, and Sergei V. Petrovskii. Dispersal, Individual Movement and Spatial Ecology: A Mathematical Perspective, volume 2071 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 385 pp. LCCN QA3 .L28 no. 2071;

QK911 .D57 2013. URL http://link.springer.com/book/
10.1007/978-3-642-35497-7;   http://www.springerlink.
com/content/978-3-642-35497-7.

Lewis:2013:FMa

[LMP13b] Mark A. Lewis, Philip K. Maini, and Sergei V. Petrovskii. Front matter. Lecture Notes in Mathematics, 2071:i–xiv, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-35497-7/1.pdf.

Lewis:2013:FMb

[LMP13c] Mark A. Lewis, Philip K. Maini, and Sergei V. Petrovskii. Front matter. Lecture Notes in Mathematics, 2071:1, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-35497-7/
1/1.pdf.

Lorenz:2010:AME

[Lor10a] Thomas Lorenz. Adapting mutational equations to examples in vector spaces: Local parameters of continuity. Lecture Notes in Mathematics, 1996:103–179, 2010. CODEN LNMAA2. ISBN 3-642-12470-4 (print), 3-642-12471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12471-6_3.pdf.

Lorenz:2010:BM

[Lor10b] Thomas Lorenz. Back matter. Lecture Notes in Mathematics, 1996:439–515, 2010. CODEN LNMAA2. ISBN 3-642-12470-4 (print), 3-642-12471-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-12471-6/1.pdf.

Lorenz:2010:EOD

[Lor10c] Thomas Lorenz. Extending ordinary differential equations to metric spaces: Aubin’s suggestion. Lecture Notes in Mathematics, 1996:31–101, 2010. CODEN LNMAA2. ISBN 3-642-12470-4 (print), 3-642-12471-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12471-6_2.pdf.

Lorenz:2010:FM

[Lor10d] Thomas Lorenz. Front matter. Lecture Notes in Mathematics, 1996:i–xiv, 2010. CODEN LNMAA2. ISBN 3-642-12470-4 (print), 3-642-12471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-12471-6/1.pdf.

Lorenz:2010:IDL

[Lor10e] Thomas Lorenz. Introducing distribution-like solutions to mutational equations. Lecture Notes in Mathematics, 1996:331–384, 2010. CODEN LNMAA2. ISBN 3-642-12470-4 (print), 3-642-12471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12471-6_5.pdf.

Lorenz:2010:I

[Lor10f] Thomas Lorenz. Introduction. Lecture Notes in Mathematics, 1996:1–29, 2010. CODEN LNMAA2. ISBN 3-642-12470-4 (print), 3-642-12471-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12471-6_1.pdf.

Lorenz:2010:LRC

[Lor10g] Thomas Lorenz. Less restrictive conditions on distance functions: Continuity instead of triangle inequality. Lecture Notes in Mathematics, 1996:181–330, 2010. CODEN LNMAA2. ISBN 3-642-12470-4 (print), 3-642-12471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12471-6_4.pdf.

Lorenz:2010:MAJ

[Lor10h] Thomas Lorenz. Mutational Analysis: A Joint Framework for Cauchy Problems in and Beyond Vector Spaces, volume 1996 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CO-DEN LNMAA2. ISBN 3-642-12470-4 (print), 3-642-12471-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 509 pp. LCCN QA3 .L28 no. 1996. URL http://link.

springer.com/book/10.1007/978-3-642-12471-6;   http:/
/www.springerlink.com/content/978-3-642-12471-6.

Lorenz:2010:MIM

[Lor10i] Thomas Lorenz. Mutational inclusions in metric spaces. Lecture Notes in Mathematics, 1996:385–438, 2010. CODEN LNMAA2. ISBN 3-642-12470-4 (print), 3-642-12471-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-12471-6_6.
pdf.

Layton:2012:FM

[LR12a] William J. Layton and Leo Rebholz. Front matter. Lecture Notes in Mathematics, 2042:i–viii, 2012. CODEN LNMAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-24409-4/1.pdf.

Layton:2012:ADM

[LR12b] William J. Layton and Leo G. Rebholz. Approximate Deconvolution Models of Turbulence: Analysis, Phenomenology and Numerical Analysis, volume 2042 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 184 pp. LCCN QA3 .L28 no.

2042. URL http://link.springer.com/book/10.1007/978-
3-642-24409-4;   http://www.springerlink.com/content/
978-3-642-24409-4.

Layton:2012:ADO

[LR12c] William J. Layton and Leo G. Rebholz. Approximate deconvolution operators and models. Lecture Notes in Mathematics, 2042: 61–88, 2012. CODEN LNMAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-24409-4_3/.

Layton:2012:BM

[LR12d] William J. Layton and Leo G. Rebholz. Back matter. Lecture Notes in Mathematics, 2042:163–184, 2012. CODEN LNMAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-24409-4/1.pdf.

Layton:2012:I

[LR12e] William J. Layton and Leo G. Rebholz. Introduction. Lecture Notes in Mathematics, 2042:1–33, 2012. CODEN LNMAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-24409-4_1/.

Layton:2012:LES

[LR12f] William J. Layton and Leo G. Rebholz. Large eddy simulation. Lecture Notes in Mathematics, 2042:35–60, 2012. CODEN LNMAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24409-4_
2/.

Layton:2012:LDR

[LR12g] William J. Layton and Leo G. Rebholz. The Leray-deconvolution regularization. Lecture Notes in Mathematics, 2042:121–144, 2012. CODEN LNMAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-24409-4_6/.

Layton:2012:NAN

[LR12h] William J. Layton and Leo G. Rebholz. NS-alpha- and NS-omega-deconvolution regularizations. Lecture Notes in Mathematics, 2042:145–162, 2012. CODEN LNMAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-24409-4_7/.

Layton:2012:PA

[LR12i] William J. Layton and Leo G. Rebholz. Phenomenology of ADMs. Lecture Notes in Mathematics, 2042:89–97, 2012. CODEN LN-MAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24409-4_
4/.

Layton:2012:TRT

[LR12j] William J. Layton and Leo G. Rebholz. Time relaxation truncates scales. Lecture Notes in Mathematics, 2042:99–120, 2012. CODEN LNMAA2. ISBN 3-642-24408-4 (print), 3-642-24409-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-24409-4_5/.

Letemplier:2014:UHT

[LS14] Julien Letemplier and Thomas Simon. Unimodality of hitting times for stable processes. Lecture Notes in Mathematics, 2123:

345–357, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-319-11970-0_13.pdf.

Magyar:2014:DSD

[Mag14] Ákos Magyar. On the distribution of solutions to Diophantine equations. Lecture Notes in Mathematics, 2107:487–538, 2014. CODEN LNMAA2. ISBN 3-319-04695-0 (print), 3-319-04696-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-04696-9_8.pdf.

Major:2013:BM

[Maj13a] Péter Major. Back matter. Lecture Notes in Mathematics, 2079: 247–290, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-37617-7/1.pdf.

Major:2013:CPT

[Maj13b] Péter Major. The completion of the proof of theorem 4.1. Lecture Notes in Mathematics, 2079:53–64, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-37617-7_7/.

Major:2013:DFP

[Maj13c] Péter Major. The diagram formula for products of degenerate U-statistics. Lecture Notes in Mathematics, 2079:121–138, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-37617-7_11/.

Major:2013:FMR

[Maj13d] Péter Major. Formulation of the main results of this work. Lecture Notes in Mathematics, 2079:65–78, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-37617-7_8/.

Major:2013:FM

[Maj13e] Péter Major. Front matter. Lecture Notes in Mathematics, 2079: i–xiii, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-37617-7/1.pdf.

Major:2013:I

[Maj13f] Péter Major. Introduction. Lecture Notes in Mathematics, 2079: 1–3, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-37617-7_1/.

Major:2013:MID

[Maj13g] Péter Major. Motivation of the investigation: Discussion of some problems. Lecture Notes in Mathematics, 2079:5–13, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-37617-7_2/.

Major:2013:MWI

[Maj13h] Péter Major. Multiple Wiener–Itô integrals and their properties. Lecture Notes in Mathematics, 2079:97–120, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-37617-7_
10/.

Major:2013:EMR

[Maj13i] Péter Major. On the Estimation of Multiple Random Integrals and U-Statistics, volume 2079 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 288 pp. LCCN QA3 .L28 no.

2079. URL http://link.springer.com/book/10.1007/978-
3-642-37617-7;   http://www.springerlink.com/content/
978-3-642-37617-7.

Major:2013:SNC

[Maj13j] Péter Major. On the supremum of a nice class of partial sums. Lecture Notes in Mathematics, 2079:21–33, 2013. CODEN LN-MAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-37617-7_
4/.

Major:2013:ORD

[Maj13k] Péter Major. An overview of the results and a discussion of the literature. Lecture Notes in Mathematics, 2079:227–245, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-37617-7_18/.

Major:2013:PDF

[Maj13l] Péter Major. The proof of the diagram formula for U-statistics. Lecture Notes in Mathematics, 2079:139–149, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-37617-7_
12/.

Major:2013:PMR

[Maj13m] Péter Major. The proof of the main result. Lecture Notes in Mathematics, 2079:209–225, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-37617-7_17/.

Major:2013:PTS

[Maj13n] Péter Major. The proof of theorems 4.1 and 4.2 on the supremum of random sums. Lecture Notes in Mathematics, 2079: 41–51, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-37617-7_6/.

Major:2013:PTE

[Maj13o] Péter Major. The proof of theorems 8.3, 8.5 and example 8.7. Lecture Notes in Mathematics, 2079:151–168, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-37617-7_
13/.

Major:2013:RMR

[Maj13p] Péter Major. Reduction of the main result in this work. Lecture Notes in Mathematics, 2079:169–179, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-37617-7_14/.

Major:2013:SEA

[Maj13q] Péter Major. Some estimates about sums of independent random variables. Lecture Notes in Mathematics, 2079:15–20, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-37617-7_3/.

Major:2013:SRA

[Maj13r] Péter Major. Some results about U-statistics. Lecture Notes in Mathematics, 2079:79–95, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-37617-7_9/.

Major:2013:SPM

[Maj13s] Péter Major. The strategy of the proof for the main result of this work. Lecture Notes in Mathematics, 2079:181–189, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-37617-7_15/.

Major:2013:SA

[Maj13t] Péter Major. A symmetrization argument. Lecture Notes in Mathematics, 2079:191–208, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-37617-7_16/.

Major:2013:VCC

[Maj13u] Péter Major. Vapnik–Červonenkis classes and L_{2}-dense classes of functions. Lecture Notes in Mathematics, 2079:35–39, 2013. CODEN LNMAA2. ISBN 3-642-37616-9 (print), 3-642-37617-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-37617-7_5/.

Major:2014:BM

[Maj14a] Péter Major. Back matter. Lecture Notes in Mathematics, 849: 123–128, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-02642-8/1.pdf.

Major:2014:FM

[Maj14b] Péter Major. Front matter. Lecture Notes in Mathematics, 849: i–xiii, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-02642-8/1.pdf.

Major:2014:HPC

[Maj14c] Péter Major. History of the problems: Comments. Lecture Notes in Mathematics, 849:113–122, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02642-8_9/.

Major:2014:MWIa

[Maj14d] Péter Major. Multiple Wiener–Itô integrals. Lecture Notes in Mathematics, 849:27–42, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02642-8_4/.

Major:2014:MWIb

[Maj14e] Péter Major. Multiple Wiener–Itô Integrals: With Applications to Limit Theorems, volume 849 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 126 pp. LCCN QA3 .L28 no. 849.

URL http://link.springer.com/book/10.1007/978-3-319-
02642-8;   http://www.springerlink.com/content/978-3-
319-02642-8.

Major:2014:NCL

[Maj14f] Péter Major. Non-central limit theorems. Lecture Notes in Mathematics, 849:87–112, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02642-8_8/.

Major:2014:LP

[Maj14g] Péter Major. On a limit problem. Lecture Notes in Mathematics, 849:1–8, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02642-8_1/.

Major:2014:OWI

[Maj14h] Péter Major. On the original Wiener–Itô integral. Lecture Notes in Mathematics, 849:81–86, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-02642-8_7/.

Major:2014:PIF

[Maj14i] Péter Major. The proof of Itô’s formula: The diagram formula and some of its consequences. Lecture Notes in Mathematics, 849: 43–64, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02642-8_5/.

Major:2014:RSM

[Maj14j] Péter Major. Random spectral measures. Lecture Notes in Mathematics, 849:15–26, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-02642-8_3/.

Major:2014:SRF

[Maj14k] Péter Major. Subordinated random fields: Construction of self-similar fields. Lecture Notes in Mathematics, 849:65–79, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02642-8_6/.

Major:2014:WP

[Maj14l] Péter Major. Wick polynomials. Lecture Notes in Mathematics, 849:9–14, 2014. CODEN LNMAA2. ISBN 3-319-02641-0 (print), 3-319-02642-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02642-8_2/.

Manfredi:2012:IRT

[Man12] Juan J. Manfredi. Introduction to random tug-of-war games and PDEs. Lecture Notes in Mathematics, 2045:133–151, 2012. CODEN LNMAA2. ISBN 3-642-27144-8 (print), 3-642-27145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27145-8_3/.

Marouby:2011:SLT

[Mar11a] Matthieu Marouby. Simulation of a local time fractional stable motion. Lecture Notes in Mathematics, 2006:221–239, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_9/.

Marrani:2011:COM

[Mar11b] Alessio Marrani. Charge orbits and moduli spaces of black hole attractors. Lecture Notes in Mathematics, 2027:155–174, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21744-9_8/.

Markowsky:2012:DIL

[Mar12] Greg Markowsky. The derivative of the intersection local time of Brownian motion through Wiener chaos. Lecture Notes in Mathematics, 2046:141–148, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27461-9_6/.

Masuda:2015:BM

[Mas15a] Hiroki Masuda. Back matter. Lecture Notes in Mathematics, 2128:287–288, 2015. CODEN LNMAA2. ISBN 3-319-12372-6 (print), 3-319-12373-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-12373-8/1.pdf.

Masuda:2015:LMI

[Mas15b] Hiroki Masuda. Lévy Matters IV: Estimation for Discretely Observed Lévy Processes, volume 2128 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LNMAA2. ISBN 3-319-12372-6 (print), 3-319-12373-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 288 (est.) pp. LCCN ???? URL

http://www.springerlink.com/content/978-3-319-12373-8.

Masuda:2015:PEL

[Mas15c] Hiroki Masuda. Parametric estimation of Lévy processes. Lecture Notes in Mathematics, 2128:179–286, 2015. CO-DEN LNMAA2. ISBN 3-319-12372-6 (print), 3-319-12373-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-12373-8_3.pdf.

Mawhin:2013:RPS

[Maw13] Jean Mawhin. Resonance problems for some non-autonomous ordinary differential equations. Lecture Notes in Mathematics, 2065:103–184, 2013. CODEN LNMAA2. ISBN 3-642-32905-5 (print), 3-642-32906-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-32906-7_3/.

Marciniak-Czochra:2015:RDO

[MC15] Anna Marciniak-Czochra. Reaction-diffusion-ODE models of pattern formation. Lecture Notes in Mathematics, 2126:387–438, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11322-7_8.pdf.

Meckes:2012:PPD

[Mec12] Elizabeth Meckes. Projections of probability distributions: a measure-theoretic Dvoretzky theorem. Lecture Notes in Mathematics, 2050:317–326, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29849-3_18/.

Mennucci:2013:BM

[Men13a] Andrea C. G. Mennucci. Back matter. Lecture Notes in Mathematics, 2090:321–322, 2013. CODEN LNMAA2. ISBN 3-319-01711-X (print), 3-319-01712-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-319-01712-9/1.pdf.

Mennucci:2013:MCS

[Men13b] Andrea C. G. Mennucci. Metrics of curves in shape optimization and analysis. Lecture Notes in Mathematics, 2090:205– 319, 2013. CODEN LNMAA2. ISBN 3-319-01711-X (print), 3-319-01712-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01712-9_4/.

Meyer:2011:UCT

[Mey11] Ralf Meyer. Universal coefficient theorems and assembly maps in KK-theory. Lecture Notes in Mathematics, 2008:45–102, 2011. CODEN LNMAA2. ISBN 3-642-15707-6 (print), 3-642-15708-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15708-0_2/.

Montaner:2013:BM

[MFR13] Josep Àlvarez Montaner and Oscar Fernández-Ramos. Back matter. Lecture Notes in Mathematics, 2083:187–196, 2013. CODEN LNMAA2. ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-642-38742-5/
1.pdf.

Mielke:2011:DEM

[Mie11] Alexander Mielke. Differential, energetic, and metric formulations for rate-independent processes. Lecture Notes in Mathematics, 2028:87–170, 2011. CODEN LNMAA2. ISBN 3-642-21718-4 (print), 3-642-21861-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21861-3_3/.

Malchow:2013:CCB

[MJB13] Horst Malchow, Alex James, and Richard Brown. Control of competitive bioinvasion. Lecture Notes in Mathematics, 2071: 293–305, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35497-7_10/.

Mokhtar-Kharroubi:2015:STN

[MK15] Mustapha Mokhtar-Kharroubi. Spectral theory for neutron transport. Lecture Notes in Mathematics, 2126:319–386, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11322-7_7.pdf.

Miclo:2013:ESM

[MM13a] Laurent Miclo and Pierre Monmarché. Étude spectrale minutieuse de processus moins indécis que les autres. (French) [A careful spectral study of processes less undecided than others]. Lecture Notes in Mathematics, 2078:459–481, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-00321-4_18/.

Mitrea:2013:BM

[MM13b] Irina Mitrea and Marius Mitrea. Back matter. Lecture Notes in Mathematics, 2063:405–424, 2013. CODEN LNMAA2. ISBN 3-642-32665-X (print), 3-642-32666-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-32666-0/1.pdf.

Mitrea:2013:DML

[MM13c] Irina Mitrea and Marius Mitrea. The double multi-layer potential operator. Lecture Notes in Mathematics, 2063:199–252, 2013. CODEN LNMAA2. ISBN 3-642-32665-X (print), 3-642-32666-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32666-0_4/.

Mitrea:2013:FM

[MM13d] Irina Mitrea and Marius Mitrea. Front matter. Lecture Notes in Mathematics, 2063:i–x, 2013. CODEN LNMAA2. ISBN 3-642-32665-X (print), 3-642-32666-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-32666-0/1.pdf.

Mitrea:2013:FSW

[MM13e] Irina Mitrea and Marius Mitrea. Function spaces of Whitney arrays. Lecture Notes in Mathematics, 2063:125–197, 2013. CODEN LNMAA2. ISBN 3-642-32665-X (print), 3-642-32666-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32666-0_3/.

Mitrea:2013:FAP

[MM13f] Irina Mitrea and Marius Mitrea. Functional analytic properties of multi-layer potentials and boundary value problems. Lecture Notes in Mathematics, 2063:293–403, 2013. CODEN LNMAA2. ISBN 3-642-32665-X (print), 3-642-32666-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32666-0_6/.

Mitrea:2013:I

[MM13g] Irina Mitrea and Marius Mitrea. Introduction. Lecture Notes in Mathematics, 2063:1–19, 2013. CODEN LNMAA2. ISBN 3-642-32665-X (print), 3-642-32666-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32666-0_1/.

Mitrea:2013:MLP

[MM13h] Irina Mitrea and Marius Mitrea. Multi-Layer Potentials and Boundary Problems: for Higher-Order Elliptic Systems in Lipschitz Domains, volume 2063 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-32665-X (print), 3-642-32666-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 424 pp. LCCN QA3 .L28 no. 2063.

URL http://link.springer.com/book/10.1007/978-3-642-
32666-0;   http://www.springerlink.com/content/978-3-
642-32666-0.

Mitrea:2013:SML

[MM13i] Irina Mitrea and Marius Mitrea. The single multi-layer potential operator. Lecture Notes in Mathematics, 2063:253–291, 2013. CODEN LNMAA2. ISBN 3-642-32665-X (print), 3-642-32666-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32666-0_5/.

Mitrea:2013:SSC

[MM13j] Irina Mitrea and Marius Mitrea. Smoothness scales and Calderón–Zygmund theory in the scalar-valued case. Lecture Notes in Mathematics, 2063:21–124, 2013. CODEN LNMAA2. ISBN 3-642-32665-X (print), 3-642-32666-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32666-0_2/.

Matsumoto:2011:BM

[MMA11a] Yukio Matsumoto and José María Montesinos-Amilibia. Back matter. Lecture Notes in Mathematics, 2030:221–238, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/bbm:978-3-
642-22534-5/1.pdf.

Matsumoto:2011:BTO

[MMA11b] Yukio Matsumoto and José María Montesinos-Amilibia. Blowing down is a topological operation. Lecture Notes in Mathematics, 2030:189–198, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-22534-5_8/.

Matsumoto:2011:CI

[MMA11c] Yukio Matsumoto and José María Montesinos-Amilibia. Conjugacy invariants. Lecture Notes in Mathematics, 2030:145– 169, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22534-5_6/.

Matsumoto:2011:FMa

[MMA11d] Yukio Matsumoto and José María Montesinos-Amilibia. Front matter. Lecture Notes in Mathematics, 2030:i–xvi, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-22534-5/
1.pdf.

Matsumoto:2011:FMb

[MMA11e] Yukio Matsumoto and José María Montesinos-Amilibia. Front matter. Lecture Notes in Mathematics, 2030:1, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-22534-5/
1/1.pdf.

Matsumoto:2011:FMc

[MMA11f] Yukio Matsumoto and José María Montesinos-Amilibia. Front matter. Lecture Notes in Mathematics, 2030:171, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-22534-5/
2/1.pdf.

Matsumoto:2011:GQ

[MMA11g] Yukio Matsumoto and José María Montesinos-Amilibia. Generalized quotient. Lecture Notes in Mathematics, 2030:53–92, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22534-5_3/.

Matsumoto:2011:PPMa

[MMA11h] Yukio Matsumoto and José María Montesinos-Amilibia. Pseudoperiodic maps. Lecture Notes in Mathematics, 2030:3–15, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22534-5_1/.

Matsumoto:2011:PPMb

[MMA11i] Yukio Matsumoto and José María Montesinos-Amilibia. Pseudoperiodic Maps and Degeneration of Riemann Surfaces, volume 2030 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 238 pp. LCCN QA3 .L28 no. 2030. URL http://link.

springer.com/book/10.1007/978-3-642-22534-5;   http:/
/www.springerlink.com/content/978-3-642-22534-5.

Matsumoto:2011:SOB

[MMA11j] Yukio Matsumoto and José María Montesinos-Amilibia. Singular open-book. Lecture Notes in Mathematics, 2030:199–220, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22534-5_9/.

Matsumoto:2011:SF

[MMA11k] Yukio Matsumoto and José María Montesinos-Amilibia. Standard form. Lecture Notes in Mathematics, 2030:17–52, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-22534-5_
2/.

Matsumoto:2011:TEN

[MMA11l] Yukio Matsumoto and José María Montesinos-Amilibia. A theorem in elementary number theory. Lecture Notes in Mathematics, 2030:131–144, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-22534-5_5/.

Matsumoto:2011:TM

[MMA11m] Yukio Matsumoto and José María Montesinos-Amilibia. Topological monodromy. Lecture Notes in Mathematics, 2030:173– 188, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22534-5_7/.

Matsumoto:2011:UMQ

[MMA11n] Yukio Matsumoto and José María Montesinos-Amilibia. Uniqueness of minimal quotient. Lecture Notes in Mathematics, 2030: 93–129, 2011. CODEN LNMAA2. ISBN 3-642-22533-0 (print), 3-642-22534-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-22534-5_4/.

Mazya:2013:BM

[MMN13a] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Back matter. Lecture Notes in Mathematics, 2077:249–260, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-00357-3/1.pdf.

Mazya:2013:FMa

[MMN13b] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Front matter. Lecture Notes in Mathematics, 2077:i–xvii, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-00357-3/1.pdf.

Mazya:2013:FMb

[MMN13c] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Front matter. Lecture Notes in Mathematics, 2077:1, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-00357-3/
1/1.pdf.

Mazya:2013:FMc

[MMN13d] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Front matter. Lecture Notes in Mathematics, 2077:95, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-00357-3/
2/1.pdf.

Mazya:2013:FMd

[MMN13e] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Front matter. Lecture Notes in Mathematics, 2077:189, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-00357-3/
3/1.pdf.

Mazya:2013:GFD

[MMN13f] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Green’s function for the Dirichlet boundary value problem in a domain with several inclusions. Lecture Notes in Mathematics, 2077:59–73, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00357-3_3/.

Mazya:2013:GKM

[MMN13g] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Green’s Kernels and Meso-Scale Approximations in Perforated Domains, volume 2077 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xviii + 258 pp. LCCN QA3 .L28 no. 2077.

URL http://link.springer.com/book/10.1007/978-3-319-
00357-3;   http://www.springerlink.com/content/978-3-
319-00357-3.

Mazya:2013:GTD

[MMN13h] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Green’s tensor for the Dirichlet boundary value problem in a domain with a single inclusion. Lecture Notes in Mathematics, 2077: 97–137, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00357-3_6/.

Mazya:2013:GTM

[MMN13i] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Green’s tensor for the mixed boundary value problem in a domain with a small hole. Lecture Notes in Mathematics, 2077: 169–188, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00357-3_8/.

Mazya:2013:GTB

[MMN13j] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Green’s tensor in bodies with multiple rigid inclusions. Lecture Notes in Mathematics, 2077:139–167, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-00357-3_7/.

Mazya:2013:MSA

[MMN13k] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Meso-scale approximations for solutions of Dirichlet problems. Lecture Notes in Mathematics, 2077:191–219, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-00357-3_
9/.

Mazya:2013:MNC

[MMN13l] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Mixed and Neumann boundary conditions for domains with small holes and inclusions: Uniform asymptotics of Green’s kernels. Lecture Notes in Mathematics, 2077:21–57, 2013. CODEN LN-MAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-00357-3_
2/.

Mazya:2013:MBV

[MMN13m] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Mixed boundary value problems in multiply-perforated domains. Lecture Notes in Mathematics, 2077:221–247, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-00357-3_
10/.

Mazya:2013:NSB

[MMN13n] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Numerical simulations based on the asymptotic approximations. Lecture Notes in Mathematics, 2077:75–81, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-00357-3_4/.

Mazya:2013:OEA

[MMN13o] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Other examples of asymptotic approximations of Green’s functions in singularly perturbed domains. Lecture Notes in Mathematics, 2077:83–94, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00357-3_5/.

Mazya:2013:UAF

[MMN13p] Vladimir Maz’ya, Alexander Movchan, and Michael Nieves. Uniform asymptotic formulae for Green’s functions for the Laplacian in domains with small perforations. Lecture Notes in Mathematics, 2077:3–19, 2013. CODEN LNMAA2. ISBN 3-319-00356-9 (print), 3-319-00357-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00357-3_1/.

Mijatovic:2012:MPG

[MNU12] Aleksandar Mijatović, Nika Novak, and Mikhail Urusov. Martingale property of generalized stochastic exponentials. Lecture Notes in Mathematics, 2046:41–59, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27461-9_2/.

Molchanov:2013:FSG

[Mol13] Ilya Molchanov. Foundations of stochastic geometry and theory of random sets. Lecture Notes in Mathematics, 2068:1–20, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-33305-7_1/;   http://link.springer.
com/content/pdf/bfm:978-3-642-33305-7/1.pdf.

Montaner:2013:LCM

[Mon13] Josep Àlvarez Montaner. Local cohomology modules supported on monomial ideals. Lecture Notes in Mathematics, 2083:109– 178, 2013. CODEN LNMAA2. ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-38742-5_5/.

Montanari:2014:BM

[Mon14a] Annamaria Montanari. Back matter. Lecture Notes in Mathematics, 2087:209–212, 2014. CODEN LNMAA2. ISBN 3-319-00941-9 (print), 3-319-00942-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-00942-1/1.pdf.

Montanari:2014:LMA

[Mon14b] Annamaria Montanari. On the Levi Monge–Ampére equation. Lecture Notes in Mathematics, 2087:151–208, 2014. CODEN LNMAA2. ISBN 3-319-00941-9 (print), 3-319-00942-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-00942-1_
4/.

Morel:2012:ATF

[Mor12a] Fabien Morel. A1-Algebraic Topology over a Field, volume 2052 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CO-DEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 259 pp. LCCN QA3 .L28 no. 2052. URL http://link.

springer.com/book/10.1007/978-3-642-29514-0;   http:/
/www.springerlink.com/content/978-3-642-29514-0.

Morel:2012:C

[Mor12b] Fabien Morel. A^{1}-coverings, π_{1}^{A^{1}}(P^{n}) and π_{1}^{A^{1}}(SL_{n}). Lecture Notes in Mathematics, 2052:177–197, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29514-0_7/.

Morel:2012:HAV

[Mor12c] Fabien Morel. A^{1}-homotopy and algebraic vector bundles. Lecture Notes in Mathematics, 2052:199–207, 2012. CODEN LN-MAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29514-0_
8/.

Morel:2012:HSH

[Mor12d] Fabien Morel. A^{1}-homotopy sheaves and A^{1}-homology sheaves. Lecture Notes in Mathematics, 2052:149–175, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29514-0_

Morel:2012:ABG

[Mor12e] Fabien Morel. The affine B.G. property for the linear groups and the Grassmannian. Lecture Notes in Mathematics, 2052:209– 226, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29514-0_9/.

Morel:2012:BM

[Mor12f] Fabien Morel. Back matter. Lecture Notes in Mathematics, 2052: 227–259, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-29514-0/1.pdf.

Morel:2012:FM

[Mor12g] Fabien Morel. Front matter. Lecture Notes in Mathematics, 2052: i–x, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-29514-0/1.pdf.

Morel:2012:GVC

[Mor12h] Fabien Morel. Geometric versus canonical transfers. Lecture Notes in Mathematics, 2052:81–112, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29514-0_4/.

Morel:2012:I

[Mor12i] Fabien Morel. Introduction. Lecture Notes in Mathematics, 2052: 1–13, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29514-0_1/.

Morel:2012:RSC

[Mor12j] Fabien Morel. The Rost–Schmid complex of a strongly A^{1}- invariant sheaf. Lecture Notes in Mathematics, 2052:113–148, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29514-0_5/.

Morel:2012:UMW

[Mor12k] Fabien Morel. Unramified Milnor–Witt K-theories. Lecture Notes in Mathematics, 2052:49–80, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29514-0_3/.

Morel:2012:USS

[Mor12l] Fabien Morel. Unramified sheaves and strongly A^{1}-invariant sheaves. Lecture Notes in Mathematics, 2052:15–48, 2012. CODEN LNMAA2. ISBN 3-642-29513-4 (print), 3-642-29514-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
642-29514-0_2/.

Morozov:2013:FM

[Mor13a] Andrew Yu. Morozov. Front matter. Lecture Notes in Mathematics, 2071:261, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-35497-7/3/1.pdf.

Morozov:2013:ICF

[Mor13b] Andrew Yu. Morozov. Incorporating complex foraging of zooplankton in models: Role of micro- and mesoscale processes in macroscale patterns. Lecture Notes in Mathematics, 2071:223– 259, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35497-7_8/.

Mourrat:2012:DPR

[Mou12] Jean-Christophe Mourrat. On the delocalized phase of the random pinning model. Lecture Notes in Mathematics, 2046:401– 407, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27461-9_18/.

Marinelli:2014:MIP

[MR14] Carlo Marinelli and Michael Röckner. On maximal inequalities for purely discontinuous martingales in infinite dimensions. Lecture Notes in Mathematics, 2123:293–315, 2014. CO-DEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_10.pdf.

Movshev:2011:MS

[MS11] M. V. Movshev and A. Schwarz. Maximal supersymmetry. Lecture Notes in Mathematics, 2027:175–193, 2011. CODEN LN-MAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21744-9_
9/.

Marty:2012:ABO

[MS12] Renaud Marty and Knut Sølna. Asymptotic behavior of oscillatory fractional processes. Lecture Notes in Mathematics, 2046: 247–269, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27461-9_12/.

Manni:2016:SNS

[MS16] Carla Manni and Hendrik Speleers. Standard and non-standard CAGD tools for isogeometric analysis: a tutorial. Lecture Notes in Mathematics, 2161:1–69, 2016. CODEN LNMAA2. ISBN 3-319-42308-8 (print), 3-319-42309-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-42309-8_1.pdf.

Mayer:2011:BM

[MSU11a] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. Back matter. Lecture Notes in Mathematics, 2036:109–112, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-23650-1/1.pdf.

Mayer:2011:CER

[MSU11b] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. Classical expanding random systems. Lecture Notes in Mathematics, 2036:75–91, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-23650-1_8/.

Mayer:2011:EM

[MSU11c] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. Expanding in the mean. Lecture Notes in Mathematics, 2036:69– 74, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23650-1_7/.

Mayer:2011:ERM

[MSU11d] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. Expanding random maps. Lecture Notes in Mathematics, 2036:5– 15, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23650-1_2/.

Mayer:2011:FSC

[MSU11e] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. Fractal structure of conformal expanding random repellers. Lecture Notes in Mathematics, 2036:47–56, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-23650-1_5/.

Mayer:2011:I

[MSU11f] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. Introduction. Lecture Notes in Mathematics, 2036:1–4, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-23650-1_
1/.

Mayer:2011:MPG

[MSU11g] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. Measurability, pressure and Gibbs condition. Lecture Notes in Mathematics, 2036:39–45, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-23650-1_4/.

Mayer:2011:MA

[MSU11h] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. Multifractal analysis. Lecture Notes in Mathematics, 2036:57–68, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23650-1_6/.

Mayer:2011:RAP

[MSU11i] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. Real analyticity of pressure. Lecture Notes in Mathematics, 2036:93– 108, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23650-1_9/.

Mayer:2011:RT

[MSU11j] Volker Mayer, Bartlomiej Skorulski, and Mariusz Urbanski. The RPF-theorem. Lecture Notes in Mathematics, 2036:17– 38, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23650-1_3/.

Macpherson:2014:MTA

[MT14] Dugald Macpherson and Carlo Toffalori. Model theory in algebra, analysis and arithmetic: a preface. Lecture Notes in Mathematics, 2111:1–11, 2014. CODEN LNMAA2. ISBN 3-642-54935-7 (print), 3-642-54936-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-54936-6_1.pdf.

Mayer:2011:DER

[MUS11a] Volker Mayer, Mariusz Urbanski, and Bartlomiej Skorulski. Distance Expanding Random Mappings, Thermodynamical Formalism, Gibbs Measures and Fractal Geometry, volume 2036 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LN-MAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 112 pp. LCCN QA3 .L28 no. 2036; QK911 .D57 2013. URL http:

//link.springer.com/book/10.1007/978-3-642-23650-1;
http://www.springerlink.com/content/978-3-642-23650-1.

Mayer:2011:FM

[MUS11b] Volker Mayer, Mariusz Urbanski, and Bartlomiej Skorulski. Front matter. Lecture Notes in Mathematics, 2036:i–x, 2011. CODEN LNMAA2. ISBN 3-642-23649-9 (print), 3-642-23650-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-23650-1/
1.pdf.

Marubayashi:2012:BM

[MV12a] Hidetoshi Marubayashi and Fred Van Oystaeyen. Back matter. Lecture Notes in Mathematics, 2059:213–218, 2012. CODEN LNMAA2. ISBN 3-642-31151-2 (print), 3-642-31152-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-642-31152-9/
1.pdf.

Marubayashi:2012:EVQ

[MV12b] Hidetoshi Marubayashi and Fred Van Oystaeyen. Extensions of valuations to quantized algebras. Lecture Notes in Mathematics, 2059:175–211, 2012. CODEN LNMAA2. ISBN 3-642-31151-2 (print), 3-642-31152-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-31152-9_3/.

Marubayashi:2012:FM

[MV12c] Hidetoshi Marubayashi and Fred Van Oystaeyen. Front matter. Lecture Notes in Mathematics, 2059:i–ix, 2012. CODEN LNMAA2. ISBN 3-642-31151-2 (print), 3-642-31152-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-31152-9/
1.pdf.

Marubayashi:2012:GTP

[MV12d] Hidetoshi Marubayashi and Fred Van Oystaeyen. General theory of primes. Lecture Notes in Mathematics, 2059:1–107, 2012. CODEN LNMAA2. ISBN 3-642-31151-2 (print), 3-642-31152-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31152-9_1/.

Marubayashi:2012:MOP

[MV12e] Hidetoshi Marubayashi and Fred Van Oystaeyen. Maximal orders and primes. Lecture Notes in Mathematics, 2059:109– 173, 2012. CODEN LNMAA2. ISBN 3-642-31151-2 (print), 3-642-31152-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31152-9_2/.

Marubayashi:2012:PDN

[MV12f] Hidetoshi Marubayashi and Fred Van Oystaeyen. Prime Divisors and Noncommutative Valuation Theory, volume 2059 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-31151-2 (print), 3-642-31152-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 218 pp. LCCN

QA3 .L28 no. 2059. URL http://link.springer.com/book/
10.1007/978-3-642-31152-9;   http://www.springerlink.
com/content/978-3-642-31152-9.

Mocha:2012:QSB

[MW12] Markus Mocha and Nicholas Westray. Quadratic semimartingale BSDEs under an exponential moments condition. Lecture Notes in Mathematics, 2046:105–139, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27461-9_5/.

Matsumoto:2011:ZOL

[MY11] Ayako Matsumoto and Kouji Yano. On a zero-one law for the norm process of transient random walk. Lecture Notes in Mathematics, 2006:105–126, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-15217-7_4/.

Naasell:2011:ASI

[Nås11a] Ingemar Nåsell. Approximation of some images under Ψ for the SIS model. Lecture Notes in Mathematics, 2022:119–139, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-20530-9_10/.

Naasell:2011:AQS

[Nås11b] Ingemar Nåsell. Approximation of the quasi-stationary distribution q of the SIS model. Lecture Notes in Mathematics, 2022:

141–147, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-20530-9_11/.

Naasell:2011:ASDb

[Nås11c] Ingemar Nåsell. Approximation of the stationary distribution p^{(0)} of the SIS model. Lecture Notes in Mathematics, 2022:115– 118, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-20530-9_9/.

Naasell:2011:ASDa

[Nås11d] Ingemar Nåsell. Approximation of the stationary distribution p^{(1)} of the SIS model. Lecture Notes in Mathematics, 2022:93– 99, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-20530-9_7/.

Naasell:2011:ATE

[Nås11e] Ingemar Nåsell. Approximation of the time to extinction for the SIS model. Lecture Notes in Mathematics, 2022:149–154, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-20530-9_12/.

Naasell:2011:BM

[Nås11f] Ingemar Nåsell. Back matter. Lecture Notes in Mathematics, 2022:183–199, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-20530-9/1.pdf.

Naasell:2011:CC

[Nås11g] Ingemar Nåsell. Concluding comments. Lecture Notes in Mathematics, 2022:177–182, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-20530-9_15/.

Naasell:2011:EQS

[Nås11h] Ingemar Nåsell. Extinction and Quasi-Stationarity in the Stochastic Logistic SIS Model, volume 2022 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 199 pp. LCCN QA3 .L28 no. 2022; QH352 .N37 2011; QA3 .L28 no. 2022. URL http:

//link.springer.com/book/10.1007/978-3-642-20530-9;
http://www.springerlink.com/content/978-3-642-20530-9.

Naasell:2011:FM

[Nås11i] Ingemar Nåsell. Front matter. Lecture Notes in Mathematics, 2022:i–xi, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-20530-9/1.pdf.

Naasell:2011:I

[Nås11j] Ingemar Nåsell. Introduction. Lecture Notes in Mathematics, 2022:1–7, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-20530-9_1/.

Naasell:2011:MF

[Nås11k] Ingemar Nåsell. Model formulation. Lecture Notes in Mathematics, 2022:9–16, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-20530-9_2/.

Naasell:2011:PSSb

[Nås11l] Ingemar Nåsell. Preparations for the study of the stationary distribution p^{(0)} of the SIS model. Lecture Notes in Mathematics, 2022:101–113, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-20530-9_8/.

Naasell:2011:PSSa

[Nås11m] Ingemar Nåsell. Preparations for the study of the stationary distribution p^{(1)} of the SIS model. Lecture Notes in Mathematics, 2022:73–91, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-20530-9_6/.

Naasell:2011:SMF

[Nås11n] Ingemar Nåsell. The SIS model: First approximations of the quasi-stationary distribution. Lecture Notes in Mathematics, 2022:41–46, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-20530-9_4/.

Naasell:2011:SAI

[Nås11o] Ingemar Nåsell. Some approximations involving the normal distribution. Lecture Notes in Mathematics, 2022:47–72, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-20530-9_5/.

Naasell:2011:SPB

[Nås11p] Ingemar Nåsell. Stochastic process background. Lecture Notes in Mathematics, 2022:17–40, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-20530-9_3/.

Naasell:2011:TSM

[Nås11q] Ingemar Nåsell. Thresholds for the SIS model. Lecture Notes in Mathematics, 2022:171–175, 2011. CODEN LNMAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-20530-9_14/.

Naasell:2011:UAS

[Nås11r] Ingemar Nåsell. Uniform approximations for the SIS model. Lecture Notes in Mathematics, 2022:155–170, 2011. CODEN LN-MAA2. ISBN 3-642-20529-1 (print), 3-642-20530-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-20530-9_

Nazarov:2012:HPB

[Naz12] Fedor Nazarov. The Hörmander proof of the Bourgain– Milman theorem. Lecture Notes in Mathematics, 2050:335– 343, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29849-3_20/.

Neuberger:2010:AAT

[Neu10a] J. W. Neuberger. An alternate approach to time-dependent PDEs. Lecture Notes in Mathematics, 1670:153–158, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_18.pdf.

Neuberger:2010:BSC

[Neu10b] J. W. Neuberger. Boundary and supplementary conditions. Lecture Notes in Mathematics, 1670:63–78, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-04041-2_8.
pdf.

Neuberger:2010:CTG

[Neu10c] J. W. Neuberger. Comparison of two gradients. Lecture Notes in Mathematics, 1670:5–13, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04041-2_2.pdf.

Neuberger:2010:CNM

[Neu10d] J. W. Neuberger. Continuous Newton’s method. Lecture Notes in Mathematics, 1670:79–83, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04041-2_9.pdf.

Neuberger:2010:CSDa

[Neu10e] J. W. Neuberger. Continuous steepest descent in Hilbert space: Linear case. Lecture Notes in Mathematics, 1670:15–17, 2010.

CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_3.pdf.

Neuberger:2010:CSDb

[Neu10f] J. W. Neuberger. Continuous steepest descent in Hilbert space: Nonlinear case. Lecture Notes in Mathematics, 1670:19–34, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_4.pdf.

Neuberger:2010:CGI

[Neu10g] J. W. Neuberger. Convexity and gradient inequalities. Lecture Notes in Mathematics, 1670:57–61, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-04041-2_7.
pdf.

Neuberger:2010:FPN

[Neu10h] J. W. Neuberger. Flow problems and non-inner product Sobolev spaces. Lecture Notes in Mathematics, 1670:147–152, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_17.pdf.

Neuberger:2010:FSCa

[Neu10i] J. W. Neuberger. Foliations and supplementary conditions I. Lecture Notes in Mathematics, 1670:159–169, 2010. CO-DEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_19.pdf.

Neuberger:2010:FSCb

[Neu10j] J. W. Neuberger. Foliations and supplementary conditions II. Lecture Notes in Mathematics, 1670:171–175, 2010. CO-DEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_20.pdf.

Neuberger:2010:FM

[Neu10k] J. W. Neuberger. Front matter. Lecture Notes in Mathematics, 1670:i–xiii, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-04041-2/1.pdf.

Neuberger:2010:ISG

[Neu10l] J. W. Neuberger. An introduction to Sobolev gradients in noninner product spaces. Lecture Notes in Mathematics, 1670:103– 107, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_12.pdf.

Neuberger:2010:MS

[Neu10m] J. W. Neuberger. Minimal surfaces. Lecture Notes in Mathematics, 1670:129–145, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-04041-2_16.pdf.

Neuberger:2010:MAF

[Neu10n] J. W. Neuberger. More about finite differences. Lecture Notes in Mathematics, 1670:85–97, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04041-2_10.pdf.

Neuberger:2010:ODE

[Neu10o] J. W. Neuberger. Ordinary differential equations and Sobolev gradients. Lecture Notes in Mathematics, 1670:53–55, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_6.pdf.

Neuberger:2010:OPA

[Neu10p] J. W. Neuberger. Orthogonal projections, adjoints and Laplacians. Lecture Notes in Mathematics, 1670:35–51, 2010. CO-DEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_5.pdf.

Neuberger:2010:SG

[Neu10q] J. W. Neuberger. Several gradients. Lecture Notes in Mathematics, 1670:1–4, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-04041-2_1.pdf.

Neuberger:2010:SSG

[Neu10r] J. W. Neuberger. Singularities and a simple Ginzburg–Landau functional. Lecture Notes in Mathematics, 1670:109–111, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_13.pdf.

Neuberger:2010:SGD

[Neu10s] J. W. Neuberger. Sobolev Gradients and Differential Equations, volume 1670 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 289 pp. LCCN QA3 .L28 no. 1670. URL http://link.

springer.com/book/10.1007/978-3-642-04041-2;   http:/
/www.springerlink.com/content/978-3-642-04041-2.

Neuberger:2010:SGV

[Neu10t] J. W. Neuberger. Sobolev gradients for variational problems. Lecture Notes in Mathematics, 1670:99–102, 2010. CO-DEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_11.pdf.

Neuberger:2010:SEG

[Neu10u] J. W. Neuberger. The superconductivity equations of Ginzburg– Landau. Lecture Notes in Mathematics, 1670:113–121, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-04041-2_14.pdf.

Neuberger:2010:TEC

[Neu10v] J. W. Neuberger. Tricomi equation: a case study. Lecture Notes in Mathematics, 1670:123–127, 2010. CODEN LNMAA2. ISBN 3-642-04040-3 (print), 3-642-04041-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-04041-2_15.pdf.

Nishitani:2014:BM

[Nis14a] Tatsuo Nishitani. Back matter. Lecture Notes in Mathematics, 2097:231–240, 2014. CODEN LNMAA2. ISBN 3-319-02272-5 (print), 3-319-02273-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-02273-4/1.pdf.

Nishitani:2014:FM

[Nis14b] Tatsuo Nishitani. Front matter. Lecture Notes in Mathematics, 2097:i–viii, 2014. CODEN LNMAA2. ISBN 3-319-02272-5 (print), 3-319-02273-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-02273-4/1.pdf.

Nishitani:2014:HSA

[Nis14c] Tatsuo Nishitani. Hyperbolic Systems with Analytic Coefficients: Well-posedness of the Cauchy Problem, volume 2097 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-02272-5 (print), 3-319-02273-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 237 pp. LCCN

QA3 .L28 no. 2097. URL http://link.springer.com/book/
10.1007/978-3-319-02273-4;   http://www.springerlink.
com/content/978-3-319-02273-4.

Nishitani:2014:I

[Nis14d] Tatsuo Nishitani. Introduction. Lecture Notes in Mathematics, 2097:1–29, 2014. CODEN LNMAA2. ISBN 3-319-02272-5 (print), 3-319-02273-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02273-4_1/.

Nishitani:2014:NCS

[Nis14e] Tatsuo Nishitani. Necessary conditions for strong hyperbolicity. Lecture Notes in Mathematics, 2097:31–84, 2014. CODEN LNMAA2. ISBN 3-319-02272-5 (print), 3-319-02273-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02273-4_
2/.

Nishitani:2014:SNC

[Nis14f] Tatsuo Nishitani. Systems with nondegenerate characteristics. Lecture Notes in Mathematics, 2097:161–229, 2014. CODEN LNMAA2. ISBN 3-319-02272-5 (print), 3-319-02273-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-02273-4_
4/.

Nishitani:2014:TTS

[Nis14g] Tatsuo Nishitani. Two by two systems with two independent variables. Lecture Notes in Mathematics, 2097:85–160, 2014. CODEN LNMAA2. ISBN 3-319-02272-5 (print), 3-319-02273-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-02273-4_3/.

Novo:2013:NAF

[NO13] Sylvia Novo and Rafael Obaya. Non-autonomous functional differential equations and applications. Lecture Notes in Mathematics, 2065:185–263, 2013. CODEN LNMAA2. ISBN 3-642-32905-5 (print), 3-642-32906-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-32906-7_4/.

Noreddine:2013:FTC

[Nor13] Salim Noreddine. Fluctuations of the traces of complex-valued random matrices. Lecture Notes in Mathematics, 2078:401– 431, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00321-4_16/.

Nourdin:2013:FM

[Nou13a] Ivan Nourdin. Front matter. Lecture Notes in Mathematics, 2078: 91, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-00321-4/2/1.pdf.

Nourdin:2013:LGA

[Nou13b] Ivan Nourdin. Lectures on Gaussian approximations with Malliavin calculus. Lecture Notes in Mathematics, 2078:3–89, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
319-00321-4_1/;   http://link.springer.com/content/pdf/
bfm:978-3-319-00321-4/1.pdf;   http://link.springer.com/
content/pdf/bfm:978-3-319-00321-4/1/1.pdf.

Niemeyer:2013:EPR

[NPS13] Alice C. Niemeyer, Cheryl E. Praeger, and Ákos Seress. Estimation problems and randomised group algorithms. Lecture Notes in Mathematics, 2070:35–82, 2013. CODEN LNMAA2. ISBN 1-4471-4813-4 (print), 1-4471-4814-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-1-4471-4814-2_2/.

Neeb:2011:LSU

[NS11] Karl-Hermann Neeb and Hadi Salmasian. Lie supergroups, unitary representations, and invariant cones. Lecture Notes in Mathematics, 2027:195–239, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21744-9_10/.

Nemethi:2012:AMS

[NS12a] András Némethi and Ágnes Szilárd. The algebraic monodromy of H_{1}(∂F): Starting point. Lecture Notes in Mathematics, 2037: 153–156, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23647-1_14/.

Nemethi:2012:CP

[NS12b] András Némethi and Ágnes Szilárd. The characteristic polynomial of ∂F via P^{♯} and P_{j}^{♯}. Lecture Notes in Mathematics, 2037: 161–166, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23647-1_16/.

Nemethi:2012:CMA

[NS12c] András Némethi and Ágnes Szilárd. The collapsing main algorithm. Lecture Notes in Mathematics, 2037:131–138, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
642-23647-1_12/.

Nemethi:2012:CCG

[NS12d] András Némethi and Ágnes Szilárd. Cyclic coverings of graphs. Lecture Notes in Mathematics, 2037:45–54, 2012. CODEN LN-MAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-23647-1_
5/.

Nemethi:2012:CPC

[NS12e] András Némethi and Ágnes Szilárd. Cylinders of plane curve singularities: f = f′(x, y). Lecture Notes in Mathematics, 2037: 201–204, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23647-1_20/.

Nemethi:2012:EFA

[NS12f] András Némethi and Ágnes Szilárd. Examples: Families associated with plane curve singularities. Lecture Notes in Mathematics, 2037:83–97, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-23647-1_9/.

Nemethi:2012:EHS

[NS12g] András Némethi and Ágnes Szilárd. Examples: Homogeneous singularities. Lecture Notes in Mathematics, 2037:79– 82, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23647-1_8/.

Nemethi:2012:FMa

[NS12h] András Némethi and Ágnes Szilárd. Front matter. Lecture Notes in Mathematics, 2037:i–xii, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-23647-1/1.pdf.

Nemethi:2012:FMb

[NS12i] András Némethi and Ágnes Szilárd. Front matter. Lecture Notes in Mathematics, 2037:9, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-23647-1/1/1.pdf.

Nemethi:2012:FMc

[NS12j] András Némethi and Ágnes Szilárd. Front matter. Lecture Notes in Mathematics, 2037:99, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-23647-1/2/1.pdf.

Nemethi:2012:FMd

[NS12k] András Némethi and Ágnes Szilárd. Front matter. Lecture Notes in Mathematics, 2037:177, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-642-23647-1/3/1.pdf.

Nemethi:2012:GPD

[NS12l] András Némethi and Ágnes Szilárd. The graph Γ_{C} of a pair (f, g): The definition. Lecture Notes in Mathematics, 2037:55– 61, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23647-1_6/.

Nemethi:2012:GP

[NS12m] András Némethi and Ágnes Szilárd. The graph Γ_{C}: Properties. Lecture Notes in Mathematics, 2037:63–77, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-23647-1_
7/.

Nemethi:2012:HS

[NS12n] András Némethi and Ágnes Szilárd. Homogeneous singularities. Lecture Notes in Mathematics, 2037:179–199, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-23647-1_
19/.

Nemethi:2012:I

[NS12o] András Némethi and Ágnes Szilárd. Introduction. Lecture Notes in Mathematics, 2037:1–7, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-23647-1_1/.

Nemethi:2012:MA

[NS12p] András Némethi and Ágnes Szilárd. The main algorithm. Lecture Notes in Mathematics, 2037:101–115, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-23647-1_10/.

Nemethi:2012:MFB

[NS12q] András Némethi and Ágnes Szilárd. Milnor Fiber Boundary of a Non-isolated Surface Singularity, volume 2037 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 240 pp. LCCN QA3 .L28 no.

2037. URL http://link.springer.com/book/10.1007/978-
3-642-23647-1;   http://www.springerlink.com/content/
978-3-642-23647-1.

Nemethi:2012:MHS

[NS12r] András Némethi and Ágnes Szilárd. The mixed Hodge structure of H_{1}(∂F). Lecture Notes in Mathematics, 2037:173–176, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23647-1_18/.

Nemethi:2012:PGO

[NS12s] András Némethi and Ágnes Szilárd. Plumbing graphs and oriented plumbed 3-manifolds. Lecture Notes in Mathematics, 2037: 25–43, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23647-1_4/.

Nemethi:2012:PCP

[NS12t] András Némethi and Ágnes Szilárd. The proof of the characteristic polynomial formulae. Lecture Notes in Mathematics, 2037: 167–172, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23647-1_17/.

Nemethi:2012:PMA

[NS12u] András Némethi and Ágnes Szilárd. Proof of the main algorithm. Lecture Notes in Mathematics, 2037:117–130, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-23647-1_
11/.

Nemethi:2012:RP

[NS12v] András Némethi and Ágnes Szilárd. The ranks of H_{1}(∂F) and H_{1}(∂F \ V g) via plumbing. Lecture Notes in Mathematics, 2037: 157–160, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-23647-1_15/.

Nemethi:2012:THG

[NS12w] András Némethi and Ágnes Szilárd. The topology of a hypersurface germ f in three variables. Lecture Notes in Mathematics, 2037:11–15, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-23647-1_2/.

Nemethi:2012:TP

[NS12x] András Némethi and Ágnes Szilárd. The topology of a pair (f, g). Lecture Notes in Mathematics, 2037:17–23, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-23647-1_
3/.

Nemethi:2012:VHM

[NS12y] András Némethi and Ágnes Szilárd. Vertical/horizontal monodromies. Lecture Notes in Mathematics, 2037:139–151, 2012. CODEN LNMAA2. ISBN 3-642-23646-4 (print), 3-642-23647-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
642-23647-1_13/.

Nayar:2012:LWT

[NT12] Piotr Nayar and Tomasz Tkocz. On a Loomis–Whitney type inequality for permutationally invariant unconditional convex bodies. Lecture Notes in Mathematics, 2050:327–333, 2012. CODEN LNMAA2. ISBN 3-642-29848-6 (print), 3-642-29849-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29849-3_19/.

Nochetto:2012:PAF

[NV12] Ricardo H. Nochetto and Andreas Veeser. Primer of adaptive finite element methods. Lecture Notes in Mathematics, 2040: 125–225, 2012. CODEN LNMAA2. ISBN 3-642-24078-X (print), 3-642-24079-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-24079-9_3/.

Ottesen:2013:DPS

[ONO13] Johnny T. Ottesen, Vera Novak, and Mette S. Olufsen. Development of patient specific cardiovascular models predicting dynamics in response to orthostatic stress challenges. Lecture Notes in Mathematics, 2064:177–213, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32882-4_10/.

Ortmann:2013:FBB

[Ort13] Janosch Ortmann. Functionals of the Brownian bridge. Lecture Notes in Mathematics, 2078:433–458, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-00321-4_17/.

Otal:2015:BM

[Ota15a] Jean-Pierre Otal. Back matter. Lecture Notes in Mathematics, 2119:415–416, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-11029-5/1.pdf.

Otal:2015:BSA

[Ota15b] Jean-Pierre Otal. Berkovich Spaces and Applications, volume 2119 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 416

(est.) pp. LCCN ???? URL http://www.springerlink.com/
content/978-3-319-11029-5.

Otal:2015:CSR

[Ota15c] Jean-Pierre Otal. Compactification of spaces of representations after Culler, Morgan and Shalen. Lecture Notes in Mathematics, 2119:367–413, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-11029-5_7.pdf.

Otway:2012:BM

[Otw12a] Thomas H. Otway. Back matter. Lecture Notes in Mathematics, 2043:169–214, 2012. CODEN LNMAA2. ISBN 3-642-24414-9 (print), 3-642-24415-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-24415-5/1.pdf.

Otway:2012:CPM

[Otw12b] Thomas H. Otway. The cold plasma model. Lecture Notes in Mathematics, 2043:87–120, 2012. CODEN LNMAA2. ISBN 3-642-24414-9 (print), 3-642-24415-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24415-5_4/.

Otway:2012:DPE

[Otw12c] Thomas H. Otway. The Dirichlet Problem for Elliptic-Hyperbolic Equations of Keldysh Type, volume 2043 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-24414-9 (print), 3-642-24415-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 214 pp. LCCN QA3 .L28 no.

2043. URL http://link.springer.com/book/10.1007/978-
3-642-24415-5;   http://www.springerlink.com/content/
978-3-642-24415-5.

Otway:2012:ECC

[Otw12d] Thomas H. Otway. The equation of Cinquini–Cibrario. Lecture Notes in Mathematics, 2043:47–85, 2012. CODEN LNMAA2. ISBN 3-642-24414-9 (print), 3-642-24415-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-24415-5_3/.

Otway:2012:FM

[Otw12e] Thomas H. Otway. Front matter. Lecture Notes in Mathematics, 2043:i–ix, 2012. CODEN LNMAA2. ISBN 3-642-24414-9 (print), 3-642-24415-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-24415-5/1.pdf.

Otway:2012:I

[Otw12f] Thomas H. Otway. Introduction. Lecture Notes in Mathematics, 2043:1–11, 2012. CODEN LNMAA2. ISBN 3-642-24414-9 (print), 3-642-24415-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-24415-5_1/.

Otway:2012:LNC

[Otw12g] Thomas H. Otway. Light near a caustic. Lecture Notes in Mathematics, 2043:121–144, 2012. CODEN LNMAA2. ISBN 3-642-24414-9 (print), 3-642-24415-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24415-5_5/.

Otway:2012:MP

[Otw12h] Thomas H. Otway. Mathematical preliminaries. Lecture Notes in Mathematics, 2043:13–45, 2012. CODEN LNMAA2. ISBN 3-642-24414-9 (print), 3-642-24415-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24415-5_2/.

Otway:2012:PG

[Otw12i] Thomas H. Otway. Projective geometry. Lecture Notes in Mathematics, 2043:145–167, 2012. CODEN LNMAA2. ISBN 3-642-24414-9 (print), 3-642-24415-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24415-5_6/.

Othmer:2013:MAB

[OX13] Hans G. Othmer and Chuan Xue. The mathematical analysis of biological aggregation and dispersal: Progress, problems and perspectives. Lecture Notes in Mathematics, 2071:79– 127, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35497-7_4/.

Obukhovskii:2013:BM

[OZLK13a] Valeri Obukhovskii, Pietro Zecca, Nguyen Van Loi, and Sergei Kornev. Back matter. Lecture Notes in Mathematics, 2076:167– 180, 2013. CODEN LNMAA2. ISBN 3-642-37069-1 (print), 3-642-37070-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-37070-0/1.pdf.

Obukhovskii:2013:B

[OZLK13b] Valeri Obukhovskii, Pietro Zecca, Nguyen Van Loi, and Sergei Kornev. Background. Lecture Notes in Mathematics, 2076:1– 24, 2013. CODEN LNMAA2. ISBN 3-642-37069-1 (print), 3-642-37070-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-37070-0_1/.

Obukhovskii:2013:FM

[OZLK13c] Valeri Obukhovskii, Pietro Zecca, Nguyen Van Loi, and Sergei Kornev. Front matter. Lecture Notes in Mathematics, 2076: i–xiii, 2013. CODEN LNMAA2. ISBN 3-642-37069-1 (print), 3-642-37070-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-37070-0/1.pdf.

Obukhovskii:2013:MGFa

[OZLK13d] Valeri Obukhovskii, Pietro Zecca, Nguyen Van Loi, and Sergei Kornev. Method of guiding functions in finite-dimensional spaces. Lecture Notes in Mathematics, 2076:25–67, 2013. CODEN LN-MAA2. ISBN 3-642-37069-1 (print), 3-642-37070-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-37070-0_
2/.

Obukhovskii:2013:MGFb

[OZLK13e] Valeri Obukhovskii, Pietro Zecca, Nguyen Van Loi, and Sergei Kornev. Method of guiding functions in Hilbert spaces. Lecture Notes in Mathematics, 2076:69–104, 2013. CODEN LNMAA2. ISBN 3-642-37069-1 (print), 3-642-37070-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-37070-0_3/.

Obukhovskii:2013:MGFc

[OZLK13f] Valeri Obukhovskii, Pietro Zecca, Nguyen Van Loi, and Sergei Kornev. Method of Guiding Functions in Problems of Nonlinear Analysis, volume 2076 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-37069-1 (print), 3-642-37070-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 177 pp. LCCN QA3 .L28 no. 2076. URL http:

//link.springer.com/book/10.1007/978-3-642-37070-0;
http://www.springerlink.com/content/978-3-642-37070-0.

Obukhovskii:2013:NFI

[OZLK13g] Valeri Obukhovskii, Pietro Zecca, Nguyen Van Loi, and Sergei Kornev. Nonlinear Fredholm inclusions and applications. Lecture Notes in Mathematics, 2076:131–165, 2013. CODEN LNMAA2. ISBN 3-642-37069-1 (print), 3-642-37070-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-37070-0_5/.

Obukhovskii:2013:SOD

[OZLK13h] Valeri Obukhovskii, Pietro Zecca, Nguyen Van Loi, and Sergei Kornev. Second-order differential inclusions. Lecture Notes in Mathematics, 2076:105–129, 2013. CODEN LNMAA2. ISBN 3-642-37069-1 (print), 3-642-37070-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-37070-0_4/.

Padula:2011:ASS

[Pad11a] Mariarosaria Padula. Asymptotic Stability of Steady Compressible Fluids, volume 2024 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-21136-4 (print), 3-642-21137-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 235 pp. LCCN QA3 .L28 no. 2024. URL http:

//link.springer.com/book/10.1007/978-3-642-21137-9;
http://www.springerlink.com/content/978-3-642-21137-9.

Padula:2011:BM

[Pad11b] Mariarosaria Padula. Back matter. Lecture Notes in Mathematics, 2024:223–235, 2011. CODEN LNMAA2. ISBN 3-642-21136-4 (print), 3-642-21137-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-21137-9/1.pdf.

Padula:2011:BFR

[Pad11c] Mariarosaria Padula. Barotropic fluids with rigid boundary. Lecture Notes in Mathematics, 2024:87–132, 2011. CODEN LN-MAA2. ISBN 3-642-21136-4 (print), 3-642-21137-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21137-9_
3/.

Padula:2011:FM

[Pad11d] Mariarosaria Padula. Front matter. Lecture Notes in Mathematics, 2024:i–xiv, 2011. CODEN LNMAA2. ISBN 3-642-21136-4 (print), 3-642-21137-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-21137-9/1.pdf.

Padula:2011:IFF

[Pad11e] Mariarosaria Padula. Isothermal fluids with free boundaries. Lecture Notes in Mathematics, 2024:133–195, 2011. CODEN LN-MAA2. ISBN 3-642-21136-4 (print), 3-642-21137-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21137-9_
4/.

Padula:2011:PFR

[Pad11f] Mariarosaria Padula. Polytropic fluids with rigid boundary. Lecture Notes in Mathematics, 2024:197–221, 2011. CODEN LN-MAA2. ISBN 3-642-21136-4 (print), 3-642-21137-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21137-9_
5/.

Padula:2011:TFM

[Pad11g] Mariarosaria Padula. Topics in fluid mechanics. Lecture Notes in Mathematics, 2024:1–52, 2011. CODEN LNMAA2. ISBN 3-642-21136-4 (print), 3-642-21137-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21137-9_1/.

Padula:2011:TS

[Pad11h] Mariarosaria Padula. Topics in stability. Lecture Notes in Mathematics, 2024:53–86, 2011. CODEN LNMAA2. ISBN 3-642-21136-4 (print), 3-642-21137-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21137-9_2/.

Pages:2013:FCM

[Pag13] Gilles Pagès. Functional co-monotony of processes with applications to peacocks and barrier options. Lecture Notes in Mathematics, 2078:365–400, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-00321-4_15/.

Plank:2013:LAP

[PAMC13] Michael J. Plank, Marie Auger-Méthé, and Edward A. Codling. Lévy or not? Analysing positional data from animal movement paths. Lecture Notes in Mathematics, 2071:33–52, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-35497-7_

Parmeggiani:2010:BM

[Par10a] Alberto Parmeggiani. Back matter. Lecture Notes in Mathematics, 1992:239–260, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-11922-4/1.pdf.

Parmeggiani:2010:ELD

[Par10b] Alberto Parmeggiani. Energy-levels, dynamics, and the Maslov index. Lecture Notes in Mathematics, 1992:161–190, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11922-4_11.pdf.

Parmeggiani:2010:FM

[Par10c] Alberto Parmeggiani. Front matter. Lecture Notes in Mathematics, 1992:i–xi, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-11922-4/1.pdf.

Parmeggiani:2010:HO

[Par10d] Alberto Parmeggiani. The harmonic oscillator. Lecture Notes in Mathematics, 1992:7–13, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11922-4_2.pdf.

Parmeggiani:2010:HSF

[Par10e] Alberto Parmeggiani. The heat-semigroup, functional calculus and kernels. Lecture Notes in Mathematics, 1992:67–77, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11922-4_5.pdf.

Parmeggiani:2010:I

[Par10f] Alberto Parmeggiani. Introduction. Lecture Notes in Mathematics, 1992:1–5, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11922-4_1.pdf.

Parmeggiani:2010:LMS

[Par10g] Alberto Parmeggiani. Localization and multiplicity of a selfadjoint elliptic 2 × 2 positive NCHO in R^{n}. Lecture Notes in Mathematics, 1992:191–238, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11922-4_12.pdf.

Parmeggiani:2010:OIG

[Par10h] Alberto Parmeggiani. On operators induced by general finite-rank orthogonal projections. Lecture Notes in Mathematics, 1992:149– 159, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11922-4_10.pdf.

Parmeggiani:2010:SPE

[Par10i] Alberto Parmeggiani. Some properties of the eigenvalues of Q_{(α,β)}^{w(x,D)}. Lecture Notes in Mathematics, 1992:111–120, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11922-4_8.pdf.

Parmeggiani:2010:STS

[Par10j] Alberto Parmeggiani. Some tools from the semiclassical calculus. Lecture Notes in Mathematics, 1992:121–147, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11922-4_9.pdf.

Parmeggiani:2010:SCFa

[Par10k] Alberto Parmeggiani. The spectral counting function N(λ) and the behavior of the eigenvalues: Part 1. Lecture Notes in Mathematics, 1992:55–66, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11922-4_4.pdf.

Parmeggiani:2010:SCFb

[Par10l] Alberto Parmeggiani. The spectral counting function N(λ) and the behavior of the eigenvalues: Part 2. Lecture Notes in Mathematics, 1992:79–92, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-11922-4_6.pdf.

Parmeggiani:2010:STN

[Par10m] Alberto Parmeggiani. Spectral Theory of Non-Commutative Harmonic Oscillators: an Introduction, volume 1992 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 254 pp. LCCN QA3 .L28 no.

1992. URL http://link.springer.com/book/10.1007/978-
3-642-11922-4;   http://www.springerlink.com/content/
978-3-642-11922-4.

Parmeggiani:2010:SZF

[Par10n] Alberto Parmeggiani. The spectral zeta function. Lecture Notes in Mathematics, 1992:93–110, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11922-4_7.pdf.

Parmeggiani:2010:WHC

[Par10o] Alberto Parmeggiani. The Weyl–Hörmander calculus. Lecture Notes in Mathematics, 1992:15–53, 2010. CODEN LNMAA2. ISBN 3-642-11921-2 (print), 3-642-11922-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-11922-4_3.
pdf.

Patrizio:2013:FM

[PBBD13a] Giorgio Patrizio, Zbigniew Błocki, François Berteloot, and Jean Pierre Demailly. Front matter. Lecture Notes in Mathematics, 2075:i–ix, 2013. CODEN LNMAA2. ISBN 3-642-36420-9 (print), 3-642-36421-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-36421-1/1.pdf.

Patrizio:2013:PTC

[PBBD13b] Giorgio Patrizio, Zbigniew Błocki, François Berteloot, and Jean-Pierre Demailly, editors. Pluripotential Theory: Cetraro, Italy 2011, volume 2075 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-36420-9 (print), 3-642-36421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2075. URL http://link.

springer.com/book/10.1007/978-3-642-36421-1;   http:/
/www.springerlink.com/content/978-3-642-36421-1.

Petrovskaya:2013:BM

[PEP13a] Natalia Petrovskaya, Nina Embleton, and Sergei V. Petrovskii. Back matter. Lecture Notes in Mathematics, 2071:387–388, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-35497-7/1.pdf.

Petrovskaya:2013:NSP

[PEP13b] Natalia Petrovskaya, Nina Embleton, and Sergei V. Petrovskii. Numerical study of pest population size at various diffusion rates. Lecture Notes in Mathematics, 2071:355–385, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35497-7_13/.

Picard:2011:FM

[Pic11a] Jean Picard. Front matter. Lecture Notes in Mathematics, 2006: 71, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-15217-7/2/1.pdf.

Picard:2011:RFF

[Pic11b] Jean Picard. Representation formulae for the fractional Brownian motion. Lecture Notes in Mathematics, 2006:3–70, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-15217-7_
1/;   http://link.springer.com/content/pdf/bfm:978-3-

642-15217-7/1.pdf;   http://link.springer.com/content/
pdf/bfm:978-3-642-15217-7/1/1.pdf.

Pillay:2014:STA

[Pil14] Anand Pillay. Some themes around first order theories without the independence property. Lecture Notes in Mathematics, 2111:13–33, 2014. CODEN LNMAA2. ISBN 3-642-54935-7 (print), 3-642-54936-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-54936-6_2.pdf.

Post:2012:BM

[Pos12a] Olaf Post. Back matter. Lecture Notes in Mathematics, 2039: 389–431, 2012. CODEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-23840-6/1.pdf.

Post:2012:FM

[Pos12b] Olaf Post. Front matter. Lecture Notes in Mathematics, 2039: i–xv, 2012. CODEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-23840-6/1.pdf.

Post:2012:FAPa

[Pos12c] Olaf Post. The functional analytic part: Scales of Hilbert spaces and boundary triples. Lecture Notes in Mathematics, 2039:97– 185, 2012. CODEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-23840-6_3.pdf.

Post:2012:FAPb

[Pos12d] Olaf Post. The functional analytic part: Two operators in different Hilbert spaces. Lecture Notes in Mathematics, 2039:187–257, 2012. CODEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-23840-6_4.pdf.

Post:2012:GCR

[Pos12e] Olaf Post. Global convergence results. Lecture Notes in Mathematics, 2039:367–388, 2012. CODEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-23840-6_7.pdf.

Post:2012:GAL

[Pos12f] Olaf Post. Graphs and associated Laplacians. Lecture Notes in Mathematics, 2039:57–96, 2012. CODEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-23840-6_2.pdf.

Post:2012:I

[Pos12g] Olaf Post. Introduction. Lecture Notes in Mathematics, 2039:1– 56, 2012. CODEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-23840-6_1.pdf.

Post:2012:MTN

[Pos12h] Olaf Post. Manifolds, tubular neighbourhoods and their perturbations. Lecture Notes in Mathematics, 2039:259–289, 2012. CODEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-23840-6_5.pdf.

Post:2012:PSE

[Pos12i] Olaf Post. Plumber’s shop: Estimates for star graphs and related spaces. Lecture Notes in Mathematics, 2039:291–366, 2012. CODEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-23840-6_6.pdf.

Post:2012:SAG

[Pos12j] Olaf Post. Spectral Analysis on Graph-like Spaces, volume 2039 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CO-DEN LNMAA2. ISBN 3-642-23839-4 (print), 3-642-23840-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 431 pp. LCCN QA3 .L28 no. 2039. URL http://link.

springer.com/book/10.1007/978-3-642-23840-6;   http:/
/www.springerlink.com/content/978-3-642-23840-6.

Potzsche:2010:BM

[Pöt10a] Christian Pötzsche. Back matter. Lecture Notes in Mathematics, 2002:345–405, 2010. CODEN LNMAA2. ISBN 3-642-14257-5 (print), 3-642-14258-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-14258-1/1.pdf.

Potzsche:2010:FM

[Pöt10b] Christian Pötzsche. Front matter. Lecture Notes in Mathematics, 2002:i–xxiv, 2010. CODEN LNMAA2. ISBN 3-642-14257-5 (print), 3-642-14258-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-14258-1/1.pdf.

Potzsche:2010:GTD

[Pöt10c] Christian Pötzsche. Geometric Theory of Discrete Nonautonomous Dynamical Systems, volume 2002 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-14257-5 (print), 3-642-14258-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxiv + 399 pp. LCCN

QA3 .L28 no. 2002. URL http://link.springer.com/book/
10.1007/978-3-642-14258-1;   http://www.springerlink.
com/content/978-3-642-14258-1.

Potzsche:2010:IFB

[Pöt10d] Christian Pötzsche. Invariant fiber bundles. Lecture Notes in Mathematics, 2002:187–316, 2010. CODEN LNMAA2. ISBN 3-642-14257-5 (print), 3-642-14258-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-14258-1_4.pdf.

Potzsche:2010:LDE

[Pöt10e] Christian Pötzsche. Linear difference equations. Lecture Notes in Mathematics, 2002:95–185, 2010. CODEN LNMAA2. ISBN 3-642-14257-5 (print), 3-642-14258-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-14258-1_3.pdf.

Potzsche:2010:L

[Pöt10f] Christian Pötzsche. Linearization. Lecture Notes in Mathematics, 2002:317–343, 2010. CODEN LNMAA2. ISBN 3-642-14257-5 (print), 3-642-14258-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-14258-1_5.pdf.

Potzsche:2010:NDE

[Pöt10g] Christian Pötzsche. Nonautonomous difference equations. Lecture Notes in Mathematics, 2002:37–94, 2010. CODEN LNMAA2. ISBN 3-642-14257-5 (print), 3-642-14258-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-14258-1_2.
pdf.

Potzsche:2010:NDS

[Pöt10h] Christian Pötzsche. Nonautonomous dynamical systems. Lecture Notes in Mathematics, 2002:1–36, 2010. CODEN LNMAA2. ISBN 3-642-14257-5 (print), 3-642-14258-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-14258-1_1.
pdf.

Popescu-Pampu:2016:GAV

[PP16a] Patrick Popescu-Pampu. The γένoς according to Aristotle. Lecture Notes in Mathematics, 2162:1, 2016. CO-DEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_1.pdf.

Popescu-Pampu:2016:ANT

[PP16b] Patrick Popescu-Pampu. Abel and the new transcendental functions. Lecture Notes in Mathematics, 2162:21–22, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_9.pdf.

Popescu-Pampu:2016:AM

[PP16c] Patrick Popescu-Pampu. Abel’s motivations. Lecture Notes in Mathematics, 2162:25–26, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-42312-8_11.pdf.

Popescu-Pampu:2016:CIP

[PP16d] Patrick Popescu-Pampu. Cauchy and the integration paths. Lecture Notes in Mathematics, 2162:27–30, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-42312-8_12.
pdf.

Popescu-Pampu:2016:CCT

[PP16e] Patrick Popescu-Pampu. Clebsch and the choice of the term “genus”. Lecture Notes in Mathematics, 2162:59–61, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_20.pdf.

Popescu-Pampu:2016:CNH

[PP16f] Patrick Popescu-Pampu. Clifford and the number of holes. Lecture Notes in Mathematics, 2162:53–57, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-42312-8_19.
pdf.

Popescu-Pampu:2016:DNW

[PP16g] Patrick Popescu-Pampu. Descartes and the new world of curves. Lecture Notes in Mathematics, 2162:5–6, 2016. CO-DEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_2.pdf.

Popescu-Pampu:2016:EAL

[PP16h] Patrick Popescu-Pampu. Euler and the addition of lemniscatic integrals. Lecture Notes in Mathematics, 2162:17–18, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_7.pdf.

Popescu-Pampu:2016:FL

[PP16i] Patrick Popescu-Pampu. Fagnano and the lemniscate. Lecture Notes in Mathematics, 2162:15–16, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-42312-8_6.
pdf.

Popescu-Pampu:2016:FMa

[PP16j] Patrick Popescu-Pampu. Front matter. Lecture Notes in Mathematics, 2162:i–xvii, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-42312-8/1.pdf.

Popescu-Pampu:2016:FMb

[PP16k] Patrick Popescu-Pampu. Front matter. Lecture Notes in Mathematics, 2162:3, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-42312-8/1/1.pdf.

Popescu-Pampu:2016:JBC

[PP16l] Patrick Popescu-Pampu. Jakob Bernoulli and the construction of curves. Lecture Notes in Mathematics, 2162:11–13, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_5.pdf.

Popescu-Pampu:2016:JTC

[PP16m] Patrick Popescu-Pampu. Jordan and the topological classification. Lecture Notes in Mathematics, 2162:51–52, 2016. CO-DEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_18.pdf.

Popescu-Pampu:2016:LEF

[PP16n] Patrick Popescu-Pampu. Legendre and elliptic functions. Lecture Notes in Mathematics, 2162:19–20, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-42312-8_8.
pdf.

Popescu-Pampu:2016:NCC

[PP16o] Patrick Popescu-Pampu. Newton and the classification of curves. Lecture Notes in Mathematics, 2162:7–8, 2016. CO-DEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_3.pdf.

Popescu-Pampu:2016:PA

[PP16p] Patrick Popescu-Pampu. A proof by Abel. Lecture Notes in Mathematics, 2162:23–24, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-42312-8_10.pdf.

Popescu-Pampu:2016:PPR

[PP16q] Patrick Popescu-Pampu. Puiseux and the permutations of roots. Lecture Notes in Mathematics, 2162:31–33, 2016. CO-DEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_13.pdf.

Popescu-Pampu:2016:RAW

[PP16r] Patrick Popescu-Pampu. A reinterpretation of Abel’s works. Lecture Notes in Mathematics, 2162:45–49, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-42312-8_17.
pdf.

Popescu-Pampu:2016:RRT

[PP16s] Patrick Popescu-Pampu. The Riemann–Roch theorem. Lecture Notes in Mathematics, 2162:43–44, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-42312-8_16.
pdf.

Popescu-Pampu:2016:RBI

[PP16t] Patrick Popescu-Pampu. Riemann and the birational invariance of genus. Lecture Notes in Mathematics, 2162:41–42, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_15.pdf.

Popescu-Pampu:2016:RCS

[PP16u] Patrick Popescu-Pampu. Riemann and the cutting of surfaces. Lecture Notes in Mathematics, 2162:35–40, 2016. CO-DEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42312-8_14.pdf.

Popescu-Pampu:2016:WG

[PP16v] Patrick Popescu-Pampu, editor. What is the Genus?, volume 2162 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 61

(est.) pp. LCCN ???? URL http://www.springerlink.com/
content/978-3-319-42312-8.

Popescu-Pampu:2016:WIH

[PP16w] Patrick Popescu-Pampu. When integrals hide curves. Lecture Notes in Mathematics, 2162:9–10, 2016. CODEN LNMAA2. ISBN 3-319-42311-8 (print), 3-319-42312-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-42312-8_4.
pdf.

Pratelli:2011:RHV

[Pra11a] Maurizio Pratelli. A remark on the 1/H-variation of the fractional Brownian motion. Lecture Notes in Mathematics, 2006: 215–219, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_8/.

Pratelli:2011:RVF

[Pra11b] Maurizio Pratelli. A remark on the 1/H-variation of the fractional Brownian motion. Lecture Notes in Mathematics, 2006: 215–219, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_8/.

Prokaj:2013:SSC

[Pro13a] Vilmos Prokaj. Some sufficient conditions for the ergodicity of the Lévy transformation. Lecture Notes in Mathematics, 2078: 93–121, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00321-4_2/.

Protter:2013:MTF

[Pro13b] Philip Protter. A mathematical theory of financial bubbles. Lecture Notes in Mathematics, 2081:1–108, 2013. CODEN LNMAA2. ISBN 3-319-00412-3 (print), 3-319-00413-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-00413-6_1/.

Prokaj:2014:ELT

[Pro14] Vilmos Prokaj. On the exactness of the Lévy-transformation. Lecture Notes in Mathematics, 2123:395–400, 2014. CO-DEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_16.pdf.

Pages:2011:CMD

[PS11] Gilles Pagès and Afef Sellami. Convergence of multi-dimensional quantized SDE’s. Lecture Notes in Mathematics, 2006:269–307, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_11/.

Patrizio:2013:BM

[PS13a] G. Patrizio and A. Spiro. Back matter. Lecture Notes in Mathematics, 2075:321–322, 2013. CODEN LNMAA2. ISBN 3-642-36420-9 (print), 3-642-36421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-36421-1/1.pdf.

Patrizio:2013:PTM

[PS13b] G. Patrizio and A. Spiro. Pluripotential theory and Monge– Ampère foliations. Lecture Notes in Mathematics, 2075:265– 319, 2013. CODEN LNMAA2. ISBN 3-642-36420-9 (print), 3-642-36421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-36421-1_4/.

Polik:2010:BM

[PT10a] Imre Pólik and Tamás Terlaky. Back matter. Lecture Notes in Mathematics, 1989:277–289, 2010. CODEN LNMAA2. ISBN 3-642-11338-9 (print), 3-642-11339-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-642-11339-0/1.pdf.

Polik:2010:IPM

[PT10b] Imre Pólik and Tamás Terlaky. Interior point methods for nonlinear optimization. Lecture Notes in Mathematics, 1989:215–276, 2010. CODEN LNMAA2. ISBN 3-642-11338-9 (print), 3-642-11339-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11339-0_4.pdf.

Queffelec:2010:BM

[Que10a] Martine Queffélec. Back matter. Lecture Notes in Mathematics, 1294:293–351, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-11212-6/1.pdf.

Queffelec:2010:BAa

[Que10b] Martine Queffélec. The Banach algebra M(T). Lecture Notes in Mathematics, 1294:1–19, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11212-6_1.pdf.

Queffelec:2010:BAb

[Que10c] Martine Queffélec. Bijective automata. Lecture Notes in Mathematics, 1294:225–242, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11212-6_9.pdf.

Queffelec:2010:CA

[Que10d] Martine Queffélec. Compact automata. Lecture Notes in Mathematics, 1294:281–291, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11212-6_12.pdf.

Queffelec:2010:DSAb

[Que10e] Martine Queffélec. Dynamical systems arising from substitutions. Lecture Notes in Mathematics, 1294:125–160, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11212-6_5.pdf.

Queffelec:2010:DSAa

[Que10f] Martine Queffélec. Dynamical systems associated with sequences. Lecture Notes in Mathematics, 1294:87–124, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11212-6_4.pdf.

Queffelec:2010:ESD

[Que10g] Martine Queffélec. Eigenvalues of substitution dynamical systems. Lecture Notes in Mathematics, 1294:161–192, 2010. CO-DEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11212-6_6.pdf.

Queffelec:2010:FM

[Que10h] Martine Queffélec. Front matter. Lecture Notes in Mathematics, 1294:I–XV, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-11212-6/1.pdf.

Queffelec:2010:MM

[Que10i] Martine Queffélec. Matrices of measures. Lecture Notes in Mathematics, 1294:193–207, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11212-6_7.pdf.

Queffelec:2010:MRP

[Que10j] Martine Queffélec. Matrix Riesz products. Lecture Notes in Mathematics, 1294:209–224, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-11212-6_8.pdf.

Queffelec:2010:MST

[Que10k] Martine Queffélec. Maximal spectral type of general automata. Lecture Notes in Mathematics, 1294:243–264, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11212-6_10.pdf.

Queffelec:2010:SMG

[Que10l] Martine Queffélec. Spectral multiplicity of general automata. Lecture Notes in Mathematics, 1294:265–280, 2010. CO-DEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-11212-6_11.pdf.

Queffelec:2010:STD

[Que10m] Martine Queffélec. Spectral theory of dynamical systems. Lecture Notes in Mathematics, 1294:49–86, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-11212-6_3.
pdf.

Queffelec:2010:STU

[Que10n] Martine Queffélec. Spectral theory of unitary operators. Lecture Notes in Mathematics, 1294:21–48, 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-11212-6_2.
pdf.

Queffelec:2010:SDS

[Que10o] Martine Queffélec. Substitution Dynamical Systems — Spectral Analysis, volume 1294 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-11211-0 (print), 3-642-11212-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 351 pp. LCCN QA3 .L28 no. 1294. URL http:

//link.springer.com/book/10.1007/978-3-642-11212-6;
http://www.springerlink.com/content/978-3-642-11212-6.

Qian:2012:MRD

[QY12] Zhongmin Qian and Jiangang Ying. Martingale representations for diffusion processes and backward stochastic differential equations. Lecture Notes in Mathematics, 2046:75–103, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27461-9_4/.

Rabin:2011:GDP

[Rab11] Jeffrey M. Rabin. Geometry of dual pairs of complex supercurves. Lecture Notes in Mathematics, 2027:241–252, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21744-9_
11/.

Reider:2013:LD

[Rei13a] Igor Reider. J(X; L, d) and the Langlands duality. Lecture Notes in Mathematics, 2072:197–212, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35662-9_12/.

Reider:2013:Sa

[Rei13b] Igor Reider. sl_{2}-structures on F′. Lecture Notes in Mathematics, 2072:99–111, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-35662-9_6/.

Reider:2013:Sb

[Rei13c] Igor Reider. sl_{2}-structures on G_{Γ}. Lecture Notes in Mathematics, 2072:113–122, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-35662-9_7/.

Reider:2013:BM

[Rei13d] Igor Reider. Back matter. Lecture Notes in Mathematics, 2072: 213–216, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-35662-9/1.pdf.

Reider:2013:CTE

[Rei13e] Igor Reider. Configurations and theirs equations. Lecture Notes in Mathematics, 2072:145–173, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35662-9_10/.

Reider:2013:FM

[Rei13f] Igor Reider. Front matter. Lecture Notes in Mathematics, 2072: i–viii, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-35662-9/1.pdf.

Reider:2013:Ia

[Rei13g] Igor Reider. Introduction. Lecture Notes in Mathematics, 2072: 1–15, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35662-9_1/.

Reider:2013:Ib

[Rei13h] Igor Reider. Involution on G_{−}. Lecture Notes in Mathematics, 2072:123–132, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-35662-9_8/.

Reider:2013:NJM

[Rei13i] Igor Reider. Nonabelian Jacobian J(X; L, d): Main properties. Lecture Notes in Mathematics, 2072:17–32, 2013. CODEN LN-MAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-35662-9_
2/.

Reider:2013:NJP

[Rei13j] Igor Reider. Nonabelian Jacobian of Projective Surfaces: Geometry and Representation Theory, volume 2072 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 214 pp. LCCN QA3 .L28 no.

2072. URL http://link.springer.com/book/10.1007/978-
3-642-35662-9;   http://www.springerlink.com/content/
978-3-642-35662-9.

Reider:2013:PMT

[Rei13k] Igor Reider. Period maps and Torelli problems. Lecture Notes in Mathematics, 2072:75–98, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35662-9_5/.

Reider:2013:RTC

[Rei13l] Igor Reider. Representation theoretic constructions. Lecture Notes in Mathematics, 2072:175–196, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35662-9_11/.

Reider:2013:SLA

[Rei13m] Igor Reider. The sheaf of Lie algebras G_{−}. Lecture Notes in Mathematics, 2072:39–73, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-35662-9_4/.

Reider:2013:SPF

[Rei13n] Igor Reider. Some properties of the filtration h̃_{−•}. Lecture Notes in Mathematics, 2072:33–38, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35662-9_3/.

Reider:2013:Sc

[Rei13o] Igor Reider. Stratification of T_{π}. Lecture Notes in Mathematics, 2072:133–144, 2013. CODEN LNMAA2. ISBN 3-642-35661-3 (print), 3-642-35662-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-35662-9_9/.

Reynolds:2013:BOS

[Rey13a] Andy Reynolds. Beyond optimal searching: Recent developments in the modelling of animal movement patterns as Lévy walks. Lecture Notes in Mathematics, 2071:53–76, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35497-7_3/.

Reynolds:2013:FM

[Rey13b] Andy Reynolds. Front matter. Lecture Notes in Mathematics, 2071:77, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-35497-7/2/1.pdf.

Riedle:2011:CWP

[Rie11] Markus Riedle. Cylindrical Wiener processes. Lecture Notes in Mathematics, 2006:191–214, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-15217-7_7/.

Rivasseau:2012:IRG

[Riv12] Vincent Rivasseau. Introduction to the renormalization group with applications to non-relativistic quantum electron gases. Lecture Notes in Mathematics, 2051:1–54, 2012. CODEN LNMAA2. ISBN 3-642-29510-X (print), 3-642-29511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29511-9_1/.

Redenbach:2015:RTT

[RL15] Claudia Redenbach and André Liebscher. Random tessellations and their application to the modelling of cellular materials. Lecture Notes in Mathematics, 2120:73–93, 2015. CO-DEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10064-7_3.pdf.

Robertz:2014:BM

[Rob14a] Daniel Robertz. Back matter. Lecture Notes in Mathematics, 2121:233–285, 2014. CODEN LNMAA2. ISBN 3-319-11444-1 (print), 3-319-11445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-11445-3/1.pdf.

Robertz:2014:DEA

[Rob14b] Daniel Robertz. Differential elimination for analytic functions. Lecture Notes in Mathematics, 2121:119–231, 2014. CO-DEN LNMAA2. ISBN 3-319-11444-1 (print), 3-319-11445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11445-3_3.pdf.

Robertz:2014:FAE

[Rob14c] Daniel Robertz. Formal Algorithmic Elimination for PDEs, volume 2121 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-11444-1 (print), 3-319-11445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 285 (est.) pp. LCCN ???? URL http://www.

springerlink.com/content/978-3-319-11445-3.

Robertz:2014:FMP

[Rob14d] Daniel Robertz. Formal methods for PDE systems. Lecture Notes in Mathematics, 2121:5–117, 2014. CODEN LNMAA2. ISBN 3-319-11444-1 (print), 3-319-11445-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11445-3_2.pdf.

Robertz:2014:FM

[Rob14e] Daniel Robertz. Front matter. Lecture Notes in Mathematics, 2121:i–viii, 2014. CODEN LNMAA2. ISBN 3-319-11444-1 (print), 3-319-11445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-11445-3/1.pdf.

Robertz:2014:I

[Rob14f] Daniel Robertz. Introduction. Lecture Notes in Mathematics, 2121:1–4, 2014. CODEN LNMAA2. ISBN 3-319-11444-1 (print), 3-319-11445-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-11445-3_1.pdf.

Rosen:2011:SCP

[Ros11] Jay Rosen. A stochastic calculus proof of the CLT for the L^{2} modulus of continuity of local time. Lecture Notes in Mathematics, 2006:95–104, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-15217-7_3/.

Rouviere:2014:BM

[Rou14a] François Rouvière. Back matter. Lecture Notes in Mathematics, 2115:177–198, 2014. CODEN LNMAA2. ISBN 3-319-09772-5 (print), 3-319-09773-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-09773-2/1.pdf.

Rouviere:2014:CHS

[Rou14b] François Rouvière. Convolution on homogeneous spaces. Lecture Notes in Mathematics, 2115:51–56, 2014. CODEN LNMAA2. ISBN 3-319-09772-5 (print), 3-319-09773-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-09773-2_2.
pdf.

Rouviere:2014:FCH

[Rou14c] François Rouvière. e-functions and the Campbell–Hausdorff formula. Lecture Notes in Mathematics, 2115:119–175, 2014.

CODEN LNMAA2. ISBN 3-319-09772-5 (print), 3-319-09773-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09773-2_4.pdf.

Rouviere:2014:FM

[Rou14d] François Rouvière. Front matter. Lecture Notes in Mathematics, 2115:i–xxi, 2014. CODEN LNMAA2. ISBN 3-319-09772-5 (print), 3-319-09773-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-09773-2/1.pdf.

Rouviere:2014:KVM

[Rou14e] François Rouvière. The Kashiwara–Vergne method for Lie groups. Lecture Notes in Mathematics, 2115:1–49, 2014. CO-DEN LNMAA2. ISBN 3-319-09772-5 (print), 3-319-09773-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09773-2_1.pdf.

Rouviere:2014:RF

[Rou14f] François Rouvière. The role of e-functions. Lecture Notes in Mathematics, 2115:57–117, 2014. CODEN LNMAA2. ISBN 3-319-09772-5 (print), 3-319-09773-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-09773-2_3.pdf.

Rouviere:2014:SSK

[Rou14g] François Rouvière. Symmetric Spaces and the Kashiwara–Vergne Method, volume 2115 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-09772-5 (print), 3-319-09773-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 198 (est.) pp. LCCN ???? URL http://www.

springerlink.com/content/978-3-319-09773-2.

Reichmann:2010:BM

[RS10a] Oleg Reichmann and Christoph Schwab. Back matter. Lecture Notes in Mathematics, 2001:197–204, 2010. CODEN LNMAA2. ISBN 3-642-14006-8 (print), 3-642-14007-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-642-14007-5/1.pdf.

Reichmann:2010:NAA

[RS10b] Oleg Reichmann and Christoph Schwab. Numerical analysis of additive, Lévy and Feller processes with applications to option pricing. Lecture Notes in Mathematics, 2001:137–196, 2010. CODEN LNMAA2. ISBN 3-642-14006-8 (print), 3-642-14007-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14007-5_3.pdf.

Rivasseau:2012:FM

[RSSS12a] Vincent Rivasseau, Robert Seiringer, Jan Philip Solovej, and Thomas Spencer. Front matter. Lecture Notes in Mathematics, 2051:i–xiii, 2012. CODEN LNMAA2. ISBN 3-642-29510-X (print), 3-642-29511-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-29511-9/1.pdf.

Rivasseau:2012:QMB

[RSSS12b] Vincent Rivasseau, Robert Seiringer, Jan Philip Solovej, and Thomas Spencer, editors. Quantum Many Body Systems: Cetraro, Italy 2010, volume 2051 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-29510-X (print), 3-642-29511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 2051;

QC174.17.P7 Q36 2012. URL http://link.springer.com/
book/10.1007/978-3-642-29511-9;   http://www.loc.gov/
catdir/enhancements/fy1306/2012941621-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1306/2012941621-t.
html;   http://www.springerlink.com/content/978-3-642-
29511-9.

Remy:2015:BTB

[RTW15a] Bertrand Rémy, Amaury Thuillier, and Annette Werner. Bruhat– Tits buildings and analytic geometry. Lecture Notes in Mathematics, 2119:141–202, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-11029-5_5.pdf.

Remy:2015:FM

[RTW15b] Bertrand Rémy, Amaury Thuillier, and Annette Werner. Front matter. Lecture Notes in Mathematics, 2119:203, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-11029-5/
3/1.pdf.

Rudnicki:2015:SOS

[Rud15] Ryszard Rudnicki. Stochastic operators and semigroups and their applications in physics and biology. Lecture Notes in Mathematics, 2126:255–318, 2015. CODEN LNMAA2. ISBN 3-319-11321-6 (print), 3-319-11322-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-11322-7_6.pdf.

Rumpf:2013:VMI

[Rum13] Martin Rumpf. Variational methods in image matching and motion extraction. Lecture Notes in Mathematics, 2090:143– 204, 2013. CODEN LNMAA2. ISBN 3-319-01711-X (print), 3-319-01712-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01712-9_3/.

Rocircuzicka:2013:AGN

[Růž13a] Michael Růžička. Analysis of generalized Newtonian fluids. Lecture Notes in Mathematics, 2073:199–238, 2013. CODEN LN-MAA2. ISBN 3-642-36296-6 (print), 3-642-36297-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-36297-2_
4/.

Ruzicka:2013:AGN

[Růž13b] Michael Růžička. Analysis of generalized Newtonian fluids. Lecture Notes in Mathematics, 2073:199–238, 2013. CODEN LN-MAA2. ISBN 3-642-36296-6 (print), 3-642-36297-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-36297-2_
4/.

Rosenbaum:2014:LTA

[RY14] Mathieu Rosenbaum and Marc Yor. On the law of a triplet associated with the pseudo-Brownian bridge. Lecture Notes in Mathematics, 2123:359–375, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-11970-0_14.pdf.

Sabbah:2013:F

[Sab13a] Claude Sabbah. I-filtrations. Lecture Notes in Mathematics, 2060:1–19, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31695-1_1/.

Sabbah:2013:AS

[Sab13b] Claude Sabbah. Abelianity and strictness. Lecture Notes in Mathematics, 2060:39–50, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31695-1_3/.

Sabbah:2013:ARH

[Sab13c] Claude Sabbah. Applications of the Riemann–Hilbert correspondence to holonomic distributions. Lecture Notes in Mathematics, 2060:79–88, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-31695-1_6/.

Sabbah:2013:BM

[Sab13d] Claude Sabbah. Back matter. Lecture Notes in Mathematics, 2060:239–249, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-31695-1/1.pdf.

Sabbah:2013:FMa

[Sab13e] Claude Sabbah. Front matter. Lecture Notes in Mathematics, 2060:i–xiv, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-31695-1/1.pdf.

Sabbah:2013:FMb

[Sab13f] Claude Sabbah. Front matter. Lecture Notes in Mathematics, 2060:21, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-31695-1/1/1.pdf.

Sabbah:2013:FMc

[Sab13g] Claude Sabbah. Front matter. Lecture Notes in Mathematics, 2060:113, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-31695-1/2/1.pdf.

Sabbah:2013:GMCb

[Sab13h] Claude Sabbah. Good meromorphic connections (analytic theory) and the Riemann–Hilbert correspondence. Lecture Notes in Mathematics, 2060:177–193, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31695-1_12/.

Sabbah:2013:GMCa

[Sab13i] Claude Sabbah. Good meromorphic connections (formal theory). Lecture Notes in Mathematics, 2060:159–175, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31695-1_
11/.

Sabbah:2013:ISS

[Sab13j] Claude Sabbah. Introduction to Stokes Structures, volume 2060 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CO-DEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 249 pp. LCCN QA3 .L28 no. 2060. URL http://link.

springer.com/book/10.1007/978-3-642-31695-1;   http:/
/www.springerlink.com/content/978-3-642-31695-1.

Sabbah:2013:INC

[Sab13k] Claude Sabbah. Irregular nearby cycles. Lecture Notes in Mathematics, 2060:207–225, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31695-1_14/.

Sabbah:2013:NCS

[Sab13l] Claude Sabbah. Nearby cycles of Stokes-filtered local systems. Lecture Notes in Mathematics, 2060:227–238, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31695-1_
15/.

Sabbah:2013:PFS

[Sab13m] Claude Sabbah. Push-forward of Stokes-filtered local systems. Lecture Notes in Mathematics, 2060:195–206, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31695-1_
13/.

Sabbah:2013:RBS

[Sab13n] Claude Sabbah. Real blow-up spaces and moderate de Rham complexes. Lecture Notes in Mathematics, 2060:115–129, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31695-1_8/.

Sabbah:2013:RHL

[Sab13o] Claude Sabbah. Riemann–Hilbert and Laplace on the affine line (the regular case). Lecture Notes in Mathematics, 2060:89– 111, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31695-1_7/.

Sabbah:2013:RHCb

[Sab13p] Claude Sabbah. The Riemann–Hilbert correspondence for good meromorphic connections (case of a smooth divisor). Lecture Notes in Mathematics, 2060:147–157, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31695-1_10/.

Sabbah:2013:RHCa

[Sab13q] Claude Sabbah. The Riemann–Hilbert correspondence for holonomic D-modules on curves. Lecture Notes in Mathematics, 2060: 65–78, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31695-1_5/.

Sabbah:2013:SFLb

[Sab13r] Claude Sabbah. Stokes-filtered local systems along a divisor with normal crossings. Lecture Notes in Mathematics, 2060:131– 146, 2013. CODEN LNMAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31695-1_9/.

Sabbah:2013:SFLa

[Sab13s] Claude Sabbah. Stokes-filtered local systems in dimension one. Lecture Notes in Mathematics, 2060:23–38, 2013. CODEN LN-MAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31695-1_
2/.

Sabbah:2013:SPS

[Sab13t] Claude Sabbah. Stokes-perverse sheaves on Riemann surfaces. Lecture Notes in Mathematics, 2060:51–64, 2013. CODEN LN-MAA2. ISBN 3-642-31694-8 (print), 3-642-31695-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31695-1_
4/.

Sainz:2014:BM

[SAC^{+}14a] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Back matter. Lecture Notes in Mathematics, 2091:307–318, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-319-01721-1/1.pdf.

Sainz:2014:ES

[SAC^{+}14b] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Equations and systems. Lecture Notes in Mathematics, 2091:143–158, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01721-1_6/.

Sainz:2014:FM

[SAC^{+}14c] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Front matter. Lecture Notes in Mathematics, 2091:i–xvi, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-319-01721-1/1.pdf.

Sainz:2014:IO

[SAC^{+}14d] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Interpretability and optimality. Lecture Notes in Mathematics, 2091:73–120, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-01721-1_
4/.

Sainz:2014:IA

[SAC^{+}14e] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Interval arithmetic. Lecture Notes in Mathematics, 2091:121–141, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01721-1_5/.

Sainz:2014:I

[SAC^{+}14f] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Intervals. Lecture Notes in Mathematics, 2091:1–16, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-01721-1_1/.

Sainz:2014:IM

[SAC^{+}14g] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Intervals of marks. Lecture Notes in Mathematics, 2091:229–264, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01721-1_9/.

Sainz:2014:M

[SAC^{+}14h] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Marks. Lecture Notes in Mathematics, 2091:185–228, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01721-1_8/.

Sainz:2014:MIA

[SAC^{+}14i] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Modal Interval Analysis: New Tools for Numerical Information, volume 2091 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 316 pp. LCCN QA3 .L28 no.

2091. URL http://link.springer.com/book/10.1007/978-
3-319-01721-1;   http://www.springerlink.com/content/
978-3-319-01721-1.

Sainz:2014:MIE

[SAC^{+}14j] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Modal interval extensions. Lecture Notes in Mathematics, 2091:39–72, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01721-1_3/.

Sainz:2014:MI

[SAC^{+}14k] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Modal intervals. Lecture Notes in Mathematics, 2091:17–37, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01721-1_2/.

Sainz:2014:SRP

[SAC^{+}14l] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Some related problems. Lecture Notes in Mathematics, 2091:265–305, 2014. CODEN LNMAA2. ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01721-1_10/.

Sainz:2014:TA

[SAC^{+}14m] Miguel A. Sainz, Joaquim Armengol, Remei Calm, Pau Herrero, Lambert Jorba, and Josep Vehi. Twins and f^{∗} algorithm. Lecture Notes in Mathematics, 2091:159–183, 2014. CODEN LNMAA2.

ISBN 3-319-01720-9 (print), 3-319-01721-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-01721-1_7/.

Salsa:2012:BM

[Sal12a] Sandro Salsa. Back matter. Lecture Notes in Mathematics, 2045: 245–247, 2012. CODEN LNMAA2. ISBN 3-642-27144-8 (print), 3-642-27145-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-27145-8/1.pdf.

Salsa:2012:POL

[Sal12b] Sandro Salsa. The problems of the obstacle in lower dimension and for the fractional Laplacian. Lecture Notes in Mathematics, 2045:153–244, 2012. CODEN LNMAA2. ISBN 3-642-27144-8 (print), 3-642-27145-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-27145-8_4/.

Sambale:2014:ADG

[Sam14a] Benjamin Sambale. Abelian defect groups. Lecture Notes in Mathematics, 2127:205–217, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-12006-5_14.pdf.

Sambale:2014:BM

[Sam14b] Benjamin Sambale. Back matter. Lecture Notes in Mathematics, 2127:229–246, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-12006-5/1.pdf.

Sambale:2014:BG

[Sam14c] Benjamin Sambale. Bicyclic groups. Lecture Notes in Mathematics, 2127:127–157, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-12006-5_10.pdf.

Sambale:2014:BFG

[Sam14d] Benjamin Sambale. Blocks of Finite Groups and Their Invariants, volume 2127 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

246 (est.) pp. LCCN ???? URL http://www.springerlink.
com/content/978-3-319-12006-5.

Sambale:2014:BFC

[Sam14e] Benjamin Sambale. Blocks with few characters. Lecture Notes in Mathematics, 2127:219–227, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-12006-5_15.pdf.

Sambale:2014:BTF

[Sam14f] Benjamin Sambale. A bound in terms of fusion systems. Lecture Notes in Mathematics, 2127:47–61, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-12006-5_5.
pdf.

Sambale:2014:CM

[Sam14g] Benjamin Sambale. The Cartan method. Lecture Notes in Mathematics, 2127:33–46, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-12006-5_4.pdf.

Sambale:2014:DGP

[Sam14h] Benjamin Sambale. Defect groups of p-rank 2. Lecture Notes in Mathematics, 2127:159–165, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-12006-5_11.pdf.

Sambale:2014:DF

[Sam14i] Benjamin Sambale. Definitions and facts. Lecture Notes in Mathematics, 2127:3–17, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-12006-5_1.pdf.

Sambale:2014:ESA

[Sam14j] Benjamin Sambale. Essential subgroups and Alperin’s fusion theorem. Lecture Notes in Mathematics, 2127:63–70, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-12006-5_6.pdf.

Sambale:2014:FMa

[Sam14k] Benjamin Sambale. Front matter. Lecture Notes in Mathematics, 2127:i–xiii, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-12006-5/1.pdf.

Sambale:2014:FMb

[Sam14l] Benjamin Sambale. Front matter. Lecture Notes in Mathematics, 2127:1, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-12006-5/1/1.pdf.

Sambale:2014:FMc

[Sam14m] Benjamin Sambale. Front matter. Lecture Notes in Mathematics, 2127:23, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-12006-5/2/1.pdf.

Sambale:2014:FMd

[Sam14n] Benjamin Sambale. Front matter. Lecture Notes in Mathematics, 2127:79, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-12006-5/3/1.pdf.

Sambale:2014:MDG

[Sam14o] Benjamin Sambale. Metacyclic defect groups. Lecture Notes in Mathematics, 2127:81–94, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-12006-5_8.pdf.

Sambale:2014:MNA

[Sam14p]  Benjamin Sambale.  Minimal non-abelian defect groups.

Lecture Notes in Mathematics, 2127:167–179, 2014. CO-DEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-12006-5_12.pdf.

Sambale:2014:OC

[Sam14q] Benjamin Sambale. Open conjectures. Lecture Notes in Mathematics, 2127:19–22, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-12006-5_2.pdf.

Sambale:2014:PMG

[Sam14r] Benjamin Sambale. Products of metacyclic groups. Lecture Notes in Mathematics, 2127:95–125, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-12006-5_9.pdf.

Sambale:2014:QF

[Sam14s] Benjamin Sambale. Quadratic forms. Lecture Notes in Mathematics, 2127:25–32, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-12006-5_3.pdf.

Sambale:2014:RQG

[Sam14t] Benjamin Sambale. Reduction to quasisimple groups and the classification. Lecture Notes in Mathematics, 2127:71–78, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-12006-5_7.pdf.

Sambale:2014:SDG

[Sam14u] Benjamin Sambale. Small defect groups. Lecture Notes in Mathematics, 2127:181–203, 2014. CODEN LNMAA2. ISBN 3-319-12005-0 (print), 3-319-12006-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-12006-5_13.pdf.

Sawatzky:2013:ETM

[SBK^{+}13] Alex Sawatzky, Christoph Brune, Thomas Kösters, Frank Wübbeling, and Martin Burger. EM–TV methods for inverse problems with Poisson noise. Lecture Notes in Mathematics, 2090: 71–142, 2013. CODEN LNMAA2. ISBN 3-319-01711-X (print), 3-319-01712-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01712-9_2/.

Schleicher:2010:BM

[Sch10a] Dierk Schleicher. Back matter. Lecture Notes in Mathematics, 1998:341–348, 2010. CODEN LNMAA2. ISBN 3-642-13170-0 (print), 3-642-13171-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-13171-4/1.pdf.

Schleicher:2010:DEF

[Sch10b] Dierk Schleicher. Dynamics of entire functions. Lecture Notes in Mathematics, 1998:295–339, 2010. CODEN LNMAA2. ISBN 3-642-13170-0 (print), 3-642-13171-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-13171-4_5.pdf.

Schoutens:2010:AHC

[Sch10c] Hans Schoutens. Asymptotic homological conjectures in mixed characteristic. Lecture Notes in Mathematics, 1999:149–169, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-13368-8_10.pdf.

Schoutens:2010:BM

[Sch10d] Hans Schoutens. Back matter. Lecture Notes in Mathematics, 1999:171–210, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-13368-8/1.pdf.

Schoutens:2010:C

[Sch10e] Hans Schoutens. Cataproducts. Lecture Notes in Mathematics, 1999:113–125, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-13368-8_8.pdf.

Schoutens:2010:F

[Sch10f] Hans Schoutens. Flatness. Lecture Notes in Mathematics, 1999: 29–50, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-13368-8_3.pdf.

Schoutens:2010:FM

[Sch10g] Hans Schoutens. Front matter. Lecture Notes in Mathematics, 1999:i–x, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-13368-8/1.pdf.

Schoutens:2010:I

[Sch10h] Hans Schoutens. Introduction. Lecture Notes in Mathematics, 1999:1–6, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-13368-8_1.pdf.

Schoutens:2010:P

[Sch10i] Hans Schoutens. Protoproducts. Lecture Notes in Mathematics, 1999:127–148, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-13368-8_9.pdf.

Schoutens:2010:TCCa

[Sch10j] Hans Schoutens. Tight closure in characteristic zero. affine case. Lecture Notes in Mathematics, 1999:81–95, 2010. CO-DEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-13368-8_6.pdf.

Schoutens:2010:TCCb

[Sch10k] Hans Schoutens. Tight closure in characteristic zero. local case. Lecture Notes in Mathematics, 1999:97–112, 2010. CO-DEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-13368-8_7.pdf.

Schoutens:2010:TCP

[Sch10l] Hans Schoutens. Tight closure in positive characteristic. Lecture Notes in Mathematics, 1999:65–80, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-13368-8_5.
pdf.

Schoutens:2010:UT

[Sch10m] Hans Schoutens. Ultraproducts and Łoś’ theorem. Lecture Notes in Mathematics, 1999:7–27, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-13368-8_2.pdf.

Schoutens:2010:UB

[Sch10n] Hans Schoutens. Uniform bounds. Lecture Notes in Mathematics, 1999:51–63, 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-13368-8_4.pdf.

Schoutens:2010:UUC

[Sch10o] Hans Schoutens. The Use of Ultraproducts in Commutative Algebra, volume 1999 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-13367-3 (print), 3-642-13368-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 204 pp. LCCN QA3 .L28 no. 1999. URL http:

//link.springer.com/book/10.1007/978-3-642-13368-8;
http://www.springerlink.com/content/978-3-642-13368-8.

Schlichting:2011:HAT

[Sch11a] Marco Schlichting. Higher algebraic K-theory (after Quillen, Thomason and others). Lecture Notes in Mathematics, 2008: 167–241, 2011. CODEN LNMAA2. ISBN 3-642-15707-6 (print), 3-642-15708-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15708-0_4/.

Schmies:2011:CPT

[Sch11b] Markus Schmies. Computing Poincaré theta series for Schottky groups. Lecture Notes in Mathematics, 2013:165–182, 2011. CODEN LNMAA2. ISBN 3-642-17412-4 (print), 3-642-17413-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-17413-1_4/.

Schotland:2012:DRM

[Sch12] John C. Schotland. Direct reconstruction methods in optical tomography. Lecture Notes in Mathematics, 2035:1–29, 2012. CODEN LNMAA2. ISBN 3-642-22989-1 (print), 3-642-22990-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
642-22990-9_1/;   http://link.springer.com/content/pdf/
bfm:978-3-642-22990-9/1.pdf.

Schachermayer:2014:ATS

[Sch14] Walter Schachermayer. Admissible trading strategies under transaction costs. Lecture Notes in Mathematics, 2123:317–331, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_11.pdf.

Schmidt:2015:FM

[Sch15a] Volker Schmidt. Front matter. Lecture Notes in Mathematics, 2120:i–xxiv, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-10064-7/1.pdf.

Schuhmacher:2015:SMA

[Sch15b] Dominic Schuhmacher. Stein’s method for approximating complex distributions, with a view towards point processes. Lecture Notes in Mathematics, 2120:1–30, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-10064-7_1.
pdf.

Swinnerton-Dyer:2010:TDE

[SD10] Sir Peter Swinnerton-Dyer. Topics in Diophantine equations. Lecture Notes in Mathematics, 2009:45–110, 2010. CO-DEN LNMAA2. ISBN 3-642-15944-3 (print), 3-642-15945-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-15945-9_2.pdf.

Skytt:2016:MIA

[SD16] Vibeke Skytt and Tor Dokken. Models for isogeometric analysis from CAD. Lecture Notes in Mathematics, 2161:71–86, 2016. CODEN LNMAA2. ISBN 3-319-42308-8 (print), 3-319-42309-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42309-8_2.pdf.

Searle:2014:IIG

[Sea14] Catherine Searle. An introduction to isometric group actions with applications to spaces with curvature bounded from below. Lecture Notes in Mathematics, 2110:21–43, 2014. CO-DEN LNMAA2. ISBN 3-319-06372-3 (print), 3-319-06373-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-06373-7_2.pdf.

Seiringer:2012:CQG

[Sei12] Robert Seiringer. Cold quantum gases and Bose–Einstein condensation. Lecture Notes in Mathematics, 2051:55–92, 2012. CODEN LNMAA2. ISBN 3-642-29510-X (print), 3-642-29511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29511-9_2/.

Serganova:2011:BM

[Ser11a] Vera Serganova. Back matter. Lecture Notes in Mathematics, 2027:275–276, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-21744-9/1.pdf.

Serganova:2011:SIR

[Ser11b] Vera Serganova. On the superdimension of an irreducible representation of a basic classical Lie superalgebra. Lecture Notes in Mathematics, 2027:253–273, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21744-9_12/.

Seregin:2013:BM

[Ser13a] Gregory Seregin. Back matter. Lecture Notes in Mathematics, 2073:315–316, 2013. CODEN LNMAA2. ISBN 3-642-36296-6 (print), 3-642-36297-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-36297-2/1.pdf.

Seregin:2013:STL

[Ser13b] Gregory Seregin. Selected topics of local regularity theory for Navier–Stokes equations. Lecture Notes in Mathematics, 2073: 239–313, 2013. CODEN LNMAA2. ISBN 3-642-36296-6 (print), 3-642-36297-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-36297-2_5/.

Serlet:2014:IPR

[Ser14] Laurent Serlet. Invariance principle for the random walk conditioned to have few zeros. Lecture Notes in Mathematics, 2123:461–472, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-11970-0_19.pdf.

Sacerdote:2013:SIF

[SG13] Laura Sacerdote and Maria Teresa Giraudo. Stochastic integrate and fire models: a review on mathematical methods and their applications. Lecture Notes in Mathematics, 2058:99–148, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32157-3_5/.

Schmidt:2015:SMM

[SGS15] Volker Schmidt, Gerd Gaiselmann, and Ole Stenzel. Stochastic 3D models for the micro-structure of advanced functional materials. Lecture Notes in Mathematics, 2120:95–141, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-10064-7_4.pdf.

Siedow:2011:RHT

[Sie11] Norbert Siedow. Radiative heat transfer and applications for glass production processes II. Lecture Notes in Mathematics, 2010: 135–171, 2011. CODEN LNMAA2. ISBN 3-642-15966-4 (print), 3-642-15967-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15967-1_3/.

Siebert:2012:BM

[Sie12a] Kunibert G. Siebert. Back matter. Lecture Notes in Mathematics, 2040:311–314, 2012. CODEN LNMAA2. ISBN 3-642-24078-X (print), 3-642-24079-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-24079-9/1.pdf.

Siebert:2012:MFD

[Sie12b] Kunibert G. Siebert. Mathematically founded design of adaptive finite element software. Lecture Notes in Mathematics, 2040: 227–309, 2012. CODEN LNMAA2. ISBN 3-642-24078-X (print), 3-642-24079-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-24079-9_4/.

Sinestrari:2016:STD

[Sin16] Carlo Sinestrari. Singularities of three-dimensional Ricci flows. Lecture Notes in Mathematics, 2166:71–104, 2016. CO-DEN LNMAA2. ISBN 3-319-42350-9 (print), 3-319-42351-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-42351-7_3.pdf.

Soicher:2013:BM

[Soi13a] Leonard H. Soicher. Back matter. Lecture Notes in Mathematics, 2070:107, 2013. CODEN LNMAA2. ISBN 1-4471-4813-4 (print), 1-4471-4814-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-1-4471-4814-2/1.pdf.

Soicher:2013:DGC

[Soi13b] Leonard H. Soicher. Designs, groups and computing. Lecture Notes in Mathematics, 2070:83–107, 2013. CODEN LNMAA2. ISBN 1-4471-4813-4 (print), 1-4471-4814-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-1-4471-4814-2_3/.

Sokol:2013:EPF

[Sok13] Alexander Sokol. An elementary proof that the first hitting time of an open set by a jump process is a stopping time. Lecture Notes in Mathematics, 2078:301–304, 2013. CODEN LNMAA2. ISBN 3-319-00320-8 (print), 3-319-00321-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00321-4_11/.

Solovej:2012:QCG

[Sol12] Jan Philip Solovej. Quantum Coulomb gases. Lecture Notes in Mathematics, 2051:93–124, 2012. CODEN LNMAA2. ISBN 3-642-29510-X (print), 3-642-29511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29511-9_3/.

Spencer:2012:BM

[Spe12a] Thomas Spencer. Back matter. Lecture Notes in Mathematics, 2051:179–180, 2012. CODEN LNMAA2. ISBN 3-642-29510-X (print), 3-642-29511-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-29511-9/1.pdf.

Spencer:2012:SSM

[Spe12b] Thomas Spencer. SUSY statistical mechanics and random band matrices. Lecture Notes in Mathematics, 2051:125–177, 2012. CODEN LNMAA2. ISBN 3-642-29510-X (print), 3-642-29511-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-29511-9_4/.

Spodarev:2013:SGS

[Spo13] Evgeny Spodarev. Stochastic Geometry, Spatial Statistics and Random Fields: Asymptotic Methods, volume 2068 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2.

ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxiv + 446 pp. LCCN QA3 .L28 no. 2068; QA3 .L28 no. 2068; QA273.5 .S745 2013.

URL http://link.springer.com/book/10.1007/978-3-642-
33305-7;   http://www.springerlink.com/content/978-3-
642-33305-7.

Surulescu:2013:BM

[SS13a] Christina Surulescu and Nicolae Surulescu. Back matter. Lecture Notes in Mathematics, 2102:309–314, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-319-03080-7/1.pdf.

Surulescu:2013:SCS

[SS13b] Christina Surulescu and Nicolae Surulescu. Some classes of stochastic differential equations as an alternative modeling approach to biomedical problems. Lecture Notes in Mathematics, 2102:269–307, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-03080-7_9/.

Sanz-Serna:2014:MCM

[SS14] J. M. Sanz-Serna. Markov chain Monte Carlo and numerical differential equations. Lecture Notes in Mathematics, 2082:39– 88, 2014. CODEN LNMAA2. ISBN 3-319-01299-1 (print), 3-319-01300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01300-8_2/.

Shchepakina:2014:AP

[SSM14a] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Appendix: Proofs. Lecture Notes in Mathematics, 2114:183–198, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09570-7_9.pdf.

Shchepakina:2014:BM

[SSM14b] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Back matter. Lecture Notes in Mathematics, 2114:199–214, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-09570-7/1.pdf.

Shchepakina:2014:BN

[SSM14c] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. The book of numbers. Lecture Notes in Mathematics, 2114:43– 80, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09570-7_3.pdf.

Shchepakina:2014:CBS

[SSM14d] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Canards and black swans. Lecture Notes in Mathematics, 2114: 141–182, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-319-09570-7_8.pdf.

Shchepakina:2014:FM

[SSM14e] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Front matter. Lecture Notes in Mathematics, 2114:i–xiii, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-09570-7/1.pdf.

Shchepakina:2014:I

[SSM14f] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Introduction. Lecture Notes in Mathematics, 2114:1–23, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09570-7_1.pdf.

Shchepakina:2014:RMC

[SSM14g] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Reduction methods for chemical systems. Lecture Notes in Mathematics, 2114:111–117, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-09570-7_6.pdf.

Shchepakina:2014:RSI

[SSM14h] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Representations of slow integral manifolds. Lecture Notes in Mathematics, 2114:81–92, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-09570-7_4.pdf.

Shchepakina:2014:SPI

[SSM14i] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Singular Perturbations: Introduction to System Order Reduction Methods with Applications, volume 2114 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 214 (est.) pp. LCCN ???? URL

http://www.springerlink.com/content/978-3-319-09570-7.

Shchepakina:2014:SSP

[SSM14j] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Singular singularly perturbed systems. Lecture Notes in Mathematics, 2114:93–110, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-09570-7_5.pdf.

Shchepakina:2014:SIM

[SSM14k] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Slow integral manifolds. Lecture Notes in Mathematics, 2114:25– 42, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09570-7_2.pdf.

Shchepakina:2014:SC

[SSM14l] Elena Shchepakina, Vladimir Sobolev, and Michael P. Mortell. Specific cases. Lecture Notes in Mathematics, 2114:119–139, 2014. CODEN LNMAA2. ISBN 3-319-09569-2 (print), 3-319-09570-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-09570-7_7.pdf.

Spodarev:2015:ESR

[SSR15] Evgeny Spodarev, Elena Shmileva, and Stefan Roth. Extrapolation of stationary random fields. Lecture Notes in Mathematics, 2120:321–368, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-10064-7_11.pdf.

Stamov:2012:APSa

[Sta12a] Gani T. Stamov. Almost periodic solutions. Lecture Notes in Mathematics, 2047:33–96, 2012. CODEN LNMAA2. ISBN 3-642-27545-1 (print), 3-642-27546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27546-3_2/.

Stamov:2012:APSb

[Sta12b] Gani T. Stamov. Almost Periodic Solutions of Impulsive Differential Equations, volume 2047 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNMAA2. ISBN 3-642-27545-1 (print), 3-642-27546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 217 pp. LCCN QA3 .L28 no. 2047.

URL http://link.springer.com/book/10.1007/978-3-642-
27546-3;   http://www.springerlink.com/content/978-3-
642-27546-3.

Stamov:2012:A

[Sta12c] Gani T. Stamov. Applications. Lecture Notes in Mathematics, 2047:151–203, 2012. CODEN LNMAA2. ISBN 3-642-27545-1 (print), 3-642-27546-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-27546-3_4/.

Stamov:2012:BM

[Sta12d] Gani T. Stamov. Back matter. Lecture Notes in Mathematics, 2047:205–217, 2012. CODEN LNMAA2. ISBN 3-642-27545-1 (print), 3-642-27546-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-27546-3/1.pdf.

Stamov:2012:FM

[Sta12e] Gani T. Stamov. Front matter. Lecture Notes in Mathematics, 2047:i–xx, 2012. CODEN LNMAA2. ISBN 3-642-27545-1 (print), 3-642-27546-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-27546-3/1.pdf.

Stamov:2012:IDE

[Sta12f] Gani T. Stamov. Impulsive differential equations and almost periodicity. Lecture Notes in Mathematics, 2047:1–32, 2012. CODEN LNMAA2. ISBN 3-642-27545-1 (print), 3-642-27546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27546-3_1/.

Stamov:2012:LMA

[Sta12g] Gani T. Stamov. Lyapunov method and almost periodicity. Lecture Notes in Mathematics, 2047:97–149, 2012. CODEN LN-MAA2. ISBN 3-642-27545-1 (print), 3-642-27546-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27546-3_
3/.

Stadtmuller:2013:SLT

[Sta13] Ulrich Stadtmüller. Strong limit theorems for increments of random fields. Lecture Notes in Mathematics, 2068:385–398, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33305-7_11/.

Stix:2013:BM

[Sti13a] Jakob Stix. Back matter. Lecture Notes in Mathematics, 2054: 233–249, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-30674-7/1.pdf.

Stix:2013:BGO

[Sti13b] Jakob Stix. Basic geometric operations in terms of sections. Lecture Notes in Mathematics, 2054:25–36, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30674-7_3/.

Stix:2013:BMD

[Sti13c] Jakob Stix. Brauer–Manin and descent obstructions. Lecture Notes in Mathematics, 2054:119–146, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30674-7_11/.

Stix:2013:CNA

[Sti13d] Jakob Stix. Continuous non-abelian H^{1} with profinite coefficients. Lecture Notes in Mathematics, 2054:3–11, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30674-7_
1/.

Stix:2013:CSB

[Sti13e] Jakob Stix. Cuspidal sections and birational analogues. Lecture Notes in Mathematics, 2054:219–231, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30674-7_18/.

Stix:2013:CCA

[Sti13f] Jakob Stix. Cycle classes in anabelian geometry. Lecture Notes in Mathematics, 2054:53–66, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30674-7_6/.

Stix:2013:EU

[Sti13g] Jakob Stix. Evaluation of units. Lecture Notes in Mathematics, 2054:45–51, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30674-7_5/.

Stix:2013:FCD

[Sti13h] Jakob Stix. Fields of cohomological dimension 1. Lecture Notes in Mathematics, 2054:213–218, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30674-7_17/.

Stix:2013:FNA

[Sti13i] Jakob Stix. Fragments of non-abelian Tate–Poitou duality. Lecture Notes in Mathematics, 2054:147–154, 2013. CODEN LN-MAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30674-7_
12/.

Stix:2013:FMa

[Sti13j] Jakob Stix. Front matter. Lecture Notes in Mathematics, 2054: i–xx, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-30674-7/1.pdf.

Stix:2013:FMb

[Sti13k] Jakob Stix. Front matter. Lecture Notes in Mathematics, 2054: 1, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-30674-7/1/1.pdf.

Stix:2013:FMc

[Sti13l] Jakob Stix. Front matter. Lecture Notes in Mathematics, 2054: 67, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-30674-7/2/1.pdf.

Stix:2013:FMd

[Sti13m] Jakob Stix. Front matter. Lecture Notes in Mathematics, 2054: 105, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-30674-7/3/1.pdf.

Stix:2013:FMe

[Sti13n] Jakob Stix. Front matter. Lecture Notes in Mathematics, 2054: 155, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-30674-7/4/1.pdf.

Stix:2013:FG

[Sti13o] Jakob Stix. The fundamental groupoid. Lecture Notes in Mathematics, 2054:13–23, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30674-7_2/.

Stix:2013:ISC

[Sti13p] Jakob Stix. Injectivity in the section conjecture. Lecture Notes in Mathematics, 2054:69–79, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30674-7_7/.

Stix:2013:LOA

[Sti13q] Jakob Stix. Local obstructions at a p-adic place. Lecture Notes in Mathematics, 2054:107–117, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30674-7_10/.

Stix:2013:NS

[Sti13r] Jakob Stix. Nilpotent sections. Lecture Notes in Mathematics, 2054:175–196, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30674-7_14/.

Stix:2013:SCT

[Sti13s] Jakob Stix. On the section conjecture for torsors. Lecture Notes in Mathematics, 2054:157–174, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30674-7_13/.

Stix:2013:SCL

[Sti13t] Jakob Stix. On the section conjecture over local fields. Lecture Notes in Mathematics, 2054:207–212, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30674-7_16/.

Stix:2013:RPA

[Sti13u] Jakob Stix. Rational Points and Arithmetic of Fundamental Groups: Evidence for the Section Conjecture, volume 2054 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CO-DEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 249 pp. LCCN QA3 .L28 no. 2054. URL http://link.

springer.com/book/10.1007/978-3-642-30674-7;   http:/
/www.springerlink.com/content/978-3-642-30674-7.

Stix:2013:RS

[Sti13v] Jakob Stix. Reduction of sections. Lecture Notes in Mathematics, 2054:81–93, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30674-7_8/.

Stix:2013:SFF

[Sti13w] Jakob Stix. Sections over finite fields. Lecture Notes in Mathematics, 2054:197–205, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-30674-7_15/.

Stix:2013:SST

[Sti13x] Jakob Stix. The space of sections as a topological space. Lecture Notes in Mathematics, 2054:37–44, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30674-7_4/.

Stix:2013:SSA

[Sti13y] Jakob Stix. The space of sections in the arithmetic case and the section conjecture in covers. Lecture Notes in Mathematics, 2054: 95–103, 2013. CODEN LNMAA2. ISBN 3-642-30673-X (print), 3-642-30674-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-30674-7_9/.

Song:2013:IKR

[SW13] Jian Song and Ben Weinkove. An introduction to the Kähler– Ricci flow. Lecture Notes in Mathematics, 2086:89–188, 2013. CODEN LNMAA2. ISBN 3-319-00818-8 (print), 3-319-00819-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
319-00819-6_3/.

Stadtmuller:2015:IFD

[SZ15] Ulrich Stadtmüller and Marta Zampiceni. An introduction to functional data analysis. Lecture Notes in Mathematics, 2120: 257–292, 2015. CODEN LNMAA2. ISBN 3-319-10063-7 (print), 3-319-10064-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-319-10064-7_9.pdf.

Tankov:2011:BM

[Tan11a] Peter Tankov. Back matter. Lecture Notes in Mathematics, 2003: 361–366, 2011. CODEN LNMAA2. ISBN 3-642-14659-7 (print), 3-642-14660-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-14660-2/1.pdf.

Tankov:2011:PHE

[Tan11b] Peter Tankov. Pricing and hedging in exponential Lévy models: Review of recent results. Lecture Notes in Mathematics, 2003: 319–359, 2011. CODEN LNMAA2. ISBN 3-642-14659-7 (print), 3-642-14660-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-14660-2_5/.

Thomaseth:2013:BM

[TBBF13a] Karl Thomaseth, Jerry J. Batzel, Mostafa Bachar, and Raffaello Furlan. Back matter. Lecture Notes in Mathematics, 2064:247– 254, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-32882-4/1.pdf.

Thomaseth:2013:PEM

[TBBF13b] Karl Thomaseth, Jerry J. Batzel, Mostafa Bachar, and Raffaello Furlan. Parameter estimation of a model for baroreflex control of unstressed volume. Lecture Notes in Mathematics, 2064:215– 246, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32882-4_11/.

Temkin:2015:IBA

[Tem15] Michael Temkin. Introduction to Berkovich analytic spaces. Lecture Notes in Mathematics, 2119:3–66, 2015. CODEN LNMAA2. ISBN 3-319-11028-4 (print), 3-319-11029-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-11029-5_1.
pdf.

Thieullen:2013:DSF

[Thi13] Michèle Thieullen. Deterministic and stochastic FitzHugh– Nagumo systems. Lecture Notes in Mathematics, 2058:175– 186, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32157-3_7/.

Tian:2016:BM

[Tia16a] Gang Tian. Back matter. Lecture Notes in Mathematics, 2166: 137–138, 2016. CODEN LNMAA2. ISBN 3-319-42350-9 (print), 3-319-42351-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-42351-7/1.pdf.

Tian:2016:NKR

[Tia16b] Gang Tian. Notes on Kähler–Ricci flow. Lecture Notes in Mathematics, 2166:105–136, 2016. CODEN LNMAA2. ISBN 3-319-42350-9 (print), 3-319-42351-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-42351-7_4.pdf.

Tian:2016:RFG

[Tia16c] Gang Tian, editor. Ricci Flow and Geometric Applications: Cetraro, Italy 2010, volume 2166 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. CODEN LNMAA2. ISBN 3-319-42350-9 (print), 3-319-42351-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 138 (est.) pp. LCCN ???? URL http://www.

springerlink.com/content/978-3-319-42351-7.

Bertrand:2011:BM

[Toë11a] Bertrand Toën. Back matter. Lecture Notes in Mathematics, 2008:303–308, 2011. CODEN LNMAA2. ISBN 3-642-15707-6 (print), 3-642-15708-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-15708-0/1.pdf.

Toen:2011:LDC

[Toë11b] Bertrand Toën. Lectures on DG-categories. Lecture Notes in Mathematics, 2008:243–302, 2011. CODEN LNMAA2. ISBN 3-642-15707-6 (print), 3-642-15708-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-15708-0_5/.

Tin:2013:IRA

[TP13] Chung Tin and Chi-Sang Poon. Integrative and reductionist approaches to modeling of control of breathing. Lecture Notes in Mathematics, 2064:89–103, 2013. CODEN LNMAA2. ISBN 3-642-32881-4 (print), 3-642-32882-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32882-4_5/.

Trevisan:2014:SPX

[Tre14a] Dario Trevisan, editor. Séminaire de Probabilités XLVI, volume 2123 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 479

(est.) pp. LCCN ???? URL http://www.springerlink.com/
content/978-3-319-11970-0.

Trevisan:2014:SPS

[Tre14b] Dario Trevisan. A short proof of Stein’s universal multiplier theorem. Lecture Notes in Mathematics, 2123:473–479, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_20.pdf.

Tuckwell:2013:BM

[Tuc13a] Henry C. Tuckwell. Back matter. Lecture Notes in Mathematics, 2058:201–206, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-32157-3/1.pdf.

Tuckwell:2013:SMS

[Tuc13b] Henry C. Tuckwell. Stochastic modeling of spreading cortical depression. Lecture Notes in Mathematics, 2058:187–200, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32157-3_8/.

Tuckwell:2013:SPD

[Tuc13c] Henry C. Tuckwell. Stochastic partial differential equations in neurobiology: Linear and nonlinear models for spiking neurons. Lecture Notes in Mathematics, 2058:149–173, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32157-3_
6/.

Tudor:2011:ACT

[Tud11] Ciprian A. Tudor. Asymptotic Cramér’s theorem and analysis on Wiener space. Lecture Notes in Mathematics, 2006:309– 325, 2011. CODEN LNMAA2. ISBN 3-642-15216-3 (print), 3-642-15217-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-15217-7_12/.

Vuolle-Apiala:2012:TIP

[VA12] Juha Vuolle-Apiala. Time inversion property for rotation invariant self-similar diffusion processes. Lecture Notes in Mathematics, 2046:271–277, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-27461-9_13/.

VanTuyl:2013:BGE

[Van13a] Adam Van Tuyl. A beginner’s guide to edge and cover ideals. Lecture Notes in Mathematics, 2083:63–94, 2013. CODEN LNMAA2.

ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-38742-5_3/.

VanTuyl:2013:EIU

[Van13b] Adam Van Tuyl. Edge ideals using Macaulay2. Lecture Notes in Mathematics, 2083:95–105, 2013. CODEN LNMAA2. ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-38742-5_4/.

VanTuyl:2013:FM

[Van13c] Adam Van Tuyl. Front matter. Lecture Notes in Mathematics, 2083:107, 2013. CODEN LNMAA2. ISBN 3-642-38741-1 (print), 3-642-38742-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-38742-5/3/1.pdf.

Varadarajan:2011:I

[Var11] V. S. Varadarajan. Introduction. Lecture Notes in Mathematics, 2027:1–15, 2011. CODEN LNMAA2. ISBN 3-642-21743-5 (print), 3-642-21744-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21744-9_1/;   http://link.springer.
com/content/pdf/bfm:978-3-642-21744-9/1.pdf.

vandenDries:2014:LMT

[vdD14] Lou van den Dries. Lectures on the model theory of valued fields. Lecture Notes in Mathematics, 2111:55–157, 2014. CODEN LNMAA2. ISBN 3-642-54935-7 (print), 3-642-54936-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-54936-6_4.pdf.

vandenDries:2014:FM

[vdDKM^{+}14] Lou van den Dries, Jochen Koenigsmann, H. Dugald Macpherson, Anand Pillay, Carlo Toffalori, and Alex J. Wilkie. Front matter. Lecture Notes in Mathematics, 2111:i–vii, 2014. CODEN LNMAA2. ISBN 3-642-54935-7 (print), 3-642-54936-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-642-54936-6/
1.pdf.

vandeGeer:2016:ALE

[vdG16a] Sara van de Geer. Asymptotically linear estimators of the precision matrix. Lecture Notes in Mathematics, 2159:215–221, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-32774-7_14.pdf.

vandeGeer:2016:BM

[vdG16b] Sara van de Geer. Back matter. Lecture Notes in Mathematics, 2159:267–276, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-32774-7/1.pdf.

vandeGeer:2016:BLW

[vdG16c] Sara van de Geer. The bias of the lasso and worst possible sub-directions. Lecture Notes in Mathematics, 2159:41–60, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-32774-7_4.pdf.

vandeGeer:2016:BFP

[vdG16d] Sara van de Geer. Brouwer’s fixed point theorem and sparsity. Lecture Notes in Mathematics, 2159:199–214, 2016. CO-DEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-32774-7_13.pdf.

vandeGeer:2016:CIC

[vdG16e] Sara van de Geer. Chaining including concentration. Lecture Notes in Mathematics, 2159:239–253, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-32774-7_17.
pdf.

vandeGeer:2016:CIU

[vdG16f] Sara van de Geer. Confidence intervals using the lasso. Lecture Notes in Mathematics, 2159:61–74, 2016. CODEN LNMAA2.

ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-32774-7_5.
pdf.

vandeGeer:2016:EPT

[vdG16g] Sara van de Geer. Empirical process theory for dual norms. Lecture Notes in Mathematics, 2159:121–131, 2016. CO-DEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-32774-7_8.pdf.

vandeGeer:2016:ETU

[vdG16h] Sara van de Geer, editor. Estimation and Testing Under Sparsity: École d’Été de Probabilités de Saint-Flour XLV – 2015, volume 2159 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 276

(est.) pp. LCCN ???? URL http://www.springerlink.com/
content/978-3-319-32774-7.

vandeGeer:2016:FM

[vdG16i] Sara van de Geer. Front matter. Lecture Notes in Mathematics, 2159:i–xiii, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-32774-7/1.pdf.

vandeGeer:2016:GLN

[vdG16j] Sara van de Geer. General loss with norm-penalty. Lecture Notes in Mathematics, 2159:103–119, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-32774-7_7.pdf.

vandeGeer:2016:ICE

[vdG16k] Sara van de Geer. Inequalities for the centred empirical risk and its derivative. Lecture Notes in Mathematics, 2159:139–150, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-32774-7_10.pdf.

vandeGeer:2016:I

[vdG16l] Sara van de Geer. Introduction. Lecture Notes in Mathematics, 2159:1–4, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-32774-7_1.pdf.

vandeGeer:2016:L

[vdG16m] Sara van de Geer. The lasso. Lecture Notes in Mathematics, 2159:5–25, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-32774-7_2.pdf.

vandeGeer:2016:LBS

[vdG16n] Sara van de Geer. Lower bounds for sparse quadratic forms. Lecture Notes in Mathematics, 2159:223–231, 2016. CO-DEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-32774-7_15.pdf.

vandeGeer:2016:MC

[vdG16o] Sara van de Geer. The margin condition. Lecture Notes in Mathematics, 2159:151–165, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-32774-7_11.pdf.

vandeGeer:2016:MSC

[vdG16p] Sara van de Geer. Metric structure of convex hulls. Lecture Notes in Mathematics, 2159:255–266, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-32774-7_18.pdf.

vandeGeer:2016:PIM

[vdG16q] Sara van de Geer. Probability inequalities for matrices. Lecture Notes in Mathematics, 2159:133–137, 2016. CO-DEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-32774-7_9.pdf.

vandeGeer:2016:SWE

[vdG16r] Sara van de Geer. Some worked-out examples. Lecture Notes in Mathematics, 2159:167–197, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-32774-7_12.pdf.

vandeGeer:2016:SRL

[vdG16s] Sara van de Geer. The square-root lasso. Lecture Notes in Mathematics, 2159:27–39, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-319-32774-7_3.pdf.

vandeGeer:2016:SS

[vdG16t] Sara van de Geer. Structured sparsity. Lecture Notes in Mathematics, 2159:75–101, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-32774-7_6.pdf.

vandeGeer:2016:SCC

[vdG16u] Sara van de Geer. Symmetrization, contraction and concentration. Lecture Notes in Mathematics, 2159:233–238, 2016. CODEN LNMAA2. ISBN 3-319-32773-9 (print), 3-319-32774-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-32774-7_16.pdf.

Veselic:2011:DOL

[Ves11a] Krešimir Veselić. Damped Oscillations of Linear Systems: A Mathematical Introduction, volume 2023 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 209 pp. LCCN QA3 .L28 no.

2023. URL http://link.springer.com/book/10.1007/978-
3-642-21335-9;   http://www.springerlink.com/content/
978-3-642-21335-9.

Veselic:2011:DS

[Ves11b] Krešimir Veselić. Definite spectra. Lecture Notes in Mathematics, 2023:73–88, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21335-9_10/.

Veselic:2011:FM

[Ves11c] Krešimir Veselić. Front matter. Lecture Notes in Mathematics, 2023:i–xv, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-21335-9/1.pdf.

Veselic:2011:GHM

[Ves11d] Krešimir Veselić. General Hermitian matrix pairs. Lecture Notes in Mathematics, 2023:89–92, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21335-9_11/.

Veselic:2011:IM

[Ves11e] Krešimir Veselić. ‘indefinite metric’. Lecture Notes in Mathematics, 2023:39–48, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21335-9_5/.

Veselic:2011:OPb

[Ves11f] Krešimir Veselić. J-orthogonal projections. Lecture Notes in Mathematics, 2023:61–65, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21335-9_8/.

Veselic:2011:MIS

[Ves11g] Krešimir Veselić. Matrices and indefinite scalar products. Lecture Notes in Mathematics, 2023:49–54, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-21335-9_6/.

Veselic:2011:ME

[Ves11h] Krešimir Veselić. The matrix exponential. Lecture Notes in Mathematics, 2023:113–120, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21335-9_13/.

Veselic:2011:MA

[Ves11i] Krešimir Veselić. Modal approximation. Lecture Notes in Mathematics, 2023:145–157, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21335-9_19/.

Veselic:2011:MAO

[Ves11j] Krešimir Veselić. Modal approximation and overdampedness. Lecture Notes in Mathematics, 2023:159–166, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21335-9_
20/.

Veselic:2011:M

[Ves11k] Krešimir Veselić. The model. Lecture Notes in Mathematics, 2023:1–13, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21335-9_1/.

Veselic:2011:OPa

[Ves11l] Krešimir Veselić. Oblique projections. Lecture Notes in Mathematics, 2023:55–60, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21335-9_7/.

Veselic:2011:PS

[Ves11m] Krešimir Veselić. Phase space. Lecture Notes in Mathematics, 2023:23–28, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21335-9_3/.

Veselic:2011:QEP

[Ves11n] Krešimir Veselić. The quadratic eigenvalue problem. Lecture Notes in Mathematics, 2023:121–127, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-21335-9_14/.

Veselic:2011:RR

[Ves11o] Krešimir Veselić. Resonances and resolvents. Lecture Notes in Mathematics, 2023:139–141, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21335-9_17/.

Veselic:2011:SEI

[Ves11p] Krešimir Veselić. Simple eigenvalue inclusions. Lecture Notes in Mathematics, 2023:129–134, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-21335-9_15/.

Veselic:2011:SDM

[Ves11q] Krešimir Veselić. Simultaneous diagonalisation (modal damping). Lecture Notes in Mathematics, 2023:15–22, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-21335-9_
2/.

Veselic:2011:SMC

[Ves11r] Krešimir Veselić. The singular mass case. Lecture Notes in Mathematics, 2023:29–37, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21335-9_4/.

Veselic:2011:SDG

[Ves11s] Krešimir Veselić. Spectral decomposition of a general J-Hermitian matrix. Lecture Notes in Mathematics, 2023:93–111, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21335-9_12/.

Veselic:2011:SPR

[Ves11t] Krešimir Veselić. Spectral properties and reduction of J-Hermitian matrices. Lecture Notes in Mathematics, 2023:67– 71, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21335-9_9/.

Veselic:2011:SS

[Ves11u] Krešimir Veselić. Spectral shift. Lecture Notes in Mathematics, 2023:135–138, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-21335-9_16/.

Veselic:2011:WP

[Ves11v] Krešimir Veselić. Well-posedness. Lecture Notes in Mathematics, 2023:143, 2011. CODEN LNMAA2. ISBN 3-642-21334-0 (print), 3-642-21335-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-21335-9_18.

Voss:2013:RTC

[VGS13] Florian Voss, Catherine Gloaguen, and Volker Schmidt. Random tessellations and Cox processes. Lecture Notes in Mathematics, 2068:151–182, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-642-33305-7_5/.

vanHandel:2014:EDP

[vH14] Ramon van Handel. Ergodicity, decisions, and partial information. Lecture Notes in Mathematics, 2123:411–459, 2014. CODEN LNMAA2. ISBN 3-319-11969-9 (print), 3-319-11970-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-11970-0_18.pdf.

Viviani:2014:BM

[Viv14a] Filippo Viviani. Back matter. Lecture Notes in Mathematics, 2108:241–242, 2014. CODEN LNMAA2. ISBN 3-319-04869-4 (print), 3-319-04870-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-04870-3/1.pdf.

Viviani:2014:CAG

[Viv14b] Filippo Viviani, editor. Combinatorial Algebraic Geometry: Levico Terme, Italy 2013, Editors: Sandra Di Rocco, Bernd Sturmfels, volume 2108 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-04869-4 (print), 3-319-04870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-319-04870-3.

Viviani:2014:THS

[Viv14c] Filippo Viviani. A tour on Hermitian symmetric manifolds. Lecture Notes in Mathematics, 2108:149–239, 2014. CO-DEN LNMAA2. ISBN 3-319-04869-4 (print), 3-319-04870-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-04870-3_5.pdf.

Vojta:2010:BM

[Voj10a] Paul Vojta. Back matter. Lecture Notes in Mathematics, 2009: 225–232, 2010. CODEN LNMAA2. ISBN 3-642-15944-3 (print), 3-642-15945-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-642-15945-9/1.pdf.

Vojta:2010:DAN

[Voj10b] Paul Vojta. Diophantine approximation and Nevanlinna theory. Lecture Notes in Mathematics, 2009:111–224, 2010. CO-DEN LNMAA2. ISBN 3-642-15944-3 (print), 3-642-15945-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-15945-9_3.pdf.

Volpert:2013:EPP

[VV13] Vitaly Volpert and Vitali Vougalter. Emergence and propagation of patterns in nonlocal reaction-diffusion equations arising in the theory of speciation. Lecture Notes in Mathematics, 2071: 331–353, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35497-7_12/.

Wainrib:2013:BIL

[Wai13a] Gilles Wainrib. A brief introduction to large deviations theory. Lecture Notes in Mathematics, 2058:57–72, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32157-3_
3/.

Wainrib:2013:FM

[Wai13b] Gilles Wainrib. Front matter. Lecture Notes in Mathematics, 2058:97, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-642-32157-3/2/1.pdf.

Wainrib:2013:SNM

[Wai13c] Gilles Wainrib. Some numerical methods for rare events simulation and analysis. Lecture Notes in Mathematics, 2058:73– 95, 2013. CODEN LNMAA2. ISBN 3-642-32156-9 (print), 3-642-32157-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32157-3_4/.

Wainwright:2015:GMM

[Wai15] Martin J. Wainwright. Graphical models and message-passing algorithms: Some introductory lectures. Lecture Notes in Mathematics, 2141:51–108, 2015. CODEN LNMAA2. ISBN 3-319-16966-1 (print), 3-319-16967-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-16967-5_3.pdf.

Weingart:2014:BM

[Wei14a] Gregor Weingart. Back matter. Lecture Notes in Mathematics, 2110:197–198, 2014. CODEN LNMAA2. ISBN 3-319-06372-3 (print), 3-319-06373-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-06373-7/1.pdf.

Weingart:2014:GMN

[Wei14b] Gregor Weingart. Geometry of Manifolds with Non-negative Sectional Curvature: Editors: Rafael Herrera, Luis Hernández-Lamoneda, volume 2110 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-06372-3 (print), 3-319-06373-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 198 (est.) pp. LCCN ???? URL http://www.

springerlink.com/content/978-3-319-06373-7.

Weingart:2014:IED

[Wei14c] Gregor Weingart. An introduction to exterior differential systems. Lecture Notes in Mathematics, 2110:117–196, 2014. CO-DEN LNMAA2. ISBN 3-319-06372-3 (print), 3-319-06373-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-06373-7_6.pdf.

Weiss:2014:BM

[Wei14d] Christian Weiß. Back matter. Lecture Notes in Mathematics, 2104:145–168, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-319-04075-2/1.pdf.

Weiss:2014:B

[Wei14e] Christian Weiß. Background. Lecture Notes in Mathematics, 2104:11–37, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-04075-2_2/.

Weiss:2014:CTT

[Wei14f] Christian Weiß. Calculations for twisted Teichmüller curves. Lecture Notes in Mathematics, 2104:85–119, 2014. CODEN LN-MAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-04075-2_
6/.

Weiss:2014:FM

[Wei14g] Christian Weiß. Front matter. Lecture Notes in Mathematics, 2104:i–xvi, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-04075-2/1.pdf.

Weiss:2014:I

[Wei14h] Christian Weiß. Introduction. Lecture Notes in Mathematics, 2104:1–10, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-04075-2_1/.

Weiss:2014:KCR

[Wei14i] Christian Weiß. Kobayashi curves revisited. Lecture Notes in Mathematics, 2104:135–144, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04075-2_9/.

Weiss:2014:LE

[Wei14j] Christian Weiß. Lyapunov exponents. Lecture Notes in Mathematics, 2104:127–133, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04075-2_8/.

Weiss:2014:PVT

[Wei14k] Christian Weiß. Prym varieties and Teichmüller curves. Lecture Notes in Mathematics, 2104:121–125, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-319-04075-2_7/.

Weiss:2014:SM

[Wei14l] Christian Weiß. Stabilizer and maximality. Lecture Notes in Mathematics, 2104:61–84, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04075-2_5/.

Weiss:2014:TC

[Wei14m] Christian Weiß. Teichmüller curves. Lecture Notes in Mathematics, 2104:39–51, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-04075-2_3/.

Weiss:2014:TTCa

[Wei14n] Christian Weiß. Twisted Teichmüller curves. Lecture Notes in Mathematics, 2104:53–59, 2014. CODEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-04075-2_4/.

Weiss:2014:TTCb

[Wei14o] Christian Weiß. Twisted Teichmüller Curves, volume 2104 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CO-DEN LNMAA2. ISBN 3-319-04074-X (print), 3-319-04075-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 166 pp. LCCN QA3 .L28 no. 2104. URL http://link.

springer.com/book/10.1007/978-3-319-04075-2;   http:/
/www.springerlink.com/content/978-3-319-04075-2.

Wilkie:2014:LMT

[Wil14] A. J. Wilkie. Lectures on the model theory of real and complex exponentiation. Lecture Notes in Mathematics, 2111:35–53, 2014. CODEN LNMAA2. ISBN 3-642-54935-7 (print), 3-642-54936-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-54936-6_3.pdf.

Witzel:2014:BM

[Wit14a] Stefan Witzel. Back matter. Lecture Notes in Mathematics, 2109: 99–116, 2014. CODEN LNMAA2. ISBN 3-319-06476-2 (print), 3-319-06477-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-319-06477-2/1.pdf.

Witzel:2014:BDP

[Wit14b] Stefan Witzel. Basic definitions and properties. Lecture Notes in Mathematics, 2109:1–44, 2014. CODEN LNMAA2. ISBN 3-319-06476-2 (print), 3-319-06477-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-06477-2_1.pdf.

Witzel:2014:FPb

[Wit14c] Stefan Witzel. Finiteness properties of G(F_{q}[t]). Lecture Notes in Mathematics, 2109:45–79, 2014. CODEN LNMAA2. ISBN 3-319-06476-2 (print), 3-319-06477-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-319-06477-2_2.pdf.

Witzel:2014:FPc

[Wit14d] Stefan Witzel. Finiteness properties of G(F_{q}[t, t^{−1}]). Lecture Notes in Mathematics, 2109:81–97, 2014. CODEN LNMAA2. ISBN 3-319-06476-2 (print), 3-319-06477-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-319-06477-2_3.
pdf.

Witzel:2014:FPA

[Wit14e] Stefan Witzel. Finiteness Properties of Arithmetic Groups Acting on Twin Buildings, volume 2109 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. CODEN LNMAA2. ISBN 3-319-06476-2 (print), 3-319-06477-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 116 (est.) pp. LCCN ???? URL http://www.

springerlink.com/content/978-3-319-06477-2.

Witzel:2014:FM

[Wit14f] Stefan Witzel. Front matter. Lecture Notes in Mathematics, 2109: i–xvi, 2014. CODEN LNMAA2. ISBN 3-319-06476-2 (print), 3-319-06477-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-319-06477-2/1.pdf.

Wei:2013:IMR

[WL13] Hsin-Hua Wei and Frithjof Lutscher. From individual movement rules to population level patterns: The case of centralplace foragers. Lecture Notes in Mathematics, 2071:159–175, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35497-7_6/.

Wechselberger:2013:CTE

[WMR13a] Martin Wechselberger, John Mitry, and John Rinzel. Canard theory and excitability. Lecture Notes in Mathematics, 2102:89– 132, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-03080-7_3/.

Wechselberger:2013:FM

[WMR13b] Martin Wechselberger, John Mitry, and John Rinzel. Front matter. Lecture Notes in Mathematics, 2102:133, 2013. CODEN LNMAA2. ISBN 3-319-03079-5 (print), 3-319-03080-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-03080-7/
2/1.pdf.

Wu:2012:OTB

[Wu12] Hao Wu. On the occupation times of Brownian excursions and Brownian loops. Lecture Notes in Mathematics, 2046:149– 166, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27461-9_7/.

Yserentant:2010:BM

[Yse10a] Harry Yserentant. Back matter. Lecture Notes in Mathematics, 2000:177–188, 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-12248-4/1.pdf.

Yserentant:2010:BQM

[Yse10b] Harry Yserentant. The basics of quantum mechanics. Lecture Notes in Mathematics, 2000:27–50, 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-12248-4_3.
pdf.

Yserentant:2010:CRC

[Yse10c] Harry Yserentant. Convergence rates and complexity bounds. Lecture Notes in Mathematics, 2000:127–140, 2010. CO-DEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12248-4_8.pdf.

Yserentant:2010:EE

[Yse10d] Harry Yserentant. Eigenfunction expansions. Lecture Notes in Mathematics, 2000:117–125, 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12248-4_7.pdf.

Yserentant:2010:ESE

[Yse10e] Harry Yserentant. The electronic schrödinger equation. Lecture Notes in Mathematics, 2000:51–58, 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-12248-4_4.
pdf.

Yserentant:2010:EDM

[Yse10f] Harry Yserentant. Existence and decay of mixed derivatives. Lecture Notes in Mathematics, 2000:87–116, 2010. CO-DEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-12248-4_6.pdf.

Yserentant:2010:FA

[Yse10g] Harry Yserentant. Fourier analysis. Lecture Notes in Mathematics, 2000:13–26, 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-12248-4_2.pdf.

Yserentant:2010:FM

[Yse10h] Harry Yserentant. Front matter. Lecture Notes in Mathematics, 2000:i–viii, 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-642-12248-4/1.pdf.

Yserentant:2010:IO

[Yse10i] Harry Yserentant. Introduction and outline. Lecture Notes in Mathematics, 2000:1–11, 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/10.1007/978-3-642-12248-4_1.pdf.

Yserentant:2010:RAD

[Yse10j] Harry Yserentant. The radial-angular decomposition. Lecture Notes in Mathematics, 2000:141–176, 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-12248-4_9.
pdf.

Yserentant:2010:RAE

[Yse10k] Harry Yserentant. Regularity and Approximability of Electronic Wave Functions, volume 2000 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 182 pp. LCCN QA3 .L28 no.

2000. URL http://link.springer.com/book/10.1007/978-
3-642-12248-4;   http://www.springerlink.com/content/
978-3-642-12248-4.

Yserentant:2010:SED

[Yse10l] Harry Yserentant. Spectrum and exponential decay. Lecture Notes in Mathematics, 2000:59–85, 2010. CODEN LNMAA2. ISBN 3-642-12247-7 (print), 3-642-12248-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-12248-4_5.
pdf.

Yuan:2010:ADO

[YSY10a] Wen Yuan, Winfried Sickel, and Dachun Yang. Almost diagonal operators and atomic and molecular decompositions. Lecture Notes in Mathematics, 2005:49–64, 2010. CODEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-14606-0_3.
pdf.

Yuan:2010:BM

[YSY10b] Wen Yuan, Winfried Sickel, and Dachun Yang. Back matter. Lecture Notes in Mathematics, 2005:271–288, 2010. CODEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-642-14606-0/
1.pdf.

Yuan:2010:FM

[YSY10c] Wen Yuan, Winfried Sickel, and Dachun Yang. Front matter. Lecture Notes in Mathematics, 2005:i–xi, 2010. CODEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-642-14606-0/1.pdf.

Yuan:2010:HS

[YSY10d] Wen Yuan, Winfried Sickel, and Dachun Yang. Homogeneous spaces. Lecture Notes in Mathematics, 2005:251–269, 2010. CODEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14606-0_8.pdf.

Yuan:2010:IBH

[YSY10e] Wen Yuan, Winfried Sickel, and Dachun Yang. Inhomogeneous Besov–Hausdorff and Triebel–Lizorkin–Hausdorff spaces. Lecture Notes in Mathematics, 2005:177–250, 2010. CO-DEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14606-0_7.pdf.

Yuan:2010:I

[YSY10f] Wen Yuan, Winfried Sickel, and Dachun Yang. Introduction. Lecture Notes in Mathematics, 2005:1–19, 2010. CO-DEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14606-0_1.pdf.

Yuan:2010:KT

[YSY10g] Wen Yuan, Winfried Sickel, and Dachun Yang. Key theorems. Lecture Notes in Mathematics, 2005:147–175, 2010. CODEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14606-0_6.pdf.

Yuan:2010:MCM

[YSY10h] Wen Yuan, Winfried Sickel, and Dachun Yang. Morrey and Campanato Meet Besov, Lizorkin and Triebel, volume 2005 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 281 pp. LCCN

QA3 .L28 no. 2005. URL http://link.springer.com/book/
10.1007/978-3-642-14606-0;   http://www.springerlink.
com/content/978-3-642-14606-0.

Yuan:2010:PDO

[YSY10i] Wen Yuan, Winfried Sickel, and Dachun Yang. Pseudodifferential operators. Lecture Notes in Mathematics, 2005:137– 146, 2010. CODEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14606-0_5.pdf.

Yuan:2010:SEC

[YSY10j] Wen Yuan, Winfried Sickel, and Dachun Yang. Several equivalent characterizations. Lecture Notes in Mathematics, 2005:65–135, 2010. CODEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14606-0_4.pdf.

Yuan:2010:S

[YSY10k] Wen Yuan, Winfried Sickel, and Dachun Yang. The spaces B_{p,q}^{s,τ} (R^{n}) and F_{p,q}^{s,τ} (R^{n}). Lecture Notes in Mathematics, 2005:21– 48, 2010. CODEN LNMAA2. ISBN 3-642-14605-8 (print), 3-642-14606-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-642-14606-0_2.pdf.

Yukich:2013:LTD

[Yuk13] Joseph Yukich. Limit theorems in discrete stochastic geometry. Lecture Notes in Mathematics, 2068:239–275, 2013. CODEN LNMAA2. ISBN 3-642-33304-4 (print), 3-642-33305-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33305-7_
8/.

Yen:2013:BM

[YY13a] Ju-Yi Yen and Marc Yor. Back matter. Lecture Notes in Mathematics, 2088:133–138, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bbm:978-3-319-01270-4/1.pdf.

Yen:2013:BET

[YY13b] Ju-Yi Yen and Marc Yor. Brownian excursion theory: a first approach. Lecture Notes in Mathematics, 2088:57–64, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01270-4_5/.

Yen:2013:ERS

[YY13c] Ju-Yi Yen and Marc Yor. The existence and regularity of semimartingale local times. Lecture Notes in Mathematics, 2088:13– 28, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01270-4_2/.

Yen:2013:FKF

[YY13d] Ju-Yi Yen and Marc Yor. The Feynman–Kac formula and excursion theory. Lecture Notes in Mathematics, 2088:107–110, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01270-4_10/.

Yen:2013:FMa

[YY13e] Ju-Yi Yen and Marc Yor. Front matter. Lecture Notes in Mathematics, 2088:i–ix, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-01270-4/1.pdf.

Yen:2013:FMb

[YY13f] Ju-Yi Yen and Marc Yor. Front matter. Lecture Notes in Mathematics, 2088:11, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-01270-4/1/1.pdf.

Yen:2013:FMc

[YY13g] Ju-Yi Yen and Marc Yor. Front matter. Lecture Notes in Mathematics, 2088:55, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-01270-4/2/1.pdf.

Yen:2013:FMd

[YY13h] Ju-Yi Yen and Marc Yor. Front matter. Lecture Notes in Mathematics, 2088:105, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-319-01270-4/3/1.pdf.

Yen:2013:IRR

[YY13i] Ju-Yi Yen and Marc Yor. Integral representations relating W and n. Lecture Notes in Mathematics, 2088:101–104, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01270-4_9/.

Yen:2013:LCR

[YY13j] Ju-Yi Yen and Marc Yor. The laws of, and conditioning with respect to, last passage times. Lecture Notes in Mathematics, 2088: 93–100, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01270-4_8/.

Yen:2013:LRR

[YY13k] Ju-Yi Yen and Marc Yor. Lévy’s representation of reflecting BM and Pitman’s representation of BES(3). Lecture Notes in Mathematics, 2088:29–41, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-01270-4_3/.

Yen:2013:LTE

[YY13l] Ju-Yi Yen and Marc Yor. Local Times and Excursion Theory for Brownian Motion: A Tale of Wiener and Itô Measures, volume 2088 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 135 pp. LCCN QA3 .L28 no. 2088. URL http://link.

springer.com/book/10.1007/978-3-319-01270-4;   http:/
/www.springerlink.com/content/978-3-319-01270-4.

Yen:2013:PLA

[YY13m] Ju-Yi Yen and Marc Yor. Paul Lévy’s arcsine laws. Lecture Notes in Mathematics, 2088:43–54, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01270-4_4/.

Yen:2013:P

[YY13n] Ju-Yi Yen and Marc Yor. Prerequisites. Lecture Notes in Mathematics, 2088:1–10, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-01270-4_1/.

Yen:2013:SPD

[YY13o] Ju-Yi Yen and Marc Yor. A simple path decomposition of Brownian motion around time t = 1. Lecture Notes in Mathematics, 2088:79–92, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-319-01270-4_7/.

Yen:2013:SIL

[YY13p] Ju-Yi Yen and Marc Yor. Some identities in law. Lecture Notes in Mathematics, 2088:111–131, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-01270-4_11/.

Yen:2013:TDI

[YY13q] Ju-Yi Yen and Marc Yor. Two descriptions of n: Itô’s and Williams’. Lecture Notes in Mathematics, 2088:65–77, 2013. CODEN LNMAA2. ISBN 3-319-01269-X (print), 3-319-01270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-01270-4_6/.

Yang:2013:AI

[YYH13a] Dachun Yang, Dongyong Yang, and Guoen Hu. Approximations of the identity. Lecture Notes in Mathematics, 2084:23– 58, 2013. CODEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00825-7_2/.

Yang:2013:BM

[YYH13b] Dachun Yang, Dongyong Yang, and Guoen Hu. Back matter. Lecture Notes in Mathematics, 2084:643–656, 2013. CODEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-319-00825-7/
1.pdf.

Yang:2013:BOa

[YYH13c] Dachun Yang, Dongyong Yang, and Guoen Hu. Boundedness of operators over (R^{D}, μ). Lecture Notes in Mathematics, 2084: 215–328, 2013. CODEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00825-7_5/.

Yang:2013:BOb

[YYH13d] Dachun Yang, Dongyong Yang, and Guoen Hu. Boundedness of operators over (X , ν). Lecture Notes in Mathematics, 2084: 483–642, 2013. CODEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00825-7_8/.

Yang:2013:FMa

[YYH13e] Dachun Yang, Dongyong Yang, and Guoen Hu. Front matter. Lecture Notes in Mathematics, 2084:i–xiii, 2013. CODEN LN-MAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-00825-7/
1.pdf.

Yang:2013:FMb

[YYH13f] Dachun Yang, Dongyong Yang, and Guoen Hu. Front matter. Lecture Notes in Mathematics, 2084:1–3, 2013. CODEN LN-MAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-00825-7/
1/1.pdf.

Yang:2013:FMc

[YYH13g] Dachun Yang, Dongyong Yang, and Guoen Hu. Front matter. Lecture Notes in Mathematics, 2084:413–415, 2013. CODEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-319-00825-7/
2/1.pdf.

Yang:2013:HSN

[YYH13h] Dachun Yang, Dongyong Yang, and Guoen Hu. The Hardy Space H^{1} with Non-doubling Measures and Their Applications, volume 2084 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. CO-DEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 653 pp. LCCN QA3 .L28 no. 2084; QA331.7 .Y36 2013. URL http:

//link.springer.com/book/10.1007/978-3-319-00825-7;
http://www.springerlink.com/content/978-3-319-00825-7.

Yang:2013:HSD

[YYH13i] Dachun Yang, Dongyong Yang, and Guoen Hu. The Hardy space H^{1}(X , ν) and its dual space RBMO(X , ν). Lecture Notes in Mathematics, 2084:417–481, 2013. CODEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-319-00825-7_7/.

Yang:2013:HS

[YYH13j] Dachun Yang, Dongyong Yang, and Guoen Hu. The Hardy space H^{1}(μ). Lecture Notes in Mathematics, 2084:59–136, 2013. CODEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00825-7_3/.

Yang:2013:LPO

[YYH13k] Dachun Yang, Dongyong Yang, and Guoen Hu. Littlewood– Paley operators and maximal operators related to approximations of the identity. Lecture Notes in Mathematics, 2084:329– 412, 2013. CODEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00825-7_6/.

Yang:2013:LAH

[YYH13l] Dachun Yang, Dongyong Yang, and Guoen Hu. The local atomic Hardy space h^{1}(μ). Lecture Notes in Mathematics, 2084:137– 214, 2013. CODEN LNMAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-319-00825-7_4/.

Yang:2013:P

[YYH13m] Dachun Yang, Dongyong Yang, and Guoen Hu. Preliminaries. Lecture Notes in Mathematics, 2084:5–22, 2013. CODEN LN-MAA2. ISBN 3-319-00824-2 (print), 3-319-00825-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-319-00825-7_
1/.

Ziller:2014:RMP

[Zil14] Wolfgang Ziller. Riemannian manifolds with positive sectional curvature. Lecture Notes in Mathematics, 2110:1–19, 2014. CODEN LNMAA2. ISBN 3-319-06372-3 (print), 3-319-06373-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-319-06373-7_1.pdf.

Zhou:2013:LMM

[ZK13] Ying Zhou and Mark Kot. Life on the move: Modeling the effects of climate-driven range shifts with integrodifference equations. Lecture Notes in Mathematics, 2071:263–292, 2013. CODEN LNMAA2. ISBN 3-642-35496-3 (print), 3-642-35497-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35497-7_9/.

Zuazua:2013:BM

[Zua13a] Enrique Zuazua. Back matter. Lecture Notes in Mathematics, 2062:495–497, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-642-32160-3/1.pdf.

Zuazua:2013:CSW

[Zua13b] Enrique Zuazua. Control and stabilization of waves on 1-D networks. Lecture Notes in Mathematics, 2062:463–493, 2013. CODEN LNMAA2. ISBN 3-642-32159-3 (print), 3-642-32160-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32160-3_9/.